A Complete Bibliography of Lecture Notes in Mathematics (1960–1969)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

13 October 2017 Version 1.11

Title word cross-reference

(Γ_{4} = 0) [Cer68g]. (L_{p}(G), L_{q}(G))(1 ≤ p, q ≤ ∞) [Lar69f]. (ω_{1}, ω) [Wei68]. 0 [Yor68d]. 1 [Cer68f, Ton65i]. 2 [Cer68f, Cer68f]. 21 [Lün69p]. 3 [Bar69a, Hes69, Lün65a]. 4 [Lün69m]. = [Dol66f]. [A, ΩX] [AC64d, AC67b]. m,2 [Gri69]. _{0} [Ros68a]. A [Lam66a, Lin69a, Nør69, Wlo69d, Wlo69e, Wlo69g, Har67c]. A^{⊥} [Wer64f, Wer64h]. ℵ [Ros68a]. B [Jen67b]. mod1 [Cha66i]. C(Z_{2}) [Con67m]. C^{n} [Hen68, Sic68, Wer68]. C_{k} [FB66d, FB66a]. χ_{o} [Ros68b]. D(∗) [Lum68e]. D^{1}(U, I) [Alb68b]. D^{2} [HM68m]. d′′ [Pol68, Pol68]. ẋ [ST69g]. ẋ = g(x) [ST69g]. E [Gro68j, Gro68k, Lum68e]. ϵ [SD69]. F [Con67o]. f(M) [Fer68a, Fer68a]. F K [Wil67b]. G [HM68d, HM68i, HM68n, Jän68a, Jän68f, Jän68h, KT68b, KT68d, KT68h, Lün65f, Ton65b, Ton65c, Ton65d]. G(R) [Swa68k]. G_{O}(R) [Swa68j, Swa68p]. G_{O}(R_{S}) [Swa68p]. Γ_{4} [Cer68c]. GF(q) [Lün65a, Lün65a]. H [Bec69b, Bor67b, DHS69]. H(R) [Swa68d]. H^{∞} [Stu69d]. H^{m} [Dol68a]. H^{P} (λ) [Wer64e]. H_{0}^{∞} [Stu69g, Stu69d]. H_{m} [Fic65j].

IR^{n} [Jän68j]. j [Her66a]. K [Buc69b, CF66d, CF66b, FW69, Hsi69, Kar69, MT69, Pio69, Swa68e, Swa68s, Bor64i]. K′(ϕ(z)) [Shi68d]. K_{0} [Swa68n]. K_{1} [Swa68n]. K_{1}(A) [Swa68a]. K_{2} [Bas69]. K_{2}(R) [Swa68b]. K_{i} [Swa68m]. K_{O} [Swa68c]. K_{O}(A) [Swa68l]. K_{O}(R) [Swa68j]. L [Kap67a, Shi68j]. L^{E} [Gro68j, Gro68j]. L^{p} [BDCK67]. L^{X} [Gro68j]. L_{1} [Wer66m, Wer66m]. L_{1}(G) [Lar69h]. L_{22} [Lün69o]. L_{γ} [Wer66d]. L_{ω_{1},ω} [Mal68]. L_{p}(G) [Lar69c]. L_{p}(Ĝ) [Lar69e]. M [Fer68b]. M∗ [Her65h, Her65h, Her65h]. M^{∗} [Her65i]. M^{φ} [Her65l]. M_{γ} [Lum68f]. R [Cer68a]. n [Lic69, vF69]. n = 1 [Pim69i]. N_{E} [Gro68k, Gro68k]. ♮ [AM69a]. O [BH69k]. O(n) [HM68a, HM68a, HM68m]. O_{∗}(H; F, F) [Con67e]. O_{∗}^{U}(Z_{2}) [Con67j]. Ω [Gri69, Tak68a]. Ω_{∗}^{SU}(BU(1)) [May69f]. Ω_{∗}^{SU}(X) [May69i]. Ω_{8n+4}^{Spin}(BSO(2k + 1)) [May69g]. P (X) [Wer64c, Wer64d]. Π_{1}^{1} [Tak68a]. P W^{m,2}(Ω, p) [Stu69e]. Q [Nor68]. q^{2} [Lün65b]. R(Z_{2}) [Con67n]. R^{3} [Cer68b, Cer68f, Cer68f]. R^{n} [Fer68a]. rmIR^{n} [Jän68j]. S [Wlo69e, Her65h]. S(m) [Wlo69c, Wlo69f]. S(q) [Lün65a, Lün65b, Lün65a, Lün65b]. S^{2} [Cer68c, Cer68d]. S^{2n−1} [Tho69]. S^{n} [HM68h]. Spec(R) [Swa68d]. SU [May69l, May69l, May69e]. SU(3) [Sim68a]. t [Dol66a, Her65h, Lün69d, Lün69i, Lün69k]. T [f] [Wer66q]. τ(f: M → R^{n}) [Fer68b, Fer68a, Fer68b]. U [CF66c]. U(k) [May69e, May69e]. V [Stu69c, Zar69c, Zar69g]. V/k(dim(V) = r ≥ 1) [Zar69f]. V = L [Jen67q].

[Vdash] [Her65k]. W [Wlo69e, Wlo69i]. W (BU(1)) [May69d]. W (X) [May69c, May69j]. W^{2n−1}(d) [HM68g]. W^{m,2}(Ω) [Gri69]. W^{m,2}(Ω; p) [Stu69e]. K̃_{0}(R) [Swa68i]. ZF [Jen67d, Jen67h]. ZT [Lün65k].

-adic [Shi68j]. -Approximation [Wer66m]. -block [Lün69d, Lün69k]. -Blockplänen [Lün69d, Lün69k]. -Bündels [May69e]. -categorical [Ros68b]. -Cohomologie [Pol68]. -cohomology [Pol68]. -comprehension [Tak68a]. -connected [Lic69]. -dimensional [Lün65a, vF69]. -dimensionalen [Lün65a]. -ellipticity [Stu69c]. -Elliptizität [Stu69c]. -entropy [SD69]. -groups [Lün65k]. -Gruppen [Lün65k]. -Homogene [Lün69i]. -Homogeneous [Lün69i]. -invariants [Her66a]. -isomorphism [AM69a]. -manifold [Cer68f, May69e]. -manifolds [CF66c, HM68d, HM68a, HM68i, HM68m, Jän68a, Jän68f, Jän68h, May69l, Ton65b]. -Mannigfaltigkeit [May69e]. -Mannigfaltigkeiten [HM68a, HM68m, HM68d, HM68i, HM68n, Jän68a, Jän68f, Jän68h, May69l]. -mappings [FB66d, FB66a]. -Modelle [Jen67h]. -modules [Har67c]. -Norm [Wer66d, Wer66d]. -objects [Ton65c]. -parameter [Ton65i]. -point [Lün69p]. -polytopes [Bar69a]. -pseudoconvex [Nor68]. -pseudoconvexes [Nor68]. -punkte [Lün69p]. -Räume [Wlo69d, Wlo69c, Wlo69i, Wlo69g, Wlo69f]. -Räumen [Wlo69e]. -rules [Tak68a]. -spaces [Bec69b, Bor67b, DHS69, Kap67a, Ton65d, Wlo69e]. -spheres [Cer68f, Cer68f]. -stability [Lin69a]. -stable [Nør69]. -step [Pio69]. -structure [KT68b, KT68h]. -structures [KT68d]. -thèorie [Kar69]. -Theories [CF66d]. -Theory [MT69, Buc69b, CF66b, FW69, Hsi69, Kar69, Swa68e]. -valued [Jen67b]. -variété [Cer68f]. -wertige [Jen67b].

/MI [CK69].

15th [AL69]. 1954 [Bor67d]. 1961 [Ada66h, Ada69g, HG67]. 1963 [Har66i]. 1963/64 [Har66i]. 1964 [Alb68d, Bor64e, Ivă65k, Wer66v]. 1965 [Fic65k, yMO68]. 1966 [Eck69, IR66f, Tar67j, Urb67b]. 1966/67 [Eck69]. 1967 [DE67, DE68, Köl68d, Tie69c]. 1968 [AL69, BKW69, CK69, DE69, Lum68a]. 1969 [Mor69d].

2nd [Fic65a].

58 [BCH^{+}66].

64 [Har66i]. 67 [Eck69]. 68 [Lel68b].

= [Dol66f, ST69g].

Abbildungen [FW68a, FW68o, HM68k, FW68i, FW68l, FW68m, FW68q, GKM68d]. Abelian [BP69a, BP69b, Ded69, Hoo69, Oor66c, Shi68a, Shi68f]. Abelsche [BP69b, BP69a]. abgeschlossene [FW68q]. Abhängigkeit [Wer66r]. Ableitungsbegriffs [Köl68g]. Abschnitt [Ber68i]. Absolute [Par69a, Fer68f, Fer68c, BKKN67a]. Absoluteness [Jen67a]. Absolutheit [Jen67a]. Absolutkrümmung [Fer68c, Fer68f]. Abstract [Lum68c, Kap67a]. abstraits [Lum68c]. AC [Zal68a]. according [Mey68c]. accroissement [Mor69c]. accroissements [Mey69a]. accuracy [Jan69l, Lin69a]. Action [Con68b]. actions [Jän68j, Jän68l, Tho69]. Acyclic [Ulm69a, DMO67]. Adams [Nør69]. Adding [Tie69a]. Addison [Jen67k]. Addition [Mül66a, Mül66b, BP66a, BP69b]. Additions [CMW68]. Additive [Gro68e, Cer68a, Del69c, Gro68f, BP66b]. Additives [Gro68e, Del69c, Dol68b, Gro68f]. adhesion [HM68b]. adic [Shi68j]. Adjoint [Dub68]. Adjointness [Lin69e]. Adjunction [Kel69]. admissible [FJ68, FB66c]. Advanced [BCH^{+}66, Löb68, Lum68a]. affin [HM68e]. affine [HM68e, Lün69f, Lün65c]. affinen [Lün69f]. after [Jän68i, Roo68]. Aktionen [Jän68j, Jän68l]. aléatoires [Del69a, Del69b, Del71, Fer67a, Fer67b, Hub69a, Mor69a, Mor69c, Mor69b]. Algebra [Ada64a, Ada66a, And67c, Qui67c, Ada64h, Ada66i, Ada69b, Ada69h, And67a, And67b, And68, Bor64b, Bor64a, Bor64f, Gro68b, Ham69, Ivă65b, Lum68b, SG65, Sim68k, Ulm69c, Wal69, Zal68g]. Algebraic [Ati69, Bor64e, MT69, Oor66a, Swa68e, Zar69a, BB69a, Buc69b, FW69, HM68e, Hsi69, Jan69j, Kar68, Law68, Mau67b, Nor68, Shi68i, Swa68s, Zar69j, Bor64j, EG69c]. algebraically [Zar69j]. algebraische [Jan69j]. algebraischer [HM68e]. Algebras [CS69c, Knu69, Bar69b, Bar69c, Cha69a, Cha69b, CS69b, Geo67, Lam66b, Lar69d, Lin69c, Man69b, Pre68, Ric69, Sim68e, Wer64c, Wer64j, Wer69b, Lum68a, Lum68j]. Algèbre [And67a, And67b, And67c, Bor64b, Bor64a, SG65, Bor64f]. Algebren [Geo67, Wer64c, Wer64j]. Algèbres [Lum68b, Lum68a, Lum68j]. algébrique [Bor64e, Mau67b]. Algébriques [EG69c, Bor64j, Nor68]. Allgemeine [Jan69a, Stu69a]. allied [Pad69]. Almost [BH67b, BH67a, May69k, May69m, Pyk69, Yos68]. almost-complex [May69k, May69m]. Alternating [McK69]. alternative [Wal69]. Analogues [Beh69]. Analyse [Lel68b]. Analysis [Kap67b, MH69, TA69, Wil67h, Fen69, Loy67, Tro69d]. Analytic [Cha66n, Nar66c, Sto66a, Sto66d, Zal68c, Zal68b, BKKN67c, BKKN67e, Blo68, Hof69, LG69, Nar66d, Nar66g, Nov68, Ram68, Ser68, Sic68, Sto66c, Tit67b, Tre69, Wer64a, Wer66g, Ram68]. Analytical [Luk67, Tit67c]. analytique [Blo68, Ram68]. analytiques [Ram68, Ser68]. Analytische [Cha66n, Wer64a, Tit67b, Tit67c]. analytischen [BKKN67c, BKKN67e]. analytischer [Wer66g]. André [Ulm69c, Ulm69b]. Anfangswertproblems [Jan69f]. Anhang [GKM68a]. Année [Lel68b]. annelés [Gro68i]. Ansatz [Wer66d]. antérieurs [Mey69b]. Anwendung [Jan69d, Sie68, Dol66a, Jan69b]. Anwendungen [BKKN67f, Dol66b]. Anzahl [Fie67]. Appendice [Mey67a]. Appendix [GKM68a, Mey67a]. Application [Gro68a, Gro68b, Har66a, Har67a, Ivă65a, Mét67, Bor64b, Del69c, Jan69d, Kni68, Roo68, Sie68, Cer68f, Dol66a, Gro68a, Gro68b, Gro68i, Jan69b]. Applications [Ada64a, Ada66a, Ada69b, BS67c, Bor64d, Bor64c, CF69a, Con67a, Gro68c, Hil69b, Hil69c, Ivă65b, Ivă65c, Ivă65d, Ivă65k, Löb69, MT69, Mül66b, Pyk69, TA69, Tro69a, Zal68d, Zal68e, Ada64c, Ada66c, Ada69c, Dol68a, Erd69, Har67f, Mey68h, Swa68o, dR67b, Bor64d, Bor64c, Dol66b, Gro68c, Roo69]. Applied [Lin69b]. approach [Wal69]. approximating [Wer66d]. Approximation [Cha66h, Dah69a, Jan69f, PP66a, Wer68, Wer66q, Wer66r, Wer66j, Wer66v, Zal68c, Ada64h, Ada66i, Ada69h, Ber68h, Fic65s, PP66b, PP66c, Wer69b, Wer69a, Wer66f, Wer66g, Wer66e, Wer66m, Wer66s, Wer66t, Zal68m, Zal68n, Ber68b, Ber68a, Wer66a, Wer66e, Wer66m, Wer66t]. Approximationen [Wer66g]. Approximationsproblem [Wer66f]. Approximationsproblems [Wer66s]. Approximationsprozesse [Ber68b]. Approximationsprozessen [Ber68h]. Approximationstheorie [Wer66v]. Approximierenden [Wer66d]. April [BKW69]. Äquivariantes [HM68b]. arbitrary [Lyc69, Zar69g]. ARF [HM68l]. ARFsche [HM68l]. arguments [Law69a]. arising [Loy67]. arithemetic [Zar69b]. arithmetic [Cha66c, Fen69, Ken67, Tro69e, Zar69h]. arithmetischen [Cha66c]. Aronszajn [Fic65t]. Artinian [Fai67i]. asintotica [ST69q]. aspect [Hub69a]. Associated [Mül66c, BM68b, Exb68, KT68b, KT68h, Mey67a, Mey68a, Mey68d].

Association [Löb68]. associé [Mey68d]. associées [Mey68a]. associés [BM68b, Exb68, Mey67a]. Asymptotic [KK69, Kat68a, Yor68a, Wer66g, ST69q, Wer66b]. asymptotische [Wer66g, Wer66b]. Atiyah [Con67b]. attainability [Alb68a, Alb68b]. Attraction [BH69a, ST69a]. Attrazione [ST69a]. Aufgaben [Jan69g]. August [AL69]. Ausarbeitung [Bau66g]. Ausnahmeorbits [Jän68g]. Aussagen [Jan69a]. Auswahlaxioms [Jen67m, Jen67p]. Auswahloperator [Her65e, Her65m, Her65q, Her65f]. auteur [Lum68a]. author [Lum68a]. automoma [ST69e]. Automorphic [HC68a, Shi68b, Shi68c]. automorphism [Lün69k]. Automorphismengruppen [Lün69k]. automorphismes [dR67c]. automorphisms [Kue68, Ton65e, dR67c]. autonoma [ST69g]. Autonomous [Tie69b, ST69e, ST69g]. aux [BM68b, Bor64c, Cer68b, Del69c, Dol68a, Gro68a, Gro68b, Gro68i, Kis68, Sch68, dR67b]. Avancée [Lum68a]. average [KK69]. axiom [Jen67p, Jen67m]. Axiomatic [Qui67a, Bau67].

Back [Ada64b, Ada66b, AC64a, Bau66a, Ber68c, Cha66a, GKM68b, Her65a, Jan69c, Köl68a, Mey67g, Ser65a]. background [Her68a, Mül66g]. Baire [FW68e]. Balayage [Bau66b]. Balayage-theorie [Bau66b]. Baltimore [Fic65k]. Banach [Ber68b, Ber68h, BM68a, FW68o, Jan69f, Pim69g, Ram68, Wer69b]. banachique [Ram68]. Banachraum [Jan69f]. Banachräumen [Ber68b, Ber68h]. Barden [Ker67]. barwise [Kar68]. base [Dol66g]. based [Csö69]. Bashforth [Nør69]. Basic [BH69j, Swa68f, Bor64g, FW68t, Jän68e, Köl68b, Wlo69b]. Basis [Köh68, Wlo69h]. Battelle [Hil69b, Hil69c, Löb69]. Baummannigfaltigkeiten [HM68f]. be [Mat68]. Beck [Dus69]. Behandlung [Ber68h]. behavior [Wer66g]. behaviour [KK69]. Beispiele [Jän68l, Dol66c, Jän68a, Wer66c, Wer66i, Wer66n]. belonging [Lün65i, Lün65j]. Bereichen [Cha66g]. Bericht [Jän68d, HM68c, Jän68b]. Berkeley [Ada66h, Ada69g, Tho66h, Tho66h]. Berlin [Ivă65k]. Beschränkte [FW68a, FW68b, Stu69b, Wer64a]. Bestimmung [Lün65e, May69i]. between [EK69, Fer68b, Fer68a, GKM68c, Her65t, Lam66i, May69e, May69n, Shi68l, Swa68j, Swa68p, Swa68s, Tar67i]. Beurling [Kur68]. Bezeichnungen [Bau66i, Köl68b, May69o, Tit67a]. Beziehungen [Fer68b, Fer68a, GKM68c]. BH [Wil67c]. bicategories [Bén67]. bifurcation [Pim69h, Pim69i, Pim69a]. Bifurcations [Pim69b]. bilinear [Gri69, Stu69b, Stu69k, Stu69c]. Bilinearformen [Gri69, Stu69b, Stu69k, Stu69c]. binary [Aig69]. binomiaux [CF69g]. bipolar [FW68g]. Bipolarensatz [FW68g]. Bivalent [IR66f, IR66a]. block [Lün69d, Lün69k]. Blockplänen [Lün69d, Lün69k]. Bochner [Bad67]. Boolean [BB69a, Geo67, Ivă65h, Ivă65j, Ivă65k, Ivă65l, IR66c, IR66d, IR66e, IR66f, Jen67f]. Booleschen [Geo67]. Booleschwertige [Jen67f]. bordism [Con67g, May69b]. Bordismusgruppen [May69b]. Borel [Par69b]. bornological [FW68a]. bornologische [FW68a]. Bott [Con67b]. bound [Sto66c]. Boundaries [yMO68, CF66c]. Boundary [Alb68a, Hei69a, Fic65m, Fic65n, Fic65o, Fic65u, Fuj68, Jan69b, Mae68a, Mat68, Mey68i, Mey68j, Pim69k, Stu69a, Stu69e, Stu69i, Tan68, Ura69, Gir68, Gri69, Jan69k, Stu69h]. Bounded [Hof69, Kou67, Tar67g]. bounds [Coo69]. Bourlet [Tar67f, Tar67i]. Branches [Pim69b, Pim69a, Pim69h, Pim69f, Pim69g]. Brouwer [Tro69b]. Buffalo [Tro69q]. Bündel [Fer68g]. Bündels [May69e]. bundle [Bor67c]. Bundles [Bor67d, Bor67e, Con68c, Eps69, KT68a, KT68c, KT68k, Tho66f]. Burkolder [dGM69]. BVP [Fic65a, Fic65c, Fic65b, Fic65p].

C [Lün69h]. C. [Lün69h]. calculi [Ros68a]. Calculus [FB66b, BKKN67c, FB66i, Her65b, Her65d, Slo68]. California [Ada66h, Ada69g]. Canada [BKW69]. Cancellation [Swa68g]. canonical [Zar69c]. Cantelli [Csö69]. Capacity [Zal68c, AW69, Zal68b, Zal68a]. caractéristique [Gro68k]. Carleman [Tar67h, Tar67k, Tar67n]. carré [Mey67a]. Cartesian [BP69f, Law69a]. Cas [Gro68d, Ser65c, Ser65i, Lum68f]. case [Pim69i, Ser65c, Gro68d, Lum68f]. cases [FB66h, Köl68f]. Categorial [Her68a]. Categorical [Eck69, Gra69, Ros68b, Tie69c]. catégorie [Gro68d]. Categories [AT69, Fit69, Gro68e, Gru69a, Lam66d, May69a, Gro68f, Gro68e, Gro68g, Gro68h, AM69g, BKKN67e, BP66b, BP69a, BP69b, BP69c, BP69i, Bri69, BP69k, Buc69b, DK69, EK69, Fre69b, Gro68f, Gro68g, Gro68h, Isb67, Kel69, Kni68, Lam66b, Lam66c, Lam69, Law69a, Lin69c, Mac69a, Mit69, Roo67, Roo69, Swa68r, Tie69b, BP66e, BP66f, Bri66, BP69g]. categorists [Mac69c]. Category [AM68, BBD^{+}69, BDD^{+}67, Swa68h, AM69j, Bar69b, FK69, GH69a, Gro68d, Har66b, Mac69b, Zis69, Hil69b, Hil69c, Löb69]. Cauchy [Hör69, Jan69f, Tre69]. Cauchyschen [Jan69f]. Center [Hil69b, Hil69c, Löb69, Bar69e]. central [SD69]. certain [D’J69, Fer68g, HM68e, Lün69l]. certains [Exb68]. Cesàro [Pey69a]. chaînes [Gir68, Gro68g]. Chain [Swa68i, Gro68g]. chains [BB69a, Gir68, Ivă65g]. Chang [Slo68]. change [Lum68d]. channels [AW69]. Chapter [HC68b, HC68c, HC68d, HC68e, HC68f]. character [KT68b]. Characteristic [BH69k, KT68a, CF66a, Gro68k, KT68d, May69e, May69n, Zar69h]. characterization [Shi68d, Wer66d, Wer66e]. characterizations [Acz69]. Characterizing [Cel68]. Charakterisierung [Wer66d, Wer66e]. charakteristischen [May69e, May69n]. Charge [Sim68a]. Chebychev [Cha66f]. Chebyshev [Cha66f, Ura69, Wer66e, Wer66q, Wer66r, Wer66t, Wer66j, Wer66u]. Chern [KT68b]. Chetaev [Yor68c]. Chicago [Bor67d]. Choice [Kre68, Tro69c, Jen67p, Jen67m]. chung [Jän68d, Jän68d]. Circuits [CF69b, CF69b]. Class [Bor66a, Bor66b, Iwa66, Cho66, FB66c, Her66b, Shi68d, Shi68h]. class-field-theoretical [Shi68d]. class-invariants [Cho66]. Classes [AC64h, AC67d, AKDR68, Hei69d, CF66a, DMO67, Hed69, KT68a, Wer64d, Röd68]. classical [Bor67a, Bre67a, Fic65a, Fic65c, Fic65b, Mar69, Shi68e, Tai68, Tho66a, Ulm69c]. Classification [Sim68b, Hei69b, HM68d, Jän68h, ST69b]. Classificazione [ST69b]. Classifying [Bor67a, Bor67e]. Claude [dC67c]. clean [Gro68i]. closed [Fai67n, FW68q, Lam69, Law69a, Sun68, Zar69j]. closed-graph [FW68q]. co [BP69h, BP69f, Ulm69c, Ulm69b]. co-Cartesian [BP69f]. co-kernels [BP69h]. coaddition [BP66a, BP66a]. Coalgebras [Bar69b]. Cobordism [CF66a, CF66d]. cocellulaire [Cer68f]. Cocellular [Cer68f]. codimension [Cer68f]. coefficients [CF69g, Dol66k, Hör69, Kis68, Tre69, vF69]. Coequalizers [Lin69c, Par69a]. Cofibered [Gro68e]. cofibrée [Gro68d]. Cofibrées [Gro68e, Gro68f, Gro68g, Gro68h]. cofibrous [Gro68d, Gro68h, Gro68f, Gro68g]. Cohen [Jen67e]. Cohensche [Jen67e]. Coherence [Nar66a]. Cohomological [AM69a]. Cohomologie [Pol68, Ser65c, Ser65d, Ser65e, Ser65f, Ser65b, Ser65k, Ver65, Bor64e, Ser65h, Ser65i, Ser65j]. Cohomology [Bar69c, Bre67b, HG67, Ada69a, And68, Bre67a, Bre67e, Ded69, Har66g, Har67a, Har67b, Hoo69, Pol68, Ser65c, Ser65d, Ser65e, Ser65f, Ser65j, Ser65k, Ver65, Bor64e, Ser65b, Ser65h, Ser65i]. cokartesische [BP69f]. Cokerne [BP69h]. colimits [Tie69a]. Collatz [Alb69]. collection [Pim69j]. Collège [SG65, SG65, Ser65d, Ser65e]. collineation [Lün65a, Lün65b, Lün69l, Lün69f]. Collineations [Lün69f]. Colloquium [Ivă65k]. combinatoires [CF69i]. Combinatorial [CF69i]. Comments [Ada64c, Ada66c, Ada69c, Cha67]. Commutatif [Ser65c, Ser65i]. commutation [CF69e, CF69i]. Commutative [And67c, Bar69c, Frö68a, Frö68b, Oor66b, Ser65f, And67a, And68, BP69j, Lar69d, Ser65c, And67a, Ser65j]. Commutativity [AC64b, AC67a, AC64h, AC67d]. Commuting [Tar67b]. Compact [BH67c, BH67a, Par69f, PP66b, PP66c, PP66d, PP66e, Bor64e, Cor68, Dol68a, Fic65l, Her68e, Man69b, Par69b, Roo69, ST69a, ST69b, ST69q, ST69r, Fer68e, FW68m]. compacte [Dol68a]. compactification [Mey68k]. Compactifications [Mey68a, Mey68e, Mey68a]. compactifié [Mey68k]. Compactness [BH69b, Bar68a, FW68r, Kar68, FW68n]. compacts [Bor64e]. Comparative [GKM68k]. Comparison [AM69b, Wil67a, Dol66r]. compatti [ST69a, ST69q, ST69r]. compatto [ST69b]. complement [dC67c, dC67c]. Compléments [CMW68, Mey68f, Bor64d, Bor64d, Mey68f]. complete [Hen68, Lam66j, Mal68, Zar69m]. completely [Urb67e, Her68e, Stu69k]. Completeness [Her65b, Fre69b, Her65r, Lam66i, Wil67d]. completion [Lam66a]. Completions [AM69c, AM69d, Lam66b, Lam66c, Lam66e, AM69h, Fre69b, Isb67, Lam66d]. Complex [BCH^{+}66, Gro68e, Blo68, Con67r, Gro68c, Gro68i, Gro68j, Gro68l, May69k, May69m, Nar66b, Ric69, Shi68e, Shi68f, dR67c, May69k, Tit67b]. Complex-analytic [Tit67b]. Complexe [Gro68e, Gro68i, Gro68j, Blo68, Gro68c, Gro68l, dR67c]. complexes [Bor64h, Gro68g, Grü69b, Har66e, Har66h, KT68h, KT68k, Swa68i]. complexité [Cer68e]. complexity [Cer68e]. Composite [Bar69d]. composition [FB66d]. compound [AW69]. comprehension [Gra69, Tak68a, Tro69m]. Computation [Her66a, Ura69]. Computational [Wer66v]. Computing [Tho66b]. con [ST69c]. concepts [Bor64j, Dah69b, FW68t, Fre69b, Jän68e, Köl68b]. concerning [Bar69a]. Concluding [Tar67a, Shi68g]. concordant [Fre69b]. condition [Fai67o, Kre69, Mat68]. conditional [PP66g]. conditions [BH69b, Fai67h, Lam66l, Pim69h, Zal68i, ST69c]. Condizioni [ST69c]. cone [ST69f]. Conference [CK69, Hil69b, Hil69c, Löb69, Tro69q, Mor69d]. Congress [AL69]. congruence [Lün65e]. Congruences [Cha66k]. conjugates [Lum68c]. Conjugués [Lum68c]. connected [Lic69]. connection [Swa68j]. cono [ST69f]. consequences [Mey68g, Mey68g]. consistency [Jen67m, Pio69]. constant [Hör69, Kis68, vF69]. constants [Hod69, Kis68]. constructible [Jen67c]. constructing [Shi68h]. Construction [Cer68a, Ehr69, Fic65d, Her66b, Fic65i, Jan69f, Man69b, Cer68a, Ehr69, Mey67a]. Constructions [Tie69c, BB69b, Con67m, Eng68, Lam69, Pyk69]. Constructive [Tro69d, Tro69k]. context [Law68]. continuation [BH69m]. continue [Bor64b]. continues [Exb68]. continuity [Hus69, Tro69a, Wer66q]. Continuous [Tar67b, Bor64b, Exb68, Fre69b, Stu69g, Stu69k, Wer66r]. continuum [Jen67o, Jen67m]. contour [ST69s]. Control [Alb68d, Kou67]. Controls [Alb68b]. conull [Wil67b]. Convergence [Cha67, Pad69, Tar67c, Cor68, Lum68d, MD69, Pio69, Wer66l]. convergence-change [Lum68d]. convergent [Pyk69]. convex [Bar69a, Cha66g, FW68j, FW68p, Pol68, Wer66s, Wer66t]. convexe [Pol68]. convolution [vB68]. Coordinate [Zar69d]. coordinates [Zar69i]. coproducts [BP66i]. Coprodukte [BP66i]. core [Her68c, Her68d]. Coreflection [Ken69]. coreflections [Her68b]. Coreflexionen [Her68b, Her68c, Her68d]. Coregular [Wil67b]. corona [Car69]. Correction [Del71]. Corrections [Mey69b]. Correctness [Her65k]. correlations [GKM68g]. correspondence [BP69a]. correspondences [BP69g, BP69j, BP69i]. cost [Kou67]. Cotangent [Gro68e, Gro68c, Gro68i, Gro68l]. cotriple [Ulm69c]. cotriples [Bar69d, Ulm69b]. coupled [Gor69]. Cours [Ser65d, Ser65e, SG65]. Courses [Ser65d, Ser65e, SG65]. covered [KT68k]. covering [Cer68a]. coverings [Bei69, Ken69]. creative [Tro69b]. criteria [AM69a, BKKN67f, FW68r, Kre68]. criterion [Dus69]. critical [Jon68a]. Critically [Lic69]. croissance [Pol68]. croissant [Mey68d]. croissants [Mey67a]. crossings [Fie67]. curvature [Fer68f, Fer68c, GKM68c]. curve [Shi68i, Shi68l, ST69s]. curves [Shi68e]. cutsets [Hed69]. Cybernetics [Ivă65k]. cycles [Fre69a, Zar69f].

d’algèbres [Gro68b]. dalla [ST69g]. d’anneaux [Gro68a]. dans [CF69b, Cer68b, Gir68, Hen68, Lel68a, Roo68, Sch68, Ser65k, Ver65, Wei67b, Wer68]. d’après [Mey68c, Nov68, Roo68]. d’arrêt [Mey69d]. Darstellbare [BP66c, Dol66d]. Darstellbarkeit [Cha66l, Dol66e]. Darstellungen [Tit67d]. decomposing [Tho66a]. decomposition [BP69j, Ivă65g, Pol68, Urb67g, Urb67j, Wer64h, Pol68, BP69k, KT66a]. decompositions [Cer68b, Cer68b]. Deductive [Lam69]. Definability [Kue68, Bar68a, Jen67a, Kre68]. defined [CF69e, Cer68f, Gro68g, Hed69, Spi69]. Definierbarkeit [Jen67a]. définies [Cer68f, Gro68g]. définis [CF69e]. definite [Urb67c]. Definition [Har67b, May69c, Swa68k, Swa68l, Jen67f, Wer66i]. definitions [BH69j, Tar67d]. deformation [FB66f]. Deformations [EE69]. degli [ST69a, ST69q]. degree [Jän68d]. Dekompositionsmethoden [KT66a]. del [ST69b]. Delivered [Urb67b, Ada66h, Ada69g, Tho66h]. Dellacherie [dC67c]. démonstration [Mey69e]. dense [Lam66f, Lam66j]. Deny [Mey67f]. dependence [Wer66r]. depending [Dun68]. depinizioni [ST69o]. depth [Har67e]. derivability [Tai68]. derivate [ST69c]. dérivation [Mét67]. derivative [Köl68g]. derivatives [Fic65o, FB66e, FB66j, FB66k, McK69, ST69c]. Derived [Sta69, Zis69, Bar69d, Har66b, Hoo69, TV69, Kar69]. derivées [Dol68a, Kis68, Sch68]. dèrivès [Kar69]. designations [Bau66i, May69o, Tit67a]. Detailed [Mey68b]. détaillé [Mey68b]. Details [Her65c]. determinant [Swa68q]. determinate [Tak68b]. determination [Lün65e, May69i]. deux [Mey67a]. Development [Pim69a, Bau66g]. developments [Bau67, Sel69]. d’Homotopie [dRMK67, Mau67b, Mau67c, dR67d]. Diagonal [Law69a]. Diagram [BP69c]. Diagrammlemmata [BP69c]. diagrams [Olu69]. difféomorphismes [Cer68g]. Diffeomorphisms [Cer68g]. difference [D’J69, Gor69, LG69, Wer66k]. different [Acz69, Hod69, Lam66i]. Differentiability [FB66e, FB66k]. Differentiable [FB66f, Cer68f, GKM68d, Jän68f]. différentiables [Cer68f]. Differential [Jon68b, Mor69d, Mül66d, Ver69, BS67e, BH69d, BM68b, BM68d, Din67, Dol68a, Fer68g, Fer68f, Fic65k, Gor69, Gou69, Gri69, Gub67, Hod69, Hör69, Jan69b, Kat68b, Kis68, LG69, Lin69a, Pio69, Sch68, Sie68, Stu69e, Stu69f, Stu69i, Tre69, Vit69, Yor68a, Zar69q, vF69, BKKN67c, BKKN67b, Blo68]. differential-difference [LG69]. Differentialgeometrie [Fer68g, Fer68f]. Differentialgleichung [Jan69b]. Differentialgleichungen [Stu69f, Sie68, Stu69i]. Differentialmoduln [BKKN67b]. Differentialoperatoren [Gri69, Din67, Stu69e]. Differentialrechnung [BKKN67c]. Differentials [Zar69e]. differentiation [Mét67, Köl68c, Köl68d]. différentiels [Blo68, BM68b, BM68d]. Differenzenquotient [Wer66k]. Differenzierbare [GKM68d, Jän68f]. dilatation [Leh69]. Dimension [Con67c, Cer68g, Frö68a, Frö68b, Geo67, Mit69, Wer64f]. dimensional [Cer68g, Ded69, Jän68j, Lün65a, Sic68, vF69]. dimensionalen [Lün65a]. Dimensionen [Jän68i]. dimensions [Jän68i, Köl68d, KMP67, McK69, Wer64h]. dinamici [ST69e, ST69p, ST69o]. dinamico [ST69f, ST69g]. Dirac [Sim68c]. direct [Fai67l, Sim68j]. directed [Tie69a]. direction [McK69]. Dirichlet [Bau66c, Cai69, Cha66c, Lum68j, Stu69e, dR67a]. Dirichletsche [Cha66c, Stu69e]. Dirichletsches [Bau66c]. discontinuous [Sel69]. Discrete [Ivă65e, Frö68b, Lum68i]. Discretization [Wer66l]. discrèts [Lum68i]. disk [Hof69, Jän68k]. Diskretisierung [Wer66l]. Displayability [Dol66e]. Displayable [BP66c, Dol66d]. distance [Csö69]. distribution [Cha66f, Cha66i, Csö69]. distributions [Fer67a, Fer67b]. Distributive [Bec69a, Roo67]. Divisorial [Zar69f]. Divisors [Zar69g]. doctrines [Law69b]. d’ordre [de 68]. double [Cer68b]. doubles [Cer68b]. dual [FW68c, KT66b, FW68i]. duale [KT66b, FW68i]. Dualität [BP66e]. Dualité [Ser65k, Ver65, Wei67a, dR67b]. Duality [Har66c, Har66i, Oor66c, BP66e, Har66e, Har66d, Har67d, Wei67a, Zis69, Ser65k, Ver65, dR67b]. Dualizing [Har66e]. Dualraum [FW68c]. d’un [AKDR68, Cai69, CF69c, Del69c, Gro68j, Gro68k, dR67c]. Dundee [Mor69d]. Dundee/Scotland [Mor69d]. d’une [Bor64b, Cer68a, Cer68b, Gro68d, Hen68, Kis68, Ram68, Roo68]. Duoplex [KT66d]. Duoplexmethode [KT66d]. dynamic [Man69a, ST69f]. Dynamical [BS67a, BS67b, BS67c, BH69i, Jon68b, BH69j, Jon68a, Sma69, ST69g, ST69p, ST69o, ST69e].

each [Mat68]. Ebene [Lün65b, Lün69m, Lün69p]. Ebenen [Lün65c, Lün65n, Lün69n]. École [Bor64e, Bor64e]. Econometric [IR66f]. Economy [Ivă65k]. ed [ST69l]. edges [HM68o, HM68p]. Edition [Bor64e]. effective [But69]. effects [Wid69]. Ehrling [Fic65r]. Eidg [Wer64j]. Eigenfunction [Pim69b]. eigenfunctions [Pim69h, Pim69f]. Eigenschaften [Lün69o]. Eigenvalue [Fic65e, Fic65k, Pim69c, Stu69h, Fic65q]. eigenvalues [Fic65s]. Eigenwertaufgaben [Stu69h]. Einbettungen [BP66d]. Einbettungssatz [Jän68c, Jen67h]. Einbettungstheoreme [Wlo69a]. Eindeutigkeit [Wer66s, Wer66m]. Einfacheit [Lün65d]. einfachen [Tit67d]. einfachsten [Jan69e]. Einführung [Ber68f, Cha66n, FW68j, Wer64g, Wer66n]. Einige [Jan69g]. Einleitung [Bau66d, BKKN67d, Dol66i, Her65g, May69h]. elasticity [Jan69k]. Elastizitätstheorie [Jan69k]. Elastostatics [Fic65c]. élémentaire [Mey67m, Mey69d]. elementaren [Cha66d]. Elementary [FB66g, Tro69e, Tro69f, Cha66d, Har67b, Mey69d, Mey67m]. Elliptic [Fic65f, Shi68e, BM68b, Fic65a, Fic65c, Fic65b, Fic65i, Fic65g, Fic65k, Fic65q, Fic65p, Gri69, KT68b, KT68h, Shi68l, Stu69e, BM68d, Stu69f]. ellipticity [Stu69c]. elliptique [BM68b]. elliptiques [BM68d]. elliptische [Gri69, Stu69f]. Elliptizität [Stu69c]. embedding [Jän68c, Wlo69a, Jen67h]. Embeddings [BP66d]. empirical [Csö69]. endlichen [Lün69f]. endlicher [Lün69j, Lün69n]. Endomorphism [Tie69f, Fai67a, Shi68f]. endomorphism-ring [Shi68f]. enforcement [Jen67e, Jen67f]. Enriched [DK69, Kel69]. Ensembles [Del69a, Del69b, Del71, Ram68, Del71]. entière [Hen68]. entières [Ser68]. entire [Ser68]. entre [CF69g]. entropies [Acz69]. entropy [SD69]. envelopes [Sic68]. enveloping [Sim68k]. environmental [HM68o, HM68p]. epireflections [Her68c]. Epireflexionen [Her68c]. Equal [Cha66i]. equality [Slo68]. equation [Jan69b, Pim69i, Sim68c, ST69e, ST69g, Tar67l, vF69, Kis68]. Equational [Dav67, Law69b]. Equations [Jon68b, Mor69d, Tar67j, Ver69, BS67e, BH69d, D’J69, Dol68a, Gor69, Gou69, Gub67, Hod69, Hör69, Ivă65l, IR66c, IR66e, Jan69d, Jan69e, Jan69j, Kat68b, Kis68, Kou67, LG69, Lin69a, McK69, Mül66d, Nob69, Pim69a, Pim69g, Pio69, Sch68, Sie68, Stu69f, Stu69i, Tre69, Vit69, Yor68a, vF69, Dol68a, Sch68]. equazione [ST69g]. équidistribuées [Hub69a]. Equilibrium [Fic65b]. equivalence [Cro68, Kat68a, Her65k]. Equivalences [AC64g, AC67f]. Equivariant [Bre67a, Bre67b, Bre67c, Bre67e, FW69, HM68b]. Erfüllbarkeit [Her65i, Her65h]. Ergebnis [May69g]. ergodic [Kre69, Urb67g]. erhöhter [Jan69l]. Erlangen [Urb67b]. Erratum [Ser65g, Ser65h, Ser65i, Ser65j, Ser65k, Ser65h, Ser65i, Ser65k]. Error [Coo69, KK69, Ura69]. Erweiterungen [Lün69l, Lün69g, Lün69k, Lün69j, Lün69d]. Erzwingungsbeziehung [Jen67f]. Erzwingungsmethode [Jen67e]. espace [Blo68, Cer68f, Kis68, Roo68, Sch68]. Espaces [BM68a, BM68b, BM68c, BDCK67, Cer68b, Dol68a, Lum68c, Lum68d, Ser68, Bor64c, Bor64a, Bor64e, Exb68, Hen68, Lel68a, Lum68a, dR67d]. especially [Shi68c]. Essential [Fai67b]. estensione [ST69t]. Estimates [Zal68f]. estimation [Ura69]. estimators [WW69]. Étale [AM69e]. etc [Ada64c, Ada66c, Ada69c]. ETH [Alb68d, Eck69, Tar67j, Tie69c]. Étude [Cai69, Lum68f, Lum68e]. Euclidean [BS67a, Fer68d]. euklidischen [Fer68d]. Euler [KT68d]. European [IR66f]. evaluation [Ber69]. evolutionary [Gou69]. exact [BP69a, BP69d, BP69i, Bri69, Dol66a, Dol66c, Dol66e, Dol66h, Dol66r, Gro68d, Gro68h, Swa68m, BP69d, BP69k, Gro68m, May69i, Dol66l, Dol66m]. exacte [Gro68d]. exactes [Gro68h, Gro68m]. Exakte [BP69k, BP69a, BP69d, BP69d, May69i]. exakten [BP69c, BP69i]. example [Pim69d, Pim69e, Yor68c]. Examples [BKKN67f, BH67d, CC68, FB66h, Qui67b, Jän68l, Swa68k, Swa68l, Wer66i, Wer66n, Dol66c, Jän68a, Wer66c]. exceptional [Cro69, Jän68g, Pim69a]. excessive [Mey68f, dC67a]. excessives [Mey68f, dC67a]. Existence [Fic65g, Fic65h, PP66f, PP66g, Jon68a, Str68b, Yor68d, Yos68]. Existenz [Lün69l]. exotic [HM68a, Jän68d]. exotische [HM68a]. exotischen [Jän68d]. Expansions [Mül66e, Cam69]. Explicit [Fic65i, Lün65e]. explizite [Lün65e]. exponential [Ton65a]. exposé [CMW68]. Exposés [Bor64e, Mey69b, dRMK67]. Extending [Yor68b]. Extension [Con67d, Kel67, Pim69f, Ric69, Sch69, Bau66j, Dol66f, Lam66j, Pim69h, Pim69g, ST69t, Yor68c, Sch69]. Extensions [Sto66d, Cha69a, Cha69b, Fai67b, Fai67n, Gro68a, Gru69a, Lün69l, Lün69g, Lün69k, Lün69j, Shi68j, Ulm69a, Ulm69b, Lün69d, Mor69b]. Extremal [Fuj68, GKM68e]. Extremaleigenschaften [GKM68e].

faces [Cro69]. Facets [CK69]. facetted [Guy69]. factoring [Olu69]. Factorization [Cas69, Hil69a, Din67]. factors [Dol66o]. facts [Zal68p]. faisceaux [Bor64i, Gro68a, Gro68b]. faits [Bor64e, dRMK67]. Faktoren [Dol66o]. Faktorisierung [Din67]. Fall [HG67]. Faltungskerne [vB68]. families [Hen68]. Famille [Hen68]. fan [Tro69a]. Fassung [Wer66q]. fast [May69k, May69m]. fast-komplexen [May69k, May69m]. Fatou [Kur68]. FD [Tie69d]. FD-spectra [Tie69d]. Federal [Wer64j]. Fédérale [Bor64e, Bor64e]. Feller [Mey67l]. Fellerian [Mey68c]. fellériennes [Mey68c]. fermé [Sun68]. février [DE67]. FH [Wil67c]. Fiber [Tho66h, Bor64c, Bor64a, Tho66g]. fibration [Tho66a]. fibrations [AM69d, Bre67d]. Fibre [Bor67d, AM69g, Bor67c]. fibres [Bor64c, Bor64a]. field [Shi68d, Shi68f]. fields [Her66b, Iwa66, Shi68h]. filters [FB66g]. final [Fic65m]. fine [Bre67f, Bre67f]. fini [Lum68f]. Finite [Con68b, And69, AC64e, AC67c, Bou69, Ivă65g, Lum68f, Lün69f, Lün69j, Lün69n, Shi68l]. First [IR66f]. fixed [Con67c]. Flat [KT68c, KT68g, KT68e, KT68f, KT68d, KT68b, KT68a, KT68h, KT68k]. Flow [BH69c, ST69b]. flows [Ivă65c, ST69d]. flussi [ST69d]. flusso [ST69b]. Folgen [BP69d]. Foncteurs [Kar69]. Fonction [CF69c, Hen68]. Fonctionnelles [Dol68b, Del69c, Kis68, Ser68]. fonctionnels [Cer68b]. Fonctions [Lel68a, Ava67, BM68a, Cer68f, Exb68, Lum68a, Lum68b, Mey68f, Ser68, dC67a]. fondamental [Bor64g]. fondamentales [EG69e]. form [Tar67i]. Formal [Frö68c, Frö68a, Frö68b, Jen67d]. Formale [Jen67d]. Formalisations [Ros68a]. formally [Gro68i]. formally-clean [Gro68i]. Formel [Wer64f]. formellement [Gro68i]. Formen [HM68l, HM68j]. Forms [HC68a, Bak69, Gri69, HM68j, HM68l, Lam66i, Ser66a, Stu69b, Stu69c, Stu69k]. formula [Con67b, Hsi69, Mül66f, Mül66m, Wer64f]. formulae [Lin69a]. Formulas [Kei68, AC64h, AC67d]. formulation [Dah69b]. Forschungsinstitut [Alb68d, Tar67j, Wer64j]. fortement [Mey68c]. fortemente [ST69b]. Fortsetzungssatz [Bau66j]. foundation [KT66c, Mac69b]. Foundations [FW69, Mac69a, FK69]. four [Cha66l]. Fourier [Ber68e, KMP67, Lar69e]. Fourierumkehrintegrals [Ber68e]. Fractional [IR66a]. fractions [Kni68]. Frage [Jen67k]. fragment [Slo68]. framed [CF66c]. France [SG65, SG65, Ser65d, Ser65e]. Franz [dR67a]. free [Ehr69, Her65n, Sim68b, Sim68h]. freies [Her65n]. French [And67a, And67b, And67c, Ava67, AKDR68, ADR69, Bad67, Blo68, BM68a, BM68b, BM68c, BM68d, BM68e, Bor64d, Bor64c, Bor64b, Bor64a, Bor64e, Bor64h, Bor64i, Bor64f, Bor64j, Bor64g, Bre67f, BDCK67, Cai69, CF69b, CF69c, CF69d, CF69e, CF69h, CF69f, CMW68, CF69i, Cat68, Cer68a, Cer68g, Cer68b, Cer68c, Cer68f, Cer68e, Cer68d, Del69c, Del69a, Del69b, Del71, Dol68a, Dol68b, EG69a, EG69b, EG69c, EG69d, EG69e, Ehr69, Exb68, Fer67a, Fer67b, FW68i, Gir68, Gro68c, Gro68a, Gro68b, Gro68f, Gro68e, Gro68d, Gro68g, Gro68h, Gro68i, Gro68j, Gro68l, Gro68k, Gro68m, dGM69, Hen68, Hub69a, Hub69b, Igo68, Kar69, Ker67, Kis68, Lel68a, Lum68a, Lum68b, Lum68f, Lum68e, Lum68c, Lum68d, Lum68g, Lum68h, Lum68i, Lum68j, Mau67a, Mau67b, Mau67c, Mét67]. French [Mey67a, Mey67b, Mey67c, Mey67d, Mey67e, Mey67i, Mey67j, Mey67k, Mey67l, Mey67f, Mey67m, Mey68a, Mey68f, Mey68e, Mey68b, Mey68g, Mey68d, Mey68k, Mey68h, Mey68i, Mey68j, Mey68c, Mey69e, Mey69a, Mey69b, Mey69d, Mey69c, Mor69a, Mor69b, Mor69c, Nor68, Pol68, Ram68, Roo68, Sch68, Sch69, SG65, Ser65c, Ser65d, Ser65e, Ser65f, Ser65b, Ser65h, Ser65i, Ser65j, Ser65k, Ser68, Sun68, ST69c, Ver65, Wei67a, Wei67b, Wer68, dC67c, dC67a, dC67b, dR67a, dR67c, dR67b, dR67d, dRMK67, de 68]. Front [Ada64d, Ada66d, AC64c, Bau66e, Ber68g, Cha66j, GKM68f, Her65j, IR66b, Jan69h, Köl68e, Mey67h, Ser65l, Wer66o]. frontière [Mey68i, Mey68j]. frontières [Gir68]. full [Fai67l, Mae68b]. full-superharmonic [Mae68b]. Fully [BP69j]. Fully-commutative [BP69j]. Function [Bre67d, Zal68g, Bor64b, CF69c, Csö69, Fic65j, FB66m, Hen68, Ric69, Shi68i, Stu69e, WW69, Wer64j, Zar69h, Lum68b]. Functional [BH69d, Tar67j, Wil67h, Del69c, Dol68b, Kat68b, Yor68a, Cer68b]. Functional-differential [BH69d]. functionals [Kis68, Ser68, Stu69g]. Functions [BH67a, Shi68b, AW69, Ava67, Bau66h, BH67b, BH69g, BM68a, Cer68f, Cha66m, Csö69, Exb68, Hei69f, Hof69, Ivă65h, IR66d, Lar69e, Lar69g, Lel68a, Lum68a, Mae68b, Mey68f, Mül66c, Mül66i, Mül66k, Nov68, Röd68, Ser68, Shi68c, Shi68a, Shi68l, Sic68, Sie68, Str68b, Stu69f, ST69d, Wer64a, Wer66a, Wer66g, Wer66r, Wlo69b, Yor68b, Yor68d, dC67a]. functor [AM69m, Ber69, Buc68, Con67e, DK69]. functorial [Law68, Lin69b, Lin69d, Lin69e]. Functors [EK69, Har67c, Bar69d, BP66c, BP66e, BP66f, Bri66, Dav67, Dol66a, Dol66c, Dol66d, Dol66e, Dol66l, Dol66m, Dol66h, Dol66r, Fre69b, Hoo69, Kar69, Lam66h, Sta69, TV69, Dol66j, Dol66k]. Fundamental [FB66i, Kat68b, AM69f, Cha66d, Shi68e, Zar69n, EG69e]. Fundamentalsatz [Cha66d]. Funk [Mül66f]. Funktion [Wer66r]. Funktionale [Stu69g]. Funktionen [Bau66h, Cha66m, Sie68, Wer64j, Stu69f, Wer64a, Wer66a, Wer66g, Wlo69b, Röd68]. Funktionen-Algebren [Wer64j]. Funktionenräume [Stu69e]. Funktoren [BP66e, Dol66a, BP66c, BP66f, Bri66, Dol66c, Dol66d, Dol66e, Dol66j, Dol66k, Dol66r]. Funzioni [ST69d]. Further [Shi68g, Swa68n, Zal68h, Shi68h].

g [ST69g]. Gäarding [Fic65q]. Galois [Ser65i, CS69a, Cha69a, Cha69b, CS69b, CS69c, Ser65c, Ser65d, Ser65e, Ser65f, Ser65j, Sha69, Shi68j]. Galoisienne [Ser65c, Ser65f, Ser65i, Ser65j, Ser65d, Ser65e]. games [Bie67]. ganzen [Cha66l]. ganzzahlige [HM68e, HM68j]. Gaps [Jän68i]. Garling [Wil67d]. gauche [Gro68d, Gro68h]. Gauss [dR67a]. Gaussian [Fie67]. Gaußschen [Fie67]. gehaltenen [Bau66g]. gehörigen [Lün65i, Lün65j].

Gelfand [FW68k]. Gelfandräume [FW68k]. Genauigkeit [Jan69l]. General [Hei69c, Mac69d, Mül66g, Csö69, Lar69a, Mey68b, Ric69, WW69, Gro68l, Jan69a, Stu69a]. générale [Mey68b]. générales [Gro68l]. Generalisation [Alb69, Cer68c]. generalised [Ada69a]. generalization [Cer68c, Shi68g]. Generalized [Bre67e, Kre68, Jon68a, Roo69, ST69f, ST69g, ST69o, Wlo69b, Her65s]. generalizzati [ST69o]. generalizzati-depinizioni [ST69o]. generalizzato [ST69f, ST69g]. generated [ST69e, ST69g]. générateurs [Lum68f]. generati [ST69e]. Generating [Lam66f]. generato [ST69g]. Generators [Nov68, Lum68f]. Generic [Jen67i]. Generische [Jen67i]. genus [Zar69b, Zar69h]. Geodätischen [GKM68e]. geodesics [GKM68e]. Geometric [MT69, Zal68i, Mau67c]. geometrical [Eng68]. Geometrie [Fer68b, Fer68a, GKM68h, Lün69p, BKKN67c, BKKN67e]. Geometrien [Lün65g]. geometries [Lün65g]. géométrique [Mau67c]. geometry [BKKN67c, BKKN67e, Fer68b, Fer68a, Fer68g, Fer68f, GKM68h, Lün69p]. German [Bau66b, Bau66c, Bau66d, Bau66f, Bau66g, Bau66h, Bau66i, Bau66j, Bau68, Ber68b, Ber68a, Ber68f, Ber68h, Ber68e, Ber68d, Ber68i, BKKN67a, BKKN67c, BKKN67b, BKKN67d, BKKN67e, BKKN67f, Bie67, BP66a, BP66b, BP66c, BP66d, BP66e, BP66f, Bri66, BP66g, BP66h, BP66i, BP69a, BP69b, BP69c, BP69d, BP69e, BP69h, BP69f, BP69i, BP69g, BP69j, BP69k, Cha66c, Cha66n, Cha66d, Cha66k, Cha66e, Cha66l, Cha66b, Cha66h, Cha66f, Cha66g, Cha66i, Cha66m, Din67, Dol66a, Dol66b, Dol66c, Dol66d, Dol66e, Dol66i, Dol66j, Dol66k, Dol66f, Dol66l, Dol66m, Dol66n, Dol66o, Dol66p, Dol66q, Dol66g, Dol66h, Dol66r, Fer68b, Fer68a, Fer68g, Fer68d, Fer68e, Fer68f, Fer68c, Fie67, FW68a, FW68b, FW68c, FW68j, FW68k, FW68l, FW68m, FW68n, FW68o, FW68p, FW68q]. German [FW68d, FW68e, FW68r, FW68f, FW68g, FW68h, FW68t, FW68s, FW68u, Geo67, Gri69, GKM68a, GKM68c, GKM68d, GKM68e, GKM68g, GKM68h, GKM68j, GKM68i, GKM68k, Her65g, Her65i, Her65h, Her65k, Her65d, Her65l, Her65m, Her65n, Her65q, Her65e, Her65f, Her65p, Her65r, Her65s, Her65t, Her68a, Her68b, Her68c, Her68d, Her68e, HM68b, HM68c, HM68e, HM68j, HM68f, HM68d, HM68a, HM68i, HM68g, HM68k, HM68l, HM68n, HM68m, HM68o, HM68p, HM68h, Jan69a, Jan69d, Jan69b, Jan69g, Jan69e, Jan69i, Jan69j, Jan69f, Jan69k, Jan69l, Jan69m, Jän68b, Jän68a, Jän68f, Jän68c, Jän68e, Jän68g, Jän68h, Jän68i, Jän68j, Jän68d, Jän68k, Jän68l, Jen67a, Jen67e, Jen67f, Jen67h, Jen67i, Jen67j, Jen67c, Jen67k, Jen67g, Jen67n, Jen67l, Jen67m, Jen67d, Jen67p, Jen67q, Jen67o]. German [Jen67b, Köh68, Köl68b, Köl68c, Köl68d, Köl68f, Köl68g, KT66a, KT66e, KT66d, KT66c, KT66b, Lün65c, Lün65e, Lün65d, Lün65f, Lün65l, Lün65a, Lün65b, Lün65k, Lün65m, Lün65n, Lün65i, Lün65j, Lün65g, Lün65o, Lün65h, Lün69d, Lün69l, Lün69a, Lün69p, Lün69i, Lün69b, Lün69e, Lün69f, Lün69c, Lün69g, Lün69h, Lün69q, Lün69k, Lün69j, Lün69n, Lün69m, Lün69o, May69b, May69c, May69h, May69g, May69i, May69d, May69j, May69k, May69l, May69m, May69e, May69n, May69f, May69o, Röd68, Sie68, Stu69a, Stu69b, Stu69c, Stu69e, Stu69f, Stu69g, Stu69h, Stu69d, Stu69i, Stu69j, Stu69k, Tit67a, Tit67b, Tit67c, Tit67d, Wer64c, Wer64a, Wer64g, Wer64f, Wer64d, Wer64i, Wer64h, Wer64e, Wer64j, Wer64b, Wer66f, Wer66a, Wer66b, Wer66g, Wer66c, Wer66d, Wer66e, Wer66k, Wer66l, Wer66i, Wer66m, Wer66n, Wer66p]. German [Wer66h, Wer66q, Wer66r, Wer66s, Wer66t, Wer66j, Wer66u, Wer66v, Wlo69a, Wlo69b, Wlo69d, Wlo69c, Wlo69e, Wlo69h, Wlo69i, Wlo69g, Wlo69f, vB68, vH68]. Gestalt [GKM68c]. gewisser [Fer68g, HM68e, Lün69l]. gewöhnlichen [Stu69i]. Gitterpunkte [Cha66g, Cha66m]. given [Alb68d, Bor67d, Har66i, HG67, Lam66d, Leh69, Tie69c]. Gleichungen [Jan69d, Jan69e, Jan69j]. Gleichverteilung [Cha66i]. Gleichwertigkeit [Her65k]. Glivenko [Csö69]. Global [Lin69a, Sma69, Str68b, ST69t]. globale [ST69t]. glossary [AM69g]. Gorenstein [Lün69h, Lün69h]. graded [Knu69, Swa68c]. Graph [And69, CK69, Ham69, CF69b, Erd69, FW68q, Sun68]. Graphe [CF69b, Sun68]. Graphen [FW68q]. Graphen-abgeschlossene [FW68q]. Graphs [Aig69, Bar69a, Bei69, Bou69, Fre69a, GH69a, GH69b, Grü69b, HW69, Hed69, Hes69, Ivă65a, Jän68k, Lic69, Man69c]. greatly [Hod69]. Green [Fic65i, Mey68h]. großen [Lün69a, GKM68h]. großer [KT66e]. Grothendieck [Har66i]. Group [AC64g, Con68b, And68, AC67f, AM69f, Gru69a, Knu69, KM68, Lün65d, Lün65f, Lün65a, Lün65b, Lün65i, Lün65j, Lün65h, Lün69b, Oor66a, Oor66b, Par69b, Swa68q, Tho69, Ton65a, Lün65g]. Groupes [Ser65b, Ser65h, Ser65k, Ver65, Mey67l, dC67a]. Groups [AC64e, AC64h, AC64d, AC67c, AC67d, Frö68c, HC68a, Par69f, Sim68f, Ton65e, Ton65f, Ton69, AC67b, AM69h, Ber68a, Bor67a, Bou69, Bri69, Ded69, Frö68a, Frö68b, Har67b, Jän68i, Kne69, Lam66e, Lün65k, Lün69l, Lün69a, Lün69i, Lün69f, Lün69c, Lün69g, Lün69q, Lün69k, Lün69j, May69b, Par69e, Par69d, Par69g, Sel69, Ser65b, Ser65h, Tho66e, Tit67b, Tit67c, Tit67d, Ton65i, Ver65, vB68, Ser65k]. growing [Mey68d]. growth [Dun68, Pol68]. Grundbegriffe [FW68t, Jän68e, Köl68b]. Grundlagen [KT66c]. Grundpunktes [Dol66g]. Grundräume [Wlo69b, Wlo69h]. Gruppe [Lün69b]. Gruppen [Lün65f, Lün69g, Tit67d, vB68, Lün65k, Tit67b, Lün69a, Tit67c]. Guide [Mey68b].

H* [Lar69d]. H*-algebras [Lar69d]. Haar [Wer66h]. Habilitation [Köl68d]. Habilitationsschrift [Köl68d]. Hahn [FW68o, FW68o]. Halbexakte [Dol66a, Dol66l, Dol66m]. halbexakten [Dol66c]. halbexakter [Dol66e, Dol66h, Dol66r]. Halbgruppen [Ber68a]. Half [Dav68, Dol66e, Dol66r, Shi68c, Dol66l, Dol66m]. half-exact [Dol66e, Dol66r, Dol66l, Dol66m]. Half-ring [Dav68]. Hall [Lün69h]. Hamburg [Bau66g, Bau66g]. Hammerstein [Pim69f, Pim69c]. Hanf [Mal68]. Hardy [Lum68a, Hei69d, Lum68a, Lum68c, Lum68d]. harmonic [Ava67, CC68, Cor68, Mat68, Sie68, vH68, Bau66f, Bau66g, BM68b, BM68c, Köh68, Mey67k, vB68]. harmonicité [Ava67]. harmonicity [Ava67].

Harmonics [Mül66n, Mül66e, Mül66h]. harmoniques [Ava67, BM68b, BM68c, Mey67k]. Harmonische [Bau66f, Bau66g, Köh68, vB68]. harmonischer [vH68]. Harnack [BM68e]. Harvard [Har66i, HG67]. Hasse [Shi68i]. Hauptorbits [Jän68g]. Hausdorff [Q. 66a, Pey69b]. Hecke [Mül66f]. HEE [Dol66f]. Held [Mor69d, BCH^{+}66, CK69, Hil69b, Hil69c, Löb69, yMO68]. HEP [Dol66f]. Herglotz [Urb67a]. Hermitean [Tar67n]. Higher [FB66j, Shi68l, de 68]. Highly [Mey68c]. Higman [Lün69b, Lün69b]. Hilbert [Ati69, FW68l, Zar69h]. Hindernistheorie [Dol66b]. Hintergrund [Her68a]. history [Dun68]. Hochschild [Mit69]. Hochschule [Wer64j]. Hölder [BM68a]. hölderiennes [BM68a]. holomorphe [Roo68]. Holomorphic [Nar66b, Roo68]. holomorphy [Sic68]. Homeomorphisms [Leh69]. Homogene [Lün69i, Jan69i]. Homogeneous [Zar69i, Gub67, Jan69i, Lün69i]. Homologic [And67b]. Homological [Ada64a, Ada66a, And67c, Ada64h, Ada66i, Ada69b, Ada69h, Bor67b, Ulm69c]. Homologie [HM68f, May69d, HM68e]. Homologique [And67c, And67b]. Homology [BB69b, Bor67d, Buc68, Eck69, Hil69b, Hil69c, Löb69, Olu69, HM68e, HM68f, May69d, Ulm69c, Ulm69b, BP69e]. Homoloqie [BP69e]. homomorphism [Gro68k]. Homomorphisms [AC64d, AC67b]. homorphisme [Gro68k]. Homotopical [Qui67c]. Homotopieerweiterungseigenschaft [Dol66f]. Homotopiefunktoren [Dol66l, Dol66m]. Homotopy [Ada64a, Ada64g, Ada64i, Ada66a, Ada66g, AC64b, AC64h, AC64g, AC67d, AM69e, AM69h, Tie69c, dRMK67, Ada66h, Ada69b, Ada69f, Ada69g, AC67a, AC67f, AM69j, Dol66f, Dol66l, Dol66m, KT68k, Kni68, Mau67b, Mau67c, Qui67a, Qui67b, Tho66e, dR67d]. Homotopy-Commutativity [AC64b, AC64h, AC67a, AC67d]. Hopf [Cha69a, Cha69b, CS69b, CS69c]. Hopkins [Fic65k]. Hörmander [Nov68]. Hsiang [Jän68d, Jän68d]. Hughes [Lün69h, Lün69h]. hull [Fai67b]. Hunt [Mey67i, vB68]. Huntsche [vB68]. Hurwitz [Cha66h]. hydrodynamic [Jan69g]. hydrodynamische [Jan69g]. Hyperarithmetic [Cle68]. hyperbolic [Jan69d, vF69]. hyperbolische [Jan69d]. Hypercoverings [AM69i]. Hyperflächen [Fer68e]. hyperfonctions [Sch68]. hyperfunctions [Sch68]. hypersurfaces [Fer68e]. Hypothese [Jen67m, Mey68g, Jen67o]. hypotheses [Pim69j]. hypothesis [Jen67o, Mey68g, Jen67m].

IB [Jen67l, Jen67l]. IB-Modelle [Jen67l]. ideals [Fai67h]. Identification [May69j]. identifications [BP66d]. Identifikation [May69j]. Identifizierungen [BP66d]. ihre [Bau66g, Lün65g]. ihren [Tit67d]. II [Del71, DE68, Hil69c, Del71, AM68, Bor66b, Del69b, Dol66j, HC68c, HM68p, Lam69, Lin69b, Mey67c]. III [Dol66k, DE69, Löb69, BBD^{+}69, Dol66k, HC68d, Mey67d, Mit69]. illustration [Tho66c]. im [Bau66g, GKM68h, Jan69f]. image [Roo68]. immersed [Cer68b]. Immersion [Fer68d]. Immersionen [Fer68d]. Implicit [Bar68a, Ver69]. Incidence [Grü69b, Lün65l, Lün69e]. inclusion [Alb69]. increase [Gro68i, Mey68d, Mor69c]. increased [Jan69l]. increasing [Mey67a]. increments [Mey69a]. indefinite [Bie67]. indépendantes [Fer67a, Fer67b, Hub69a]. independants [Mor69c, Mey69a]. independence [Jen67m, Jen67p, Jen67q, Jen67o]. independent [Fer67a, Fer67b, Hub69a, Mey69a, Mor69c]. indetermined [Urb67e]. Index [Con67g, Con67f, KT68h]. Induced [Sim68d]. Inductive [Wil67e, Wlo69d, Wlo69c]. Induktive [Wlo69d, Wlo69c]. inégalité [BM68e]. inégalités [dGM69]. inequalities [dGM69, Ivă65l, IR66c, IR66e]. inequality [BM68e, Fic65q]. Infinitary [Bar68b, Bar68a, Cha68, Kre68, Kue68]. Infinite [Shi68j, Bec69b, May69a, Spi69]. infinitely [KMP67]. infinitesimal [Gro68b, Gro68i, Gro68i]. infinitésimales [Gro68b]. inflections [Her68d]. Information [BKW69, KK69, Kem69]. initial [Fit69, Jan69f]. Injective [Fai67c, Fai67e, Fai67k, Fai67b, Fai67a]. injectives [Sta69]. Inner [Jen67j]. innere [Jen67l, Jen67j]. insieme [ST69b]. Insiemi [ST69h, ST69i, ST69a, ST69l, ST69q, ST69r]. Instability [Nob69, Yor68c]. Institut [Lum68a, Wer66v]. Institute [BCH^{+}66, Hil69b, Hil69c, IR66f, Löb68, Löb69, Lum68a, yMO68, Wer64j, Wer66v]. Instrumental [Wer66v]. Instrumentelle [Wer66v]. integer [Cha66l, HM68e, HM68j]. integrable [Mey67a, Mey67a]. Integral [Mül66h, Tar67d, Ber68e, Ber68d, Jan69j, Kne69, Nob69, Sie68]. Integraldarstellung [Sie68]. Intégrales [Mey67b, Mey67c, Mey67d, Mey67e]. Integralgleichungen [Jan69j]. integrals [Mey67b, Mey67c, Mey67d, Mey67e, Zal68f]. Integration [Jan69e, Hod69, Lin69a, Pio69]. Integrodifferentialgleichungen [Jan69j]. interaction [Spi69]. Interior [Fic65f]. intermediate [Fic65d, Jan69d, Jan69b, Jan69e, Jan69m]. intermediate-step [Jan69d, Jan69b]. internal [Jen67l]. International [BKW69]. Interpolation [Ber68h]. Interpolationsmethoden [Ber68h]. Intersection [Zar69j]. Introduction [Ada64e, Ada66e, Ada69d, ABB^{+}69, AC64f, AC67e, Bén67, BH67e, BH69e, CF69d, Con67h, Con68a, EG69a, EG69b, Har66f, Ivă65f, KT68i, Lam66g, Mae68a, Nar66c, Par69c, PP66h, Shi68k, Sto66b, Tar67e, Tho66d, Tie69e, Ton65f, Ton69, Tro69g, Zal68j, Zar69a, Mac69d, Bau66d, Ber68f, BKKN67d, CF69d, Cha66n, Dol66i, EG69a, FW68j, Her65g, May69h, ST69j, Wer64g, Wer66n]. Introductory [Pey69c]. Introduzione [ST69j]. Intuitionism [Tro69q]. Invariance [BH69f, Sim68i, ST69k]. Invariant [Jac67, BH69c, Con67g, Con67i, Con67j, Con67q, Par69g, ST69b, Yor68c, ADR69]. invariante [ADR69, ST69b, HM68l]. Invarianten [Dol66p]. invariants [Bor66a, Bor66b, Cho66, Dol66p, Fic65l, Fic65s, Her66a, HM68l, Tho66b]. Invarianza [ST69k]. inversion [Ber68e]. Inverting [Tie69f]. involutions [Con67r, Con68d]. Inzidenzstrukturen [Lün65l, Lün69e]. irrational [Cha66h]. irrationalen [Cha66h]. irreducible [Zar69d]. isomorphism [AM69a, CF66b, Her65p]. isospin [Sim68a]. Italian [ST69a, ST69b, ST69f, ST69d, ST69j, ST69k, ST69h, ST69i, ST69l, ST69m, ST69q, ST69s, ST69e, ST69g, ST69p, ST69r, ST69o, ST69n, ST69t, ST69u]. iterated [Hub69a]. iteration [Tar67b]. iterative [D’J69]. itéré [Hub69a]. IV [HC68e, Mey67e].

Jacobian [Mar69]. Johns [Fic65k]. Johnson [Fai67d]. Jordan [Lün69h, Lün69h]. Juin [DE69]. July [Hil69b, Hil69c, Löb69, Tar67j]. June [Hil69b, Hil69c, Löb69, Mor69d].

K-W [Mey68g]. Kac [Csö69]. Kalamazoo [CK69]. Kalkül [Her65d]. Kan [Ulm69a, Ulm69b]. Kartesische [BP69f]. Kategorieller [Her68a]. Kategorien [BP69a, BKKN67e, BP66b, BP66e, BP66f, Bri66, BP69b, BP69c, BP69i, BP69g, BP69k]. Kennzeichnung [Lün65k]. Kerne [BP69h, vH68]. kernels [BP69h, Pim69c, Tar67o, vB68, vH68, BP69h]. Killing [Tho66e]. kind [Nob69]. Klassen [Wer64d, Röd68]. Klassifikation [Jän68h]. Klassifikationssatz [HM68d]. kleinen [Lün69c]. knot [Jän68b]. Knoten [Jän68b]. Knoten-Mannigfaltigkeiten [Jän68b]. knots [dR67b]. Koeffizienten [Dol66k]. Kollineationen [Lün69f]. Kollineationsgruppe [Lün65b, Lün65a]. Kollineationsgruppen [Lün69f, Lün69l]. Kolmogorov [Urb67h]. Kompakte [Her68e, Fer68e, FW68m]. Kompaktheit [FW68n]. Kompaktheitskriterien [FW68r]. Komplex [Tit67b]. Komplex-analytische [Tit67b]. Komplexe [May69k]. komplexen [May69k, May69m]. Kongruenz [Lün65e]. Kongruenzen [Cha66k]. konstruktible [Jen67c]. Konstruktion [Jan69f]. Kontinuum [Jen67m, Jen67o]. Kontinuum-Hypothese [Jen67m, Jen67o]. Konvergenz [Wer66l]. konvexe [Wer66s, Wer66t]. konvexen [Cha66g]. Korrektheit [Her65k]. Korrespondenzen [BP69a, BP69i, BP69g, BP69j]. Krein [Urb67h]. Kronecker [Cha66i]. Krull [Swa68o]. Krümmung [GKM68c]. Künneth [Hsi69]. Kuramochi [Fuj68, Mae68a, yMO68, Mat68, Oht68, Tan68]. Kutta [But69]. Kyoto [yMO68].

L [Jän68i]. L. [Alb69, Jän68i, Nov68]. language [Jen67d]. Languages [Bar68b, Bar68a, Cha68, Dah69b, Kre68, Kue68]. Laplace [Jan69b]. Laplaceschen [Jan69b]. Large [KT66e, GKM68h, Lün69a, Pio69]. lattice [Cha66g, Cha66m]. lattices [Cor68, Fai67n]. Lausanne [dRMK67]. law [Cha66b, Hub69a]. Lawless [Tro69h]. lawlike [Tro69d, Tro69k]. laws [Bec69a, BDCK67]. Lebesgue [Wer64h]. Lebesguesche [Wer64h]. Lecture [Bau66g, Har66i, IR66f]. Lectures [Ada69a, Bor67d, Buc69a, Con68b, Fai67e, Fef68, Lün65g, MH69, Pey69d, TA69, Tho66h, Tro69q, Lam66d, Ada66h, Ada69g, Urb67b, Wer66v]. Leeds [Löb68]. left [Gro68d, Gro68h]. left-exact [Gro68d, Gro68h]. Legendre [Mül66c, Mül66i, Mül66k]. lemma [Hub69b, Zar69n]. lemmas [BS67d, BP69c, Fic65n, Fic65r]. lemme [Hub69b]. length [Fuj68]. lens [dR67d]. lenticulaires [dR67d]. level [Fie67, Shi68l]. Lévy [Bad67].

Liapunov [BS67e, BH69g, Str68b, ST69d, Yor68b, Yor68d]. libres [Ehr69]. Lie [Tit67d, Frö68d, Ham69, HC68a, Par69e, Sim68e, Sim68f, Tit67d, Ton65a, Ton65f, Ton69]. liés [Cer68b]. Lifting [Sim68g]. liftings [Köl68c, Köl68c]. Limes [FW68d]. Limit [BH69h, Lam66h, FW68d, SD69, ST69l, ST69h]. limite [ST69h, ST69l]. Limited [FW68b, Stu69b]. Limits [Sto66d, Fre69b, Wil67e]. Line [Con68c, Tro69j]. linéaires [BM68b]. Linear [Fic65c, IR66c, Zar69k, BM68b, Fai67l, Fic65a, Fic65f, Gub67, Hod69, Kat68b, Kou67, Lar69g, Tre69, Wer66d, Wer66i, Wer66s, Wer66t, Wil67g, Fic65k, FW68o, GKM68g, Stu69g]. lineare [Wer66t, FW68o, GKM68g, KT66e, Stu69g]. linearem [Wer66d]. linearen [Wer66i, Wer66s]. Linearitätsbeweises [Jän68j]. Linearitätssätze [Jän68l]. linearity [Jän68j, Jän68l]. linearly [Kei68, Roo69]. lines [Hed69, ST69s]. Linnik [Igo68]. Lipschitz [Yor68b]. Literatur [Wer64i]. livello [ST69s]. Local [BH69j, BH69i, Har66g, Har67d, Nar66d, Tre69, Buc69a, Con67i, Con67j, Fic65g, Fie67, Har66e, Har67a, Har67b, Kre69, Sto66c, FW68p, HG67, SG65, ST69m, ST69t]. locale [SG65, ST69t]. localement [Bor64e]. locali [ST69m]. Localization [Swa68p]. Locally [Par69f, Roo67, Roo69, Bor64e, FW68j, Par69b, Tar67g]. log [Lum68j]. logarithm [Hub69a]. logarithme [Hub69a]. Logic [Löb68, Tro69i, Her65m, Her65q, Her65e, Her65f, Her65t, Tai68, Tak68b, BP66g, Löb68]. logical [Ivă65e]. Logik [BP66g]. logmodulaires [Lum68j]. loi [Hub69a]. Lois [BDCK67]. lokalen [Fie67]. Lokalkonvexe [FW68p]. lokalkonvexen [FW68j]. loop [Bec69b, May69a]. loss [WW69]. Lösung [Jan69f, Jan69m, Jen67k]. low [KK69]. lower [Sic68, Sto66c]. Lucid [Fre69b]. Lücken [Jän68i]. Łukasiewicz [Ros68a]. l’Université [dRMK67].

M [CMW68, Lün69h]. M. [Csö69, Lün69h]. Mackey [FW68f]. Macmahon [CF69h]. made [Bor64e, dRMK67]. Main [Jän68g]. Mainz [Lün65g]. majoration [Mey68d]. Management [IR66f]. manifold [Cer68f, HM68g, May69e, Ram68]. Manifolds [KT68f, CF66c, Dol68a, GKM68d, GKM68j, GKM68i, Hil69a, HM68e, HM68f, HM68d, HM68a, HM68i, HM68n, HM68m, Jän68b, Jän68a, Jän68f, Jän68h, KT68g, KT68e, KT68d, May69k, May69l, May69m, Ros69, Ton65b]. Mann [Jän68i, Jän68i]. Mannigfaltigkeit [May69e]. Mannigfaltigkeiten [GKM68d, GKM68i, HM68a, HM68g, HM68m, GKM68j, HM68e, HM68d, HM68i, HM68n, Jän68b, Jän68a, Jän68f, Jän68h, May69k, May69l, May69m]. Many [CK69, Guy69, KMP67, Shi68f]. many-facetted [Guy69]. map [Roo68, Ton65a]. mapping [Q. 66b]. mappings [FB66d, FB66f, FB66a, Nar66b]. Maps [Con67p, Con67k, FW68a, FW68i, FW68l, FW68m, FW68o, FW68q, GKM68d, HM68k, Ken69]. March [Fic65k]. marking [Lün65k]. Markov [Mey68j, AKDR68, ADR69, BB69a, Gir68, Hus69, Mey67j, Mey67m, Mey68i, Mey68j, Wei67b]. markoviens [Wei67b]. mars [DE68]. Martin [Mey68e, Mey68k, Mey68i, Mey68j, Tan68]. Martingale [Cha67, dGM69, Hub69b, Mey67a]. Martingales [Mét67, dGM69, Hub69b]. Masse [Geo67, Wer64f]. Massen [Wer64h, Köl68d]. Master [CF69h]. Mathematical [Fic65a, Fic65c, Fic65b, yMO68, Jan69m, Kac69, EG69b, KT66c]. Mathematics [Ivă65k, Lum68a, Wer66v, Wer64j]. Mathematik [Alb68d, Tar67j, Wer66v, Wer64j]. Mathématique [Lum68a, EG69b]. Mathematische [KT66c]. mathematischen [Jan69m]. Mathieu [Lün69a, Lün69c, Lün69q]. Mathieu-gruppen [Lün69a]. Mathieugruppen [Lün69c]. Mathieurgruppen [Lün69q]. Matrix [Pey69e, Fic65i, Kat68b]. Matroids [HW69]. Matter [Ada64b, Ada64d, Ada66b, Ada66d, AC64a, AC64c, Bau66a, Bau66e, Ber68c, Ber68g, Cha66a, Cha66j, GKM68b, GKM68f, Her65a, Her65j, IR66b, Jan69c, Jan69h, Köl68a, Köl68e, Mey67g, Mey67h, Ser65a, Ser65l, Wer66o]. Maxima [Bil69]. Maximal [Fai67f, Fai67g, Jen67l]. Maximale [Jen67l]. Maximalisierung [Her65l]. Maximalstellen [Fie67]. Maximization [Her65l]. Maximum [WW69, Cel68, Fai67h, Fai67o, Fie67]. May [Fic65k]. Mazur [Ker67, Mau67a]. McMaster [BKW69]. Md [Fic65k]. means [Kre68, Pey69a]. measure [ADR69, Q. 66a, Kri69, Lum68j, Mat68, Par69g]. Measurement [Lum68h]. measures [FW68b, Jac67, Köh68, Lar69g, Lum68g, Mey67k, Mor69a, Mor69b, Mor69c, PP66f, PP66g, PP66i, Urb67i]. Mécanique [EG69c, EG69e]. Mechanics [EG69c, Sim68f, EG69e, Kac69]. medium [Ber68e]. Meeting [IR66f, Löb68]. mehrdimensionaler [Jan69m]. mehreren [Wer66j]. Mehrphasen [KT66d]. Mehrphasen- [KT66d]. Melnikov [Zal68k]. Memorial [Hil69b, Hil69c, Löb69]. Mengen [FW68b, Köh68]. Mengenlehre [Jen67m, BP66g]. Mergelyan [Wer64b]. Mesure [ADR69, Lum68h]. Mesures [Lum68g, Mor69a, Mey67k, Mor69b, Mor69c]. metalanguage [Jen67g]. Metasprache [Jen67g]. Method [And67c, BS67e, Fic65e, Fic65t, Hod69, Jan69d, Jan69b, Jan69f, Jan69m, Jen67e, KT66d, KT66b, Shi68h, Tho66a, Ura69, Wid69, Yor68b]. Methode [Jan69f, And67c]. Méthodes [EG69c]. Methods [EG69c, IR66f, Kap67b, Ver69, BB69a, Ber68h, But69, Coo69, D’J69, KT66a, Luk67, Lyc69, McK69, MD69, Nob69, Nør69, Pio69, Zal68g]. metric [BS67b, Q. 66a, GKM68i, Tro69o]. metrische [GKM68i]. MI [CK69]. Michigan [CK69]. Midwest [AM68, BBD^{+}69, BDD^{+}67]. Miltipliers [Par69d]. Minimal [Ivă65g, Man69a, ST69i]. Minimality [BH69k]. minimally [Lic69]. minimi [ST69i]. Minimization [Ivă65h, IR66d]. minimize [Sto66a]. Minkowski [Cha66g, Cha66g]. Minlos [Sch69]. mittel [Ber68e]. Mixed [Wil67f, McK69]. mixing [Kri69]. Möbius [CF69c, Lün65m, Lün65i]. Möbiusebenen [Lün65m, Lün65i]. model [Cha68, Dun68, Jen67c]. Modèles [EG69d]. Modell [Jen67c]. Modelle [Jen67l, Jen67h, Jen67i, Jen67j, Jen67n, Jen67m, Jen67b]. models [AT69, DMO67, Fen69, Fit69, Jen67h, Jen67i, Jen67j, Jen67l, Jen67b, Mor68, Tie69b, Ulm69a, Wei68, EG69d, Jen67n, Jen67m]. Modern [TA69]. modification [Lum68d, Nør69]. modified [Csö69]. Modular [Ser66a, Shi68c, Shi68a, Shi68l]. modulars [Lum68j]. module [Fai67a]. Modules [Fai67e, Bak69, BKKN67b, Cat68, Fai67c, Fai67i, Fai67k, Har67c, Lam66c]. moduli [Shi68f]. Mokobodzki [Del69c, Mey68c]. moments [de 68]. monadic [Slo68]. Mono [Her68c]. Mono- [Her68c]. monoid [CF69c, Dol66a, Dol66j, Dol66k, Dol66h]. Monoïde [CF69c]. Monoïdes [CF69e]. Monoidfunktoren [Dol66h]. Monoids [CF69e]. Monoidstruktur [Dol66a, Dol66j, Dol66k]. Monotonic [Gor69]. Montgomery [Jän68j, Jän68j]. Moore [Tho66a]. Morphismen [BP69k, BP69b]. morphismes [Gro68i]. morphisms [BP66h, BP69b, BP69k, Gro68i, Har66c]. morphologies [Dav68]. motions [Sel69]. Mr. [CMW68]. Muenster [Wer66v]. multi [ST69o, ST69n]. multi-vocal [ST69n]. multi-vocal-depinizioni [ST69o]. multidimensional [Jan69m]. Multiphase [KT66d]. multiples [HM68h]. Multiplication [BCH^{+}66, Tar67f, Shi68e]. multiplications [Shi68f]. Multiplicative [Dol66n]. multiplicités [SG65]. multiplicities [SG65]. Multiplier [Lar69b]. Multipliers [Lar69g, Par69f, Par69e, Lar69a, Lar69c, Lar69d, Lar69e, Lar69f, Lar69h, Par69b]. Multiplikative [Dol66n]. multistep [Lyc69, Nob69]. Multivoche [ST69n, ST69o]. Münster [Wer66v]. Musterbeispiel [Jän68j].

N [Jän68i]. N.A.T.O [Löb68]. N.J [BCH^{+}66]. N.Y. [Tro69q]. nach [Jän68i]. Natural [Eps69, Mey68d, Tro69h]. naturalness [BP66e]. naturel [Mey68d]. Natürlichkeit [BP66e]. near [BH69c, Pim69a]. Nearly [Cro69]. necessary [Kre69]. Negative [BH69m]. neighborhood [ST69b]. nell’intorno [ST69b]. nets [Gro68i]. networks [Ivă65c]. Neumann [Tar67n]. Nicht [Lün69l]. Nicht-Existenz [Lün69l]. Nichtexistenz [Lün69g]. nichtlinearer [Jän68l]. Nikodym [Geo67]. Nil [Fai67h]. nilpotence [Swa68i]. Niveaudurchgänge [Fie67]. Noetherian [Fai67i, Fai67p, Roo69]. noeuds [dR67b]. nombre [Lum68f]. Non [Fen69, Hoo69, MH69, Ded69, Hod69, Jac67, Ser65f, Ser65j, Yor68b, Zar69i]. Non-Abelian [Hoo69, Ded69]. non-homogeneous [Zar69i]. non-invariant [Jac67]. non-linear [Hod69]. non-Lipschitz [Yor68b]. Non-Standard [MH69, Fen69]. Noncommutative [Ser65f, Ser65i, Ser65j]. nonexistence [Lün69l, Lün69g]. Nonlinear [IR66e, Pim69b, D’J69, Dah69a, Jän68l, Pim69a, Pim69g, KT66e]. nonsteppoints [Lyc69]. Nörlund [Pey69b]. Norm [FB66b, Wer66d, Wil67f]. Normal [Isb67, Kne69, Tai68, Zar69l, Zar69f]. Normalität [Wer66p]. Normality [Wer66p]. Normalization [Zar69m, Nar66e]. normed [Wer66i]. Normen [Wer66s]. normierten [Wer66i]. norms [Wer66s]. Notation [BS67d, Mül66g]. Notations [Ivă65i]. Note [FJ68, GH69a]. Notes [Har66i]. Notions [Bor64j]. nouvelle [Mey69e]. November [CK69]. novembre [DE67]. ntegrodifferential [Jan69j]. Null [BP66h]. nullité [Cer68c]. nullity [Cer68c]. Nullmorphismen [BP66h]. Number [Cha66n, Shi68b, Cha66d, Cha66e, Cha66h, Erd69, Fie67, Fre69a, Lum68f, Lyc69, Mal68, Cha66m]. numbers [Beh69, Cha66c, Cha66i, LG69, May69e, May69n, Tro69f, Tro69h, Tro69n]. Numerical [LG69, Mor69d, Ura69, Gou69, Hod69, Pio69, Vit69, vF69]. Numerische [Wer66v]. Nürnberg [Urb67b].

o [ST69o]. object [Gro68j, Gro68k]. objects [AM69j, CS69a, Cha69a, Cha69b, Fit69, Mit69, Ton65c, Tro69k]. objet [Gro68j, Gro68k]. observables [Sim68e]. observations [Hei69c]. obstruction [Bre67c, Dol66b]. Obstructions [Tho69, Bar69c]. occurrences [Her65n]. October [CK69, Ivă65k, yMO68, Tar67j]. Octobre [DE69, DE68]. Offene [FW68q]. One [Mac69b, Frö68a, Frö68b, Lin69a, Hil69b]. one- [Lin69a]. Open [Fai67j, Gro68c, Zal68o, FW68q, Tar67a]. opening [Pol68]. Opérateurs [Blo68, BM68d, BM68b]. Operationen [Wlo69e]. operations [Ada64f, Ada66f, Ada69e, Wlo69e]. operator [Ber68a, Fic65r, Her65m, Her65q, Her65e, Her65f]. Operatoren [Stu69c, Ber68a, Stu69b, Stu69k]. Operators [Pim69b, Tar67g, Tar67j, Ati69, Blo68, BM68b, BM68d, Con68c, Dah69a, Din67, Fic65d, Fic65l, Fic65q, Gri69, Pim69f, Pim69c, Stu69b, Stu69c, Stu69e, Stu69k, Tar67b, Tar67d, Tar67f, Tar67i, Tar67h, Tar67k, Tar67l, Tar67m, Tar67n]. Optimal [Kou67, Lyc69]. Optimierung [KT66e]. Optimierungstheorie [KT66c]. Optimization [KT66e, KT66c]. optimum [Kem69]. orbit [Jän68j]. orbital [BH69n]. Orbitraum [Jän68j]. Orbits [Jän68g]. order [BM68b, BM68d, But69, Fic65a, Gub67, Lün65b, Lün69m, Lyc69, Shi68l, Tro69j, de 68]. ordered [Ivă65g, Kei68]. Ordering [Tro69j]. orderings [Tro69p]. Ordinal [Law69b]. ordinals [FJ68, Tro69p]. ordinari [ST69p]. ordinary [BS67e, Hod69, Lin69a, Pio69, Stu69i, ST69p]. Ordnung [Lün65b, Lün69m]. ordre [BM68b, BM68d]. Orientation [Con68d]. Oriented [Tho66f, May69k, May69m]. orientierten [May69k, May69m]. origin [Pim69i]. Orthogonal [Fic65l, Mül66j, Fic65s, Kne69, Lum68g]. orthogonales [Lum68g]. oscillation [Pim69c]. Oslo [AL69]. Other [Lün69o]. outline [Lin69d]. ouvert [Pol68]. ouverts [Gro68c]. Overview [Her65r].

P. [Bad67]. packings [Bei69]. pair [Lar69f]. paper [Oht68]. parabolic [Gor69, Gub67, Jan69e, McK69, Sie68]. parabolische [Sie68]. parabolischer [Jan69e]. Paradigm [Jän68j]. parallel [KT68g]. parameter [Ton65i]. parfaites [Dol68b]. Partial [FB66k, Bil69, Dol68a, Gou69, Gub67, Kis68, Stu69f, Tre69, Sch68]. partially [Ivă65g]. particle [Sim68h, Spi69]. particles [Sim68b]. particularly [vF69]. partieller [Stu69f]. partielles [Dol68a, Kis68, Sch68]. Partitions [Mor68]. past [Dun68]. patterns [Gre67, Grü69b]. PCO [Fic65s]. PDE [Fic65a]. Peaceman [Wid69]. Peak [Zal68l]. Perfect [PP66i, Dol68b]. Periodic [BH67a, Con67k, Con67p, BH67b, Stu69f, Yos68, HM68k]. periodicity [Ada64h, Ada66i, Ada69h]. periodische [Stu69f, HM68k]. permutation [Bou69, Fre69a, Lün69i, Lün69j]. Permutationsgruppen [Lün69i, Lün69j]. Perspectives [Lün65n]. Perspektivitäten [Lün65n]. perturbation [Str68a]. petite [Cer68e]. physical [Sim68e]. Physics [Fic65c, Fic65b, Fic65a, Jan69m]. Physik [Jan69m]. Picard [Swa68q]. plainness [Lün65d]. planar [Bar69a]. plane [Lün65b, Lün65m, Lün65n, Lün69p, Lün69m, Shi68c]. planes [And69, Lün65c, Lün65i, Lün65j, Lün65o, Lün69d, Lün69k, Lün69n]. plates [Fic65b]. plongée [Cer68b]. plurisousharmonique [Hen68]. plurisousharmoniques [Lel68a]. plurisubharmonic [Hen68, Lel68a]. poids [Exb68]. Poincaré [Zis69]. point [BH69o, BM68e, Con67c, Dol66g, Lün69p, Mat68, Pim69a, Pim69k, Zar69i]. points [Alb68a, Cha66g, Cha66m, Fie67, Jon68a, Shi68l, Tan68, Zal68l]. Poisson [Jan69b]. Poissonschen [Jan69b]. polaritäten [Lün69n]. polarities [Lün69n]. polyhedra [Cro69]. Polynomial [Wer69a, Ivă65e]. polynomial-logical [Ivă65e]. polynomials [Cas69, Zar69p]. Polytechnique [Bor64e]. polytopes [Bar69a]. ponctuel [BM68e]. Pontriagin [Cel68]. population [Dun68]. Pósa [Kro69]. posed [Fic65u]. positifs [Mey69a]. Positive [Urb67c, BH69l, Cha66l, Fic65l, Mey69a, Mor69c]. positiven [Cha66l]. positives [Mor69c]. possède [Lum68f]. Possible [Mac69c]. Postnikov [Dol66o, Dol66p, Tho66a, Tho66b]. Postnikov-Faktoren [Dol66o]. Postnikov-Invarianten [Dol66p]. potential [Bau66g, Bau67, Bau68, Bre67f, Gub67, BM68e, vH68]. Potentiale [Bau66h]. Potentials [Oht68, Bau66h, Mey68h]. Potentialtheorie [Bau66g, Bau68, vH68]. Potentiel [Bre67f, Mey69c]. Potentiels [BM68e, Mey68h]. pour [Cer68a, Cer68c, Cer68d, Hub69a, dR67b]. Poussin [Ber68d]. Prädikatenlogik [Her65t, Her65m]. pre [Fre69b]. pre-completeness [Fre69b]. pre-limits [Fre69b]. precise [BP69c]. predicate [Her65t, Slo68, Her65m]. Prediction [Urb67b, Urb67d, Lum68i, Lum68i]. Predictors [Urb67e]. Preliminaries [Alb68c, BH67f, Frö68e, Nar66f, Oor66d, Hei69c]. preliminary [BS67d, Fic65n]. preparation [Nar66h]. Preparations [Bau66i, May69o]. Preparatory [Ber68i]. preschemes [Har66a, Har67a]. presentation [CMW68]. Presentations [Bor64e, dRMK67, Mey69b]. Presented [Ivă65k, Tro69q]. preserving [Con68d, Lam66h, Lam66j]. preventing [Pim69h]. previous [Mey69b]. Primary [Ada64f, Ada66f, Ada69e]. prime [Cha66c, Cha66e, Fai67p]. primes [Cha66f]. Primzahlen [Cha66c, Cha66f]. Primzahlsatz [Cha66e]. Princeton [BCH^{+}66]. Principal [Tho66g]. principle [Cel68, Fic65h, Tro69m]. principles [Tro69a, Tro69q]. printemps [Bor64e]. Pro [AM69j]. Pro-objects [AM69j]. probabiliste [Cai69]. probabilistes [CF69a]. Probabilistic [Cai69]. Probabilités [DE67, DE68, DE69, dC67b]. Probability [Kap67b, KK69, Luk67, Pad69, PP66g, WW69, BKW69, dC67b]. Problem [Lar69b, Wer66h, Cai69, Guy69, Hei69b, Hör69, Ivă65d, Jan69f, Pim69c, Pim69k, Tre69, Wer66f, Wer66s, Zal68m, Bau66c]. Probleme [Jan69m, Cai69]. problèmes [Gro68c, CF69i]. Problems [KT68j, Bau66c, CF69i, Fai67j, Fic65e, Fic65k, Fic65q, Fic65u, Gri69, Gro68c, Hei69a, Jan69b, Jan69g, Jan69k, Jan69m, Kac69, KM68, Law68, Loy67, Stu69a, Stu69e, Stu69h, Stu69i, Ura69, Urb67d]. Proceedings [BKW69, CK69, Hil69b, Hil69c, Löb69, AL69, Löb68]. process [ADR69, Csö69, Mey68d, Mey68i]. processes [AKDR68, Ber68b, Ber68h, Cam69, Fie67, Hus69, Kel67, Lum68i, Mey67a, Mey67i, Mey67j, Mey67m, Mey68b, Mey68j, Pyk69, Spi69, Wei67b, Mey68k, Mey69a]. Processus [Mey67i, Mey67j, Mey68k, Mey68i, Mey69a, AKDR68, ADR69, Lum68i, Mey67a, Mey67m, Mey68b, Mey68d, Wei67b, Mey68j]. Procomplexe [Gro68j]. product [PP66f]. products [Fai67l, Sim68j, BP66i, Dol66q]. Produkte [BP66i, Dol66q]. Profinis [Ser65b, Ser65k, Ver65, Ser65h]. profinite [Ser65b, Ser65h, Ser65k, Ver65]. profiniteness [AM69k]. Programmierung [Wer66t]. Programming [Dah69b, Ivă65k, IR66f, Ivă65e, Ivă65j, IR66a, Wer66t]. programs [Mac69c]. progress [Ada64c, Ada66c, Ada69c]. progression [Cha66c, Cha66c]. project [Bau66g]. projectif [Roo68]. Projective [DMO67, And69, FW68d, Har66c, Lün65c, Lün65a, Lün65n, Lün69f, Lün69n, Lün69m, Roo68, Sim68g, Zar69f, Wlo69h, Wlo69i, Wlo69g, Wlo69f]. projektive [FW68d, Lün65c, Wlo69h, Wlo69i, Wlo69g, Wlo69f]. projektiven [Lün65a, Lün65n, Lün69f, Lün69m]. projektiver [Lün69n]. Prologue [Lar69h]. prolongated [ST69l]. prolongation [BH69l, ST69l]. prolongazionali [ST69l]. Prolongazioni [ST69l]. Prolongements [Mor69b]. proobjet [Gro68k]. Proof [Tro69q, Zar69n, Fef68, Her65c, Her65o, Her65r, Jän68j, Kar68, Mey69e]. proper [Q. 66b]. properness [Lam66l]. Properties [Mül66k, Bar69a, BH69j, Bor67b, FB66g, GKM68e, Gro68l, Har67b, Lün69o, Nar66d, Wei68, Wil67g, dR67a]. property [Dol66f]. propositional [Ros68a]. Propriétés [Gro68l, dR67a]. prospective [Ada64c, Ada66c, Ada69c]. Prozessen [Fie67]. Pseudo [FB66m, FB66l, Ivă65j, Ivă65h, Ivă65l, IR66c, IR66d, IR66e, Pol68, Ivă65k, IR66f]. Pseudo-Boolean [Ivă65j, Ivă65h, Ivă65l, IR66c, IR66d, IR66e, Ivă65k, IR66f]. pseudo-convex [Pol68]. pseudo-convexe [Pol68]. Pseudo-topological [FB66l]. Pseudo-topologies [FB66m]. pseudoconvex [Nor68, Ros69]. pseudoconvexes [Nor68]. punkte [Lün69p].

Quadrate [BP69j, BP69d, BP69f]. Quadraten [Cha66l]. quadratic [Bak69, Cha66b, HM68j, Kou67, Cha66l, HM68l]. quadratische [Cha66b, HM68j, Cha66l, HM68l]. Quantified [Pre68]. quantifier [Slo68]. quantifiers [Kei68]. Quantitative [Wer66q]. Quantum [Sim68f]. Quasi [Fai67k, Fai67a]. Quasi-Injective [Fai67k, Fai67a]. Quelques [Mey68g]. question [Jen67k]. questions [Pim69j, Sma69, Tar67a, Zal68o]. Quotient [Fai67e, Fai67l, Swa68r, Fai67f, Fai67g, Wer66k].

Rachford [Wid69]. Radical [Fai67m]. Radon [Geo67]. Radon-Nikodym [Geo67]. Ramseyxo [Beh69]. Rand [Stu69h]. Rand- [Stu69h]. Random [Spi69, Cam69, Csö69, Fer67a, Fer67b, Geo67, Hub69a, Mor69b, Mor69c, Del69a, Del69b, Mor69a, Del71]. Randwertaufgaben [Jan69b, Stu69e, Stu69i, Gri69, Jan69k, Stu69a]. Rang [Her65n]. range [Tar67g]. Rank [AC64h, AC64g, AC67d, Hes69, AC64e, AC67c, AC67f, Her65n]. rapport [Lum68h]. rate [Kem69]. rates [KK69, MD69]. Rational [Fai67n, Zal68c, Cha66h, Dol66k, Wer66a, Zal68m]. rationale [Cha66h]. rationalen [Dol66k, Wer66a]. Raum [Fer68d]. Räume [Bau66g, Köh68, Stu69d, Wer64e, Bau66f, FW68a, FW68j, FW68p, FW68s, GKM68i, Her68e, Stu69j, Wlo69d, Wlo69c, Wlo69i, Wlo69g, Wlo69f]. Räumen [Gri69, Lün69f, Wlo69e, Stu69h]. Raumes [Lün65a, Wer66i]. Rayleigh [Fic65e]. Real [Con67l, Nar66g, Cer68f, Her68e, Tro69f, Tro69j, Tro69n, May69l, May69m, Tit67c]. Réarrangements [CF69f, CF69i, CF69i, CF69f]. Recherche [Lum68a]. reciprocity [Cha66b]. reconsidered [Pim69e]. Reconstructing [GH69b]. reconstruction [Man69c]. Rectifications [Mey69b]. Recurrent [AKDR68]. récurrentes [AKDR68]. rècurrents [ADR69]. recurring [ADR69]. recursion [Kre68]. Recursive [Cro68, Röd68]. Rédigé [Lum68a]. reduced [Mey67k]. reduction [Con67d]. Réduites [Mey67k]. Reell [Tit67c]. Reell-analytische [Tit67c]. Reelle [May69l, May69m]. réelles [Cer68f]. Reellkompakte [Her68e]. References [Wer64i]. reflections [Her68c, Her68d, Her68b]. Reflexionen [Her68d, Her68b]. Regeln [Her65k]. regions [Cha66g]. regular [Buc69a, Cro69, Her68e, PP66g, Roo67, Stu69i]. Reguläre [Her68e, Stu69i]. Regularität [BKKN67a]. Regularitätskriterien. [BKKN67f]. Regularity [Fic65m, Fic65n, Fic65o, BKKN67a, Fic65f, BKKN67f]. rekursiver [Röd68]. related [Cer68b, Cho66]. Relatif [Gro68e, Gro68c, Gro68i, Gro68l]. Relation [CF66d, Har67e, Shi68l, Swa68p, Her65t]. Relationen [May69e, May69n]. Relations [Bri69, CF69g, Swa68s, Tar67i, Aig69, CF69e, May69e, Tro69j, GKM68c, May69n]. relationship [Jen67f, Lam66i, Ulm69c]. Relationships [Fer68b, Fer68a]. relativ [Her65h]. Relative [Gro68e, Lin69e, Buc68, Gro68c, Gro68l, Her65h, Gro68i]. Relativistic [Sim68i, Sim68b, Sim68h]. relèvement [Gro68i]. remark [Kat68c]. Remarks [Cho66, Eng68, Tar67d, Cha68, Pey69c, Shi68g, Swa68f, Tar67a, Bad67]. Remarques [Bad67]. Renewal [Ken67]. report [Jän68d, HM68c, Jän68b]. Reports [AM68, BBD^{+}69, BDD^{+}67]. Reprensentation [Mül66l]. représentabilité [Gro68g]. representability [Cha66l, Gro68g]. Representation [Kap67a, Mor69c, Sie68, Urb67f, Fic65s, Mor69c]. Representations [Sim68j, Con67l, Mül66h, Shi68j, Sim68d, Sim68g, Tit67d]. representative [Lum68j]. Research [Hil69b, Hil69c, Löb69, yMO68, Wer64j, Lum68a]. resemble [Ken69]. Residual [Har66h]. residues [Cha66l, Har66i]. resolutions [AM69g]. résolvante [Mey68a]. Résolvantes [Wei67a, Mey68c]. resolvent [Mey68a]. resolvents [Mey68c, Wei67a]. respect [Lum68h]. Reste [Cha66l]. restricted [Her65b, FW68a, Wer64a]. result [Kat68c, May69g, Mey69d, Mey69c]. résultat [Mey69d, Mey69c]. Resultate [Jän68d]. results [AM69l, Fic65m, Jän68d, Kri69, Lin69e, Pad69, Ric69, Ser66b, Str68a, Swa68n, ST69m, Tar67h, Zal68h]. Retournement [Wei67b, CMW68]. revêtement [Cer68a]. revidierten [KT66b]. revised [KT66b]. revisited [Tro69m]. Reziprozitätsgesetz [Cha66b]. Riemann [yMO68, Oht68, Con67k, EE69, Fuj68, Hei69b, Hei69d, Tan68, Zar69o]. Riemannian [GKM68h, GKM68j, GKM68i]. Riemannsche [GKM68h, GKM68j, GKM68i]. Riesz [Ber68e]. Rieszschen [Ber68e]. Rigen [BKKN67e, BKKN67e]. right [Par69g]. ring [Con67m, Con67o, Con67n, Dav68, Fai67a, Frö68b, Gro68a, Shi68f]. ringed [Gro68i]. Rings [Fai67e, Buc69a, Fai67d, Fai67f, Fai67g, Fai67i, Fai67l, Fai67m, Fai67p, Fai67o, Nov68, Roo67, Roo69, Swa68c, Zar69d]. Risultati [ST69m]. Ritz [Fic65e]. Roch [Zar69o]. Rodrigues [Mül66m]. role [Dol66g, Tro69m]. Rolle [Dol66g, ST69u, ST69u]. Rolle-type [ST69u]. rotations [dR67d]. rules [Her65k, Tak68a]. Runge [But69].

Saarbrücken [Köl68d]. Samelson [Jän68j, Jän68j]. sample [Csö69]. satisfiability [Her65o, Her65i, Her65h]. Satz [Cha66c, Cha66h, Cha66f, Cha66g, Cha66i, FW68o, FW68e, FW68f, Geo67, Jän68j, Wer64b]. Sätze [Lün69h]. Sazonov [Sch69]. scaling [Wid69]. Scan [Bau66b]. Scandinavian [AL69]. Scheibendiagramme [Jän68k]. schema [Tak68a]. Schemata [Jan69e, Jan69i, Jan69i, Jan69l]. scheme [Gra69, Zal68n]. schemes [Gor69, Jan69e, Oor66a, Oor66b, Oor66c, Jan69l]. Schmidt [FW68l, Swa68o]. Schmidt-Abbildungen [FW68l]. Schönflies [Cer68c, Cer68d]. School [Löb68]. Schröder [Tar67l]. schwach [May69k, May69m]. schwache [FW68g]. schwachen [Jan69f]. Sciences [IR66f, yMO68]. Scotland [Mor69d]. Seattle [Hil69b, Hil69c, Löb69]. second [And68, BS67e, BM68b, BM68d, Gub67, Nob69, Yor68b]. second-order [BM68b]. secondary [Pim69h]. Section [Bou69, Ber68i, Cer68a, Mey69e, Pim69e]. selection [Her65m, Her65q, Her65e, Her65f]. Semantics [Bar68b, Law68, Lin69b, Lin69d, Lin69e]. Semester [Wer66v, Bau66g, Lün65g]. Semi [BH69i, Dol66a, Mey67l, dC67a, Ber68a, BH69j, Dol66c, Dol66h, Sim68j, Stu69e]. semi-direct [Sim68j]. Semi-Dynamical [BH69i, BH69j]. semi-elliptic [Stu69e]. Semi-exact [Dol66a, Dol66c, Dol66h]. Semi-groupes [Mey67l, dC67a]. semi-groups [Ber68a]. semidefinite [ST69c]. semielliptische [Stu69e]. Semigroups [BH67a, BH67b, BH67c, Mey67l, dC67a, ST69o, ST69n]. Semigruppi [ST69o, ST69n]. Séminaire [Bor64e, dC67b, DE67, DE68, DE69, Lel68b, dRMK67]. Seminar [Alb68d, AM68, BBD^{+}69, BCH^{+}66, Con67p, Jon68b, Lam66d, Tie69c, dRMK67, Har66i, HG67, dC67b, Bau68, BDD^{+}67, BCH^{+}66, Eck69, Tar67j, Tho66h, Wer64j]. Seminary [Bor64e]. Semiprime [Fai67o]. semiprimitivity [Fai67m]. semiregular [Köh68]. semiregulärer [Köh68]. Semisimple [HC68a, Fai67i]. Semitopological [BH67a, BH67b, BH67c]. Semiweak [Fic65p]. sentences [Mal68]. senza [ST69d]. separable [Tro69o]. Separately [Sic68, Ava67]. separately-harmonic [Ava67]. Separation [Bau66j]. séparément [Ava67]. September [IR66f]. sequence [Bor67c, Dol66q, Dol66h, Swa68m]. Sequences [Tro69k, BP69d, CF69f, Gro68m, Ken67, May69i, Tro69c, Tro69h, Urb67e, Urb67c]. Sequenzen [May69i]. Series [Cam69, Urb67f, Fer67a, Fer67b, Ura69, dR67a, dR67a, Fer67a, Fer67b]. Serre [Swa68t]. Set [FK69, BH69c, BH69o, BP66g, Cha66c, Con67c, FW68e, Jän68j, Jen67m, ST69b, Wer64b]. Set-Theoretical [FK69]. Sets [PP66e, Alb68a, BH69h, Cor68, Del69a, Del69b, Ivă65g, Jän68l, Mac69a, Nar66d, Nar66g, PP66d, ST69a, ST69h, ST69i, ST69l, ST69q, ST69r, Yor68c]. Several [Fre69b, Cas69, Ric69, Wer66j]. shall [AM69g]. shape [GKM68c, Wer66d]. sheaf [HM68h, Hoo69, May69l, May69m, May69e]. sheaves [Bor64i, Fer68g, Gro68a, Gro68b, May69k]. Siegel [Shi68a]. Simple [dRMK67, Lum68i, Mau67b, Mau67c, Tit67d, Mau67b, Mau67c]. simple-stationary [Lum68i]. simplement [Lum68i]. simplest [Jan69e]. simplex [KT66b]. Simplexverfahren [KT66b, KT66b]. Simplicial [TV69, Tie69g, Qui67b, And67c]. Simpliciale [And67c]. Sims [Lün69b, Lün69b]. Single [Dun68, Coo69]. singular [Ber68d, Fai67h, Her66a, Jän68g]. Singuläre [Ber68d, Jän68g]. Singularitäten [HM68o, HM68p, HM68a]. Singularités [Cer68f]. singularities [HM68a, HM68o, HM68p, Cer68f]. sistema [ST69f, ST69g]. Sistemi [ST69p, ST69e, ST69o]. size [Csö69]. small [Cer68e, Lam66a, Lün69c]. Sobolev [Fic65r, Gri69, Stu69e, Stu69h, Stu69j]. Sobolewsche [Stu69e, Stu69j]. Sobolewschen [Gri69, Stu69h]. sollen [Her65h]. Solomentsev [Hei69e]. Solution [Mor69d, Pim69i, Gou69, Jan69f, Jan69m, LG69, Ura69, Vit69, Jen67k]. Solutions [BH69m, Cel68, Fic65g, Fic65p, Kis68, Pim69a, Pim69g, Yor68d, Yos68]. solving [D’J69, Nob69]. Some [Cha68, Erd69, Har67f, Kac69, KM68, Law68, Loy67, Wil67g, Zal68o, Zal68p, Bar69a, Coo69, Exb68, FB66m, Nov68, Par69d, Ros68a, Sic68, Swa68l, Jan69g, Mey68g]. Sommer [Wer66v]. Sommer-Semester [Wer66v]. Sommersemester [Bau66g, Lün65g]. sommes [dR67a]. Sonderfälle [Köl68f]. Sous [Ram68, Cer68e, Hen68]. Sous-ensembles [Ram68]. sous-espaces [Hen68]. sous-variétés [Cer68e]. space [Alb68b, Ati69, BS67a, Blo68, Cer68f, Fer68d, FW68c, Jan69f, Jän68j, Kis68, Köh68, Lün65a, McK69, Roo68, Sch68, Stu69g, Wer64e, Wer66i]. Spaces [FB66b, Nar66c, Tho66h, Bau66f, Bau66g, Bec69b, Ber68b, Ber68h, BS67b, BM68a, BM68b, BM68c, Bor64c, Bor64a, Bor64e, Bor67a, Bor67b, Bor67e, Bre67d, BDCK67, Cer68b, CC68, Cor68, Dol68a, DHS69, EK69, Exb68, Fic65j, FW68a, FW68j, FW68k, FW68p, FW68s, FW68u, FB66c, FB66m, FB66l, Gri69, GKM68i, Her68e, Kap67a, Kri69, Lar69g, Lel68a, Lum68a, Lum68c, Lum68d, Lün69f, May69a, Nar66b, PP66c, Pim69g, Sel69, Stu69e, Stu69h, Stu69d, Stu69j, Tho66g, Ton65d, Tro69o, Wil67b, Wil67c, Wil67f, Wil67g, Wlo69b, Wlo69d, Wlo69c, Wlo69e, Wlo69h, Wlo69i, Wlo69g, Wlo69f, dR67d, Ser68]. special [FB66h, Hed69, HM68d, Jän68h, Par69d, FW68r, HM68n, HM68m, Köl68f]. Species [Tro69l, Tro69m, Dun68]. specific [Jän68a, Tar67i]. spectra [Bre67d, May69a, Tie69d, Tie69g]. Spectral [Bor67c, Urb67g, Dol66q, Dol66h, Loy67, Roo67, de 68, Bor64b, Bor64a, Bor64f]. spectrale [Bor64b, Bor64a, Bor64f]. spectraux [de 68]. Spektralsequenz [Dol66h, Dol66q]. spezielle [HM68d, FW68r, HM68n, HM68m]. speziellen [Jän68h]. spezieller [Jän68a]. Sphären [Jän68d, HM68c, HM68a, HM68k, HM68o, HM68p]. sphere [Cer68g, Cer68b, KT68k, Tho66f, Cer68g, Cer68b]. spheres [Cer68f, HM68c, HM68a, HM68k, Jän68d, Cer68f, HM68o, HM68p]. Spherical [Mül66n, Mül66e, Mül66h]. Spiele [Bie67]. splitting [Hsi69]. Sprache [Jen67d]. Spreads [Tro69n]. spring [Bor64e]. square [Mey67a, Tar67f]. squares [BP69d, BP69f, BP69j, Cha66l, Olu69]. Stabilità [ST69q, ST69r]. Stability [BS67c, BH69n, Pio69, Yos68, Lin69a, Sma69, ST69q, Vit69, Yor68a, ST69r]. Stable [Ada64a, Ada64g, Ada64i, Ada66a, Ada66g, Ada66h, Ada69f, Ada69g, AM69l, Tie69c, Ada69b, Nør69, BDCK67]. Stables [BDCK67]. Stably [KT68k]. Stallings [Ker67]. Standard [MH69, Par69g, BB69b, Fen69, Mey67i, Lam69]. starke [FW68h]. start [BH69o]. Statement [Ser66b]. statements [Jan69a]. stationary [Kel67, Lum68i]. stationnaires [Lum68i]. Statistical [EG69c, EG69e, Kac69]. Statistique [EG69c, EG69e]. Steinitz [Bar69a]. step [Coo69, Jan69d, Jan69b, Jan69m, Lin69a, Pio69]. steps [Jan69e]. stetige [Stu69g, Wer66r]. Stetigkeit [Wer66q]. stiff [Lin69a]. stochastic [Kou67, Mey67b, Mey67c, Mey67d, Mey67e]. stochastiques [Mey67b, Mey67c, Mey67d, Mey67e]. Stoll [Roo68]. Stone [dC67c]. stopping [Mey69d]. Strasbourg [DE67, DE68, DE69, Lum68a]. Strauss [Kat68c]. strictement [Nor68]. strictly [Roo69, Nor68, Wer66s]. Strikt [Wer66s]. Strong [Tar67k, Tar67o, FW68h]. Strongly [Fic65q, Ros69, Roo67, ST69b]. strongly-invariant [ST69b]. Structure [Fai67p, AW69, Dol66a, Dol66j, Dol66k, KT68b, KT68h, Lün69q, ST69s]. Structures [EG69e, Dol66n, Ehr69, Hir69, KT68d, Lün65l, Lün69e]. Struktur [Lün69q]. Strukturen [Dol66n]. Struttura [ST69s]. studies [Wer66b]. Study [BCH^{+}66, Löb68, Cai69, Lum68f, Lum68e]. Subadjoint [Zar69p]. subalgebras [Man69a, Ton65g]. subcategories [Lam66f]. subcategory [Tie69a]. Subdivision [Cer68f]. Subgroups [Ton65g, Kne69, Lün65h]. subject [Tro69b]. sublinear [Pim69c]. submodules [Fai67n]. subregion [Fuj68]. subsets [Ram68, Sic68]. subspaces [Hen68]. Substitution [Tar67l, Her65n, Her65s, Tar67b, Tar67m, Her65n, Her65s]. Subvarieties [Cer68e]. sufficient [Kre69]. Suites [CF69f, Gro68m]. sum [Cha66l]. Summability [Pey69d, Pey69b]. Summary [Pim69j, Tar67e]. Summe [Cha66l]. Summer [Löb68, Wer66v, Bau66g, Lün65g, Tro69q]. sums [Bil69, Law69b, dR67a]. SUNY [Tro69q]. sup [Lam66f, Lam66j]. sup-complete [Lam66j]. sup-dense [Lam66f, Lam66j]. sup-preserving [Lam66j]. Super [Sie68]. superharmonic [Mae68b, Bau66h]. Superharmonische [Bau66h]. superharmonischer [Sie68]. superieur [de 68]. superlinear [Pim69c]. supermartingale [Mey68d]. Superposition [Tar67m]. support [BM68e]. Supports [Kis68, Kis68]. surely [Pyk69]. surface [Con67k, Fuj68]. Surfaces [Hei69d, yMO68, EE69, Hei69b, Oht68, Tan68, Zar69j, Zar69a]. surmartingale [Mey68d]. Survey [Her65o, Bei69, Cro68]. Suzuki [Lün65d, Lün65i, Lün65j, Lün65g, Lün65h, Lün69g]. Suzuki-Typ [Lün69g]. Suzuki-type [Lün69g]. Suzukigruppen [Lün65i, Lün65j, Lün65g, Lün65d, Lün65h]. Swiss [Wer64j]. switching [CF69i, Ivă65b]. Symbolic [Löb68]. symbols [Bas69, KT68b]. symmetric [Sel69]. Symmetriegrad [Jän68d]. symmetry [Jän68d]. Symposium [BKW69, Kap67b, yMO68]. Syntax [Bar68b]. System [Jen67d, Fic65i, Lum68h, Spi69, ST69f, ST69g, Zar69c]. Systeme [Stu69f, Lum68h, KT66e, Wer66u]. Systems [BS67c, BH69i, Ivă65l, Jon68b, KT66e, PP66e, BS67a, BS67b, BH69j, Dav67, Fic65f, Fic65g, Fic65k, Fic65p, Gor69, Jon68a, Lam69, PP66d, Pio69, Sma69, Stu69f, ST69e, ST69p, ST69o, Wer66u, Zar69k, Zar69m]. Szego [Hei69e, Urb67h].

Tabellen [Tit67d]. Tables [Tit67d]. tangent [HM68h]. tangential [Fic65o]. Tangentialbündels [HM68h]. Tauberian [Pey69f]. Technische [Wer64j]. Technology [Wer64j]. temps [CMW68, Mey69d, Wei67b]. tending [Yor68d]. Teoremi [ST69t, ST69u]. term [Her65b, Her65d, Her65e, Her65f, Her65t, Köl68g, Her65q, Her65p]. Terme [Her65h]. Terminology [Lam66k, BS67d, Ivă65i]. Termisomorphismen [Her65p]. Termlogik [Her65e, Her65f, Her65d, Her65q, Her65t]. terms [Her65h, Köl68b]. test [Stu69g]. Testraum [Stu69g]. tge [Gro68c]. th [Wei67b]. Their [AC64d, Löb69, AC67b, Bau66g, Hil69b, Hil69c, Ken67, Lün65g, Pim69b, Tit67d, Ulm69c]. Theorem [Cha67, Hei69e, Her65o, Alb69, AM69k, Bar69a, Bau66j, Bor64g, Car69, CF69h, Cer68c, Cer68d, Cha66d, Cha66e, Cha66h, Cha66f, Cha66g, Cha66i, Del69c, FW68o, FW68f, FW68g, FB66i, Geo67, Har66d, HM68d, Hsi69, Igo68, Jän68c, Jen67h, Kar68, Ker67, Kre69, Kro69, Lum68d, Mau67a, Mey67f, Mül66b, Mül66a, Mül66l, Nar66e, Nar66h, Pol68, Sch69, SD69, Sun68, Swa68t, Tro69a, Urb67a, Urb67g, Urb67h, Wil67d, Wlo69a, Yor68c, Zal68e, Zal68k, Zal68q, Zar69o, dC67c, dR67a, dR67b]. théorème [Bor64g, Cer68c, Cer68d, Del69c, Igo68, Ker67, Lum68d, Mau67a, Mey67f, Pol68, Sch69, Sun68, dC67c, dR67a, dR67b]. théorèmes [Bad67, Gro68g, Mey69e]. Theorems [Ada64h, Ada66i, Ada69h, Lam66l, AM69b, Bad67, Cor68, Csö69, GKM68k, Gro68g, Kur68, Mey69e, Nar66a, Oor66c, Pey69f, Shi68e, Swa68g, Swa68o, ST69t, ST69u, Tar67c, Tar67f, Tar67n, Lün69h]. theoretic [Cha66m, Man69b]. Theoretical [FK69, Shi68d]. Theorie [FW68j, Stu69a, Bre67f, Gir68, Lum68h, Lum68i, Mau67c, Mey67m, dGM69, Hub69b, Mau67b, Mey68b, Mey69c, Kar69, Bau66b]. Theories [Bre67b, CF66d, Ros68b, Law68, Qui67b]. Theory [Ada64a, Ada64g, Ada64i, Ada66a, Ada66g, Alb68d, BKW69, BS67c, Bor67d, Cha66n, CK69, CS69c, Eck69, Hil69b, Hil69c, Löb69, MT69, Nar66c, Shi68b, Swa68h, Tie69c, Urb67b, Wer66v, Zar69a, Ada66h, Ada69b, Ada69f, Ada69g, And69, Bau66b, Bau66g, Bau67, Bau68, Bre67c, Bre67f, BP66g, Buc69b, Cha66d, Cha68, CS69b, CF66b, CC68, Dol66b, Eng68, Erd69, Fef68, FK69, FW68j, Frö68d, FW69, Gir68, Gre67, Gub67, dGM69, Ham69, Hsi69, Hub69b, Ivă65c, Ivă65a, Jan69k, Jen67m, Kar69, Kni68, KM68, Kre68, KT66c, LG69, Lar69a, Luk67, Lum68h, Lum68i, Mac69b, Mar69, Mau67b, Mau67c, Mey67m, Mey68b, Mey69c, Pim69a, Qui67a, Sel69, Sha69, Shi68e, Str68a, Stu69a, Swa68e, Swa68s, Tro69b, Tro69f, Tro69q, Tro69n, Zar69j, vH68]. thesis [Köl68d]. thin [Fic65b]. Thom [CF66b]. Three [Ded69, Cer68g, McK69, Löb69]. three-dimensional [Cer68g]. Tightening [Köl68g]. time [CMW68, Hod69, Mey69d, Wei67b]. tipo [ST69u]. togic [Her65d]. Tonnelierte [FW68s]. tools [Dah69b]. Topics [Wil67h, Cho66, Alb68d, Bor67d]. topological [Cat68, FB66l, GKM68c, Kri69, Lar69g, Lel68a, PP66c, Swa68f, Swa68s, Wil67g, FW68t, FW68u, Her68c, Her68d]. Topologie [Bor64e, Bre67f, FW68g, dRMK67, FW68h, Fer68f]. topologies [FB66m, Wil67a, Wil67f]. topologiques [Cat68, Lel68a]. Topologische [FW68t, FW68u, Her68c, Her68d]. topologischer [GKM68c]. Topology [Tro69o, Ati69, Bre67f, Fer68f, FW68g, FW68h, Bor64e, dRMK67]. topos [Gro68i]. Torsion [May69f, dR67c, dRMK67, KT68g, dR67b, dR67c]. Total [Fer68f, Fer68c]. totale [Fer68c, Fer68f]. Trace [Zar69q, Con67q, Fic65r, Hen68]. traces [Hen68]. traiettorie [ST69f]. trajectories [ST69f]. transform [KMP67]. Transformation [Tar67o, Ton65f, Ton69, Jän68i]. transformations [Mül66j, Pey69e, ST69o, ST69n, Ton65i]. Transformationsgruppen [Jän68i]. transforms [Lar69e]. transition [dC67a, dC67a]. transitive [Lün69l, Lün69g, Lün69k, Lün69j]. transitiver [Lün69l, Lün69g]. transitivité [Gro68m]. transitivity [Gro68m]. translation [Lün65j, Lün65o]. Translationsebenen [Lün65j, Lün65o]. transmission [KK69]. transmitting [Kem69]. transportation [Ivă65d]. Trasformazioni [ST69n, ST69o]. treatment [Ber68h, vF69]. tree [HM68f]. triangles [Dub68]. Triangulated [Buc69b]. triple [Man69b]. tripleability [Dus69].

Triples [Eck69, Man69a]. trois [Cer68g]. Troisième [Bor64e]. truth [Jen67f]. Tschebyscheff [Wer66q, Wer66r, Wer66j, Wer66e, Wer66t, Wer66u]. Tschebyscheff-Approximation [Wer66q, Wer66r, Wer66j, Wer66e, Wer66t]. Tschebyscheff-Systeme [Wer66u]. turnaround [CMW68, Wei67b]. Twist [dRMK67]. two [Cro69, Jän68j, LG69, Lin69a, McK69, Mey67a, Pim69k, Wil67f, Hil69c]. two-dimensional [Jän68j]. two-point [Pim69k]. two-step [Lin69a]. Typ [Lün69g]. Type [Mau67b, Mau67c, dR67d, dRMK67, BM68b, Csö69, Lün69g, ST69u]. Typical [Gro68j]. typique [Gro68j].

Übersicht [Her65r]. ultrafilters [Cle68]. Umgebungsränder [HM68o, HM68p]. Unabhängigkeit [Jen67m, Jen67p, Jen67q, Jen67o]. und-topologie [Fer68f]. un’equazione [ST69e]. unicità [ST69d]. Unicité [Lum68j]. Uniform [Hir69, Wer69b, Wer68]. uniforme [Wer68]. unions [Wei68]. uniqueness [ST69d, Wer66s, Lum68j, Wer66m]. unit [Hof69]. Unitale [Lün69m]. Unitals [Lün69m]. Unitäre [Lün69n]. Unitary [Lün69n]. Universal [Bor67e, Sim68k, Ken69, Wal69]. universality [Buc68]. universe [Mac69b]. Universität [Bau66g, Wer66v]. Université [DE67, DE68, DE69, Lum68a]. University [BKW69, Bor67d, CK69, Fic65k, HG67, Lum68a, yMO68, Urb67b, dRMK67, Ada66h, Ada69g, Bau66g, Wer66v]. unsettled [Pim69j]. Untergruppen [Lün65h]. Untersuchungen [Wer66b]. Upper [Fic65s, Shi68c]. use [Q. 66b, Str68b]. useful [Zal68p]. using [Yor68c].

V [HC68f]. Valeurs [Mét67]. validity [Kre69]. Vallée [Ber68d]. valuation [Frö68b]. value [Fic65u, Gri69, Jan69b, Jan69f, Jan69k, Pim69k, Stu69a, Stu69e, Stu69i, Ura69]. valued [Geo67, Hei69f, Jen67f, Jen67b, Mét67, Urb67i]. vanishing [And68]. variable [Geo67, Her65n, Pio69]. Variablen [Her65n]. variables [Cas69, Hub69a, Ric69, Wer66j]. variational [MD69]. Variations [Dus69, Kro69, Gro68b]. variété [Cer68f, Dol68a, Ram68, Roo68]. Variétés [Nor68, Cer68e, Dol68a]. Varieties [Sto66d, Mar69, Nor68, Shi68a, Sto66a, Zar69d, Zar69l]. variety [Dol68a, Roo68, Shi68f, Sto66c, Zar69c, Zar69g, Zar69f, Zar69h]. Vector [FB66b, Hei69f, Urb67i, Cor68, EK69, Eps69, FW68u, FB66c, FB66l, KT68k, Lel68a, May69k, May69l, May69m, Geo67, Mét67]. Vector-valued [Hei69f, Geo67, Mét67]. Vectorfields [Ton65h, Ton65i]. Vectorielles [Mét67]. vectoriels [Lel68a]. Vektorraumbündel [May69k, May69l, May69m]. Vektorräume [FW68u]. Vektorwertige [Geo67]. verallgemeinerte [Wlo69b, Her65s]. Veränderlichen [Wer66j]. Verdier [AM69m]. Vergleich [Dol66r]. Vergleichssätze [GKM68k]. Verhalten [Wer66g]. Verkleben [HM68b]. Verschärfungen [Köl68g]. version [Wer66q]. versus [HW69]. Verteilung [Cha66f]. very [Bie67].

Vielfachen [HM68h]. vier [Cha66l]. Vitushkin [Zal68e, Zal68q]. vocal [ST69o, ST69n]. Vollkommutative [BP69j]. Vollständig [Her68e]. Vollständigkeitsbeweis [Her65r]. Vollstetige [Stu69k]. Volterra [Nob69]. Volume [Hil69b, Hil69c, Löb69, Sto66a, Sto66c]. Volumes [Sto66d]. vom [Lün69g]. Vorbereitender [Ber68i]. Vorbereitungen [Bau66i, May69o]. Vorkommen [Her65n]. Vorlesung [Lün65g, Bau66g, Wer66v].

W [Gri69, Mey68g]. W. [Roo68]. Wahrheitsdefinition [Jen67f]. Warsaw [IR66f]. wave [vF69]. Weak [Cor68, Jan69f, May69m, FW68g]. Weakly [BH67a, Con67r, BH67b, Gor69, May69k, Pyk69]. Weierstrass [Nar66h, dC67c]. weight [Exb68]. Weil [CMW68]. Weinstein [Fic65t]. Weitere [Jän68l, Lün69o]. Well [Fic65u, Tro69p]. Well-orderings [Tro69p]. wertige [Jen67b]. wesentlich [Bie67]. Western [CK69]. Weyl [Cha66i, Tar67n]. Weylsche [Cha66i]. where [Lum68f]. which [AM69g, Ken69, Ros68b]. Widerspruchsfreiheit [Jen67m]. without [FB66b, Lam66l, Sta69, ST69d]. Wold [Urb67j]. work [Ada64c, Ada66c, Ada69c, AM69g, Har66i]. Written [Lum68a]. Wu [Jän68d, Jän68d]. Wu-chung [Jän68d, Jän68d]. Wu-yi [Jän68d, Jän68d].

x [ST69g].

Yang [Jän68j, Jän68j]. yi [Jän68d, Jän68d].

Zahl [Cha66l, Cha66h]. Zahlen [Cha66i, May69e, May69n]. Zahlentheoretische [Cha66m]. Zahlentheorie [Cha66n, Cha66d]. Zarankiewicz [Guy69]. Zerlegung [BP69j, Wer64h, BP69k]. Zerlegungs [Bau66j]. Zerlegungs- [Bau66j]. zero [Mat68]. zeta [Shi68i]. Zippin [Jän68j, Jän68j]. Zufallsvariable [Geo67]. zur [Ber68h, Jan69e, Jan69m, KT66c, May69i, Fer68g, Lün69p, Lün69q]. Zurich [Wer64j, Tar67j, Tho66h, Tie69c, Lam66d, Wer64j]. Zusammenhang [Her65t]. Zusammenhänge [GKM68g]. zweidimensionalem [Jän68j]. zwischen [Fer68b, Fer68a, GKM68c, Her65t, May69e, May69n]. Zwischenschritten [Jan69e]. Zwischenschrittmethode [Jan69d, Jan69b, Jan69m].

References

Appelgate:1969:I

[ABB^{+}69] H. Appelgate, M. Barr, J. Beck, F. W. Lawvere, F. E. J. Linton, E. Manes, M. Tierney, and F. Ulmer. Introduction. Lecture Notes in Mathematics, 80:1–6A, 1969. CODEN LNMAA2. ISBN 3-540-04601-1 (print), 3-540-36091-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

           com/chapter/10.1007/BFb0083079/.

Arkowitz:1964:BM

[AC64a] M. Arkowitz and C. R. Curjel. Back matter. Lecture Notes in Mathematics, 4:35–37, 1964. CODEN LNMAA2. ISBN 3-662-15915-5 (print), 3-662-15913-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-662-15913-2/1.pdf.

Arkowitz:1964:CHC

[AC64b] M. Arkowitz and C. R. Curjel. Commutativity and homotopycommutativity. Lecture Notes in Mathematics, 4:20–26, 1964. CO-DEN LNMAA2. ISBN 3-662-15915-5 (print), 3-662-15913-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-662-15913-2_
4/.

Arkowitz:1964:FM

[AC64c] M. Arkowitz and C. R. Curjel. Front matter. Lecture Notes in Mathematics, 4:N2–iii, 1964. CODEN LNMAA2. ISBN 3-662-15915-5 (print), 3-662-15913-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-662-15913-2/1.pdf.

Arkowitz:1964:GTH

[AC64d] M. Arkowitz and C. R. Curjel. The groups [A, ΩX] and their homomorphisms. Lecture Notes in Mathematics, 4:10–19, 1964. CODEN LNMAA2. ISBN 3-662-15915-5 (print), 3-662-15913-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-662-15913-2_3/.

Arkowitz:1964:GFR

[AC64e] M. Arkowitz and C. R. Curjel. Groups of finite rank. Lecture Notes in Mathematics, 4:3–9, 1964. CODEN LNMAA2. ISBN 3-662-15915-5 (print), 3-662-15913-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-15913-2_2/.

Arkowitz:1964:I

[AC64f] M. Arkowitz and C. R. Curjel. Introduction. Lecture Notes in Mathematics, 4:1–2, 1964. CODEN LNMAA2. ISBN 3-662-15915-5 (print), 3-662-15913-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-15913-2_1.

Arkowitz:1964:RGH

[AC64g] M. Arkowitz and C. R. Curjel. The rank of the group of homotopy equivalences. Lecture Notes in Mathematics, 4:27– 34, 1964. CODEN LNMAA2. ISBN 3-662-15915-5 (print), 3-662-15913-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-15913-2_5/.

Arkowitz:1964:GHC

[AC64h] Martin Arkowitz and Caspar R. Curjel. Groups of Homotopy Classes: Rank formulas and homotopy-commutativity, volume 4 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1964. CO-DEN LNMAA2. ISBN 3-662-15915-5 (paperback), 3-662-15913-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 36

pp. LCCN QA3 .L28 no. 4. URL http://link.springer.com/
book/10.1007/978-3-662-15913-2.

Arkowitz:1967:CHC

[AC67a] M. Arkowitz and C. R. Curjel. Commutativity and homotopycommutativity. Lecture Notes in Mathematics, 4:19–25, 1967. CO-DEN LNMAA2. ISBN 3-540-03900-7 (print), 3-540-34968-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0099651/.

Arkowitz:1967:GTH

[AC67b] M. Arkowitz and C. R. Curjel. The groups [A, ΩX] and their homomorphisms. Lecture Notes in Mathematics, 4:10–18, 1967. CODEN LNMAA2. ISBN 3-540-03900-7 (print), 3-540-34968-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099650/.

Arkowitz:1967:GFR

[AC67c] M. Arkowitz and C. R. Curjel. Groups of finite rank. Lecture Notes in Mathematics, 4:3–10, 1967. CODEN LNMAA2. ISBN 3-540-03900-7 (print), 3-540-34968-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099649/.

Arkowitz:1967:GHC

[AC67d] M. Arkowitz and C. R. Curjel. Groups of Homotopy Classes: Rank formulas and homotopy-commutativity, volume 4 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1967. CODEN LNMAA2. ISBN 3-540-03900-7 (paperback), 3-540-34968-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 35 pp. LCCN QA3 .L28 no. 4; QA3 .L28 no. 4, 1967. URL http://

link.springer.com/book/10.1007/BFb0099647.

Arkowitz:1967:I

[AC67e] M. Arkowitz and C. R. Curjel. Introduction. Lecture Notes in Mathematics, 4:1–2, 1967. CODEN LNMAA2. ISBN 3-540-03900-7 (print), 3-540-34968-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099648.

Arkowitz:1967:RGH

[AC67f] M. Arkowitz and C. R. Curjel. The rank of the group of homotopy equivalences. Lecture Notes in Mathematics, 4:26–33, 1967. CODEN LNMAA2. ISBN 3-540-03900-7 (print), 3-540-34968-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0099652/.

Aczel:1969:DCE

[Acz69] J. Aczel. On different characterizations of entropies. Lecture Notes in Mathematics, 89:1–11, 1969. CODEN LNMAA2. ISBN 3-540-04608-9 (print), 3-540-36098-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079114/.

Adams:1964:AHA

[Ada64a] J. Frank Adams. Applications of homological algebra to stable homotopy theory. Lecture Notes in Mathematics, 3:38– 57, 1964. CODEN LNMAA2. ISBN 3-662-15944-9 (print), 3-662-15942-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-15942-2_4/.

Adams:1964:BM

[Ada64b] J. Frank Adams. Back matter. Lecture Notes in Mathematics, 3:74–77, 1964. CODEN LNMAA2. ISBN 3-662-15944-9 (print), 3-662-15942-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-662-15942-2/1.pdf.

Adams:1964:CPA

[Ada64c] J. Frank Adams. Comments on prospective applications of (5), work in progress, etc. Lecture Notes in Mathematics, 3:69– 73, 1964. CODEN LNMAA2. ISBN 3-662-15944-9 (print), 3-662-15942-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-15942-2_6/.

Adams:1964:FM

[Ada64d] J. Frank Adams. Front matter. Lecture Notes in Mathematics, 3:N2–iii, 1964. CODEN LNMAA2. ISBN 3-662-15944-9 (print), 3-662-15942-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-662-15942-2/1.pdf.

Adams:1964:I

[Ada64e] J. Frank Adams. Introduction. Lecture Notes in Mathematics, 3: 1–3, 1964. CODEN LNMAA2. ISBN 3-662-15944-9 (print), 3-662-15942-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-15942-2_1/.

Adams:1964:PO

[Ada64f] J. Frank Adams. Primary operations. Lecture Notes in Mathematics, 3:4–21, 1964. CODEN LNMAA2. ISBN 3-662-15944-9 (print), 3-662-15942-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-15942-2_2/.

Adams:1964:SHTa

[Ada64g] J. Frank Adams. Stable homotopy theory. Lecture Notes in Mathematics, 3:22–37, 1964. CODEN LNMAA2. ISBN 3-662-15944-9 (print), 3-662-15942-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-15942-2_3/.

Adams:1964:TPA

[Ada64h] J. Frank Adams. Theorems of periodicity and approximation in homological algebra. Lecture Notes in Mathematics, 3:58– 68, 1964. CODEN LNMAA2. ISBN 3-662-15944-9 (print), 3-662-15942-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-15942-2_5/.

Adams:1964:SHTb

[Ada64i] J. Frank (John Frank) Adams. Stable Homotopy Theory, volume 3 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1964. CO-DEN LNMAA2. ISBN 3-662-15944-9 (paperback), 3-662-15942-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 77

pp. LCCN QA3 .L28 no. 3. URL http://link.springer.com/
book/10.1007/978-3-662-15942-2. Notes by A. T. Vasquez.

Adams:1966:AHA

[Ada66a] J. Frank Adams. Applications of homological algebra to stable homotopy theory. Lecture Notes in Mathematics, 3:38– 57, 1966. CODEN LNMAA2. ISBN 3-662-15907-4 (print), 3-662-15905-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-15905-7_4/.

Adams:1966:BM

[Ada66b] J. Frank Adams. Back matter. Lecture Notes in Mathematics, 3:74–81, 1966. CODEN LNMAA2. ISBN 3-662-15907-4 (print), 3-662-15905-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-662-15905-7/1.pdf.

Adams:1966:CPA

[Ada66c] J. Frank Adams. Comments on prospective applications of (5), work in progress, etc. Lecture Notes in Mathematics, 3:69– 73, 1966. CODEN LNMAA2. ISBN 3-662-15907-4 (print), 3-662-15905-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-15905-7_6/.

Adams:1966:FM

[Ada66d] J. Frank Adams. Front matter. Lecture Notes in Mathematics, 3:ii–v, 1966. CODEN LNMAA2. ISBN 3-662-15907-4 (print), 3-662-15905-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-662-15905-7/1.pdf.

Adams:1966:I

[Ada66e] J. Frank Adams. Introduction. Lecture Notes in Mathematics, 3: 1–3, 1966. CODEN LNMAA2. ISBN 3-662-15907-4 (print), 3-662-15905-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-15905-7_1/.

Adams:1966:PO

[Ada66f] J. Frank Adams. Primary operations. Lecture Notes in Mathematics, 3:4–21, 1966. CODEN LNMAA2. ISBN 3-662-15907-4 (print), 3-662-15905-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-15905-7_2/.

Adams:1966:SHTa

[Ada66g] J. Frank Adams. Stable homotopy theory. Lecture Notes in Mathematics, 3:22–37, 1966. CODEN LNMAA2. ISBN 3-662-15907-4 (print), 3-662-15905-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-15905-7_3/.

Adams:1966:SHTb

[Ada66h] J. Frank Adams. Stable homotopy theory. Lectures delivered at the University of California at Berkeley 1961, volume 3 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1966. CO-DEN LNMAA2. ISBN 3-662-15907-4 (paperback), 3-662-15905-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 78 pp. LCCN QA3 .L28 no. 3; QA3 .L28 no. 3 1966. URL http:/

/link.springer.com/book/10.1007/978-3-662-15905-7. Notes

by A. T. Vasquez.

Adams:1966:TPA

[Ada66i] J. Frank Adams. Theorems of periodicity and approximation in homological algebra. Lecture Notes in Mathematics, 3:58– 68, 1966. CODEN LNMAA2. ISBN 3-662-15907-4 (print), 3-662-15905-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-15905-7_5/.

Adams:1969:LGC

[Ada69a] J. F. Adams. Lectures on generalised cohomology. Lecture Notes in Mathematics, 99:1–138, 1969. CODEN LNMAA2. ISBN 3-540-04618-6 (print), 3-540-36140-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081960/.

Adams:1969:AHA

[Ada69b] J. Frank Adams. Applications of homological algebra to stable homotopy theory. Lecture Notes in Mathematics, 3:38–57, 1969. CODEN LNMAA2. ISBN 3-540-04598-8 (print), 3-540-36088-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0103454/.

Adams:1969:CPA

[Ada69c] J. Frank Adams. Comments on prospective applications of (5), work in progress, etc. Lecture Notes in Mathematics, 3:69–73, 1969. CODEN LNMAA2. ISBN 3-540-04598-8 (print), 3-540-36088-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0103456/.

Adams:1969:I

[Ada69d] J. Frank Adams. Introduction. Lecture Notes in Mathematics, 3: 1–3, 1969. CODEN LNMAA2. ISBN 3-540-04598-8 (print), 3-540-36088-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103451/.

Adams:1969:PO

[Ada69e] J. Frank Adams. Primary operations. Lecture Notes in Mathematics, 3:4–21, 1969. CODEN LNMAA2. ISBN 3-540-04598-8 (print), 3-540-36088-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103452/.

Adams:1969:SHTa

[Ada69f] J. Frank Adams. Stable homotopy theory. Lecture Notes in Mathematics, 3:22–37, 1969. CODEN LNMAA2. ISBN 3-540-04598-8 (print), 3-540-36088-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103453/.

Adams:1969:SHTb

[Ada69g] J. Frank Adams. Stable homotopy theory. Lectures delivered at the University of California at Berkeley 1961, volume 3 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1969. CO-DEN LNMAA2. ISBN 3-540-04598-8 (paperback), 3-540-36088-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 78 pp. LCCN QA3 .L28 no. 3; QA3 .L28 no. 3, 1969. URL http://

link.springer.com/book/10.1007/BFb0103450. Notes by A. T.

Vasquez.

Adams:1969:TPA

[Ada69h] J. Frank Adams. Theorems of periodicity and approximation in homological algebra. Lecture Notes in Mathematics, 3:58–68, 1969. CODEN LNMAA2. ISBN 3-540-04598-8 (print), 3-540-36088-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0103455/.

Azema:1969:MIP

[ADR69] J. Azéma, M. Duflo, and D. Revuz. Mesure invariante des processus de Markov rècurrents. (French) [Invariant measure of the recurring Markov process]. Lecture Notes in Mathematics, 88:24–33, 1969. CODEN LNMAA2. ISBN 3-540-04607-0 (print), 3-540-36097-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0081625/.

Aigner:1969:GBR

[Aig69] Martin Aigner. Graphs and binary relations. Lecture Notes in Mathematics, 110:1–21, 1969. CODEN LNMAA2. ISBN 3-540-04629-1 (print), 3-540-36161-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060100/.

Azema:1968:CRP

[AKDR68] J. Azéma, M. Kaplan-Duflo, and D. Revuz. Classes récurrentes d’un processus de Markov. (French) [Recurrent classes of Markov processes]. Lecture Notes in Mathematics, 51:1–21, 1968. CO-DEN LNMAA2. ISBN 3-540-04221-0 (print), 3-540-35827-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0059176/.

Aubert:1969:PSC

[AL69] Karl Egil Aubert and Wilhelm Ljunggren, editors. Proceedings of the 15th Scandinavian Congress Oslo [August 12–16] 1968, volume 118 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CO-DEN LNMAA2. ISBN 3-540-04907-X (paperback), 3-540-36246-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN

QA3 .L28 no. 118. URL http://link.springer.com/book/10.
1007/BFb0060247.

Albrecht:1968:BPS

[Alb68a] Felix Albrecht. Boundary points of sets of attainability. Lecture Notes in Mathematics, 63:46–64, 1968. CODEN LNMAA2. ISBN 3-540-04233-4 (print), 3-540-35867-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076925/.

Albrecht:1968:CAS

[Alb68b] Felix Albrecht. Controls and attainability. The space D^{1}(U, I). Lecture Notes in Mathematics, 63:27–45, 1968. CODEN LN-MAA2. ISBN 3-540-04233-4 (print), 3-540-35867-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0076924/.

Albrecht:1968:P

[Alb68c] Felix Albrecht. Preliminaries. Lecture Notes in Mathematics, 63: 1–26, 1968. CODEN LNMAA2. ISBN 3-540-04233-4 (print), 3-540-35867-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076923/.

Albrecht:1968:TCT

[Alb68d] Felix Albrecht, editor. Topics in Control Theory: A Seminar given at the Forschungsinstitut für Mathematik, ETH, in 1964, volume 63 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. CO-DEN LNMAA2. ISBN 3-540-04233-4 (paperback), 3-540-35867-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN

QA3 .L28 no. 63. URL http://link.springer.com/book/10.
1007/BFb0076922.

Albrecht:1969:GIT

[Alb69] J. Albrecht. Generalisation of an inclusion theorem of L. Collatz. Lecture Notes in Mathematics, 109:1–6, 1969. CODEN LN-MAA2. ISBN 3-540-04628-3 (print), 3-540-36158-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0060013/.

Andre:1968:RMC

[AM68] M. (Michel) André and Saunders Mac Lane, editors. Reports of the Midwest Category Seminar II, volume 61 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. CODEN LNMAA2. ISBN 3-540-04231-8 (paperback), 3-540-35863-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 61. URL

http://link.springer.com/book/10.1007/BFb0077113.

Artin:1969:CCI

[AM69a] M. Artin and B. Mazur. Cohomological criteria for ♮-isomorphism. Lecture Notes in Mathematics, 100:35–59, 1969. CODEN LN-MAA2. ISBN 3-540-04619-4 (print), 3-540-36142-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080961/.

Artin:1969:CT

[AM69b] M. Artin and B. Mazur. Comparison theorems. Lecture Notes in Mathematics, 100:129–146, 1969. CODEN LNMAA2. ISBN 3-540-04619-4 (print), 3-540-36142-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080969/.

Artin:1969:C

[AM69c] M. Artin and B. Mazur. Completions. Lecture Notes in Mathematics, 100:25–34, 1969. CODEN LNMAA2. ISBN 3-540-04619-4 (print), 3-540-36142-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080960/.

Artin:1969:CF

[AM69d] M. Artin and B. Mazur. Completions and fibrations. Lecture Notes in Mathematics, 100:60–69, 1969. CODEN LNMAA2. ISBN 3-540-04619-4 (print), 3-540-36142-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080962/.

Artin:1969:EH

[AM69e] M. Artin and B. Mazur. Étale Homotopy, volume 100 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CODEN LNMAA2. ISBN 3-540-04619-4 (paperback), 3-540-36142-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 169 pp. LCCN

QA3 .L28 no. 100. URL http://link.springer.com/book/10.
1007/BFb0080957.

Artin:1969:FG

[AM69f] M. Artin and B. Mazur. The fundamental group. Lecture Notes in Mathematics, 100:117–123, 1969. CODEN LNMAA2. ISBN 3-540-04619-4 (print), 3-540-36142-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080967/.

Artin:1969:GCW

[AM69g] M. Artin and B. Mazur. A glossary of the categories in which we shall work, and fibre resolutions. Lecture Notes in Mathematics, 100:6–19, 1969. CODEN LNMAA2. ISBN 3-540-04619-4 (print), 3-540-36142-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080958/.

Artin:1969:HGC

[AM69h] M. Artin and B. Mazur. Homotopy groups of completions. Lecture Notes in Mathematics, 100:70–74, 1969. CODEN LNMAA2. ISBN 3-540-04619-4 (print), 3-540-36142-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080963/.

Artin:1969:H

[AM69i] M. Artin and B. Mazur. Hypercoverings. Lecture Notes in Mathematics, 100:93–110, 1969. CODEN LNMAA2. ISBN 3-540-04619-4 (print), 3-540-36142-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080965/.

Artin:1969:POH

[AM69j] M. Artin and B. Mazur. Pro-objects in the homotopy category. Lecture Notes in Mathematics, 100:20–24, 1969. CODEN LN-MAA2. ISBN 3-540-04619-4 (print), 3-540-36142-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080959/.

Artin:1969:PT

[AM69k] M. Artin and B. Mazur. A profiniteness theorem. Lecture Notes in Mathematics, 100:124–128, 1969. CODEN LNMAA2. ISBN 3-540-04619-4 (print), 3-540-36142-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080968/.

Artin:1969:SR

[AM69l] M. Artin and B. Mazur. Stable results. Lecture Notes in Mathematics, 100:75–92, 1969. CODEN LNMAA2. ISBN 3-540-04619-4 (print), 3-540-36142-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080964/.

Artin:1969:VF

[AM69m] M. Artin and B. Mazur. The Verdier functor. Lecture Notes in Mathematics, 100:111–116, 1969. CODEN LNMAA2. ISBN 3-540-04619-4 (print), 3-540-36142-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080966/.

Andre:1967:ACF

[And67a] Michel André. Algèbre commutative. (French) [Commutative algebra]. Lecture Notes in Mathematics, 32:59–119, 1967. CODEN LNMAA2. ISBN 3-540-03903-1 (print), 3-540-34971-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077201/.

Andre:1967:AHF

[And67b] Michel André. Algèbre homologique. (French) [Homologic algebra]. Lecture Notes in Mathematics, 32:1–58, 1967. CODEN LN-MAA2. ISBN 3-540-03903-1 (print), 3-540-34971-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077200/.

Andre:1967:MSA

[And67c] Michel André. Méthode Simpliciale en Algèbre Homologique et Algèbre Commutative. (French) [Simplicial Method in Homological Algebra and Commutative Algebra], volume 32 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. CODEN LNMAA2. ISBN 3-540-03903-1 (paperback), 3-540-34971-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 122 pp. LCCN QA3

.L28 no. 32. URL http://link.springer.com/book/10.1007/
BFb0077199.

Andre:1968:VSC

[And68] Michel André. On the vanishing of the second cohomology group of a commutative algebra. Lecture Notes in Mathematics, 61:1– 27, 1968. CODEN LNMAA2. ISBN 3-540-04231-8 (print), 3-540-35863-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077114/.

Anderson:1969:GTF

[And69] Sabra S. Anderson. Graph theory and finite projective planes. Lecture Notes in Mathematics, 110:23–26, 1969. CODEN LN-MAA2. ISBN 3-540-04629-1 (print), 3-540-36161-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0060101/.

Appelgate:1969:CM

[AT69] H. Appelgate and M. Tierney. Categories with models. Lecture Notes in Mathematics, 80:156–244, 1969. CODEN LNMAA2. ISBN 3-540-04601-1 (print), 3-540-36091-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083086/.

Atiyah:1969:ATO

[Ati69] M. F. Atiyah. Algebraic topology and operators in Hilbert space. Lecture Notes in Mathematics, 103:101–121, 1969. CODEN LN-MAA2. ISBN 3-540-04622-4 (print), 3-540-36145-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0099987/.

Avanissian:1967:HFS

[Ava67] V. Avanissian. Sur l’harmonicité des fonctions séparément harmoniques. (French) [On the harmonicity of separately-harmonic functions]. Lecture Notes in Mathematics, 39:3–17, 1967. CO-DEN LNMAA2. ISBN 3-540-03910-4 (print), 3-540-34978-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0080542/.

Ahlswede:1969:SCF

[AW69] R. Ahlswede and J. Wolfowitz. The structure of capacity functions for compound channels. Lecture Notes in Mathematics, 89: 12–54, 1969. CODEN LNMAA2. ISBN 3-540-04608-9 (print), 3-540-36098-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079115/.

Badrikian:1967:RTB

[Bad67] A. Badrikian. Remarques sur les théorèmes de Bochner et P. Lévy. (French) [Remarks on the theorems of Bochner and P. Lévy]. Lecture Notes in Mathematics, 31:1–19, 1967. CODEN LNMAA2. ISBN 3-540-03902-3 (print), 3-540-34970-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061102/.

Bak:1969:MQF

[Bak69] Anthony Bak. On modules with quadratic forms. Lecture Notes in Mathematics, 108:55–66, 1969. CODEN LNMAA2. ISBN 3-540-04627-5 (print), 3-540-36156-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059990/.

Barwise:1968:IDC

[Bar68a] Jon Barwise. Implicit definability and compactness in infinitary languages. Lecture Notes in Mathematics, 72:1–35, 1968. CODEN LNMAA2. ISBN 3-540-04242-3 (print), 3-540-35900-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079680/.

Barwise:1968:SSI

[Bar68b] Jon Barwise. The Syntax and Semantics of Infinitary Languages, volume 72 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. CODEN LNMAA2. ISBN 3-540-04242-3 (paperback), 3-540-35900-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 268 pp. LCCN QA3 .L28 no. 72. URL http://link.

springer.com/book/10.1007/BFb0079679.

Barnette:1969:STC

[Bar69a] David W. Barnette. On Steinitz’s theorem concerning convex 3polytopes and on some properties of planar graphs. Lecture Notes in Mathematics, 110:27–40, 1969. CODEN LNMAA2. ISBN 3-540-04629-1 (print), 3-540-36161-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060102/.

Barr:1969:CCA

[Bar69b] Michael Barr. Coalgebras in a category of algebras. Lecture Notes in Mathematics, 86:1–12, 1969. CODEN LNMAA2. ISBN 3-540-04605-4 (print), 3-540-36095-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079381/.

Barr:1969:COC

[Bar69c] Michael Barr. Cohomology and obstructions: Commutative algebras. Lecture Notes in Mathematics, 80:357–375, 1969. CODEN LNMAA2. ISBN 3-540-04601-1 (print), 3-540-36091-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083089/.

Barr:1969:CCD

[Bar69d] Michael Barr. Composite cotriples and derived functors. Lecture Notes in Mathematics, 80:336–356, 1969. CODEN LNMAA2. ISBN 3-540-04601-1 (print), 3-540-36091-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083088/.

Barr:1969:WC

[Bar69e] Michael Barr. What is the center? Lecture Notes in Mathematics, 106:1–12, 1969. CODEN LNMAA2. ISBN 3-540-04625-9 (print), 3-540-36150-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059140/.

Bass:1969:S

[Bas69] Hyman Bass. K_{2} and symbols. Lecture Notes in Mathematics, 108:1–11, 1969. CODEN LNMAA2. ISBN 3-540-04627-5 (print), 3-540-36156-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059987/.

Bauer:1966:BM

[Bau66a] Heinz Bauer. Back matter. Lecture Notes in Mathematics, 22: 165–176, 1966. CODEN LNMAA2. ISBN 3-540-03605-9 (print), 3-540-34803-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-34803-0/1.pdf.

Bauer:1966:BTG

[Bau66b] Heinz Bauer. Balayage-theorie. (German) [Scan theory]. Lecture Notes in Mathematics, 22:88–119, 1966. CODEN LNMAA2. ISBN 3-540-03605-9 (print), 3-540-34803-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075365/.

Bauer:1966:DPG

[Bau66c] Heinz Bauer. Dirichletsches Problem. (German) [Dirichlet’s problems]. Lecture Notes in Mathematics, 22:120–151, 1966. CODEN LNMAA2. ISBN 3-540-03605-9 (print), 3-540-34803-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075366/.

Bauer:1966:EGI

[Bau66d] Heinz Bauer. Einleitung. (German) [Introduction]. Lecture Notes in Mathematics, 22:1–2, 1966. CODEN LNMAA2. ISBN 3-540-03605-9 (print), 3-540-34803-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075361.

Bauer:1966:FM

[Bau66e] Heinz Bauer. Front matter. Lecture Notes in Mathematics, 22: i–iv, 1966. CODEN LNMAA2. ISBN 3-540-03605-9 (print), 3-540-34803-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-34803-0/1.pdf.

Bauer:1966:HRG

[Bau66f] Heinz Bauer. Harmonische Räume. (German) [Harmonic spaces]. Lecture Notes in Mathematics, 22:9–44, 1966. CODEN LN-MAA2. ISBN 3-540-03605-9 (print), 3-540-34803-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075363/.

Bauer:1966:HRI

[Bau66g] Heinz Bauer, editor. Harmonische Räume und ihre Potentialtheorie: Ausarbeitung einer im Sommersemester 1965 an der Universität Hamburg gehaltenen Vorlesung. (German) [Harmonic spaces and their potential theory: Lecture on the development of a project at the University of Hamburg in the summer semester 1965], volume 22 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1966. CO-DEN LNMAA2. ISSN 0075-8434 (print), 1617-9692 (electronic).

LCCN QA3 .L28 no. 22. URL http://link.springer.com/book/
10.1007/BFb0075360.

Bauer:1966:SFP

[Bau66h] Heinz Bauer. Superharmonische Funktionen und Potentiale. (German) [Superharmonic functions and potentials]. Lecture Notes in Mathematics, 22:45–87, 1966. CODEN LNMAA2. ISBN 3-540-03605-9 (print), 3-540-34803-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075364/.

Bauer:1966:VBG

[Bau66i] Heinz Bauer. Vorbereitungen und Bezeichnungen. (German) [Preparations and designations]. Lecture Notes in Mathematics, 22:3–8, 1966. CODEN LNMAA2. ISBN 3-540-03605-9 (print), 3-540-34803-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075362/.

Bauer:1966:ZFG

[Bau66j] Heinz Bauer. Zerlegungs- und Fortsetzungssatz. (German) [Separation and extension theorem]. Lecture Notes in Mathematics, 22:152–164, 1966. CODEN LNMAA2. ISBN 3-540-03605-9 (print), 3-540-34803-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075367/.

Bauer:1967:RDA

[Bau67] Heinz Bauer. Recent developments in axiomatic potential theory. Lecture Notes in Mathematics, 31:20–27, 1967. CODEN LNMAA2. ISBN 3-540-03902-3 (print), 3-540-34970-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061103/.

Bauer:1968:SPG

[Bau68] H. Bauer. Seminar über Potentialtheorie. (German) [Seminar on potential theory], volume 69 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. CODEN LNMAA2. ISBN 3-540-04239-3 (paperback), 3-540-35884-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 180 pp. LCCN QA3 .L28 no. 69. URL

http://link.springer.com/book/10.1007/BFb0080473.

Barahamihir:1969:BAM

[BB69a] M. Barahamihir and D. Behara. Boolean algebraic methods in Markov chains. Lecture Notes in Mathematics, 89:55–63, 1969.

CODEN LNMAA2. ISBN 3-540-04608-9 (print), 3-540-36098-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0079116/.

Barr:1969:HSC

[BB69b] Michael Barr and Jon Beck. Homology and standard constructions. Lecture Notes in Mathematics, 80:245–335, 1969. CODEN LNMAA2. ISBN 3-540-04601-1 (print), 3-540-36091-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083087/.

Barr:1969:RMC

[BBD^{+}69] M. Barr, P. Berthiaume, B. J. Day, J. Duskin, S. Feferman, G. M. Kelly, S. Mac Lane, M. Tierney, and R. F. C. Walters, editors. Reports of the Midwest Category Seminar III, volume 106 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CODEN LNMAA2. ISBN 3-540-04625-9 (paperback), 3-540-36150-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 106.

URL http://link.springer.com/book/10.1007/BFb0059139.

Borel:1966:SCM

[BCH^{+}66] A. Borel, S. Chowla, C. S. Herz, K. Iwasawa, and J.-P. Serre, editors. Seminar on Complex Multiplication: Seminar held at the Institute for Advanced Study, Princeton, N.J., 1957–58, volume 21 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1966. CODEN LN-MAA2. ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN

QA3 .L28 no. 21. URL http://link.springer.com/book/10.
1007/BFb0091550.

Bretagnolle:1967:LSE

[BDCK67] J. Bretagnolle, D. Dacunha-Castelle, and J. L. Krivine. Lois stables et espaces L^{p}. (French) [Stable laws and L^{p} spaces]. Lecture Notes in Mathematics, 31:48–54, 1967. CODEN LNMAA2. ISBN 3-540-03902-3 (print), 3-540-34970-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061106/.

Benabou:1967:RMC

[BDD^{+}67] J. Bénabou, R. Davis, A. Dold, J. Isbell, S. Mac Lane, U. Oberst, and J.-E. Roos, editors. Reports of the Midwest Category Seminar, volume 47 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. CODEN LNMAA2. ISBN 3-540-03918-X (paperback), 3-540-35545-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

LCCN QA3 .L28 no. 47. URL http://link.springer.com/book/
10.1007/BFb0074298.

Beck:1969:DL

[Bec69a] Jon Beck. Distributive laws. Lecture Notes in Mathematics, 80: 119–140, 1969. CODEN LNMAA2. ISBN 3-540-04601-1 (print), 3-540-36091-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083084/.

Beck:1969:SIL

[Bec69b] Jon Beck. On H-spaces and infinite loop spaces. Lecture Notes in Mathematics, 99:139–153, 1969. CODEN LNMAA2. ISBN 3-540-04618-6 (print), 3-540-36140-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081961/.

Behzad:1969:ARN

[Beh69] Mehdi Behzad. Analogues of Ramseyxo numbers. Lecture Notes in Mathematics, 110:41–43, 1969. CODEN LNMAA2. ISBN 3-540-04629-1 (print), 3-540-36161-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060103/.

Beineke:1969:SPC

[Bei69] Lowell W. Beineke. A survey of packings and coverings of graphs. Lecture Notes in Mathematics, 110:45–53, 1969. CODEN LN-MAA2. ISBN 3-540-04629-1 (print), 3-540-36161-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0060104/.

Benabou:1967:IB

[Bén67] Jean Bénabou. Introduction to bicategories. Lecture Notes in Mathematics, 47:1–77, 1967. CODEN LNMAA2. ISBN 3-540-03918-X (print), 3-540-35545-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074299/.

Berens:1968:AHO

[Ber68a] Hubert Berens. Approximation und Halbgruppen von Operatoren. (German) [Approximation and operator semi-groups]. Lecture Notes in Mathematics, 64:37–55, 1968. CODEN LNMAA2. ISBN 3-540-04234-2 (print), 3-540-35871-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077260/.

Berens:1968:ABG

[Ber68b] Hubert Berens. Approximationsprozesse auf Banachräumen. (German) [Approximation processes on Banach spaces]. Lecture Notes in Mathematics, 64:22–36, 1968. CODEN LNMAA2. ISBN 3-540-04234-2 (print), 3-540-35871-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077259/.

Berens:1968:BM

[Ber68c] Hubert Berens. Back matter. Lecture Notes in Mathematics, 64: 85–90, 1968. CODEN LNMAA2. ISBN 3-540-04234-2 (print), 3-540-35871-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-35871-8/1.pdf.

Berens:1968:SIV

[Ber68d] Hubert Berens. Das Singuläre integral von de la Vallée Poussin. (German) [The de la Vallée Poussin singular integral]. Lecture Notes in Mathematics, 64:56–65, 1968. CODEN LNMAA2. ISBN 3-540-04234-2 (print), 3-540-35871-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077261/.

Berens:1968:RMF

[Ber68e] Hubert Berens. Die Rieszschen mittel des Fourierumkehrintegrals. (German) [The Riesz medium of the Fourier inversion integral]. Lecture Notes in Mathematics, 64:66–84, 1968. CODEN LNMAA2. ISBN 3-540-04234-2 (print), 3-540-35871-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077262/.

Berens:1968:EGI

[Ber68f] Hubert Berens. Einführung. (German) [Introduction]. Lecture Notes in Mathematics, 64:1–5, 1968. CODEN LNMAA2. ISBN 3-540-04234-2 (print), 3-540-35871-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077257/.

Berens:1968:FM

[Ber68g] Hubert Berens. Front matter. Lecture Notes in Mathematics, 64: I–V, 1968. CODEN LNMAA2. ISBN 3-540-04234-2 (print), 3-540-35871-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-35871-8/1.pdf.

Berens:1968:IBA

[Ber68h] Hubert Berens. Interpolationsmethoden zur Behandlung von Approximationsprozessen auf Banachräumen. (German) [Interpolation methods for the treatment of approximation processes on Banach spaces], volume 64 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. CODEN LNMAA2. ISBN 3-540-04234-2 (paperback), 3-540-35871-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 90 pp. LCCN QA3 .L28 no. 64. URL http://link.

springer.com/book/10.1007/BFb0077256.

Berens:1968:VAG

[Ber68i] Hubert Berens. Vorbereitender Abschnitt. (German) [Preparatory section]. Lecture Notes in Mathematics, 64:6–21, 1968. CODEN LNMAA2. ISBN 3-540-04234-2 (print), 3-540-35871-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077258/.

Berthiaume:1969:FE

[Ber69] P. Berthiaume. The functor evaluation. Lecture Notes in Mathematics, 106:13–63, 1969. CODEN LNMAA2. ISBN 3-540-04625-9 (print), 3-540-36150-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059141/.

Berglund:1967:CSSb

[BH67a] J. F. Berglund and K. H. Hofmann. Compact Semitopological Semigroups and Weakly Almost Periodic Functions, volume 42 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. CODEN LN-MAA2. ISBN 3-540-03913-9 (paperback), 3-540-35184-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 160 pp. LCCN

QA3 .L28 no. 42. URL http://link.springer.com/book/10.
1007/BFb0073920.

Berglund:1967:APW

[BH67b] John F. Berglund and Karl Heinrich Hofmann. Almost periodic and weakly almost periodic functions on semitopological semigroups. Lecture Notes in Mathematics, 42:112–145, 1967. CODEN LN-MAA2. ISBN 3-540-03913-9 (print), 3-540-35184-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0073924/.

Berglund:1967:CSSa

[BH67c] John F. Berglund and Karl Heinrich Hofmann. Compact semitopological semigroups. Lecture Notes in Mathematics, 42:44–111, 1967. CODEN LNMAA2. ISBN 3-540-03913-9 (print), 3-540-35184-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0073923/.

Berglund:1967:E

[BH67d] John F. Berglund and Karl Heinrich Hofmann. Examples. Lecture Notes in Mathematics, 42:146–157, 1967. CODEN LNMAA2. ISBN 3-540-03913-9 (print), 3-540-35184-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073925/.

Berglund:1967:I

[BH67e] John F. Berglund and Karl Heinrich Hofmann. Introduction. Lecture Notes in Mathematics, 42:1–11, 1967. CODEN LN-MAA2. ISBN 3-540-03913-9 (print), 3-540-35184-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0073921/.

Berglund:1967:P

[BH67f] John F. Berglund and Karl Heinrich Hofmann. Preliminaries. Lecture Notes in Mathematics, 42:12–43, 1967. CODEN LN-MAA2. ISBN 3-540-03913-9 (print), 3-540-35184-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0073922/.

Bhatia:1969:A

[BH69a] N. P. Bhatia and O. Hajek. Attraction. Lecture Notes in Mathematics, 90:79–93, 1969. CODEN LNMAA2. ISBN 3-540-04609-7 (print), 3-540-36099-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079594/.

Bhatia:1969:CC

[BH69b] N. P. Bhatia and O. Hajek. Compactness conditions. Lecture Notes in Mathematics, 90:42–48, 1969. CODEN LNMAA2. ISBN 3-540-04609-7 (print), 3-540-36099-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079590/.

Bhatia:1969:FNI

[BH69c] N. P. Bhatia and O. Hajek. Flow near an invariant set. Lecture Notes in Mathematics, 90:94–99, 1969. CODEN LNMAA2. ISBN 3-540-04609-7 (print), 3-540-36099-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079595/.

Bhatia:1969:FDE

[BH69d] N. P. Bhatia and O. Hajek. Functional-differential equations. Lecture Notes in Mathematics, 90:141–153, 1969. CODEN LN-MAA2. ISBN 3-540-04609-7 (print), 3-540-36099-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079599/.

Bhatia:1969:Ia

[BH69e] N. P. Bhatia and O. Hajek. Introduction. Lecture Notes in Mathematics, 90:2–11, 1969. CODEN LNMAA2. ISBN 3-540-04609-7 (print), 3-540-36099-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079586/.

Bhatia:1969:Ib

[BH69f] N. P. Bhatia and O. Hajek. Invariance. Lecture Notes in Mathematics, 90:35–41, 1969. CODEN LNMAA2. ISBN 3-540-04609-7 (print), 3-540-36099-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079589/.

Bhatia:1969:LF

[BH69g] N. P. Bhatia and O. Hajek. Liapunov functions. Lecture Notes in Mathematics, 90:100–119, 1969. CODEN LNMAA2. ISBN 3-540-04609-7 (print), 3-540-36099-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079596/.

Bhatia:1969:LS

[BH69h] N. P. Bhatia and O. Hajek. Limit sets. Lecture Notes in Mathematics, 90:49–58, 1969. CODEN LNMAA2. ISBN 3-540-04609-7 (print), 3-540-36099-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079591/.

Bhatia:1969:LSDb

[BH69i] N. P. Bhatia and O. Hajek. Local Semi-Dynamical Systems, volume 90 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CO-DEN LNMAA2. ISBN 3-540-04609-7 (paperback), 3-540-36099-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 157

pp. LCCN QA3 .L28 no. 90. URL http://link.springer.com/
book/10.1007/BFb0079585.

Bhatia:1969:LSDa

[BH69j] N. P. Bhatia and O. Hajek. Local semi-dynamical systems: Basic definitions and properties. Lecture Notes in Mathematics, 90: 12–26, 1969. CODEN LNMAA2. ISBN 3-540-04609-7 (print), 3-540-36099-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079587/.

Bhatia:1969:MCI

[BH69k] N. P. Bhatia and O. Hajek. Minimality; characteristic O. Lecture Notes in Mathematics, 90:132–140, 1969. CODEN LNMAA2. ISBN 3-540-04609-7 (print), 3-540-36099-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079598/.

Bhatia:1969:PP

[BH69l] N. P. Bhatia and O. Hajek. The positive prolongation. Lecture Notes in Mathematics, 90:59–73, 1969. CODEN LNMAA2. ISBN 3-540-04609-7 (print), 3-540-36099-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079592/.

Bhatia:1969:SNC

[BH69m] N. P. Bhatia and O. Hajek. Solutions: Negative continuation. Lecture Notes in Mathematics, 90:27–34, 1969. CODEN LN-MAA2. ISBN 3-540-04609-7 (print), 3-540-36099-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079588/.

Bhatia:1969:SOS

[BH69n] N. P. Bhatia and O. Hajek. Stability and orbital stability. Lecture Notes in Mathematics, 90:74–78, 1969. CODEN LNMAA2. ISBN 3-540-04609-7 (print), 3-540-36099-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079593/.

Bhatia:1969:SPS

[BH69o] N. P. Bhatia and O. Hajek. The start point set. Lecture Notes in Mathematics, 90:120–131, 1969. CODEN LNMAA2. ISBN 3-540-04609-7 (print), 3-540-36099-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079597/.

Bierlein:1967:WIS

[Bie67] D. Bierlein. Über wesentlich indefinite Spiele. (German) [On very indefinite games]. Lecture Notes in Mathematics, 31:28–35, 1967. CODEN LNMAA2. ISBN 3-540-03902-3 (print), 3-540-34970-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0061104/.

Billingsley:1969:MPS

[Bil69] Patrick Billingsley. Maxima of partial sums. Lecture Notes in Mathematics, 89:64–76, 1969. CODEN LNMAA2. ISBN 3-540-04608-9 (print), 3-540-36098-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079117/.

Berger:1967:ARG

[BKKN67a] Professor Dr. R. Berger, Dr. R. Kiehl, Dr. E. Kunz, and Professor Dr. H.-J. Nastold. Absolute Regularität. (German) [Absolute regularity]. Lecture Notes in Mathematics, 38:110–131, 1967. CO-DEN LNMAA2. ISBN 3-540-03909-0 (print), 3-540-34977-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0073992/.

Berger:1967:DGD

[BKKN67b] Professor Dr. R. Berger, Dr. R. Kiehl, Dr. E. Kunz, and Professor Dr. H.-J. Nastold. Differentialmoduln. (German) [Differential modules]. Lecture Notes in Mathematics, 38:41–82, 1967. CO-DEN LNMAA2. ISBN 3-540-03909-0 (print), 3-540-34977-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0073990/.

Berger:1967:DAG

[BKKN67c] Professor Dr. R. Berger, Dr. R. Kiehl, Dr. E. Kunz, and Professor Dr. H.-J. Nastold. Differentialrechnung in der analytischen Geometrie. (German) [Differential calculus in analytic geometry], volume 38 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. CO-DEN LNMAA2. ISBN 3-540-03909-0 (paperback), 3-540-34977-4 (electronic). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 134 pp. LCCN QA3 .L28 no. 38. URL http://link.springer.

com/book/10.1007/BFb0073987.

Berger:1967:EGI

[BKKN67d] Professor Dr. R. Berger, Dr. R. Kiehl, Dr. E. Kunz, and Professor Dr. H.-J. Nastold. Einleitung. (German) [Introduction]. Lecture Notes in Mathematics, 38:1–11, 1967. CODEN LNMAA2. ISBN 3-540-03909-0 (print), 3-540-34977-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073988/.

Berger:1967:KRA

[BKKN67e] Professor Dr. R. Berger, Dr. R. Kiehl, Dr. E. Kunz, and Professor Dr. H.-J. Nastold. Kategorien von Rigen in der analytischen Geometrie. (German) [Rigen categories in analytic geometry]. Lecture Notes in Mathematics, 38:12–40, 1967. CODEN LNMAA2. ISBN 3-540-03909-0 (print), 3-540-34977-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073989/.

Berger:1967:RAG

[BKKN67f] Professor Dr. R. Berger, Dr. R. Kiehl, Dr. E. Kunz, and Professor Dr. H.-J. Nastold. Regularitätskriterien. Anwendungen. (German) [Regularity criteria. examples]. Lecture Notes in Mathematics, 38:83–109, 1967. CODEN LNMAA2. ISBN 3-540-03909-0 (print), 3-540-34977-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073991/.

Behara:1969:PIT

[BKW69] M. Behara, K. Krickeberg, and J. Wolfowitz, editors. Probability and Information Theory: Proceedings of the International Symposium at McMaster University, Canada, April, 1968, volume 89 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CO-DEN LNMAA2. ISBN 3-540-04608-9 (paperback), 3-540-36098-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN

QA3 .L28 no. 89. URL http://link.springer.com/book/10.
1007/BFb0079113.

Bloom:1968:ODE

[Blo68] Thomas Bloom. Opérateurs différentiels sur un espace analytique complexe. (French) [Differential operators on a complex analytic space]. Lecture Notes in Mathematics, 71:1–20, 1968. CODEN LNMAA2. ISBN 3-540-04241-5 (print), 3-540-35889-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080092/.

Boboc:1968:EBF

[BM68a] N. Boboc and P. Mustaţă. Espaces de Banach des fonctions hölderiennes. (French) [Banach spaces of Hölder functions]. Lecture Notes in Mathematics, 68:14–35, 1968. CODEN LNMAA2. ISBN 3-540-04238-5 (print), 3-540-35882-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097560/.

Boboc:1968:EHA

[BM68b] N. Boboc and P. Mustaţă. Espaces harmoniques associés aux opérateurs différentiels linéaires du second ordre de type elliptique. (French) [Harmonic spaces associated with linear second-order elliptic differential operators], volume 68 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. CODEN LNMAA2. ISBN 3-540-04238-5 (paperback), 3-540-35882-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 95 pp. LCCN QA3 .L28 no. 68. URL

http://link.springer.com/book/10.1007/BFb0097558.

Boboc:1968:EHF

[BM68c] N. Boboc and P. Mustaţă. Espaces harmoniques. (French) [Harmonic spaces]. Lecture Notes in Mathematics, 68:1–13, 1968. CO-DEN LNMAA2. ISBN 3-540-04238-5 (print), 3-540-35882-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0097559/.

Boboc:1968:ODE

[BM68d] N. Boboc and P. Mustaţă. Opérateurs différentiels elliptiques du second ordre. (French) [Elliptic differential operators of second order]. Lecture Notes in Mathematics, 68:36–75, 1968. CODEN LNMAA2. ISBN 3-540-04238-5 (print), 3-540-35882-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097561/.

Boboc:1968:PSP

[BM68e] N. Boboc and P. Mustaţă. Potentiels à support ponctuel et l’inégalité de Harnack. (French) [Potential at point support and the Harnack inequality]. Lecture Notes in Mathematics, 68:76–94, 1968. CODEN LNMAA2. ISBN 3-540-04238-5 (print), 3-540-35882-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097562/.

Borel:1964:ASE

[Bor64a] Armand Borel. Algèbre spectrale des espaces fibres. (French) [Spectral algebra of fiber spaces]. Lecture Notes in Mathematics, 2: VIII–1–VIII–8, 1964. CODEN LNMAA2. ISBN 3-540-03179-0 (print), 3-540-36803-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097859/.

Borel:1964:ASA

[Bor64b] Armand Borel. Algèbre spectrale d’une application continue. (French) [Spectral algebra of a continuous function]. Lecture Notes in Mathematics, 2:VII–1–VII–8, 1964. CODEN LNMAA2. ISBN 3-540-03179-0 (print), 3-540-36803-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097858/.

Borel:1964:AEF

[Bor64c] Armand Borel. Applications aux espaces fibres. (French) [Applications to fiber spaces]. Lecture Notes in Mathematics, 2:IX–1– IX–10, 1964. CODEN LNMAA2. ISBN 3-540-03179-0 (print), 3-540-36803-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097860/.

Borel:1964:ACF

[Bor64d] Armand Borel. Applications et compléments. (French) [Applications and complements]. Lecture Notes in Mathematics, 2:IV–1– IV–12, 1964. CODEN LNMAA2. ISBN 3-540-03179-0 (print), 3-540-36803-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097855/.

Borel:1964:CEL

[Bor64e] Armand Borel, editor. Cohomologie des espaces localement compacts: Exposés faits au Séminaire de Topologie algébrique de l’École Polytechnique Fédérale au printemps 1951 Troisième Edition, 1964. (French) [Cohomology of locally compact spaces: Presentations made at the Seminary of Algebraic Topology from the École Polytechnique Fédérale in spring 1951], volume 2 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1964. CODEN LNMAA2. ISBN 3-540-03179-0 (paperback), 3-540-36803-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 2;

QA3 .L28 no. 2, 1964. URL http://link.springer.com/book/
10.1007/BFb0097851.

Borel:1964:LSF

[Bor64f] Armand Borel. L’algébre spectrale. (French) [Spectral algebra]. Lecture Notes in Mathematics, 2:VI–1–VI–12, 1964. CODEN LN-MAA2. ISBN 3-540-03179-0 (print), 3-540-36803-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0097857/.

Borel:1964:TFF

[Bor64g] Armand Borel. Le théorème fondamental. (French) [The basic theorem]. Lecture Notes in Mathematics, 2:III–1–III–9, 1964. CO-DEN LNMAA2. ISBN 3-540-03179-0 (print), 3-540-36803-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0097854/.

Borel:1964:CFC

[Bor64h] Armand Borel. Les complexes. (French) [The complexes]. Lecture Notes in Mathematics, 2:II–1–II–14, 1964. CODEN LN-MAA2. ISBN 3-540-03179-0 (print), 3-540-36803-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0097853/.

Borel:1964:FFS

[Bor64i] Armand Borel. Les faisceaux k. (French) [The k sheaves]. Lecture Notes in Mathematics, 2:V–1–V–12, 1964. CODEN LN-MAA2. ISBN 3-540-03179-0 (print), 3-540-36803-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0097856/.

Borel:1964:NAF

[Bor64j] Armand Borel. Notions algébriques. (French) [Algebraic concepts]. Lecture Notes in Mathematics, 2:I–1–I–8, 1964. CODEN LN-MAA2. ISBN 3-540-03179-0 (print), 3-540-36803-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0097852/.

Borel:1966:CI

[Bor66a] A. Borel. Class invariants I. Lecture Notes in Mathematics, 21:III–1–III–8, 1966. CODEN LNMAA2. ISBN 3-540-03604-0 (print), 3-540-34798-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091553/.

Borel:1966:CII

[Bor66b] A. Borel. Class invariants II. Lecture Notes in Mathematics, 21:IV–1–IV–10, 1966. CODEN LNMAA2. ISBN 3-540-03604-0 (print), 3-540-34798-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091554/.

Borel:1967:CSC

[Bor67a] Armand Borel. Classifying spaces of the classical groups. Lecture Notes in Mathematics, 36:71–92, 1967. CODEN LNMAA2. ISBN 3-540-03907-4 (print), 3-540-34975-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096871/.

Borel:1967:HPS

[Bor67b] Armand Borel. The homological properties of H-spaces. Lecture Notes in Mathematics, 36:1–25, 1967. CODEN LNMAA2. ISBN 3-540-03907-4 (print), 3-540-34975-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096868/.

Borel:1967:SSF

[Bor67c] Armand Borel. Spectral sequence of a fibre bundle. Lecture Notes in Mathematics, 36:26–51, 1967. CODEN LNMAA2. ISBN 3-540-03907-4 (print), 3-540-34975-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096869/.

Borel:1967:THT

[Bor67d] Armand Borel, editor. Topics in the Homology Theory of Fibre Bundles: Lectures given at the University of Chicago, 1954, volume 36 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. CO-DEN LNMAA2. ISBN 3-540-03907-4 (paperback), 3-540-34975-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN

QA3 .L28 no. 36. URL http://link.springer.com/book/10.

1007/BFb0096867. Notes by Edward Halpern.

Borel:1967:UBC

[Bor67e] Armand Borel. Universal bundles and classifying spaces. Lecture Notes in Mathematics, 36:52–70, 1967. CODEN LNMAA2. ISBN 3-540-03907-4 (print), 3-540-34975-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096870/.

Bouwer:1969:SGF

[Bou69] I. Z. Bouwer. Section graphs for finite permutation groups. Lecture Notes in Mathematics, 110:55–61, 1969. CODEN LNMAA2. ISBN 3-540-04629-1 (print), 3-540-36161-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060105/.

Brinkmann:1966:ACG

[BP66a] H.-B. Brinkmann and D. Puppe. Addition und Coaddition. (German) [Addition and coaddition]. Lecture Notes in Mathematics, 18:71–91, 1966. CODEN LNMAA2. ISBN 3-540-03601-6 (print), 3-540-37191-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079030/.

Brinkmann:1966:AKG

[BP66b] H.-B. Brinkmann and D. Puppe. Additive Kategorien. (German) [Additive categories]. Lecture Notes in Mathematics, 18:

92–100, 1966. CODEN LNMAA2. ISBN 3-540-03601-6 (print), 3-540-37191-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079031/.

Brinkmann:1966:DFG

[BP66c] H.-B. Brinkmann and D. Puppe. Darstellbare Funktoren. (German) [Displayable functors]. Lecture Notes in Mathematics, 18: 33–37, 1966. CODEN LNMAA2. ISBN 3-540-03601-6 (print), 3-540-37191-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079026/.

Brinkmann:1966:EIG

[BP66d] H.-B. Brinkmann and D. Puppe. Einbettungen und Identifizierungen. (German) [Embeddings and identifications]. Lecture Notes in Mathematics, 18:38–53, 1966. CODEN LNMAA2. ISBN 3-540-03601-6 (print), 3-540-37191-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079027/.

Brinkmann:1966:KDF

[BP66e] H.-B. Brinkmann and D. Puppe. Kategorien, Dualität, Funktoren, Natürlichkeit. (German) [Categories, duality functors, naturalness]. Lecture Notes in Mathematics, 18:18–32, 1966. CODEN LNMAA2. ISBN 3-540-03601-6 (print), 3-540-37191-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079025/.

Brinkmann:1966:KFGa

[BP66f] H.-B. Brinkmann and D. Puppe. Kategorien und Funktoren. (German) [Categories and functors]. Lecture Notes in Mathematics, 18:1–5, 1966. CODEN LNMAA2. ISBN 3-540-03601-6 (print), 3-540-37191-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079023/.

Brinkmann:1966:LMG

[BP66g] H.-B. Brinkmann and D. Puppe. Logik und Mengenlehre. (German) [Logic and set theory]. Lecture Notes in Mathematics, 18: 6–17, 1966. CODEN LNMAA2. ISBN 3-540-03601-6 (print), 3-540-37191-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079024/.

Brinkmann:1966:NGN

[BP66h] H.-B. Brinkmann and D. Puppe. Nullmorphismen. (German) [Null morphisms]. Lecture Notes in Mathematics, 18:66–70, 1966. CO-DEN LNMAA2. ISBN 3-540-03601-6 (print), 3-540-37191-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0079029/.

Brinkmann:1966:PCG

[BP66i] H.-B. Brinkmann and D. Puppe. Produkte und Coprodukte. (German) [Products and coproducts]. Lecture Notes in Mathematics, 18:54–65, 1966. CODEN LNMAA2. ISBN 3-540-03601-6 (print), 3-540-37191-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079028/.

Brinkmann:1969:AEK

[BP69a] Hans-Berndt Brinkmann and Dieter Puppe. Abelsche und exakte Kategorien, Korrespondenzen. (German) [Abelian and exact categories, correspondence], volume 96 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CODEN LNMAA2. ISBN 3-540-04615-1 (paperback), 3-540-36132-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 141 pp. LCCN QA3 .L28 no. 96. URL

http://link.springer.com/book/10.1007/BFb0079036.

Brinkmann:1969:AMA

[BP69b] Hans-Berndt Brinkmann and Dieter Puppe. Addition von morphismen. Abelsche Kategorien. (German) [Addition of morphisms. Abelian categories]. Lecture Notes in Mathematics, 96:21–26, 1969. CODEN LNMAA2. ISBN 3-540-04615-1 (print), 3-540-36132-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0079039/.

Brinkmann:1969:DEK

[BP69c] Hans-Berndt Brinkmann and Dieter Puppe. Diagrammlemmata in exakten Kategorien. (German) [Diagram lemmas in precise categories]. Lecture Notes in Mathematics, 96:118–122, 1969. CODEN LNMAA2. ISBN 3-540-04615-1 (print), 3-540-36132-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079046/.

Brinkmann:1969:EFE

[BP69d] Hans-Berndt Brinkmann and Dieter Puppe. Exakte Folgen und exakte Quadrate. (German) [Exact sequences and exact squares]. Lecture Notes in Mathematics, 96:36–51, 1969. CODEN LN-MAA2. ISBN 3-540-04615-1 (print), 3-540-36132-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079041/.

Brinkmann:1969:HGH

[BP69e] Hans-Berndt Brinkmann and Dieter Puppe. Homoloqie. (German) [Homology]. Lecture Notes in Mathematics, 96:102–117, 1969. CO-DEN LNMAA2. ISBN 3-540-04615-1 (print), 3-540-36132-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0079045/.

Brinkmann:1969:KCQ

[BP69f] Hans-Berndt Brinkmann and Dieter Puppe. Kartesische und cokartesische Quadrate. (German) [Cartesian and co-Cartesian squares]. Lecture Notes in Mathematics, 96:27–35, 1969. CO-DEN LNMAA2. ISBN 3-540-04615-1 (print), 3-540-36132-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0079040/.

Brinkmann:1969:KKG

[BP69g] Hans-Berndt Brinkmann and Dieter Puppe. Kategorien von Korrespondenzen. (German) [Categories of correspondences]. Lecture Notes in Mathematics, 96:52–70, 1969. CODEN LNMAA2. ISBN 3-540-04615-1 (print), 3-540-36132-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079042/.

Brinkmann:1969:KCG

[BP69h] Hans-Berndt Brinkmann and Dieter Puppe. Kerne und Cokerne. (German) [Kernels and co-kernels]. Lecture Notes in Mathematics, 96:1–11, 1969. CODEN LNMAA2. ISBN 3-540-04615-1 (print), 3-540-36132-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079037/.

Brinkmann:1969:KEK

[BP69i] Hans-Berndt Brinkmann and Dieter Puppe. Korrespondenzen über exakten Kategorien. (German) [Correspondences on exact categories]. Lecture Notes in Mathematics, 96:84–101, 1969. CODEN LNMAA2. ISBN 3-540-04615-1 (print), 3-540-36132-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079044/.

Brinkmann:1969:VQZ

[BP69j] Hans-Berndt Brinkmann and Dieter Puppe. Vollkommutative Quadrate, Zerlegung von Korrespondenzen. (German) [Fullycommutative squares, decomposition of correspondences]. Lecture Notes in Mathematics, 96:71–83, 1969. CODEN LNMAA2. ISBN 3-540-04615-1 (print), 3-540-36132-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079043/.

Brinkmann:1969:ZME

[BP69k] Hans-Berndt Brinkmann and Dieter Puppe. Zerlegung von Morphismen. Exakte Kategorien. (German) [Decomposition of morphisms. Exact categories]. Lecture Notes in Mathematics, 96:12– 20, 1969. CODEN LNMAA2. ISBN 3-540-04615-1 (print), 3-540-36132-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079038/.

Bredon:1967:ECC

[Bre67a] Glen E. Bredon. Equivariant classical cohomology. Lecture Notes in Mathematics, 34:I.1–I.27, 1967. CODEN LNMAA2. ISBN 3-540-03905-8 (print), 3-540-34973-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082691/.

Bredon:1967:ECT

[Bre67b] Glen E. Bredon. Equivariant Cohomology Theories, volume 34 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. CODEN LN-MAA2. ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 64

pp. LCCN QA3 .L28 no. 34. URL http://link.springer.com/
book/10.1007/BFb0082690.

Bredon:1967:EOT

[Bre67c] Glen E. Bredon. Equivariant obstruction theory. Lecture Notes in Mathematics, 34:II.1–II.18, 1967. CODEN LNMAA2. ISBN 3-540-03905-8 (print), 3-540-34973-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082692/.

Bredon:1967:FSF

[Bre67d] Glen E. Bredon. Function spaces, fibrations and spectra. Lecture Notes in Mathematics, 34:III.1–III.8, 1967. CODEN LN-MAA2. ISBN 3-540-03905-8 (print), 3-540-34973-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082693/.

Bredon:1967:GEC

[Bre67e] Glen E. Bredon. Generalized equivariant cohomology. Lecture Notes in Mathematics, 34:IV.1–IV.11, 1967. CODEN LN-MAA2. ISBN 3-540-03905-8 (print), 3-540-34973-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082694/.

Brelot:1967:TFT

[Bre67f] Marcel Brelot. La topologie fine en théorie du potentiel. (French) [Fine topology in potential theory]. Lecture Notes in Mathematics, 31:36–47, 1967. CODEN LNMAA2. ISBN 3-540-03902-3 (print), 3-540-34970-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061105/.

Brinkmann:1966:KFGb

[Bri66] Hans-Bernd Brinkmann. Kategorien und Funktoren. (German) [Categories and functors], volume 18 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1966. CODEN LNMAA2. ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 107 pp. LCCN QA3

.L28 no. 18. URL http://link.springer.com/book/10.1007/

BFb0079022. With an introduction by Dieter Puppe.

Brinkmann:1969:RGE

[Bri69] Hans-Berndt Brinkmann. Relations for groups and for exact categories. Lecture Notes in Mathematics, 92:1–9, 1969. CODEN LNMAA2. ISBN 3-540-04611-9 (print), 3-540-36101-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080762/.

Bhatia:1967:DSE

[BS67a] N. P. Bhatia and G. P. Szegő. Dynamical systems in a Euclidean space. Lecture Notes in Mathematics, 35:9–113, 1967. CODEN LNMAA2. ISBN 3-540-03906-6 (print), 3-540-34974-X (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080632/.

Bhatia:1967:DSM

[BS67b] N. P. Bhatia and G. P. Szegő. Dynamical systems in metric spaces. Lecture Notes in Mathematics, 35:114–245, 1967. CODEN LN-MAA2. ISBN 3-540-03906-6 (print), 3-540-34974-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080633/.

Bhatia:1967:DSS

[BS67c] N. P. Bhatia and G. P. Szegő. Dynamical Systems: Stability Theory and Applications, volume 35 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. CODEN LNMAA2. ISBN 3-540-03906-6 (paperback), 3-540-34974-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 416 pp. LCCN QA3 .L28 no. 35. URL

http://link.springer.com/book/10.1007/BFb0080630.

Bhatia:1967:NTP

[BS67d] N. P. Bhatia and G. P. Szegő. Notation, terminology and preliminary lemmas. Lecture Notes in Mathematics, 35:1–8, 1967. CODEN LNMAA2. ISBN 3-540-03906-6 (print), 3-540-34974-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080631/.

Bhatia:1967:SML

[BS67e] N. P. Bhatia and G. P. Szegő. The second method of Liapunov for ordinary differential equations. Lecture Notes in Mathematics, 35:246–367, 1967. CODEN LNMAA2. ISBN 3-540-03906-6 (print), 3-540-34974-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080634/.

Buchsbaum:1968:HUR

[Buc68] David A. Buchsbaum. Homology and universality relative to a functor. Lecture Notes in Mathematics, 61:28–40, 1968. CODEN LNMAA2. ISBN 3-540-04231-8 (print), 3-540-35863-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077115/.

Buchsbaum:1969:LRL

[Buc69a] D. A. Buchsbaum. Lectures on regular local rings. Lecture Notes in Mathematics, 86:13–32, 1969. CODEN LNMAA2. ISBN 3-540-04605-4 (print), 3-540-36095-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079382/.

Bucur:1969:TCA

[Buc69b] I. Bucur. Triangulated categories and algebraic K-theory. Lecture Notes in Mathematics, 108:28–54, 1969. CODEN LNMAA2. ISBN 3-540-04627-5 (print), 3-540-36156-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059989/.

Butcher:1969:EOR

[But69] J. C. Butcher. The effective order of Runge–Kutta methods. Lecture Notes in Mathematics, 109:133–139, 1969. CODEN LN-MAA2. ISBN 3-540-04628-3 (print), 3-540-36158-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0060019/.

Cairoli:1969:EPP

[Cai69] R. Cairoli. Étude probabiliste d’un problème de Dirichlet. (French) [Probabilistic study of a problem of Dirichlet]. Lecture Notes in Mathematics, 88:34–92, 1969. CODEN LNMAA2. ISBN 3-540-04607-0 (print), 3-540-36097-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081626/.

Campbell:1969:SER

[Cam69] L. L. Campbell. Series expansions for random processes. Lecture Notes in Mathematics, 89:77–95, 1969. CODEN LNMAA2. ISBN 3-540-04608-9 (print), 3-540-36098-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079118/.

Carleson:1969:CT

[Car69] Lennart Carleson. The corona theorem. Lecture Notes in Mathematics, 118:121–132, 1969. CODEN LNMAA2. ISBN 3-540-04907-X (print), 3-540-36246-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060255/.

Cassels:1969:FPS

[Cas69] J. W. S. Cassels. Factorization of polynomials in several variables. Lecture Notes in Mathematics, 118:1–17, 1969. CODEN LNMAA2. ISBN 3-540-04907-X (print), 3-540-36246-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060248/.

Cathelineau:1968:MTF

[Cat68] J.-L. Cathelineau. Sur les modules topologiques. (French) [On topological modules]. Lecture Notes in Mathematics, 71:99– 112, 1968. CODEN LNMAA2. ISBN 3-540-04241-5 (print), 3-540-35889-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080101/.

Constantinescu:1968:ETH

[CC68] C. Constantinescu and A. Cornea. Examples in the theory of harmonic spaces. Lecture Notes in Mathematics, 69:161–171, 1968. CODEN LNMAA2. ISBN 3-540-04239-3 (print), 3-540-35884-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080478/.

Cellina:1968:CSP

[Cel68] Arrigo Cellina. Characterizing solutions of the Pontriagin maximum principle. Lecture Notes in Mathematics, 60:37–47, 1968. CODEN LNMAA2. ISBN 3-540-04230-X (print), 3-540-35862-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077126/.

Cerf:1968:CSA

[Cer68a] Jean Cerf. Construction d’une section additive pour le revêtement R. (French) [Construction of an additive section to the covering R]. Lecture Notes in Mathematics, 53:94–112, 1968. CODEN LNMAA2. ISBN 3-540-04223-7 (print), 3-540-35831-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060401/.

Cerf:1968:EFL

[Cer68b] Jean Cerf. Espaces fonctionnels liés aux doubles décompositions d’une sphère plongée dans R^{3}. (French) [Functional spaces related to double decompositions of a sphere immersed in R^{3}]. Lecture Notes in Mathematics, 53:51–75, 1968. CODEN LNMAA2. ISBN 3-540-04223-7 (print), 3-540-35831-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060399/.

Cerf:1968:NGD

[Cer68c] Jean Cerf. La nullité de Γ_{4}, généralisation du théorème de Schönflies pour S^{2}. (French) [The nullity of Γ_{4}, generalization of the Schönflies theorem for S^{2}]. Lecture Notes in Mathematics, 53:1–10, 1968. CODEN LNMAA2. ISBN 3-540-04223-7 (print), 3-540-35831-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060396/.

Cerf:1968:TSP

[Cer68d] Jean Cerf. Le théorème de Schönflies pour S^{2}. (French) [The Schönflies theorem for S^{2}]. Lecture Notes in Mathematics, 53: 39–50, 1968. CODEN LNMAA2. ISBN 3-540-04223-7 (print), 3-540-35831-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060398/.

Cerf:1968:SVP

[Cer68e] Jean Cerf. Les sous-variétés de petite complexité. (French) [Subvarieties of small complexity]. Lecture Notes in Mathematics, 53: 76–93, 1968. CODEN LNMAA2. ISBN 3-540-04223-7 (print), 3-540-35831-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060400/.

Cerf:1968:SCF

[Cer68f] Jean Cerf. Singularités de codimension 1 des fonctions différentiables réelles définies sur une 2-variété. Application: Subdivision cocellulaire de l’espace des 2-sphères de R^{3}. (French) [Singularities of codimension 1 of functions defined on a real differentiable 2manifold. Application: Cocellular subdivision of space of 2-spheres in R^{3}]. Lecture Notes in Mathematics, 53:11–38, 1968. CODEN LNMAA2. ISBN 3-540-04223-7 (print), 3-540-35831-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060397/.

Cerf:1968:DSD

[Cer68g] Jean Cerf. Sur les difféomorphismes de la sphère de dimension trois (Γ_{4} = 0). (French) [Diffeomorphisms on the sphere of threedimensional (Γ_{4} = 0)], volume 53 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. CODEN LNMAA2. ISBN 3-540-04223-7 (paperback), 3-540-35831-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 133 pp. LCCN QA3 .L28 no. 53.

URL http://link.springer.com/book/10.1007/BFb0060395.

Conner:1966:CCC

[CF66a] P. E. Conner and E. E. Floyd. Cobordism characteristic classes. Lecture Notes in Mathematics, 28:38–68, 1966. CODEN LN-MAA2. ISBN 3-540-03610-5 (print), 3-540-69974-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0071093/.

Conner:1966:TIT

[CF66b] P. E. Conner and E. E. Floyd. The Thom isomorphism in Ktheory. Lecture Notes in Mathematics, 28:1–37, 1966. CODEN LNMAA2. ISBN 3-540-03610-5 (print), 3-540-69974-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071092/.

Conner:1966:MFB

[CF66c] P. E. Conner and E. E. Floyd. U-manifolds with framed boundaries. Lecture Notes in Mathematics, 28:69–110, 1966. CODEN LNMAA2. ISBN 3-540-03610-5 (print), 3-540-69974-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071094/.

Conner:1966:RCT

[CF66d] P. E. (Pierre E.) Conner and E. E. Floyd. The Relation of Cobordism to K-Theories, volume 28 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1966. CODEN LNMAA2. ISSN 0075-8434 (print), 1617-9692 (electronic). 110 + 2 pp. LCCN QA3 .L28 no. 28. URL

http://link.springer.com/book/10.1007/BFb0071091.

Carter:1969:AP

[CF69a] P. Carter and D. Foata. Applications probabilistes. Lecture Notes in Mathematics, 85:72–83, 1969. CODEN LNMAA2. ISBN 3-540-04604-6 (print), 3-540-36094-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079476/.

Carter:1969:CGF

[CF69b] P. Carter and D. Foata. Circuits dans un graphe. (French) [Circuits in a graph]. Lecture Notes in Mathematics, 85:24–38, 1969. CODEN LNMAA2. ISBN 3-540-04604-6 (print), 3-540-36094-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0079472/.

Carter:1969:FMM

[CF69c] P. Carter and D. Foata. Fonction de Möbius d’un monoïde. (French) [Möbius function of a monoid]. Lecture Notes in Mathematics, 85:18–23, 1969. CODEN LNMAA2. ISBN 3-540-04604-6 (print), 3-540-36094-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079471/.

Carter:1969:IFI

[CF69d] P. Carter and D. Foata. Introduction. (French) [Introduction]. Lecture Notes in Mathematics, 85:1–7, 1969. CODEN LNMAA2. ISBN 3-540-04604-6 (print), 3-540-36094-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079469/.

Carter:1969:MDP

[CF69e] P. Carter and D. Foata. Monoïdes définis par des relations de commutation. (French) [Monoids defined by commutation relations]. Lecture Notes in Mathematics, 85:8–17, 1969. CODEN LNMAA2. ISBN 3-540-04604-6 (print), 3-540-36094-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079470/.

Carter:1969:RSF

[CF69f] P. Carter and D. Foata. Réarrangements de suites. (French) [Rearrangements of sequences]. Lecture Notes in Mathematics, 85: 39–53, 1969. CODEN LNMAA2. ISBN 3-540-04604-6 (print), 3-540-36094-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079473/.

Carter:1969:REC

[CF69g] P. Carter and D. Foata. Relations entre coefficients binomiaux. Lecture Notes in Mathematics, 85:61–71, 1969. CODEN LN-MAA2. ISBN 3-540-04604-6 (print), 3-540-36094-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079475/.

Carter:1969:MTM

[CF69h] P. Carter and D. Foata. Sur le “Master theorem” de Macmahon. (French) [On the “Master theorem” of Macmahon]. Lecture Notes in Mathematics, 85:54–60, 1969. CODEN LNMAA2. ISBN 3-540-04604-6 (print), 3-540-36094-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079474/.

Cartier:1969:PCC

[CF69i] P. Cartier and D. Foata. Problèmes combinatoires de commutation et réarrangements. (French) [Combinatorial problems of switching and rearrangements], volume 85 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CODEN LNMAA2. ISBN 3-540-04604-6 (paperback), 3-540-36094-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 88 pp. LCCN QA3 .L28 no. 85. URL

http://link.springer.com/book/10.1007/BFb0079468.

Chandrasekharan:1966:BM

[Cha66a] K. Chandrasekharan. Back matter. Lecture Notes in Mathematics, 29:??, 1966. CODEN LNMAA2. ISBN 3-540-03611-3 (print), 3-540-34855-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-34855-9/1.pdf.

Chandrasekharan:1966:QRG

[Cha66b] K. Chandrasekharan. Das quadratische Reziprozitätsgesetz. (German) [The law of quadratic reciprocity]. Lecture Notes in Mathematics, 29:48–71, 1966. CODEN LNMAA2. ISBN 3-540-03611-3 (print), 3-540-34855-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082660/.

Chandrasekharan:1966:DSP

[Cha66c] K. Chandrasekharan. Der Dirichletsche Satz von den Primzahlen in einer arithmetischen Progression. (German) [The Dirichlet set of prime numbers in an arithmetic progression]. Lecture Notes in Mathematics, 29:160–187, 1966. CODEN LNMAA2. ISBN 3-540-03611-3 (print), 3-540-34855-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082665/.

Chandrasekharan:1966:FEZ

[Cha66d] K. Chandrasekharan. Der Fundamentalsatz der elementaren Zahlentheorie. (German) [The fundamental theorem of elementary number theory]. Lecture Notes in Mathematics, 29:1–13, 1966. CODEN LNMAA2. ISBN 3-540-03611-3 (print), 3-540-34855-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082656/.

Chandrasekharan:1966:PGP

[Cha66e] K. Chandrasekharan. Der Primzahlsatz. (German) [The prime number theorem]. Lecture Notes in Mathematics, 29:188–199, 1966. CODEN LNMAA2. ISBN 3-540-03611-3 (print), 3-540-34855-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0082666/.

Chandrasekharan:1966:SCV

[Cha66f] K. Chandrasekharan. Der Satz von Chebychev über die Verteilung der Primzahlen. (German) [The Chebyshev theorem on the distribution of primes]. Lecture Notes in Mathematics, 29:99–130, 1966. CODEN LNMAA2. ISBN 3-540-03611-3 (print), 3-540-34855-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0082662/.

Chandrasekharan:1966:SMG

[Cha66g] K. Chandrasekharan. Der Satz von Minkowski über Gitterpunkte in konvexen Bereichen. (German) [The Minkowski theorem on lattice points in convex regions]. Lecture Notes in Mathematics, 29:147–159, 1966. CODEN LNMAA2. ISBN 3-540-03611-3 (print), 3-540-34855-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082664/.

Chandrasekharan:1966:RAI

[Cha66h] K. Chandrasekharan. Die rationale Approximation einer irrationalen Zahl. Der Satz von Hurwitz. (German) [The rational approximation of an irrational number. The Hurwitz theorem]. Lecture Notes in Mathematics, 29:23–37, 1966. CODEN LN-MAA2. ISBN 3-540-03611-3 (print), 3-540-34855-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082658/.

Chandrasekharan:1966:WGZ

[Cha66i] K. Chandrasekharan. Die Weylsche “Gleichverteilung von Zahlen mod1” und der Satz von Kronecker. (German) [The Weyl “Equal distribution of numbers mod1” and the Kronecker theorem]. Lecture Notes in Mathematics, 29:131–146, 1966. CODEN LN-MAA2. ISBN 3-540-03611-3 (print), 3-540-34855-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082663/.

Chandrasekharan:1966:FM

[Cha66j] K. Chandrasekharan. Front matter. Lecture Notes in Mathematics, 29:??, 1966. CODEN LNMAA2. ISBN 3-540-03611-3 (print), 3-540-34855-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-34855-9/1.pdf.

Chandrasekharan:1966:KGC

[Cha66k] K. Chandrasekharan. Kongruenzen. (German) [Congruences]. Lecture Notes in Mathematics, 29:14–22, 1966. CODEN LN-MAA2. ISBN 3-540-03611-3 (print), 3-540-34855-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082657/.

Chandrasekharan:1966:QRD

[Cha66l] K. Chandrasekharan. Quadratische Reste, und die Darstellbarkeit einer positiven ganzen Zahl als Summe von vier Quadraten. (German) [Quadratic residues and the representability of a positive integer as a sum of four squares]. Lecture Notes in Mathematics, 29:38–47, 1966. CODEN LNMAA2. ISBN 3-540-03611-3 (print), 3-540-34855-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082659/.

Chandrasekharan:1966:ZFG

[Cha66m] K. Chandrasekharan. Zahlentheoretische Funktionen und Gitterpunkte. (German) [Number theoretic functions and lattice points]. Lecture Notes in Mathematics, 29:72–98, 1966. CODEN LN-MAA2. ISBN 3-540-03611-3 (print), 3-540-34855-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082661/.

Chandrasekharan:1966:EAZ

[Cha66n] K. (Komaravolu) Chandrasekharan. Einführung in die Analytische Zahlentheorie. (German) [Introduction to Analytic Number Theory], volume 29 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1966. CODEN LNMAA2. ISBN 3-540-03611-3 (paperback), 3-540-34855-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 199 pp. LCCN QA3 .L28 no. 29. URL http://link.

springer.com/book/10.1007/BFb0082655.

Chatterji:1967:CMC

[Cha67] S. D. Chatterji. Comments on the martingale convergence theorem. Lecture Notes in Mathematics, 31:55–61, 1967. CODEN LNMAA2. ISBN 3-540-03902-3 (print), 3-540-34970-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061107/.

Chang:1968:SRM

[Cha68] C. C. Chang. Some remarks on the model theory of infinitary languages. Lecture Notes in Mathematics, 72:36–63, 1968. CO-DEN LNMAA2. ISBN 3-540-04242-3 (print), 3-540-35900-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0079681/.

Chase:1969:GOEa

[Cha69a] Stephen U. Chase. Galois objects and extensions of Hopf algebras. Lecture Notes in Mathematics, 92:10–31, 1969. CODEN LNMAA2. ISBN 3-540-04611-9 (print), 3-540-36101-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080763/.

Chase:1969:GOEb

[Cha69b] Stephen U. Chase. Galois objects and extensions of Hopf algebras. Lecture Notes in Mathematics, 97:84–126, 1969. CODEN LNMAA2. ISBN 3-540-04616-X (print), 3-540-36134-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101436/.

Chowla:1966:RCI

[Cho66] S. Chowla. Remarks on class-invariants and related topics. Lecture Notes in Mathematics, 21:VI–1–VII–15, 1966. CODEN LN-MAA2. ISBN 3-540-03604-0 (print), 3-540-34798-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0091556/.

Chartrand:1969:MFG

[CK69] G. Chartrand and S. F. Kapoor, editors. The Many Facets of Graph Theory: Proceedings of the Conference held at Western Michigan University, Kalamazoo /MI, October 31–November 2, 1968, volume 110 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CO-DEN LNMAA2. ISBN 3-540-04629-1 (paperback), 3-540-36161-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN

QA3 .L28 no. 110. URL http://link.springer.com/book/10.
1007/BFb0060099.

Cleave:1968:HU

[Cle68] J. P. Cleave. Hyperarithmetic ultrafilters. Lecture Notes in Mathematics, 70:223–240, 1968. CODEN LNMAA2. ISBN 3-540-04240-7 (print), 3-540-35886-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079097/.

Cartier:1968:RDT

[CMW68] P. Cartier, P.-A. Meyer, and M. Weil. Le retournement du temps: Compléments à l’exposé de m. weil. (French) [The turnaround time: Additions to the presentation by Mr. Weil]. Lecture Notes in Mathematics, 51:22–33, 1968. CODEN LNMAA2. ISBN 3-540-04221-0 (print), 3-540-35827-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059177/.

Conner:1967:A

[Con67a] P. E. Conner. Applications. Lecture Notes in Mathematics, 46: 100–102, 1967. CODEN LNMAA2. ISBN 3-540-03917-1 (print), 3-540-35543-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077385/.

Conner:1967:ABF

[Con67b] P. E. Conner. The Atiyah–Bott formula. Lecture Notes in Mathematics, 46:60–61, 1967. CODEN LNMAA2. ISBN 3-540-03917-1 (print), 3-540-35543-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077380.

Conner:1967:DFP

[Con67c] P. E. Conner. Dimension of the fixed point set. Lecture Notes in Mathematics, 46:103–106, 1967. CODEN LNMAA2. ISBN 3-540-03917-1 (print), 3-540-35543-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077386/.

Conner:1967:ER

[Con67d] P. E. Conner. Extension and reduction. Lecture Notes in Mathematics, 46:30–32, 1967. CODEN LNMAA2. ISBN 3-540-03917-1 (print), 3-540-35543-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077374/.

Conner:1967:FHF

[Con67e] P. E. Conner. The functor O_{∗}(H; F, F). Lecture Notes in Mathematics, 46:33–35, 1967. CODEN LNMAA2. ISBN 3-540-03917-1 (print), 3-540-35543-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077375/.

Conner:1967:Ib

[Con67f] P. E. Conner. The index. Lecture Notes in Mathematics, 46: 19–24, 1967. CODEN LNMAA2. ISBN 3-540-03917-1 (print), 3-540-35543-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077372/.

Conner:1967:IBI

[Con67g] P. E. Conner. Index as a bordism invariant. Lecture Notes in Mathematics, 46:36–40, 1967. CODEN LNMAA2. ISBN 3-540-03917-1 (print), 3-540-35543-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077376/.

Conner:1967:Ia

[Con67h] P. E. Conner. Introduction. Lecture Notes in Mathematics, 46: 1–18, 1967. CODEN LNMAA2. ISBN 3-540-03917-1 (print), 3-540-35543-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077371/.

Conner:1967:LI

[Con67i] P. E. Conner. The local invariant. Lecture Notes in Mathematics, 46:44–49, 1967. CODEN LNMAA2. ISBN 3-540-03917-1 (print), 3-540-35543-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077378/.

Conner:1967:LIZ

[Con67j] P. E. Conner. A local invariant for O_{∗}^{U}(Z_{2}). Lecture Notes in Mathematics, 46:107–115, 1967. CODEN LNMAA2. ISBN 3-540-03917-1 (print), 3-540-35543-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077387/.

Conner:1967:PMR

[Con67k] P. E. Conner. Periodic maps on a Riemann surface. Lecture Notes in Mathematics, 46:50–59, 1967. CODEN LNMAA2. ISBN 3-540-03917-1 (print), 3-540-35543-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077379/.

Conner:1967:RR

[Con67l] P. E. Conner. Real representations. Lecture Notes in Mathematics, 46:25–29, 1967. CODEN LNMAA2. ISBN 3-540-03917-1 (print), 3-540-35543-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077373/.

Conner:1967:RC

[Con67m] P. E. Conner. The ring C(Z_{2}) of constructions. Lecture Notes in Mathematics, 46:92–99, 1967. CODEN LNMAA2. ISBN 3-540-03917-1 (print), 3-540-35543-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077384/.

Conner:1967:RRZ

[Con67n] P. E. Conner. The ring R(Z_{2}). Lecture Notes in Mathematics, 46:80–91, 1967. CODEN LNMAA2. ISBN 3-540-03917-1 (print), 3-540-35543-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077383/.

Conner:1967:RF

[Con67o] P. E. Conner. The ring F. Lecture Notes in Mathematics, 46: 68–79, 1967. CODEN LNMAA2. ISBN 3-540-03917-1 (print), 3-540-35543-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077382/.

Conner:1967:SPM

[Con67p] P. E. Conner. Seminar on Periodic Maps, volume 46 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. CODEN LNMAA2. ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 116 pp. LCCN

QA3 .L28 no. 46. URL http://link.springer.com/book/10.
1007/BFb0077370.

Conner:1967:TI

[Con67q] P. E. Conner. The trace invariant. Lecture Notes in Mathematics, 46:41–43, 1967. CODEN LNMAA2. ISBN 3-540-03917-1 (print), 3-540-35543-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077377/.

Conner:1967:WCI

[Con67r] P. E. Conner. Weakly complex involutions. Lecture Notes in Mathematics, 46:62–67, 1967. CODEN LNMAA2. ISBN 3-540-03917-1 (print), 3-540-35543-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077381/.

Conner:1968:I

[Con68a] Pierre E. Conner. Introduction. Lecture Notes in Mathematics, 73:1–18, 1968. CODEN LNMAA2. ISBN 3-540-04243-1 (print), 3-540-35915-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079603/.

Conner:1968:LAF

[Con68b] Pierre E. Conner. Lectures on the Action of a Finite Group, volume 73 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. CO-DEN LNMAA2. ISBN 3-540-04243-1 (paperback), 3-540-35915-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 123

pp. LCCN QA3 .L28 no. 73. URL http://link.springer.com/
book/10.1007/BFb0079602.

Conner:1968:LBO

[Con68c] Pierre E. Conner. Line bundles with operators. Lecture Notes in Mathematics, 73:19–81, 1968. CODEN LNMAA2. ISBN 3-540-04243-1 (print), 3-540-35915-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079604/.

Conner:1968:OPI

[Con68d] Pierre E. Conner. Orientation preserving involutions. Lecture Notes in Mathematics, 73:82–122, 1968. CODEN LNMAA2. ISBN 3-540-04243-1 (print), 3-540-35915-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079605/.

Cooper:1969:EBS

[Coo69] G. J. Cooper. Error bounds for some single step methods. Lecture Notes in Mathematics, 109:140–147, 1969. CODEN LN-MAA2. ISBN 3-540-04628-3 (print), 3-540-36158-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0060020/.

Cornea:1968:WCS

[Cor68] Aurel Cornea. Weak compact sets in vector lattices and convergence theorems in harmonic spaces. Lecture Notes in Mathematics, 69:173–180, 1968. CODEN LNMAA2. ISBN 3-540-04239-3 (print), 3-540-35884-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080479/.

Crossley:1968:RES

[Cro68] John N. Crossley. Recursive equivalence: a survey. Lecture Notes in Mathematics, 70:241–251, 1968. CODEN LNMAA2. ISBN 3-540-04240-7 (print), 3-540-35886-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079098/.

Crowe:1969:NRP

[Cro69] D. W. Crowe. Nearly regular polyhedra with two exceptional faces. Lecture Notes in Mathematics, 110:63–76, 1969. CODEN LNMAA2. ISBN 3-540-04629-1 (print), 3-540-36161-8 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060106/.

Chase:1969:GO

[CS69a] Stephen U. Chase and Moss E. Sweedler. Galois objects. Lecture Notes in Mathematics, 97:4–51, 1969. CODEN LNMAA2. ISBN 3-540-04616-X (print), 3-540-36134-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101434/.

Chase:1969:HAGa

[CS69b] Stephen U. Chase and Moss E. Sweedler. Hopf algebras and Galois theory. Lecture Notes in Mathematics, 97:52–83, 1969. CO-DEN LNMAA2. ISBN 3-540-04616-X (print), 3-540-36134-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0101435/.

Chase:1969:HAGb

[CS69c] Stephen U. Chase and Moss E. Sweedler. Hopf Algebras and Galois Theory, volume 97 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CODEN LNMAA2. ISBN 3-540-04616-X (paperback), 3-540-36134-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 133 pp. LCCN QA3 .L28 no. 97. URL http://link.

springer.com/book/10.1007/BFb0101433.

Csorgo:1969:GCT

[Csö69] Miklós Csörgő. Glivenko–Cantelli type theorems for distance functions based on the modified empirical distribution function of M. Kac and for the empirical process with random sample size in general. Lecture Notes in Mathematics, 89:96–98, 1969. CODEN LNMAA2. ISBN 3-540-04608-9 (print), 3-540-36098-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079119/.

Dahl:1969:ANO

[Dah69a] Olav Dahl. Approximation of nonlinear operators. Lecture Notes in Mathematics, 109:148–153, 1969. CODEN LNMAA2. ISBN 3-540-04628-3 (print), 3-540-36158-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060021/.

Dahl:1969:PLT

[Dah69b] Ole-Johan Dahl. Programming languages as tools for the formulation of concepts. Lecture Notes in Mathematics, 118:18–29, 1969. CODEN LNMAA2. ISBN 3-540-04907-X (print), 3-540-36246-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060249/.

Davis:1967:ESF

[Dav67] Robert Davis. Equational systems of functors. Lecture Notes in Mathematics, 47:92–109, 1967. CODEN LNMAA2. ISBN 3-540-03918-X (print), 3-540-35545-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074301/.

Davis:1968:HRM

[Dav68] A. S. Davis. Half-ring morphologies. Lecture Notes in Mathematics, 70:253–268, 1968. CODEN LNMAA2. ISBN 3-540-04240-7 (print), 3-540-35886-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079099/.

deCairoli:1967:SGT

[dC67a] R. de Cairoli. Semi-groupes de transition et fonctions excessives. (French) [Transition semigroups and excessive functions]. Lecture Notes in Mathematics, 39:18–33, 1967. CODEN LNMAA2. ISBN 3-540-03910-4 (print), 3-540-34978-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080543/.

deCairoli:1967:SPF

[dC67b] R. de Cairoli. Séminaire de probabilités. (French) [Probability seminar]. Lecture Notes in Mathematics, 39:34–51, 1967. CODEN LNMAA2. ISBN 3-540-03910-4 (print), 3-540-34978-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080544/.

deCairoli:1967:CAT

[dC67c] R. de Cairoli. Un complément au théorème de Weierstrass– Stone par Claude Dellacherie. (French) [A complement to the Weierstrass–Stone theorem by Claude Dellacherie]. Lecture Notes in Mathematics, 39:52–53, 1967. CODEN LNMAA2. ISBN 3-540-03910-4 (print), 3-540-34978-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080545.

deSamLazaro:1968:MSO

[de 68] José de Sam Lazaro. Sur les moments spectraux d’ordre superieur. (French) [On higher order spectral moments]. Lecture Notes in Mathematics, 51:123–139, 1968. CODEN LNMAA2. ISBN 3-540-04221-0 (print), 3-540-35827-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059182/.

Dold:1967:SPU

[DE67] A. Dold and B. Eckmann, editors. Séminaire de Probabilités I Université de Strasbourg: novembre 1966–février 1967, volume 39 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. CO-DEN LNMAA2. ISBN 3-540-03910-4 (paperback), 3-540-34978-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN

QA3 .L28 no. 39. URL http://link.springer.com/book/10.
1007/BFb0080541.

Dold:1968:SPI

[DE68] A. Dold and B. Eckmann, editors. Séminaire de Probabilités II. Université de Strasbourg mars 1967–octobre 1967, volume 51 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. CO-DEN LNMAA2. ISBN 3-540-04221-0 (paperback), 3-540-35827-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN

QA3 .L28 no. 51. URL http://link.springer.com/book/10.
1007/BFb0059175.

Dold:1969:SPI

[DE69] A. Dold and B. Eckmann, editors. Séminaire de Probabilités III Université de Strasbourg: Octobre 1967–Juin 1968, volume 88 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CO-DEN LNMAA2. ISBN 3-540-04607-0 (paperback), 3-540-36097-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN

QA3 .L28 no. 88. URL http://link.springer.com/book/10.
1007/BFb0081623.

Dedecker:1969:TDN

[Ded69] Paul Dedecker. Three dimensional non-Abelian cohomology for groups. Lecture Notes in Mathematics, 92:32–64, 1969. CODEN LNMAA2. ISBN 3-540-04611-9 (print), 3-540-36101-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080764/.

Dellacherie:1969:EAF

[Del69a] C. Dellacherie. Ensembles aléatoires I. (French) [Random sets I]. Lecture Notes in Mathematics, 88:97–114, 1969. CODEN LN-MAA2. ISBN 3-540-04607-0 (print), 3-540-36097-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081628/.

Dellacherie:1969:EAI

[Del69b] C. Dellacherie. Ensembles aléatoires II. (French) [Random sets II]. Lecture Notes in Mathematics, 88:115–136, 1969. CODEN LNMAA2. ISBN 3-540-04607-0 (print), 3-540-36097-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081629/. See correc-

tion [Del71].

Dellacherie:1969:AFA

[Del69c] Claude Dellacherie. Une application aux fonctionnelles additives d’un théorème de Mokobodzki. (French) [An application of a functional additive theorem of Mokobodzki]. Lecture Notes in Mathematics, 88:93–96, 1969. CODEN LNMAA2. ISBN 3-540-04607-0 (print), 3-540-36097-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081627/.

Dellacherie:1971:CEA

[Del71] C. Dellacherie. Correction à ‘Ensembles aléatoires II’. (French) [correction to ‘Random ensembles II”]. Lecture Notes in Mathematics, 191:86, 1971. CODEN LNMAA2. ISBN 3-540-05397-2 (print), 3-540-36517-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058850. See [Del69b].

Gundy:1969:IBT

[dGM69] d’après Gundy and P.-A. Meyer. Les inégalités de Burkolder en théorie des martingales. (French) [Burkolder inequalities in martingale theory]. Lecture Notes in Mathematics, 88:163–174, 1969. CODEN LNMAA2. ISBN 3-540-04607-0 (print), 3-540-36097-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0081636/.

Douglas:1969:S

[DHS69] R. R. Douglas, P. J. Hilton, and F. Sigrist. H-spaces. Lecture Notes in Mathematics, 92:65–73, 1969. CODEN LNMAA2. ISBN 3-540-04611-9 (print), 3-540-36101-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080765/.

Dinges:1967:FDG

[Din67] Hermann Dinges. Faktorisierung von Differentialoperatoren. (German) [Factorization of differential operators]. Lecture Notes in Mathematics, 31:62, 1967. CODEN LNMAA2. ISBN 3-540-03902-3 (print), 3-540-34970-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061108.

DJakonov:1969:CIM

[D’J69] E. G. D’Jakonov. On certain iterative methods for solving nonlinear difference equations. Lecture Notes in Mathematics, 109: 7–22, 1969. CODEN LNMAA2. ISBN 3-540-04628-3 (print), 3-540-36158-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060014/.

Day:1969:EFC

[DK69] B. J. Day and G. M. Kelly. Enriched functor categories. Lecture Notes in Mathematics, 106:178–191, 1969. CODEN LN-MAA2. ISBN 3-540-04625-9 (print), 3-540-36150-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0059146/.

Dold:1967:PCA

[DMO67] A. Dold, S. Mac Lane, and U. Oberst. Projective classes and acyclic models. Lecture Notes in Mathematics, 47:78–91, 1967. CODEN LNMAA2. ISBN 3-540-03918-X (print), 3-540-35545-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074300/.

Dold:1966:AHF

[Dol66a] Albrecht Dold. Anwendung: Halbexakte Funktoren t mit Monoidstruktur I. (German) [Application: Semi-exact functors t with monoid structure I]. Lecture Notes in Mathematics, 12:8.1– 8.12, 1966. CODEN LNMAA2. ISBN 3-540-03595-8 (print), 3-540-37166-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084865.

Dold:1966:AHG

[Dol66b] Albrecht Dold. Anwendungen; Hindernistheorie. (German) [Applications; obstruction theory]. Lecture Notes in Mathematics, 12:13.1–13.9, 1966. CODEN LNMAA2. ISBN 3-540-03595-8 (print), 3-540-37166-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084870.

Dold:1966:BHF

[Dol66c] Albrecht Dold. Beispiele von halbexakten Funktoren. (German) [Examples of semi-exact functors]. Lecture Notes in Mathematics, 12:6.1–6.7, 1966. CODEN LNMAA2. ISBN 3-540-03595-8 (print), 3-540-37166-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084863.

Dold:1966:DFG

[Dol66d] Albrecht Dold. Darstellbare Funktoren. (German) [Displayable functors]. Lecture Notes in Mathematics, 12:1.1–1.9, 1966. CO-DEN LNMAA2. ISBN 3-540-03595-8 (print), 3-540-37166-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0084858.

Dold:1966:DHF

[Dol66e] Albrecht Dold. Darstellbarkeit halbexakter Funktoren. (German) [Displayability of half-exact functors]. Lecture Notes in Mathematics, 12:16.1–16.8, 1966. CODEN LNMAA2. ISBN 3-540-03595-8 (print), 3-540-37166-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084873.

Dold:1966:HHGa

[Dol66f] Albrecht Dold. Die Homotopieerweiterungseigenschaft (= HEE). (German) [The homotopy extension property (= HEP)]. Lecture Notes in Mathematics, 12:3.1–3.9, 1966. CODEN LNMAA2. ISBN 3-540-03595-8 (print), 3-540-37166-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084860.

Dold:1966:RGG

[Dol66g] Albrecht Dold. Die Rolle des Grundpunktes. (German) [The role of the base point]. Lecture Notes in Mathematics, 12:4.1–4.7, 1966. CODEN LNMAA2. ISBN 3-540-03595-8 (print), 3-540-37166-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0084861.

Dold:1966:SHM

[Dol66h] Albrecht Dold. Die Spektralsequenz halbexakter Monoidfunktoren. (German) [The spectral sequence of semi-exact monoid functors]. Lecture Notes in Mathematics, 12:14.1–14.8, 1966. CODEN LN-MAA2. ISBN 3-540-03595-8 (print), 3-540-37166-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0084871.

Dold:1966:EGI

[Dol66i] Albrecht Dold. Einleitung. (German) [Introduction]. Lecture Notes in Mathematics, 12:0.1–0.3, 1966. CODEN LNMAA2. ISBN 3-540-03595-8 (print), 3-540-37166-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084857.

Dold:1966:FMIa

[Dol66j] Albrecht Dold. Funktoren mit Monoidstruktur II. (German) [Functors with monoid structure II]. Lecture Notes in Mathematics, 12:9.1–9.6, 1966. CODEN LNMAA2. ISBN 3-540-03595-8 (print), 3-540-37166-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084866.

Dold:1966:FMIb

[Dol66k] Albrecht Dold. Funktoren mit Monoidstruktur (III) und rationalen Koeffizienten. (German) [Functors with monoid structure III and rational coefficients]. Lecture Notes in Mathematics, 12:10.1– 10.11, 1966. CODEN LNMAA2. ISBN 3-540-03595-8 (print), 3-540-37166-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084867.

Dold:1966:HHGb

[Dol66l] Albrecht Dold. Halbexakte Homotopiefunktoren. (German) [Halfexact homotopy functors]. Lecture Notes in Mathematics, 12:5.1– 5.11, 1966. CODEN LNMAA2. ISBN 3-540-03595-8 (print), 3-540-37166-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084862.

Dold:1966:HHGc

[Dol66m] Albrecht Dold. Halbexakte Homotopiefunktoren. (German) [Halfexact homotopy functors], volume 12 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1966. CODEN LNMAA2. ISBN 3-540-03595-8 (paperback), 3-540-37166-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 156 pp. LCCN QA3 .L28 no. 12. URL

http://link.springer.com/book/10.1007/BFb0084856.

Dold:1966:MSG

[Dol66n] Albrecht Dold. Multiplikative Strukturen. (German) [Multiplicative structures]. Lecture Notes in Mathematics, 12:2.1–2.7, 1966. CODEN LNMAA2. ISBN 3-540-03595-8 (print), 3-540-37166-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0084859.

Dold:1966:PFG

[Dol66o] Albrecht Dold. Postnikov-Faktoren. (German) [Postnikov factors]. Lecture Notes in Mathematics, 12:11.1–11.8, 1966. CODEN LN-MAA2. ISBN 3-540-03595-8 (print), 3-540-37166-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0084868.

Dold:1966:PIG

[Dol66p] Albrecht Dold. Postnikov-Invarianten. (German) [Postnikov invariants]. Lecture Notes in Mathematics, 12:12.1–12.11, 1966. CO-DEN LNMAA2. ISBN 3-540-03595-8 (print), 3-540-37166-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0084869.

Dold:1966:PSG

[Dol66q] Albrecht Dold. Produkte in der Spektralsequenz. (German) [Products in the spectral sequence]. Lecture Notes in Mathematics, 12:15.1–15.10, 1966. CODEN LNMAA2. ISBN 3-540-03595-8 (print), 3-540-37166-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084872.

Dold:1966:VHF

[Dol66r] Albrecht Dold. Vergleich halbexakter Funktoren. (German) [Comparison of half-exact functors]. Lecture Notes in Mathematics, 12: 7.1–7.3, 1966. CODEN LNMAA2. ISBN 3-540-03595-8 (print), 3-540-37166-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084864.

Doleans:1968:EVA

[Dol68a] Catherine Doléans. Espaces H^{m} sur les variétés, et applications aux équations aux derivées partielles sur une variété compacte. (French) [H^{m} spaces on manifolds and applications to partial differential equations on a compact variety]. Lecture Notes in Mathematics, 51:43–74, 1968. CODEN LNMAA2. ISBN 3-540-04221-0 (print), 3-540-35827-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059179/.

Doleans:1968:FAP

[Dol68b] Catherine Doleans. Fonctionnelles additives parfaites. (French) [Perfect functional additives]. Lecture Notes in Mathematics, 51: 34–42, 1968. CODEN LNMAA2. ISBN 3-540-04221-0 (print), 3-540-35827-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059178/.

deRham:1967:PSG

[dR67a] G. de Rham. Propriètés des sommes de Gauss et des séries de Dirichlet théorème de Franz. (French) [Properties of Gauss sums and Dirichlet series [by the] theorem of Franz]. Lecture Notes in Mathematics, 48:1–12, 1967. CODEN LNMAA2. ISBN 3-540-03919-8 (print), 3-540-35547-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073929/.

deRham:1967:TDP

[dR67b] G. de Rham. théorème de dualité pour la torsion et applications aux noeuds. (French) [Duality theorem for torsion and applications to knots]. Lecture Notes in Mathematics, 48:74–82, 1967. CODEN LNMAA2. ISBN 3-540-03919-8 (print), 3-540-35547-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073934/.

deRham:1967:TCA

[dR67c] G. de Rham. Torsion d’un complexe à automorphismes. (French) [Torsion of a complex of automorphisms]. Lecture Notes in Mathematics, 48:13–36, 1967. CODEN LNMAA2. ISBN 3-540-03919-8 (print), 3-540-35547-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073930/.

deRham:1967:THR

[dR67d] G. de Rham. Type d’homotopie des rotations et des espaces lenticulaires. (French) [Homotopy type of rotations and lens spaces]. Lecture Notes in Mathematics, 48:96–101, 1967. CODEN LN-MAA2. ISBN 3-540-03919-8 (print), 3-540-35547-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0073936/.

deRham:1967:TTS

[dRMK67] G. de Rham, S. Maumary, and M. A. Kervaire, editors. Torsion et Type Simple d’Homotopie: Exposés faits au Séminaire de Topologie de l’Université de Lausanne. (French) [Twist and Simple Type Homotopy: Presentations made at the Topology Seminar of the University of Lausanne], volume 48 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. CODEN LNMAA2. ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 48. URL http://

link.springer.com/book/10.1007/BFb0073928.

Dubuc:1968:AT

[Dub68] Eduardo Dubuc. Adjoint triangles. Lecture Notes in Mathematics, 61:69–91, 1968. CODEN LNMAA2. ISBN 3-540-04231-8 (print), 3-540-35863-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077118/.

Dunkel:1968:SSM

[Dun68] Gregory Dunkel. Single species model for population growth depending on past history. Lecture Notes in Mathematics, 60:92– 99, 1968. CODEN LNMAA2. ISBN 3-540-04230-X (print), 3-540-35862-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077132/.

Duskin:1969:VBT

[Dus69] J. Duskin. Variations on Beck’s tripleability criterion. Lecture Notes in Mathematics, 106:74–129, 1969. CODEN LNMAA2. ISBN 3-540-04625-9 (print), 3-540-36150-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059143/.

Eckmann:1969:STC

[Eck69] B. Eckmann, editor. Seminar on Triples and Categorical Homology Theory: ETH 1966/67, volume 80 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CODEN LNMAA2. ISBN 3-540-04601-1 (paperback), 3-540-36091-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 80. URL http://

link.springer.com/book/10.1007/BFb0083078.

Earle:1969:DRS

[EE69] Clifford J. Earle and James Eells. Deformations of Riemann surfaces. Lecture Notes in Mathematics, 103:122–149, 1969. CODEN LNMAA2. ISBN 3-540-04622-4 (print), 3-540-36145-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099988/.

Eckmann:1969:IFI

[EG69a] J.-P. Eckmann and M. Guenin. Introduction. (French) [Introduction]. Lecture Notes in Mathematics, 81:1, 1969. CODEN LN-MAA2. ISBN 3-540-04249-0 (print), 3-540-35927-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0075598.

Eckmann:1969:IMF

[EG69b] J.-P. Eckmann and M. Guenin. Introduction mathématique. (French) [Mathematical introduction]. Lecture Notes in Mathematics, 81:2–44, 1969. CODEN LNMAA2. ISBN 3-540-04249-0 (print), 3-540-35927-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075599/.

Eckmann:1969:MAM

[EG69c] J.-P. Eckmann and M. Guenin. Méthodes Algébriques en Mécanique Statistique. (French) [Algebraic Methods in Statistical Mechanics], volume 81 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CODEN LNMAA2. ISBN 3-540-04249-0 (paperback), 3-540-35927-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 131 pp. LCCN QA3 .L28 no. 81. URL http://link.springer.

com/book/10.1007/BFb0075597.

Eckmann:1969:MFM

[EG69d] J.-P. Eckmann and M. Guenin. Modèles. (French) [Models]. Lecture Notes in Mathematics, 81:85–117, 1969. CODEN LNMAA2. ISBN 3-540-04249-0 (print), 3-540-35927-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075601/.

Eckmann:1969:SFM

[EG69e] J.-P. Eckmann and M. Guenin. Structures fondamentales de la mècanique statistique. (French) [Fundamental structures of statistical mechanics]. Lecture Notes in Mathematics, 81:45–84, 1969. CODEN LNMAA2. ISBN 3-540-04249-0 (print), 3-540-35927-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075600/.

Ehresmann:1969:CSL

[Ehr69] Charles Ehresmann. Construction de structures libres. (French) [Construction of free structures]. Lecture Notes in Mathematics, 92:74–104, 1969. CODEN LNMAA2. ISBN 3-540-04611-9 (print), 3-540-36101-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080766/.

Epstein:1969:FBC

[EK69] D. B. A. Epstein and M. Kneser. Functors between categories of vector spaces. Lecture Notes in Mathematics, 99:154–170, 1969. CODEN LNMAA2. ISBN 3-540-04618-6 (print), 3-540-36140-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0081962/.

Engeler:1968:RTG

[Eng68] Erwin Engeler. Remarks on the theory of geometrical constructions. Lecture Notes in Mathematics, 72:64–76, 1968. CODEN LNMAA2. ISBN 3-540-04242-3 (print), 3-540-35900-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079682/.

Epstein:1969:NVB

[Eps69] D. B. A. Epstein. Natural vector bundles. Lecture Notes in Mathematics, 99:171–195, 1969. CODEN LNMAA2. ISBN 3-540-04618-6 (print), 3-540-36140-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081963/.

Erdos:1969:SAG

[Erd69] Paul Erdős. Some applications of graph theory to number theory. Lecture Notes in Mathematics, 110:77–82, 1969. CODEN LNMAA2. ISBN 3-540-04629-1 (print), 3-540-36161-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060107/.

Exbrayat:1968:CEF

[Exb68] Madame F. Exbrayat. Sur certains espaces de fonctions continues associés á des poids. (French) [On some spaces of continuous functions associated with weight]. Lecture Notes in Mathematics, 71:46–56, 1968. CODEN LNMAA2. ISBN 3-540-04241-5 (print), 3-540-35889-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080096/.

Faith:1967:ERQ

[Fai67a] Carl Faith. The endomorphism ring of a quasi-injective module. Lecture Notes in Mathematics, 49:44–50, 1967. CODEN LNMAA2. ISBN 3-540-03920-1 (print), 3-540-35551-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074324/.

Faith:1967:EEI

[Fai67b] Carl Faith. Essential extensions and the injective hull. Lecture Notes in Mathematics, 49:13–21, 1967. CODEN LNMAA2. ISBN 3-540-03920-1 (print), 3-540-35551-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074321/.

Faith:1967:IM

[Fai67c] Carl Faith. Injective modules. Lecture Notes in Mathematics, 49: 1–12, 1967. CODEN LNMAA2. ISBN 3-540-03920-1 (print), 3-540-35551-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074320/.

Faith:1967:JR

[Fai67d] Carl Faith. Johnson rings. Lecture Notes in Mathematics, 49: 123–126, 1967. CODEN LNMAA2. ISBN 3-540-03920-1 (print), 3-540-35551-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074333/.

Faith:1967:LIM

[Fai67e] Carl Faith. Lectures on Injective Modules and Quotient Rings, volume 49 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. CO-DEN LNMAA2. ISBN 3-540-03920-1 (paperback), 3-540-35551-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xv + 140

pp. LCCN QA3 .L28 no. 49. URL http://link.springer.com/
book/10.1007/BFb0074319.

Faith:1967:MQRa

[Fai67f] Carl Faith. Maximal quotient rings. Lecture Notes in Mathematics, 49:64–75, 1967. CODEN LNMAA2. ISBN 3-540-03920-1 (print), 3-540-35551-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074327/.

Faith:1967:MQRb

[Fai67g] Carl Faith. Maximal quotient rings. Lecture Notes in Mathematics, 49:96–104, 1967. CODEN LNMAA2. ISBN 3-540-03920-1 (print), 3-540-35551-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074331/.

Faith:1967:NSI

[Fai67h] Carl Faith. Nil and singular ideals under maximum conditions. Lecture Notes in Mathematics, 49:82–85, 1967. CODEN LN-MAA2. ISBN 3-540-03920-1 (print), 3-540-35551-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074329/.

Faith:1967:NAS

[Fai67i] Carl Faith. Noetherian, Artinian, and semisimple modules and rings. Lecture Notes in Mathematics, 49:51–57, 1967. CODEN LNMAA2. ISBN 3-540-03920-1 (print), 3-540-35551-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074325/.

Faith:1967:OP

[Fai67j] Carl Faith. Open problems. Lecture Notes in Mathematics, 49: 127–131, 1967. CODEN LNMAA2. ISBN 3-540-03920-1 (print), 3-540-35551-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074334/.

Faith:1967:QIM

[Fai67k] Carl Faith. Quasi-injective modules. Lecture Notes in Mathematics, 49:22–34, 1967. CODEN LNMAA2. ISBN 3-540-03920-1 (print), 3-540-35551-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074322/.

Faith:1967:QRD

[Fai67l] Carl Faith. Quotient rings and direct products of full linear rings. Lecture Notes in Mathematics, 49:105–122, 1967. CODEN LN-MAA2. ISBN 3-540-03920-1 (print), 3-540-35551-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074332/.

Faith:1967:RSR

[Fai67m] Carl Faith. Radical and semiprimitivity in rings. Lecture Notes in Mathematics, 49:35–43, 1967. CODEN LNMAA2. ISBN 3-540-03920-1 (print), 3-540-35551-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074323/.

Faith:1967:REL

[Fai67n] Carl Faith. Rational extensions and lattices of closed submodules. Lecture Notes in Mathematics, 49:58–63, 1967. CODEN LNMAA2. ISBN 3-540-03920-1 (print), 3-540-35551-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074326/.

Faith:1967:SRM

[Fai67o] Carl Faith. Semiprime rings with maximum condition. Lecture Notes in Mathematics, 49:76–81, 1967. CODEN LNMAA2. ISBN 3-540-03920-1 (print), 3-540-35551-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074328/.

Faith:1967:SNP

[Fai67p] Carl Faith. Structure of Noetherian prime rings. Lecture Notes in Mathematics, 49:86–95, 1967. CODEN LNMAA2. ISBN 3-540-03920-1 (print), 3-540-35551-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074330/.

Frolicher:1966:M

[FB66a] A. Frölicher and W. Bucher. C_{k}-mappings. Lecture Notes in Mathematics, 30:99–109, 1966. CODEN LNMAA2. ISBN 3-540-03612-1 (print), 3-540-34862-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068442/.

Frolicher:1966:CVS

[FB66b] A. Frölicher and W. Bucher. Calculus in Vector Spaces without Norm, volume 30 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1966. CODEN LNMAA2. ISSN 0075-8434 (print), 1617-9692 (electronic). x + 146 pp. LCCN QA3 .L28 no. 30. URL http://

link.springer.com/book/10.1007/BFb0068432.

Frolicher:1966:CAV

[FB66c] A. Frölicher and W. Bucher. The class of admissible vector spaces. Lecture Notes in Mathematics, 30:82–89, 1966. CODEN LN-MAA2. ISBN 3-540-03612-1 (print), 3-540-34862-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068439/.

Frolicher:1966:CM

[FB66d] A. Frölicher and W. Bucher. The composition of C_{k}-mappings. Lecture Notes in Mathematics, 30:110–130, 1966. CODEN LN-MAA2. ISBN 3-540-03612-1 (print), 3-540-34862-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068443/.

Frolicher:1966:DD

[FB66e] A. Frölicher and W. Bucher. Differentiability and derivatives. Lecture Notes in Mathematics, 30:32–41, 1966. CODEN LN-MAA2. ISBN 3-540-03612-1 (print), 3-540-34862-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068435/.

Frolicher:1966:DDD

[FB66f] A. Frölicher and W. Bucher. Differentiable deformation of differentiable mappings. Lecture Notes in Mathematics, 30:131–136, 1966. CODEN LNMAA2. ISBN 3-540-03612-1 (print), 3-540-34862-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068444/.

Frolicher:1966:EPF

[FB66g] A. Frölicher and W. Bucher. Elementary properties of filters. Lecture Notes in Mathematics, 30:1–5, 1966. CODEN LNMAA2. ISBN 3-540-03612-1 (print), 3-540-34862-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068433/.

Frolicher:1966:ESC

[FB66h] A. Frölicher and W. Bucher. Examples and special cases. Lecture Notes in Mathematics, 30:42–49, 1966. CODEN LNMAA2. ISBN 3-540-03612-1 (print), 3-540-34862-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068436/.

Frolicher:1966:FTC

[FB66i] A. Frölicher and W. Bucher. Fundamental theorem of calculus. Lecture Notes in Mathematics, 30:50–64, 1966. CODEN LNMAA2. ISBN 3-540-03612-1 (print), 3-540-34862-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068437/.

Frolicher:1966:HD

[FB66j] A. Frölicher and W. Bucher. Higher derivatives. Lecture Notes in Mathematics, 30:93–98, 1966. CODEN LNMAA2. ISBN 3-540-03612-1 (print), 3-540-34862-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068441/.

Frolicher:1966:PDD

[FB66k] A. Frölicher and W. Bucher. Partial derivatives and differentiability. Lecture Notes in Mathematics, 30:90–92, 1966. CODEN LNMAA2. ISBN 3-540-03612-1 (print), 3-540-34862-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068440/.

Frolicher:1966:PTV

[FB66l] A. Frölicher and W. Bucher. Pseudo-topological vector spaces. Lecture Notes in Mathematics, 30:6–31, 1966. CODEN LN-MAA2. ISBN 3-540-03612-1 (print), 3-540-34862-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068434/.

Frolicher:1966:PTS

[FB66m] A. Frölicher and W. Bucher. Pseudo-topologies on some function spaces. Lecture Notes in Mathematics, 30:65–81, 1966. CODEN LNMAA2. ISBN 3-540-03612-1 (print), 3-540-34862-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068438/.

Feferman:1968:LPT

[Fef68] Solomon Feferman. Lectures on proof theory. Lecture Notes in Mathematics, 70:1–107, 1968. CODEN LNMAA2. ISBN 3-540-04240-7 (print), 3-540-35886-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079094/.

Fenstad:1969:NSM

[Fen69] Jens Erik Fenstad. Non-standard models for arithmetic and analysis. Lecture Notes in Mathematics, 118:30–47, 1969. CODEN LNMAA2. ISBN 3-540-04907-X (print), 3-540-36246-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060250/.

Fernique:1967:SDAa

[Fer67a] X. Fernique. Séries de distributions aléatoires indépendantes. (French) [Series of independent random distributions]. Lecture Notes in Mathematics, 39:54–64, 1967. CODEN LNMAA2. ISBN 3-540-03910-4 (print), 3-540-34978-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080546/.

Fernique:1967:SDAb

[Fer67b] X. Fernique. Séries de distributions aléatoires indépendantes. (French) [Series of independent random distributions]. Lecture Notes in Mathematics, 39:65–71, 1967. CODEN LNMAA2. ISBN 3-540-03910-4 (print), 3-540-34978-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080547/.

Ferus:1968:BZGb

[Fer68a] Dirk Ferus. Beziehungen zwischen τ(f: M → R^{n}) und der Geometrie von f(M) in R^{n}. (German) [Relationships between τ(f: M → R^{n}) and the geometry of f(M) in R^{n}]. Lecture Notes in Mathematics, 66:40–58, 1968. CODEN LNMAA2. ISBN 3-540-04236-9 (print), 3-540-35877-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076985/.

Ferus:1968:BZGa

[Fer68b] Dirk Ferus. Beziehungen zwischen τ(f: M → R^{n}) und der Geometrie von M. (German) [Relationships between τ(f: M → R^{n}) and the geometry of M]. Lecture Notes in Mathematics, 66:38–39, 1968. CODEN LNMAA2. ISBN 3-540-04236-9 (print), 3-540-35877-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0076984.

Ferus:1968:TAG

[Fer68c] Dirk Ferus. Die totale Absolutkrümmung. (German) [Total absolute curvature]. Lecture Notes in Mathematics, 66:9–22, 1968. CODEN LNMAA2. ISBN 3-540-04236-9 (print), 3-540-35877-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076982/.

Ferus:1968:IER

[Fer68d] Dirk Ferus. Immersionen in den euklidischen Raum. (German) [Immersion in Euclidean space]. Lecture Notes in Mathematics, 66:23–37, 1968. CODEN LNMAA2. ISBN 3-540-04236-9 (print), 3-540-35877-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076983/.

Ferus:1968:KHG

[Fer68e] Dirk Ferus. Kompakte Hyperflächen. (German) [Compact hypersurfaces]. Lecture Notes in Mathematics, 66:59–70, 1968. CODEN LNMAA2. ISBN 3-540-04236-9 (print), 3-540-35877-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076986/.

Ferus:1968:TAD

[Fer68f] Dirk Ferus. Totale absolutkrümmung in Differentialgeometrie undtopologie. (German) [Total absolute curvature in differential geometry and topology], volume 66 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. CODEN LNMAA2. ISBN 3-540-04236-9 (paperback), 3-540-35877-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 85 pp. LCCN QA3 .L28 no. 66. URL

http://link.springer.com/book/10.1007/BFb0076980.

Ferus:1968:DGB

[Fer68g] Dirk Ferus. Zur Differentialgeometrie gewisser Bündel. (German) [On the differential geometry of certain sheaves]. Lecture Notes in Mathematics, 66:1–8, 1968. CODEN LNMAA2. ISBN 3-540-04236-9 (print), 3-540-35877-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076981/.

Fichera:1965:CEBa

[Fic65a] Gaetano Fichera. The classical elliptic BVP of mathematical physics: 2nd order linear PDE. Lecture Notes in Mathematics, 8:80–87, 1965. CODEN LNMAA2. ISBN 3-540-03351-3 (print), 3-540-37134-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079970/.

Fichera:1965:CEBc

[Fic65b] Gaetano Fichera. The classical elliptic BVP of mathematical physics: Equilibrium of thin plates. Lecture Notes in Mathematics, 8:95–100, 1965. CODEN LNMAA2. ISBN 3-540-03351-3 (print), 3-540-37134-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079972/.

Fichera:1965:CEBb

[Fic65c] Gaetano Fichera. The classical elliptic BVP of mathematical physics: Linear elastostatics. Lecture Notes in Mathematics, 8: 88–94, 1965. CODEN LNMAA2. ISBN 3-540-03351-3 (print), 3-540-37134-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079971/.

Fichera:1965:CIO

[Fic65d] Gaetano Fichera. Construction of the intermediate operators. Lecture Notes in Mathematics, 8:130–138, 1965. CODEN LN-MAA2. ISBN 3-540-03351-3 (print), 3-540-37134-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079976/.

Fichera:1965:EPR

[Fic65e] Gaetano Fichera. Eigenvalue problems. The Rayleigh–Ritz method. Lecture Notes in Mathematics, 8:112–119, 1965. CO-DEN LNMAA2. ISBN 3-540-03351-3 (print), 3-540-37134-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0079974/.

Fichera:1965:ELSa

[Fic65f] Gaetano Fichera. Elliptic linear systems. Interior regularity. Lecture Notes in Mathematics, 8:30–38, 1965. CODEN LN-MAA2. ISBN 3-540-03351-3 (print), 3-540-37134-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079964/.

Fichera:1965:ELSb

[Fic65g] Gaetano Fichera. Existence of local solutions for elliptic systems. Lecture Notes in Mathematics, 8:39–43, 1965. CODEN LNMAA2. ISBN 3-540-03351-3 (print), 3-540-37134-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079965/.

Fichera:1965:EP

[Fic65h] Gaetano Fichera. Existence principle. Lecture Notes in Mathematics, 8:11–16, 1965. CODEN LNMAA2. ISBN 3-540-03351-3 (print), 3-540-37134-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079961/.

Fichera:1965:ECG

[Fic65i] Gaetano Fichera. Explicit construction of the Green’s matrix for an elliptic system. Lecture Notes in Mathematics, 8:164–174, 1965. CODEN LNMAA2. ISBN 3-540-03351-3 (print), 3-540-37134-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0079979/.

Fichera:1965:FSH

[Fic65j] Gaetano Fichera. The function spaces and H_{m}. Lecture Notes in Mathematics, 8:17–23, 1965. CODEN LNMAA2. ISBN 3-540-03351-3 (print), 3-540-37134-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079962/.

Fichera:1965:LED

[Fic65k] Gaetano Fichera, editor. Linear elliptic differential systems and eigenvalue problems: The Johns Hopkins University, Baltimore Md, March–May 1965, volume 8 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1965. CODEN LNMAA2. ISBN 3-540-03351-3 (paperback), 3-540-37134-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 8. URL http://

link.springer.com/book/10.1007/BFb0079959.

Fichera:1965:OIP

[Fic65l] Gaetano Fichera. Orthogonal invariants of positive compact operators. Lecture Notes in Mathematics, 8:139–151, 1965. CODEN LNMAA2. ISBN 3-540-03351-3 (print), 3-540-37134-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079977/.

Fichera:1965:RBF

[Fic65m] Gaetano Fichera. Regularity at the boundary: final results. Lecture Notes in Mathematics, 8:69–79, 1965. CODEN LNMAA2. ISBN 3-540-03351-3 (print), 3-540-37134-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079969/.

Fichera:1965:RBP

[Fic65n] Gaetano Fichera. Regularity at the boundary: preliminary lemmas. Lecture Notes in Mathematics, 8:52–60, 1965. CODEN LN-MAA2. ISBN 3-540-03351-3 (print), 3-540-37134-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079967/.

Fichera:1965:RBT

[Fic65o] Gaetano Fichera. Regularity at the boundary: tangential derivatives. Lecture Notes in Mathematics, 8:61–68, 1965. CODEN LNMAA2. ISBN 3-540-03351-3 (print), 3-540-37134-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079968/.

Fichera:1965:SSB

[Fic65p] Gaetano Fichera. Semiweak solutions of BVP for elliptic systems. Lecture Notes in Mathematics, 8:44–51, 1965. CODEN LNMAA2. ISBN 3-540-03351-3 (print), 3-540-37134-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079966/.

Fichera:1965:SEO

[Fic65q] Gaetano Fichera. Strongly elliptic operators. Gäarding inequality. Eigenvalue problems. Lecture Notes in Mathematics, 8:101– 111, 1965. CODEN LNMAA2. ISBN 3-540-03351-3 (print), 3-540-37134-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079973/.

Fichera:1965:TOS

[Fic65r] Gaetano Fichera. The trace operator. Sobolev and Ehrling lemmas. Lecture Notes in Mathematics, 8:24–29, 1965. CODEN LN-MAA2. ISBN 3-540-03351-3 (print), 3-540-37134-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079963/.

Fichera:1965:UAE

[Fic65s] Gaetano Fichera. Upper approximation of the eigenvalues of a PCO. Representation of orthogonal invariants. Lecture Notes in Mathematics, 8:152–163, 1965. CODEN LNMAA2. ISBN 3-540-03351-3 (print), 3-540-37134-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079978/.

Fichera:1965:WAM

[Fic65t] Gaetano Fichera. The Weinstein–Aronszajn method. Lecture Notes in Mathematics, 8:120–129, 1965. CODEN LNMAA2. ISBN 3-540-03351-3 (print), 3-540-37134-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079975/.

Fichera:1965:WPB

[Fic65u] Gaetano Fichera. “well posed” boundary value problems. Lecture Notes in Mathematics, 8:1–10, 1965. CODEN LNMAA2. ISBN 3-540-03351-3 (print), 3-540-37134-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079960/.

Fieger:1967:ANL

[Fie67] Werner Fieger. Die Anzahl der Niveaudurchgänge und der lokalen Maximalstellen von Gaußschen Prozessen. (German) [The number of level crossings and the local maximum points of Gaussian processes]. Lecture Notes in Mathematics, 31:63–68, 1967. CODEN LNMAA2. ISBN 3-540-03902-3 (print), 3-540-34970-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061109/.

Fittler:1969:CMI

[Fit69] Dr. R. Fittler. Categories of models with initial objects. Lecture Notes in Mathematics, 86:33–45, 1969. CODEN LNMAA2. ISBN 3-540-04605-4 (print), 3-540-36095-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079383/.

Friedman:1968:NAO

[FJ68] Harvey Friedman and Ronald Jensen. Note on admissible ordinals. Lecture Notes in Mathematics, 72:77–79, 1968. CODEN LNMAA2. ISBN 3-540-04242-3 (print), 3-540-35900-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079683/.

Feferman:1969:STF

[FK69] Solomon Feferman and G. Kreisel. Set-theoretical foundations of category theory. Lecture Notes in Mathematics, 106:201–247, 1969. CODEN LNMAA2. ISBN 3-540-04625-9 (print), 3-540-36150-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0059148/.

Frechen:1969:NCP

[Fre69a] Joseph B. Frechen. On the number of cycles in permutation graphs. Lecture Notes in Mathematics, 110:83–87, 1969. CODEN LN-MAA2. ISBN 3-540-04629-1 (print), 3-540-36161-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0060108/.

Freyd:1969:SNC

[Fre69b] Peter Freyd. Several new concepts: Lucid and concordant functors, pre-limits, pre-completeness, the continuous and concordant completions of categories. Lecture Notes in Mathematics, 99:196– 241, 1969. CODEN LNMAA2. ISBN 3-540-04618-6 (print), 3-540-36140-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081964/.

Frohlich:1968:CFGa

[Frö68a] A. Fröhlich. Commutative formal groups of dimension one. Lecture Notes in Mathematics, 74:57–95, 1968. CODEN LNMAA2. ISBN 3-540-04244-X (print), 3-540-35916-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074376/.

Frohlich:1968:CFGb

[Frö68b] A. Fröhlich. Commutative formal groups of dimension one over a discrete valuation ring. Lecture Notes in Mathematics, 74:96– 138, 1968. CODEN LNMAA2. ISBN 3-540-04244-X (print), 3-540-35916-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074377/.

Frohlich:1968:FG

[Frö68c] A. Fröhlich. Formal Groups, volume 74 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. CODEN LNMAA2. ISBN 3-540-04244-X (paperback), 3-540-35916-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 140 pp. LCCN QA3 .L28 no. 74. URL

http://link.springer.com/book/10.1007/BFb0074373.

Frohlich:1968:LT

[Frö68d] A. Fröhlich. Lie theory. Lecture Notes in Mathematics, 74:29– 56, 1968. CODEN LNMAA2. ISBN 3-540-04244-X (print), 3-540-35916-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074375/.

Frohlich:1968:P

[Frö68e] A. Fröhlich. Preliminaries. Lecture Notes in Mathematics, 74: 1–28, 1968. CODEN LNMAA2. ISBN 3-540-04244-X (print), 3-540-35916-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074374/.

Fujiie:1968:ELK

[Fuj68] Tatsuo Fuji’i’e. Extremal length and Kuramochi boundary of a subregion of a Riemann surface. Lecture Notes in Mathematics, 58:97–102, 1968. CODEN LNMAA2. ISBN 3-540-04228-8 (print), 3-540-35859-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098448/.

Floret:1968:BAB

[FW68a] Klaus Floret and Joseph Wloka. Beschränkte Abbildungen und bornologische Räume. (German) [Restricted maps and bornological spaces]. Lecture Notes in Mathematics, 56:54–57, 1968. CO-DEN LNMAA2. ISBN 3-540-04226-1 (print), 3-540-35855-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0098560/.

Floret:1968:BMG

[FW68b] Klaus Floret and Joseph Wloka. Beschränkte Mengen. (German) [Limited measures]. Lecture Notes in Mathematics, 56:43–46, 1968. CODEN LNMAA2. ISBN 3-540-04226-1 (print), 3-540-35855-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0098557/.

Floret:1968:DGD

[FW68c] Klaus Floret and Joseph Wloka. Der Dualraum. (German) [The dual space]. Lecture Notes in Mathematics, 56:57–61, 1968. CO-DEN LNMAA2. ISBN 3-540-04226-1 (print), 3-540-35855-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0098561/.

Floret:1968:PLG

[FW68d] Klaus Floret and Joseph Wloka. Der projektive Limes. (German) [The projective limit]. Lecture Notes in Mathematics, 56: 34–39, 1968. CODEN LNMAA2. ISBN 3-540-04226-1 (print), 3-540-35855-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098555/.

Floret:1968:SBG

[FW68e] Klaus Floret and Joseph Wloka. Der Satz von Baire. (German) [The Baire set]. Lecture Notes in Mathematics, 56:6–10, 1968. CODEN LNMAA2. ISBN 3-540-04226-1 (print), 3-540-35855-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098551/.

Floret:1968:SMG

[FW68f] Klaus Floret and Joseph Wloka. Der Satz von Mackey. (German) [The Mackey theorem]. Lecture Notes in Mathematics, 56: 72–74, 1968. CODEN LNMAA2. ISBN 3-540-04226-1 (print), 3-540-35855-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098564/.

Floret:1968:STB

[FW68g] Klaus Floret and Joseph Wloka. Die schwache Topologie und der Bipolarensatz. (German) [Weak topology and the bipolar theorem]. Lecture Notes in Mathematics, 56:66–72, 1968. CODEN LNMAA2. ISBN 3-540-04226-1 (print), 3-540-35855-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098563/.

Floret:1968:STG

[FW68h] Klaus Floret and Joseph Wloka. Die starke Topologie. (German) [Strong topology]. Lecture Notes in Mathematics, 56:61–66, 1968. CODEN LNMAA2. ISBN 3-540-04226-1 (print), 3-540-35855-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0098562/.

Floret:1968:DAF

[FW68i] Klaus Floret and Joseph Wloka. Duale Abbildungen. (French) [Dual maps]. Lecture Notes in Mathematics, 56:81–85, 1968. CO-DEN LNMAA2. ISBN 3-540-04226-1 (print), 3-540-35855-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0098567/.

Floret:1968:ETL

[FW68j] Klaus Floret and Joseph Wloka. Einführung in die Theorie der lokalkonvexen Räume. (German) [Introduction to the theory of locally convex spaces], volume 56 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. CODEN LNMAA2. ISBN 3-540-04226-1 (paperback), 3-540-35855-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 194 pp. LCCN QA3 .L28 no. 56.

URL http://link.springer.com/book/10.1007/BFb0098549.

Floret:1968:GGG

[FW68k] Klaus Floret and Joseph Wloka. Gelfandräume. (German) [Gelfand spaces]. Lecture Notes in Mathematics, 56:46–49, 1968. CODEN LNMAA2. ISBN 3-540-04226-1 (print), 3-540-35855-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098558/.

Floret:1968:HSA

[FW68l] Klaus Floret and Joseph Wloka. Hilbert–Schmidt-Abbildungen. (German) [Hilbert–Schmidt maps]. Lecture Notes in Mathematics, 56:93–99, 1968. CODEN LNMAA2. ISBN 3-540-04226-1 (print), 3-540-35855-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098569/.

Floret:1968:KAG

[FW68m] Klaus Floret and Joseph Wloka. Kompakte Abbildungen. (German) [Compact maps]. Lecture Notes in Mathematics, 56:85– 93, 1968. CODEN LNMAA2. ISBN 3-540-04226-1 (print), 3-540-35855-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098568/.

Floret:1968:KGC

[FW68n] Klaus Floret and Joseph Wloka. Kompaktheit. (German) [Compactness]. Lecture Notes in Mathematics, 56:74–77, 1968. CODEN LNMAA2. ISBN 3-540-04226-1 (print), 3-540-35855-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098565/.

Floret:1968:LAS

[FW68o] Klaus Floret and Joseph Wloka. Lineare Abbildungen und der Satz von Hahn–Banach. (German) [Linear maps and the Hahn– Banach theorem]. Lecture Notes in Mathematics, 56:27–34, 1968. CODEN LNMAA2. ISBN 3-540-04226-1 (print), 3-540-35855-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0098554/.

Floret:1968:LRG

[FW68p] Klaus Floret and Joseph Wloka. Lokalkonvexe Räume. (German) [Local convex spaces]. Lecture Notes in Mathematics, 56: 19–27, 1968. CODEN LNMAA2. ISBN 3-540-04226-1 (print), 3-540-35855-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098553/.

Floret:1968:OGA

[FW68q] Klaus Floret and Joseph Wloka. Offene und Graphen-abgeschlossene Abbildungen. (German) [Open and closed-graph maps]. Lecture Notes in Mathematics, 56:39–43, 1968. CODEN LNMAA2. ISBN 3-540-04226-1 (print), 3-540-35855-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098556/.

Floret:1968:SKG

[FW68r] Klaus Floret and Joseph Wloka. Spezielle Kompaktheitskriterien. (German) [Special compactness criteria]. Lecture Notes in Mathematics, 56:78–81, 1968. CODEN LNMAA2. ISBN 3-540-04226-1 (print), 3-540-35855-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098566/.

Floret:1968:TRG

[FW68s] Klaus Floret and Joseph Wloka. Tonnelierte Räume. (German) [???? spaces]. Lecture Notes in Mathematics, 56:49–54, 1968. CODEN LNMAA2. ISBN 3-540-04226-1 (print), 3-540-35855-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098559/.

Floret:1968:TGG

[FW68t] Klaus Floret and Joseph Wloka. Topologische Grundbegriffe. (German) [Topological basic concepts]. Lecture Notes in Mathematics, 56:1–6, 1968. CODEN LNMAA2. ISBN 3-540-04226-1 (print), 3-540-35855-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098550/.

Floret:1968:TVG

[FW68u] Klaus Floret and Joseph Wloka. Topologische Vektorräume. (German) [Topological vector spaces]. Lecture Notes in Mathematics, 56:11–19, 1968. CODEN LNMAA2. ISBN 3-540-04226-1 (print), 3-540-35855-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098552/.

Frohlich:1969:FEA

[FW69] A. Fröhlich and C. T. C. Wall. Foundations of equivariant algebraic K-theory. Lecture Notes in Mathematics, 108:12–27, 1969.

CODEN LNMAA2. ISBN 3-540-04627-5 (print), 3-540-36156-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059988/.

Georgiou:1967:VMZ

[Geo67] Paul Georgiou. Vektorwertige Masse und Zufallsvariable auf Booleschen Algebren und der Satz von Radon-Nikodym. (German) [Vector-valued dimension and random variable on Boolean algebras and the Radon–Nikodym theorem]. Lecture Notes in Mathematics, 31:69–78, 1967. CODEN LNMAA2. ISBN 3-540-03902-3 (print), 3-540-34970-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061110/.

Geller:1969:NCG

[GH69a] Dennis P. Geller and Stephen T. Hedetniemi. A note on a category of graphs. Lecture Notes in Mathematics, 110:89–90, 1969. CODEN LNMAA2. ISBN 3-540-04629-1 (print), 3-540-36161-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060109.

Greenwell:1969:RG

[GH69b] D. L. Greenwell and R. L. Hemminger. Reconstructing graphs. Lecture Notes in Mathematics, 110:91–114, 1969. CODEN LN-MAA2. ISBN 3-540-04629-1 (print), 3-540-36161-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0060110/.

Giroux:1968:TFC

[Gir68] G. Giroux. Théorie des frontières dans les chaînes de Markov. (French) [Boundary theory in Markov chains]. Lecture Notes in Mathematics, 51:75–110, 1968. CODEN LNMAA2. ISBN 3-540-04221-0 (print), 3-540-35827-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059180/.

Gromoll:1968:AGA

[GKM68a] Dr. Detlef Gromoll, Professor Dr. Wilhelm Klingenberg, and Dr. Wolfgang Meyer. Anhang. (German) [Appendix]. Lecture Notes in Mathematics, 55:272–282, 1968. CODEN LNMAA2. ISBN 3-540-04225-3 (print), 3-540-35901-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-35901-2_8/.

Gromoll:1968:BM

[GKM68b] Dr. Detlef Gromoll, Professor Dr. Wilhelm Klingenberg, and Dr. Wolfgang Meyer. Back matter. Lecture Notes in Mathematics, 55:283–290, 1968. CODEN LNMAA2. ISBN 3-540-04225-3 (print), 3-540-35901-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-35901-2/1.pdf.

Gromoll:1968:BZK

[GKM68c] Dr. Detlef Gromoll, Professor Dr. Wilhelm Klingenberg, and Dr. Wolfgang Meyer. Beziehungen zwischen Krümmung und topologischer Gestalt. (German) [Relations between curvature and topological shape]. Lecture Notes in Mathematics, 55:197– 271, 1968. CODEN LNMAA2. ISBN 3-540-04225-3 (print), 3-540-35901-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-35901-2_7/.

Gromoll:1968:DMA

[GKM68d] Dr. Detlef Gromoll, Professor Dr. Wilhelm Klingenberg, and Dr. Wolfgang Meyer. Differenzierbare Mannigfaltigkeiten und Abbildungen. (German) [Differentiable manifolds and maps]. Lecture Notes in Mathematics, 55:1–34, 1968. CODEN LNMAA2. ISBN 3-540-04225-3 (print), 3-540-35901-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-35901-2_1/.

Gromoll:1968:EGG

[GKM68e] Dr. Detlef Gromoll, Professor Dr. Wilhelm Klingenberg, and Dr. Wolfgang Meyer. Extremaleigenschaften von Geodätischen. (German) [Extremal properties of geodesics]. Lecture Notes in Mathematics, 55:121–155, 1968. CODEN LNMAA2. ISBN 3-540-04225-3 (print), 3-540-35901-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-35901-2_4/.

Gromoll:1968:FM

[GKM68f] Dr. Detlef Gromoll, Professor Dr. Wilhelm Klingenberg, and Dr. Wolfgang Meyer. Front matter. Lecture Notes in Mathematics, 55:I–VI, 1968. CODEN LNMAA2. ISBN 3-540-04225-3 (print), 3-540-35901-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-35901-2/1.pdf.

Gromoll:1968:LZG

[GKM68g] Dr. Detlef Gromoll, Professor Dr. Wilhelm Klingenberg, and Dr. Wolfgang Meyer. Lineare Zusammenhänge. (German) [Linear correlations]. Lecture Notes in Mathematics, 55:35–68, 1968. CODEN LNMAA2. ISBN 3-540-04225-3 (print), 3-540-35901-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-35901-2_2/.

Gromoll:1968:RGI

[GKM68h] Dr. Detlef Gromoll, Professor Dr. Wilhelm Klingenberg, and Dr. Wolfgang Meyer. Riemannsche Geometrie im Großen. (German) [Riemannian geometry in the large], volume 55 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. CODEN LNMAA2. ISBN 3-540-04225-3 (paperback), 3-540-35901-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 287 pp. LCCN

QA3 .L28 no. 55. URL http://link.springer.com/book/10.
1007/978-3-540-35901-2.

Gromoll:1968:RMM

[GKM68i] Dr. Detlef Gromoll, Professor Dr. Wilhelm Klingenberg, and Dr. Wolfgang Meyer. Riemannsche Mannigfaltigkeiten als metrische Räume. (German) [Riemannian manifolds as metric spaces]. Lecture Notes in Mathematics, 55:156–173, 1968. CODEN LNMAA2. ISBN 3-540-04225-3 (print), 3-540-35901-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-540-35901-2_5/.

Gromoll:1968:RMG

[GKM68j] Dr. Detlef Gromoll, Professor Dr. Wilhelm Klingenberg, and Dr. Wolfgang Meyer. Riemannsche Mannigfaltigkeiten. (German) [Riemannian manifolds]. Lecture Notes in Mathematics, 55:69– 120, 1968. CODEN LNMAA2. ISBN 3-540-04225-3 (print), 3-540-35901-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-35901-2_3/.

Gromoll:1968:VGC

[GKM68k] Dr. Detlef Gromoll, Professor Dr. Wilhelm Klingenberg, and Dr. Wolfgang Meyer. Vergleichssätze. (German) [Comparative theorems]. Lecture Notes in Mathematics, 55:174–196, 1968. CODEN LNMAA2. ISBN 3-540-04225-3 (print), 3-540-35901-X (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-540-35901-2_6/.

Gorenflo:1969:MDS

[Gor69] Rudolf Gorenflo. Monotonic difference schemes for weakly coupled systems of parabolic differential equations. Lecture Notes in Mathematics, 109:160–167, 1969. CODEN LNMAA2. ISBN 3-540-04628-3 (print), 3-540-36158-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060023/.

Gourlay:1969:NSE

[Gou69] A. R. Gourlay. The numerical solution of evolutionary partial differential equations. Lecture Notes in Mathematics, 109:168– 171, 1969. CODEN LNMAA2. ISBN 3-540-04628-3 (print), 3-540-36158-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060024/.

Gray:1969:CCS

[Gra69] John W. Gray. The categorical comprehension scheme. Lecture Notes in Mathematics, 99:242–312, 1969. CODEN LNMAA2. ISBN 3-540-04618-6 (print), 3-540-36140-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081965/.

Grenander:1967:TTP

[Gre67] Ulf Grenander. Toward a theory of patterns. Lecture Notes in Mathematics, 31:79–111, 1967. CODEN LNMAA2. ISBN 3-540-03902-3 (print), 3-540-34970-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061111/.

Grigorieff:1969:RED

[Gri69] Rolf Dieter Grigorieff. Randwertaufgaben für elliptische Differentialoperatoren und Bilinearformen in den Sobolewschen Räumen W^{m,2} (Ω). (German) [Boundary value problems for elliptic differential operators and bilinear forms in the Sobolev spaces W^{m,2}(Ω)]. Lecture Notes in Mathematics, 102:339–381, 1969. CODEN LN-MAA2. ISBN 3-540-04621-6 (print), 3-540-36144-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0059071/.

Grothendieck:1968:AEF

[Gro68a] A. Grothendieck. Application aux extensions de faisceaux d’anneaux. (French) [Application to extensions of ring sheaves]. Lecture Notes in Mathematics, 79:69–85, 1968. CODEN LN-MAA2. ISBN 3-540-04248-2 (print), 3-540-35922-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082444/.

Grothendieck:1968:AVI

[Gro68b] A. Grothendieck. Application aux “variations infinitésimales” de faisceaux d’algèbres. (French) [Application to “infinitesimal” variations of algebra sheaves]. Lecture Notes in Mathematics, 79:85– 92, 1968. CODEN LNMAA2. ISBN 3-540-04248-2 (print), 3-540-35922-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082445/.

Grothendieck:1968:ADC

[Gro68c] A. Grothendieck. Applications du complexe cotangent relatif, et problèmes ouverts. (French) [Applications of tge relative cotangent complex, and open problems]. Lecture Notes in Mathematics, 79:154–164, 1968. CODEN LNMAA2. ISBN 3-540-04248-2 (print), 3-540-35922-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082449/.

Grothendieck:1968:CCC

[Gro68d] A. Grothendieck. Cas d’une catégorie cofibrée exacte à gauche. (French) [Case of a left-exact cofibrous category]. Lecture Notes in Mathematics, 79:40–48, 1968. CODEN LNMAA2. ISBN 3-540-04248-2 (print), 3-540-35922-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082442/.

Grothendieck:1968:CCAb

[Gro68e] A. Grothendieck. Catégories Cofibrées Additives et Complexe Cotangent Relatif. (French) [Cofibered Additive Categories and Relative Complex Cotangent], volume 79 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. CODEN LNMAA2. ISBN 3-540-04248-2 (paperback), 3-540-35922-2 (electronic). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 167 pp. LCCN QA3 .L28 no. 79. URL

http://link.springer.com/book/10.1007/BFb0082437.

Grothendieck:1968:CCAa

[Gro68f] A. Grothendieck. Catégories cofibrées additives. (French) [Cofibrous additive categories]. Lecture Notes in Mathematics, 79:5– 14, 1968. CODEN LNMAA2. ISBN 3-540-04248-2 (print), 3-540-35922-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082438/.

Grothendieck:1968:CCD

[Gro68g] A. Grothendieck. Catégories cofibrées définies par des complexes de chaînes, et théorèmes de représentabilité. (French) [Cofibrous defined categories of chain complexes, and representability theorems]. Lecture Notes in Mathematics, 79:48–68, 1968. CODEN LNMAA2. ISBN 3-540-04248-2 (print), 3-540-35922-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082443/.

Grothendieck:1968:CCE

[Gro68h] A. Grothendieck. Catégories cofibrées exactes à gauche. (French) [Left-exact cofibrous categories]. Lecture Notes in Mathematics, 79:14–26, 1968. CODEN LNMAA2. ISBN 3-540-04248-2 (print), 3-540-35922-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082439/.

Grothendieck:1968:CCR

[Gro68i] A. Grothendieck. Complexe cotangent relatif et relèvement infinitésimal de morphismes de topos annelés. Application aux morphismes formellement nets. (French) [Relative cotangent complex and infinitesimal increase of morphisms of ringed topos. Application to formally-clean morphisms]. Lecture Notes in Mathematics, 79:128–154, 1968. CODEN LNMAA2. ISBN 3-540-04248-2 (print), 3-540-35922-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082448/.

Grothendieck:1968:CTO

[Gro68j] A. Grothendieck. Complexe typique L^{X} d’un objet de E. procomplexe typique L^{E} de E. (French) [Typical complex L^{X} of an object E. Typical procomplexe L^{E} of E]. Lecture Notes in Mathematics, 79:27–36, 1968. CODEN LNMAA2. ISBN 3-540-04248-2 (print), 3-540-35922-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082440/.

Grothendieck:1968:PHC

[Gro68k] A. Grothendieck. Le proobjet N_{E} et l’homorphisme caractéristique d’un objet de E. (French) [The proobjet N_{E} and the characteristic homomorphism of an object E]. Lecture Notes in Mathematics, 79:36–40, 1968. CODEN LNMAA2. ISBN 3-540-04248-2 (print), 3-540-35922-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082441/.

Grothendieck:1968:PGD

[Gro68l] A. Grothendieck. Propriétés générales du complexe cotangent relatif. (French) [General properties of the relative cotangent complex]. Lecture Notes in Mathematics, 79:93–111, 1968. CODEN LNMAA2. ISBN 3-540-04248-2 (print), 3-540-35922-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082446/.

Grothendieck:1968:SET

[Gro68m] A. Grothendieck. Suites exactes de transitivité. (French) [Exact sequences of transitivity]. Lecture Notes in Mathematics, 79:112– 128, 1968. CODEN LNMAA2. ISBN 3-540-04248-2 (print), 3-540-35922-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082447/.

Gruenberg:1969:CGE

[Gru69a] K. W. Gruenberg. Categories of group extensions. Lecture Notes in Mathematics, 92:105–116, 1969. CODEN LNMAA2. ISBN 3-540-04611-9 (print), 3-540-36101-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080767/.

Grunbaum:1969:IPG

[Grü69b] Branko Grünbaum. Incidence patterns of graphs and complexes. Lecture Notes in Mathematics, 110:115–128, 1969. CODEN LN-MAA2. ISBN 3-540-04629-1 (print), 3-540-36161-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0060111/.

Guber:1967:PTL

[Gub67] Siegfried Guber. On the potential theory of linear, homogeneous parabolic partial differential equations of second order. Lecture Notes in Mathematics, 31:112–117, 1967. CODEN LNMAA2. ISBN 3-540-03902-3 (print), 3-540-34970-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061112/.

Guy:1969:MFP

[Guy69] Richard K. Guy. A many-facetted problem of Zarankiewicz. Lecture Notes in Mathematics, 110:129–148, 1969. CODEN LN-MAA2. ISBN 3-540-04629-1 (print), 3-540-36161-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0060112/.

Hamelink:1969:GTL

[Ham69] Ronald C. Hamelink. Graph theory and Lie algebra. Lecture Notes in Mathematics, 110:149–153, 1969. CODEN LNMAA2. ISBN 3-540-04629-1 (print), 3-540-36161-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060113/.

Hartshorne:1966:AP

[Har66a] Robin Hartshorne. Application to preschemes. Lecture Notes in Mathematics, 20:85–136, 1966. CODEN LNMAA2. ISBN 3-540-03603-2 (print), 3-540-34794-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080485/.

Hartshorne:1966:DC

[Har66b] Robin Hartshorne. The derived category. Lecture Notes in Mathematics, 20:19–84, 1966. CODEN LNMAA2. ISBN 3-540-03603-2 (print), 3-540-34794-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080484/.

Hartshorne:1966:DPM

[Har66c] Robin Hartshorne. Duality for projective morphisms. Lecture Notes in Mathematics, 20:137–214, 1966. CODEN LNMAA2. ISBN 3-540-03603-2 (print), 3-540-34794-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080486/.

Hartshorne:1966:DT

[Har66d] Robin Hartshorne. The duality theorem. Lecture Notes in Mathematics, 20:357–393, 1966. CODEN LNMAA2. ISBN 3-540-03603-2 (print), 3-540-34794-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080490/.

Hartshorne:1966:DCL

[Har66e] Robin Hartshorne. Dualizing complexes and local duality. Lecture Notes in Mathematics, 20:252–301, 1966. CODEN LNMAA2. ISBN 3-540-03603-2 (print), 3-540-34794-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080488/.

Hartshorne:1966:I

[Har66f] Robin Hartshorne. Introduction. Lecture Notes in Mathematics, 20:1–18, 1966. CODEN LNMAA2. ISBN 3-540-03603-2 (print), 3-540-34794-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080483/.

Hartshorne:1966:LC

[Har66g] Robin Hartshorne. Local cohomology. Lecture Notes in Mathematics, 20:215–251, 1966. CODEN LNMAA2. ISBN 3-540-03603-2 (print), 3-540-34794-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080487/.

Hartshorne:1966:RC

[Har66h] Robin Hartshorne. Residual complexes. Lecture Notes in Mathematics, 20:302–356, 1966. CODEN LNMAA2. ISBN 3-540-03603-2 (print), 3-540-34794-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080489/.

Hartshorne:1966:RDL

[Har66i] Robin Hartshorne, editor. Residues and Duality: Lecture Notes of a seminar on the work of A. Grothendieck, given at Harvard 1963/ 64, volume 20 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1966. CODEN LNMAA2. ISBN 3-540-03603-2 (paperback), 3-540-34794-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

LCCN QA3 .L28 no. 20. URL http://link.springer.com/book/

10.1007/BFb0080482. With an appendix by P. Deligne.

Hartshorne:1967:ALC

[Har67a] Robin Hartshorne. Application of local cohomology to preschemes. Lecture Notes in Mathematics, 41:16–35, 1967. CODEN LN-MAA2. ISBN 3-540-03912-0 (print), 3-540-35183-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0073973/.

Hartshorne:1967:DEP

[Har67b] Robin Hartshorne. Definition and elementary properties of the local cohomology groups. Lecture Notes in Mathematics, 41:1– 15, 1967. CODEN LNMAA2. ISBN 3-540-03912-0 (print), 3-540-35183-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073972/.

Hartshorne:1967:FM

[Har67c] Robin Hartshorne. Functors on a-modules. Lecture Notes in Mathematics, 41:48–68, 1967. CODEN LNMAA2. ISBN 3-540-03912-0 (print), 3-540-35183-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073975/.

Hartshorne:1967:LD

[Har67d] Robin Hartshorne. Local duality. Lecture Notes in Mathematics, 41:81–104, 1967. CODEN LNMAA2. ISBN 3-540-03912-0 (print), 3-540-35183-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073977/.

Hartshorne:1967:RD

[Har67e] Robin Hartshorne. Relation to depth. Lecture Notes in Mathematics, 41:36–47, 1967. CODEN LNMAA2. ISBN 3-540-03912-0 (print), 3-540-35183-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073974/.

Hartshorne:1967:SA

[Har67f] Robin Hartshorne. Some applications. Lecture Notes in Mathematics, 41:69–80, 1967. CODEN LNMAA2. ISBN 3-540-03912-0 (print), 3-540-35183-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073976/.

Harish-Chandra:1968:AFS

[HC68a] Harish-Chandra. Automorphic Forms on Semisimple Lie Groups, volume 62 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. CODEN LNMAA2. ISBN 3-540-04232-6 (paperback), 3-540-35865-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 138 pp. LCCN QA3 .L28 no. 62. URL http://

link.springer.com/book/10.1007/BFb0098434. Notes by J. G.

M. Mars.

Harish-Chandra:1968:C

[HC68b] Harish-Chandra. Chapter I. Lecture Notes in Mathematics, 62: 1–23, 1968. CODEN LNMAA2. ISBN 3-540-04232-6 (print), 3-540-35865-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098435/.

Harish-Chandra:1968:CIa

[HC68c] Harish-Chandra. Chapter II. Lecture Notes in Mathematics, 62: 24–53, 1968. CODEN LNMAA2. ISBN 3-540-04232-6 (print), 3-540-35865-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098436/.

Harish-Chandra:1968:CIb

[HC68d] Harish-Chandra. Chapter III. Lecture Notes in Mathematics, 62: 54–68, 1968. CODEN LNMAA2. ISBN 3-540-04232-6 (print), 3-540-35865-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098437/.

Harish-Chandra:1968:CIc

[HC68e] Harish-Chandra. Chapter IV. Lecture Notes in Mathematics, 62: 69–108, 1968. CODEN LNMAA2. ISBN 3-540-04232-6 (print), 3-540-35865-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098438/.

Harish-Chandra:1968:CV

[HC68f] Harish-Chandra. Chapter V. Lecture Notes in Mathematics, 62: 109–138, 1968. CODEN LNMAA2. ISBN 3-540-04232-6 (print), 3-540-35865-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098439/.

Hedetniemi:1969:CGD

[Hed69] Stephen Hedetniemi. On classes of graphs defined by special cutsets of lines. Lecture Notes in Mathematics, 110:171–189, 1969. CODEN LNMAA2. ISBN 3-540-04629-1 (print), 3-540-36161-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060115/.

Heins:1969:BP

[Hei69a] Maurice Heins. Boundary problems. Lecture Notes in Mathematics, 98:52–87, 1969. CODEN LNMAA2. ISBN 3-540-04617-8 (print), 3-540-36139-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080779/.

Heins:1969:CPR

[Hei69b] Maurice Heins. A classification problem for Riemann surfaces. Lecture Notes in Mathematics, 98:34–51, 1969. CODEN LN-MAA2. ISBN 3-540-04617-8 (print), 3-540-36139-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080778/.

Heins:1969:GOP

[Hei69c] Maurice Heins. General observations and preliminaries. Lecture Notes in Mathematics, 98:2–12, 1969. CODEN LNMAA2. ISBN 3-540-04617-8 (print), 3-540-36139-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080776/.

Heins:1969:HCR

[Hei69d] Maurice Heins. Hardy Classes on Riemann Surfaces, volume 98 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CO-DEN LNMAA2. ISBN 3-540-04617-8 (paperback), 3-540-36139-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 106

pp. LCCN QA3 .L28 no. 98. URL http://link.springer.com/
book/10.1007/BFb0080775.

Heins:1969:TSS

[Hei69e] Maurice Heins. The theorem of Szegő–Solomentsev. Lecture Notes in Mathematics, 98:13–33, 1969. CODEN LNMAA2. ISBN 3-540-04617-8 (print), 3-540-36139-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080777/.

Heins:1969:VVF

[Hei69f] Maurice Heins. Vector-valued functions. Lecture Notes in Mathematics, 98:88–104, 1969. CODEN LNMAA2. ISBN 3-540-04617-8 (print), 3-540-36139-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080780/.

Hengartner:1968:FTS

[Hen68] Walter Hengartner. Famille de traces sur les sous-espaces d’une fonction plurisousharmonique ou entière dans C^{n}. (French) [Trace families on the subspaces of a plurisubharmonic or complete function on C^{n}]. Lecture Notes in Mathematics, 71:81–93, 1968. CO-DEN LNMAA2. ISBN 3-540-04241-5 (print), 3-540-35889-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0080099/.

Hermes:1965:BM

[Her65a] Hans Hermes. Back matter. Lecture Notes in Mathematics, 6: 40–42, 1965. CODEN LNMAA2. ISBN 3-540-04899-5 (print), 3-540-36203-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-36203-6/1.pdf.

Hermes:1965:CRT

[Her65b] Hans Hermes. Completeness of restricted term calculus. Lecture Notes in Mathematics, 6:52, 1965. CODEN LNMAA2. ISBN 3-540-04899-5 (print), 3-540-36203-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36203-6_15.

Hermes:1965:DP

[Her65c] Hans Hermes. Details of the proof. Lecture Notes in Mathematics, 6:47–51, 1965. CODEN LNMAA2. ISBN 3-540-04899-5 (print), 3-540-36203-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36203-6_14/.

Hermes:1965:KTG

[Her65d] Hans Hermes. Ein Kalkül für die Termlogik. (German) [A calculus for term togic]. Lecture Notes in Mathematics, 6:17–20, 1965. CODEN LNMAA2. ISBN 3-540-04899-5 (print), 3-540-36203-7 (e-book), 3-540-03350-5 (print), 3-540-69980-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-540-36203-6_6/;
 http://link.springer.com/chapter/10.1007/BFb0097727/.

Hermes:1965:TAGb

[Her65e] Hans Hermes. Eine Termlogik mit Auswahloperator. (German) [A term logic with selection operator], volume 6 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1965. CODEN LNMAA2. ISBN 3-540-04899-5 (paperback), 3-540-36203-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 42 pp. LCCN QA3 .L28 no.

6. URL http://link.springer.com/book/10.1007/978-3-540-
36203-6.

Hermes:1965:TAGc

[Her65f] Hans Hermes. Eine Termlogik mit Auswahloperator. (German) [A term logic with selection operator], volume 6 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1965. CODEN LNMAA2. ISBN 3-540-03350-5 (paperback), 3-540-69980-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 55 pp. LCCN QA3

.L28 no. 6. URL http://link.springer.com/book/10.1007/
BFb0097721.

Hermes:1965:EGI

[Her65g] Hans Hermes. Einleitung. (German) [Introduction]. Lecture Notes in Mathematics, 6:1–3, 1965. CODEN LNMAA2. ISBN 3-540-04899-5 (print), 3-540-36203-7 (e-book), 3-540-03350-5 (print), 3-540-69980-5 (e-book). ISSN 0075-8434 (print), 1617-9692

(electronic). URL http://link.springer.com/chapter/10.
1007/978-3-540-36203-6_1/;   http://link.springer.com/
chapter/10.1007/BFb0097722/.

Hermes:1965:ETS

[Her65h] Hans Hermes. Erfüllbarkeit von M∗. Die Terme s und t sollen (relativ zu M∗). (German) [Satisfiability of M∗. The terms s and t are (relative to M∗)]. Lecture Notes in Mathematics, 6:

35–39, 1965. CODEN LNMAA2. ISBN 3-540-04899-5 (print), 3-540-36203-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36203-6_12/.

Hermes:1965:EGS

[Her65i] Hans Hermes. Erfüllbarkeit von M^{∗}. (German) [Satisfiability of M^{∗}]. Lecture Notes in Mathematics, 6:35–39, 1965. CODEN LN-MAA2. ISBN 3-540-03350-5 (print), 3-540-69980-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0097733/.

Hermes:1965:FM

[Her65j] Hans Hermes. Front matter. Lecture Notes in Mathematics, 6: i–iii, 1965. CODEN LNMAA2. ISBN 3-540-04899-5 (print), 3-540-36203-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-36203-6/1.pdf.

Hermes:1965:GKR

[Her65k] Hans Hermes. Gleichwertigkeit von [Vdash] und ⊢. Korrektheit der Regeln. (German) [Equivalence of [Vdash] and ⊢. Correctness of the rules]. Lecture Notes in Mathematics, 6:21–22, 1965. CODEN LNMAA2. ISBN 3-540-04899-5 (print), 3-540-36203-7 (e-book), 3-540-03350-5 (print), 3-540-69980-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36203-6_7;   http://link.
springer.com/chapter/10.1007/BFb0097728.

Hermes:1965:MGM

[Her65l] Hans Hermes. Maximalisierung von M^{φ}. (German) [Maximization of M^{φ}]. Lecture Notes in Mathematics, 6:26–27, 1965. CODEN LNMAA2. ISBN 3-540-03350-5 (print), 3-540-69980-5 (e-book), 3-540-04899-5 (print), 3-540-36203-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-540-36203-6_10;
 http://link.springer.com/chapter/10.1007/BFb0097731.

Hermes:1965:PAG

[Her65m] Hans Hermes. Prädikatenlogik mit Auswahloperator. (German) [Predicate logic with selection operator]. Lecture Notes in Mathematics, 6:4–6, 1965. CODEN LNMAA2. ISBN 3-540-04899-5 (print), 3-540-36203-7 (e-book), 3-540-03350-5 (print), 3-540-69980-5 (e-book). ISSN 0075-8434 (print), 1617-9692

(electronic). URL http://link.springer.com/chapter/10.
1007/978-3-540-36203-6_2/;   http://link.springer.com/
chapter/10.1007/BFb0097723/.

Hermes:1965:RFV

[Her65n] Hans Hermes. Rang, freies Vorkommen einer Variablen, Substitution. (German) [Rank, free occurrences of a variable, substitution]. Lecture Notes in Mathematics, 6:15–16, 1965. CODEN LNMAA2. ISBN 3-540-04899-5 (print), 3-540-36203-7 (e-book), 3-540-03350-5 (print), 3-540-69980-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36203-6_5;   http://link.
springer.com/chapter/10.1007/BFb0097726.

Hermes:1965:SPT

[Her65o] Hans Hermes. Survey of the proof of the theorem on satisfiability. Lecture Notes in Mathematics, 6:42–46, 1965. CODEN LN-MAA2. ISBN 3-540-04899-5 (print), 3-540-36203-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-36203-6_13/.

Hermes:1965:TGT

[Her65p] Hans Hermes. Termisomorphismen. (German) [Term isomorphism]. Lecture Notes in Mathematics, 6:24–25, 1965. CODEN LNMAA2. ISBN 3-540-04899-5 (print), 3-540-36203-7 (e-book), 3-540-03350-5 (print), 3-540-69980-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36203-6_9;   http://link.
springer.com/chapter/10.1007/BFb0097730.

Hermes:1965:TAGa

[Her65q] Hans Hermes. Termlogik mit Auswahloperator. (German) [Term logic with selection operator]. Lecture Notes in Mathematics, 6: 7–9, 1965. CODEN LNMAA2. ISBN 3-540-04899-5 (print), 3-540-36203-7 (e-book), 3-540-03350-5 (print), 3-540-69980-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-540-36203-6_3/;
 http://link.springer.com/chapter/10.1007/BFb0097724/.

Hermes:1965:UVG

[Her65r] Hans Hermes. Übersicht über den Vollständigkeitsbeweis. (German) [Overview of the completeness proof]. Lecture Notes in Mathematics, 6:23, 1965. CODEN LNMAA2. ISBN 3-540-04899-5 (print), 3-540-36203-7 (e-book), 3-540-03350-5 (print), 3-540-69980-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36203-6_8;   http://link.springer.com/chapter/
10.1007/BFb0097729.

Hermes:1965:VSG

[Her65s] Hans Hermes. Verallgemeinerte Substitution. (German) [Generalized substitution]. Lecture Notes in Mathematics, 6:28–34, 1965. CODEN LNMAA2. ISBN 3-540-03350-5 (print), 3-540-69980-5 (e-book), 3-540-04899-5 (print), 3-540-36203-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-540-36203-6_11/
;   http://link.springer.com/chapter/10.1007/BFb0097732/

.

Hermes:1965:ZZP

[Her65t] Hans Hermes. Zusammenhang zwischen der Prädikatenlogik und der Termlogik. (German) [Relation between the predicate logic and term logic]. Lecture Notes in Mathematics, 6:10–14, 1965. CODEN LNMAA2. ISBN 3-540-04899-5 (print), 3-540-36203-7 (e-book), 3-540-03350-5 (print), 3-540-69980-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-540-36203-6_4/;
 http://link.springer.com/chapter/10.1007/BFb0097725/.

Herz:1966:CSI

[Her66a] C. Herz. Computation of singular j-invariants. Lecture Notes in Mathematics, 21:VIII–1–VIII–11, 1966. CODEN LNMAA2. ISBN 3-540-03604-0 (print), 3-540-34798-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091558/.

Herz:1966:CCF

[Her66b] Carl S. Herz. Construction of class fields. Lecture Notes in Mathematics, 21:VII–1–VII–21, 1966. CODEN LNMAA2. ISBN 3-540-03604-0 (print), 3-540-34798-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091557/.

Herrlich:1968:KHG

[Her68a] Horst Herrlich. Kategorieller Hintergrund. (German) [Categorial background]. Lecture Notes in Mathematics, 78:32–75, 1968. CO-DEN LNMAA2. ISBN 3-540-04247-4 (print), 3-540-35920-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0074314/.

Herrlich:1968:RCG

[Her68b] Horst Herrlich. Reflexionen und Coreflexionen. (German) [Reflections and coreflections]. Lecture Notes in Mathematics, 78:76– 113, 1968. CODEN LNMAA2. ISBN 3-540-04247-4 (print), 3-540-35920-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074315/.

Herrlich:1968:TEM

[Her68c] Horst Herrlich. Topologische Epireflexionen und (Mono-) Coreflexionen. (German) [Topological epireflections and (mono-) core reflections]. Lecture Notes in Mathematics, 78:114–152, 1968. CO-DEN LNMAA2. ISBN 3-540-04247-4 (print), 3-540-35920-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0074316/.

Herrlich:1968:TRC

[Her68d] Horst Herrlich. Topologische Reflexionen und Coreflexionen. (German) [Topological reflections and core inflections], volume 78 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. CODEN LN-MAA2. ISBN 3-540-04247-4 (paperback), 3-540-35920-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 166 pp. LCCN QA3 .L28 no. 78; QA3 .L28 no.78. URL http://link.

springer.com/book/10.1007/BFb0074312.

Herrlich:1968:VRK

[Her68e] Horst Herrlich. Vollständig Reguläre, Kompakte und Reellkompakte Räume. (German) [Completely regular, compact and real compact spaces]. Lecture Notes in Mathematics, 78:1–31, 1968. CODEN LNMAA2. ISBN 3-540-04247-4 (print), 3-540-35920-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074313/.

Hestenes:1969:RG

[Hes69] Marshall D. Hestenes. Rank 3 graphs. Lecture Notes in Mathematics, 110:191–192, 1969. CODEN LNMAA2. ISBN 3-540-04629-1 (print), 3-540-36161-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060116.

Hartshorne:1967:LCS

[HG67] Robin Hartshorne and A. (Alexandre) Grothendieck, editors. Local Cohomology: A seminar given by A. Grothendieck Harvard University Fall, 1961, volume 41 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. CODEN LNMAA2. ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 41. URL http://

link.springer.com/book/10.1007/BFb0073971.

Hilton:1969:FM

[Hil69a] Peter Hilton. On factorization of manifolds. Lecture Notes in Mathematics, 118:48–57, 1969. CODEN LNMAA2. ISBN 3-540-04907-X (print), 3-540-36246-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060251/.

Hilton:1969:CTHa

[Hil69b] Peter John Hilton, editor. Category Theory, Homology Theory and their Applications I: Proceedings of the Conference held at the Seattle Research Center of the Battelle Memorial Institute, June 24– July 19, 1968. Volume One, volume 70 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CODEN LNMAA2. ISBN 3-540-04605-4 (paperback), 3-540-36095-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 70; BC135 .S66 1967;

QA3 .L28 no. 70. URL http://link.springer.com/book/10.
1007/BFb0079380.

Hilton:1969:CTHb

[Hil69c] Peter John Hilton, editor. Category Theory, Homology Theory and their Applications II: Proceedings of the Conference held at the Seattle Research Center of the Battelle Memorial Institute, June 24–July 19, 1968 Volume Two, volume 92 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CODEN LNMAA2. ISBN 3-540-04611-9 (paperback), 3-540-36101-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 92. URL

http://link.springer.com/book/10.1007/BFb0080761.

Hirschfelt:1969:US

[Hir69] J. Hirschfelt. Uniform structures. Lecture Notes in Mathematics, 94:48–79, 1969. CODEN LNMAA2. ISBN 3-540-04613-5 (print), 3-540-36129-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101449/.

Hirzebruch:1968:MES

[HM68a] F. Hirzebruch and K. H. Mayer. O(n)-Mannigfaltigkeiten, exotische Sphären und Singularitäten. (German) [O(n)-manifolds, exotic spheres and singularities], volume 57 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. CODEN LNMAA2. ISBN 3-540-04227-X (paperback), 3-540-35858-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 132 pp. LCCN QA3 .L28 no. 57. URL

http://link.springer.com/book/10.1007/BFb0074355.

Hirzebruch:1968:AVG

[HM68b] F. Hirzebruch and K. H. Mayer. Äquivariantes Verkleben. (German) [Equivariant adhesion]. Lecture Notes in Mathematics, 57: 42–48, 1968. CODEN LNMAA2. ISBN 3-540-04227-X (print), 3-540-35858-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074362/.

Hirzebruch:1968:BSG

[HM68c] F. Hirzebruch and K. H. Mayer. Bericht über Sphären. (German) [Report on spheres]. Lecture Notes in Mathematics, 57:67–75, 1968. CODEN LNMAA2. ISBN 3-540-04227-X (print), 3-540-35858-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074365/.

Hirzebruch:1968:KSM

[HM68d] F. Hirzebruch and K. H. Mayer. Der Klassifikationssatz für spezielle G-Mannigfaltigkeiten. (German) [The classification theorem for special G-manifolds]. Lecture Notes in Mathematics, 57: 15–21, 1968. CODEN LNMAA2. ISBN 3-540-04227-X (print), 3-540-35858-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074358/.

Hirzebruch:1968:GHG

[HM68e] F. Hirzebruch and K. H. Mayer. Die ganzzahlige homologie gewisser affin algebraischer Mannigfaltigkeiten. (German) [The integer homology of certain affine algebraic manifolds]. Lecture Notes in Mathematics, 57:79–88, 1968. CODEN LNMAA2. ISBN 3-540-04227-X (print), 3-540-35858-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074367/.

Hirzebruch:1968:HBG

[HM68f] F. Hirzebruch and K. H. Mayer. Die Homologie von Baummannigfaltigkeiten. (German) [The homology of tree manifolds]. Lecture Notes in Mathematics, 57:49–59, 1968. CODEN LNMAA2. ISBN 3-540-04227-X (print), 3-540-35858-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074363/.

Hirzebruch:1968:MGMb

[HM68g] F. Hirzebruch and K. H. Mayer. Die Mannigfaltigkeiten W^{2n−1}(d). (German) [The manifold W^{2n−1}(d)]. Lecture Notes in Mathematics, 57:29–36, 1968. CODEN LNMAA2. ISBN 3-540-04227-X (print), 3-540-35858-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074360/.

Hirzebruch:1968:VTG

[HM68h] F. Hirzebruch and K. H. Mayer. Die Vielfachen des Tangentialbündels von S^{n}. (German) [The multiples of the tangent sheaf of S^{n}]. Lecture Notes in Mathematics, 57:37–41, 1968. CODEN LNMAA2. ISBN 3-540-04227-X (print), 3-540-35858-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074361/.

Hirzebruch:1968:MGMa

[HM68i] F. Hirzebruch and K. H. Mayer. G-Mannigfaltigkeiten. (German) [G-manifolds]. Lecture Notes in Mathematics, 57:1–9, 1968. CO-DEN LNMAA2. ISBN 3-540-04227-X (print), 3-540-35858-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0074356/.

Hirzebruch:1968:GQF

[HM68j] F. Hirzebruch and K. H. Mayer. Ganzzahlige quadratische Formen. (German) [Integer quadratic forms]. Lecture Notes in Mathematics, 57:89–101, 1968. CODEN LNMAA2. ISBN 3-540-04227-X (print), 3-540-35858-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074368/.

Hirzebruch:1968:PAS

[HM68k] F. Hirzebruch and K. H. Mayer. Periodische Abbildungen von Sphären. (German) [Periodic maps on spheres]. Lecture Notes in Mathematics, 57:113–127, 1968. CODEN LNMAA2. ISBN 3-540-04227-X (print), 3-540-35858-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074370/.

Hirzebruch:1968:QFA

[HM68l] F. Hirzebruch and K. H. Mayer. Quadratische Formen, ARFsche Invariante. (German) [Quadratic forms, ARF invariants]. Lecture Notes in Mathematics, 57:60–66, 1968. CODEN LNMAA2. ISBN 3-540-04227-X (print), 3-540-35858-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074364/.

Hirzebruch:1968:SMGb

[HM68m] F. Hirzebruch and K. H. Mayer. Spezielle O(n)-Mannigfaltigkeiten über D^{2}. (German) [Special O(n)-manifolds on D^{2}]. Lecture Notes in Mathematics, 57:22–28, 1968. CODEN LNMAA2. ISBN 3-540-04227-X (print), 3-540-35858-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074359/.

Hirzebruch:1968:SMGa

[HM68n] F. Hirzebruch and K. H. Mayer. Spezielle G-Mannigfaltigkeiten. (German) [Special G manifolds]. Lecture Notes in Mathematics, 57:10–14, 1968. CODEN LNMAA2. ISBN 3-540-04227-X (print), 3-540-35858-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074357/.

Hirzebruch:1968:SUSa

[HM68o] F. Hirzebruch and K. H. Mayer. Sphären als Umgebungsränder von Singularitäten I. (German) [Spheres as environmental edges of singularities I]. Lecture Notes in Mathematics, 57:76–78, 1968. CODEN LNMAA2. ISBN 3-540-04227-X (print), 3-540-35858-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074366/.

Hirzebruch:1968:SUSb

[HM68p] F. Hirzebruch and K. H. Mayer. Sphären als Umgebungsränder von Singularitäten II. (German) [Spheres as environmental edges of singularities II]. Lecture Notes in Mathematics, 57:102–112, 1968. CODEN LNMAA2. ISBN 3-540-04227-X (print), 3-540-35858-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074369/.

Hodgkins:1969:MNI

[Hod69] W. R. Hodgkins. A method for the numerical integration of nonlinear ordinary differential equations with greatly different time constants. Lecture Notes in Mathematics, 109:172–177, 1969. CO-DEN LNMAA2. ISBN 3-540-04628-3 (print), 3-540-36158-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0060025/.

Hoffman:1969:BAF

[Hof69] Kenneth Hoffman. Bounded analytic functions in the unit disk. Lecture Notes in Mathematics, 103:1–9, 1969. CODEN LN-MAA2. ISBN 3-540-04622-4 (print), 3-540-36145-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0099981/.

Hoobler:1969:NAS

[Hoo69] R. T. Hoobler. Non-Abelian sheaf cohomology by derived functors. Lecture Notes in Mathematics, 99:313–364, 1969. CODEN LNMAA2. ISBN 3-540-04618-6 (print), 3-540-36140-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081966/.

Hormander:1969:CPD

[Hör69] Lars Hörmander. The Cauchy problem for differential equations with constant coefficients. Lecture Notes in Mathematics, 103: 60–71, 1969. CODEN LNMAA2. ISBN 3-540-04622-4 (print), 3-540-36145-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099985/.

Hsiang:1969:STK

[Hsi69] W. C. Hsiang. A splitting theorem and the Künneth formula in algebraic K-theory. Lecture Notes in Mathematics, 108:72–77, 1969. CODEN LNMAA2. ISBN 3-540-04627-5 (print), 3-540-36156-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0059992/.

Huber:1969:ALD

[Hub69a] Catherine Huber. Un aspect de la loi du logarithme itéré pour des variables aléatoires indépendantes et équidistribuées. (French) [An aspect of the law of the iterated logarithm for independent random variables]. Lecture Notes in Mathematics, 88:137–142, 1969. CODEN LNMAA2. ISBN 3-540-04607-0 (print), 3-540-36097-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0081630/.

Huber:1969:LTM

[Hub69b] Catherine Huber. Un lemme de théorie des martingales. (French) [A lemma on martingale theory]. Lecture Notes in Mathematics, 88:143, 1969. CODEN LNMAA2. ISBN 3-540-04607-0 (print), 3-540-36097-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081631.

Husain:1969:CMP

[Hus69] Taqdir Husain. On the continuity of Markov processes. Lecture Notes in Mathematics, 89:99–105, 1969. CODEN LNMAA2. ISBN 3-540-04608-9 (print), 3-540-36098-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079120/.

Harary:1969:MVG

[HW69] Frank Harary and Dominic Welsh. Matroids versus graphs. Lecture Notes in Mathematics, 110:155–170, 1969. CODEN LN-MAA2. ISBN 3-540-04629-1 (print), 3-540-36161-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0060114/.

Igot:1968:TLF

[Igo68] J. P. Igot. Un théorème de Linnik. (French) [A theorem of Linnik]. Lecture Notes in Mathematics, 51:111–122, 1968. CODEN LNMAA2. ISBN 3-540-04221-0 (print), 3-540-35827-7 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059181/.

Ivanescu:1966:FBP

[IR66a] P. L. Ivănescu and S. Rudeanu. Fractional bivalent programming. Lecture Notes in Mathematics, 23:115–120, 1966. CODEN LN-MAA2. ISBN 3-540-03606-7 (print), 3-540-34835-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077208/.

Ivanescu:1966:FM

[IR66b] P. L. Ivănescu and S. Rudeanu. Front matter. Lecture Notes in Mathematics, 23:??, 1966. CODEN LNMAA2. ISBN 3-540-03606-7 (print), 3-540-34835-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-34835-1/1.pdf.

Ivanescu:1966:LPB

[IR66c] P. L. Ivănescu and S. Rudeanu. Linear pseudo-Boolean equations and inequalities. Lecture Notes in Mathematics, 23:7–56, 1966. CODEN LNMAA2. ISBN 3-540-03606-7 (print), 3-540-34835-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0077205/.

Ivanescu:1966:MPB

[IR66d] P. L. Ivănescu and S. Rudeanu. Minimization of pseudo-Boolean functions. Lecture Notes in Mathematics, 23:85–114, 1966. CO-DEN LNMAA2. ISBN 3-540-03606-7 (print), 3-540-34835-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0077207/.

Ivanescu:1966:NPB

[IR66e] P. L. Ivănescu and S. Rudeanu. Nonlinear pseudo-Boolean equations and inequalities. Lecture Notes in Mathematics, 23:57– 84, 1966. CODEN LNMAA2. ISBN 3-540-03606-7 (print), 3-540-34835-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077206/.

Ivanescu:1966:PBM

[IR66f] P. L. Ivănescu and S. Rudeanu, editors. Pseudo-Boolean Methods for Bivalent Programming: Lecture at the First European Meeting of the Institute of Management Sciences and of the Econometric Institute, Warsaw, September 2–7, 1966, volume 23 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1966. CODEN LNMAA2. ISBN 3-540-03606-7 (paperback), 3-540-34835-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 23.

URL http://link.springer.com/book/10.1007/BFb0077204.

Isbell:1967:NCC

[Isb67] John R. Isbell. Normal completions of categories. Lecture Notes in Mathematics, 47:110–155, 1967. CODEN LNMAA2. ISBN 3-540-03918-X (print), 3-540-35545-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074302/.

Ivanescu:1965:ATG

[Ivă65a] Petru L. Ivănescu. Application to the theory of graphs. Lecture Notes in Mathematics, 9:23–27, 1965. CODEN LNMAA2. ISBN 3-540-03352-1 (print), 3-540-37136-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073945/.

Ivanescu:1965:ASA

[Ivă65b] Petru L. Ivănescu. Applications to switching algebra. Lecture Notes in Mathematics, 9:32–34, 1965. CODEN LNMAA2. ISBN 3-540-03352-1 (print), 3-540-37136-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073948/.

Ivanescu:1965:ATF

[Ivă65c] Petru L. Ivănescu. Applications to the theory of flows in networks. Lecture Notes in Mathematics, 9:27–30, 1965. CODEN LNMAA2. ISBN 3-540-03352-1 (print), 3-540-37136-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073946/.

Ivanescu:1965:ATP

[Ivă65d] Petru L. Ivănescu. Applications to the transportation problem. Lecture Notes in Mathematics, 9:30–32, 1965. CODEN LN-MAA2. ISBN 3-540-03352-1 (print), 3-540-37136-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0073947/.

Ivanescu:1965:DPL

[Ivă65e] Petru L. Ivănescu. Discrete polynomial-logical programming. Lecture Notes in Mathematics, 9:20–23, 1965. CODEN LN-MAA2. ISBN 3-540-03352-1 (print), 3-540-37136-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0073944/.

Ivanescu:1965:I

[Ivă65f] Petru L. Ivănescu. Introduction. Lecture Notes in Mathematics, 9:1–2, 1965. CODEN LNMAA2. ISBN 3-540-03352-1 (print), 3-540-37136-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073939.

Ivanescu:1965:MDF

[Ivă65g] Petru L. Ivănescu. Minimal decomposition of finite partially ordered sets in chains. Lecture Notes in Mathematics, 9:34–37, 1965. CODEN LNMAA2. ISBN 3-540-03352-1 (print), 3-540-37136-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0073949/.

Ivanescu:1965:MPB

[Ivă65h] Petru L. Ivănescu. Minimization of pseudo-Boolean functions. Lecture Notes in Mathematics, 9:7–13, 1965. CODEN LNMAA2. ISBN 3-540-03352-1 (print), 3-540-37136-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073941/.

Ivanescu:1965:NT

[Ivă65i] Petru L. Ivănescu. Notations and terminology. Lecture Notes in Mathematics, 9:3–7, 1965. CODEN LNMAA2. ISBN 3-540-03352-1 (print), 3-540-37136-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073940/.

Ivanescu:1965:PBPa

[Ivă65j] Petru L. Ivănescu. Pseudo-Boolean programming. Lecture Notes in Mathematics, 9:18–20, 1965. CODEN LNMAA2. ISBN 3-540-03352-1 (print), 3-540-37136-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073943/.

Ivanescu:1965:PBPb

[Ivă65k] Petru L. Ivănescu, editor. Pseudo-Boolean Programming and Applications: Presented at the Colloquium on Mathematics and Cybernetics in the Economy, Berlin, October 1964, volume 9 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1965. CODEN LNMAA2. ISBN 3-540-03352-1 (paperback), 3-540-37136-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN

QA3 .L28 no. 9. URL http://link.springer.com/book/10.
1007/BFb0073938.

Ivanescu:1965:SPB

[Ivă65l] Petru L. Ivănescu. Systems of pseudo-Boolean equations and inequalities. Lecture Notes in Mathematics, 9:13–18, 1965. CODEN LNMAA2. ISBN 3-540-03352-1 (print), 3-540-37136-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073942/.

Iwasawa:1966:CF

[Iwa66] K. Iwasawa. Class fields. Lecture Notes in Mathematics, 21:V– 1–V–13, 1966. CODEN LNMAA2. ISBN 3-540-03604-0 (print), 3-540-34798-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091555/.

Jacobs:1967:INI

[Jac67] Konrad Jacobs. Invariant and non-invariant measures. Lecture Notes in Mathematics, 31:118–135, 1967. CODEN LNMAA2. ISBN 3-540-03902-3 (print), 3-540-34970-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061113/.

Janich:1968:BSM

[Jän68a] Klaus Jänich. Beispiele spezieller G-Mannigfaltigkeiten. (German) [Examples of specific G-manifolds]. Lecture Notes in Mathematics, 59:36–43, 1968. CODEN LNMAA2. ISBN 3-540-04229-6 (print), 3-540-35861-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098478/.

Janich:1968:BKM

[Jän68b] Klaus Jänich. Bericht über Knoten-Mannigfaltigkeiten. (German) [Report on knot manifolds]. Lecture Notes in Mathematics, 59:43–47, 1968. CODEN LNMAA2. ISBN 3-540-04229-6 (print), 3-540-35861-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098479/.

Janich:1968:EGE

[Jän68c] Klaus Jänich. Der Einbettungssatz. (German) [The embedding theorem]. Lecture Notes in Mathematics, 59:13–20, 1968. CO-DEN LNMAA2. ISBN 3-540-04229-6 (print), 3-540-35861-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0098475/.

Janich:1968:SES

[Jän68d] Klaus Jänich. Der Symmetriegrad der exotischen Sphären (Bericht über Resultate von Wu-chung Hsiang und Wu-yi Hsiang). (German) [The degree of symmetry of exotic spheres (report on results of Wu-chung Hsiang and Wu-yi Hsiang)]. Lecture Notes in Mathematics, 59:72–84, 1968. CODEN LNMAA2. ISBN 3-540-04229-6 (print), 3-540-35861-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098483/.

Janich:1968:GGB

[Jän68e] Klaus Jänich. Die Grundbegriffe. (German) [The basic concepts]. Lecture Notes in Mathematics, 59:1–9, 1968. CODEN LNMAA2. ISBN 3-540-04229-6 (print), 3-540-35861-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098473/.

Janich:1968:DMG

[Jän68f] Klaus Jänich. Differenzierbare G-Mannigfaltigkeiten. (German) [Differentiable G-manifolds], volume 59 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. CODEN LNMAA2. ISBN 3-540-04229-6 (paperback), 3-540-35861-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 89 pp. LCCN QA3 .L28 no. 59. URL

http://link.springer.com/book/10.1007/BFb0098472.

Janich:1968:HSO

[Jän68g] Klaus Jänich. Hauptorbits, singuläre Orbits und Ausnahmeorbits. (German) [Main orbits, singular orbits and exceptional orbits]. Lecture Notes in Mathematics, 59:9–13, 1968. CODEN LNMAA2. ISBN 3-540-04229-6 (print), 3-540-35861-7 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098474/.

Janich:1968:KSM

[Jän68h] Klaus Jänich. Klassifikation der “speziellen” G-Mannigfaltigkeiten. (German) [Classification of “special” G-manifolds]. Lecture Notes in Mathematics, 59:26–35, 1968. CODEN LNMAA2. ISBN 3-540-04229-6 (print), 3-540-35861-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098477/.

Janich:1968:LDT

[Jän68i] Klaus Jänich. Lücken in den Dimensionen der Transformationsgruppen (nach L. N. Mann). (German) [Gaps in the dimensions of the transformation groups (after L. N. Mann)]. Lecture Notes in Mathematics, 59:66–72, 1968. CODEN LNMAA2. ISBN 3-540-04229-6 (print), 3-540-35861-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098482/.

Janich:1968:MLS

[Jän68j] Klaus Jänich. Musterbeispiel eines Linearitätsbeweises: Der Satz von Montgomery, Samelson, Yang und Zippin über Aktionen auf IR^{n} mit zweidimensionalem Orbitraum. (German) [Paradigm of linearity proof: The set of Montgomery, Samelson, Yang and Zippin on actions on rmIR^{n} with a two-dimensional orbit space]. Lecture Notes in Mathematics, 59:48–59, 1968. CODEN LNMAA2. ISBN 3-540-04229-6 (print), 3-540-35861-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098480/.

Janich:1968:SGD

[Jän68k] Klaus Jänich. Scheibendiagramme. (German) [Disk graphs]. Lecture Notes in Mathematics, 59:21–25, 1968. CODEN LN-MAA2. ISBN 3-540-04229-6 (print), 3-540-35861-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0098476/.

Janich:1968:WLB

[Jän68l] Klaus Jänich. Weitere Linearitätssätze und Beispiele nichtlinearer Aktionen. (German) [More linearity sets and examples of nonlinear actions]. Lecture Notes in Mathematics, 59:60–65, 1968. CODEN LNMAA2. ISBN 3-540-04229-6 (print), 3-540-35861-7 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098481/.

Janenko:1969:AAG

[Jan69a] N. N. Janenko. Allgemeine Aussagen. (German) [General statements]. Lecture Notes in Mathematics, 91:142–161, 1969. CODEN LNMAA2. ISBN 3-540-04610-0 (print), 3-540-36100-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098299/.

Janenko:1969:AZR

[Jan69b] N. N. Janenko. Anwendung der Zwischenschrittmethode auf Randwertaufgaben der Laplaceschen und Poissonschen Differentialgleichung. (German) [Application of the intermediate-step method to boundary value problems of the Laplace and Poisson differential equation]. Lecture Notes in Mathematics, 91:59–97, 1969. CODEN LNMAA2. ISBN 3-540-04610-0 (print), 3-540-36100-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0098294/.

Janenko:1969:BM

[Jan69c] N. N. Janenko. Back matter. Lecture Notes in Mathematics, 91:184–197, 1969. CODEN LNMAA2. ISBN 3-540-04610-0 (print), 3-540-36100-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-36100-8/1.pdf.

Janenko:1969:AZH

[Jan69d] N. N. Janenko. Die Anwendung der Zwischenschrittmethode auf hyperbolische Gleichungen. (German) [The application of the intermediate-step method to hyperbolic equations]. Lecture Notes in Mathematics, 91:45–59, 1969. CODEN LNMAA2. ISBN 3-540-04610-0 (print), 3-540-36100-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098293/.

Janenko:1969:ESZ

[Jan69e] N. N. Janenko. Die einfachsten Schemata mit Zwischenschritten zur Integration parabolischer Gleichungen. (German) [The simplest schemes with intermediate steps for integration of parabolic equations]. Lecture Notes in Mathematics, 91:16–45, 1969. CODEN LNMAA2. ISBN 3-540-04610-0 (print), 3-540-36100-6 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098292/.

Janenko:1969:MSA

[Jan69f] N. N. Janenko. Die Methode der schwachen Approximation und die Konstruktion einer Lösung des Cauchyschen Anfangswertproblems im Banachraum. (German) [The method of weak approximation and the construction of a solution of the Cauchy initial value problem in Banach space]. Lecture Notes in Mathematics, 91:161–183, 1969. CODEN LNMAA2. ISBN 3-540-04610-0 (print), 3-540-36100-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098300/.

Janenko:1969:EHA

[Jan69g] N. N. Janenko. Einige hydrodynamische Aufgaben. (German) [Some hydrodynamic problems]. Lecture Notes in Mathematics, 91:123–142, 1969. CODEN LNMAA2. ISBN 3-540-04610-0 (print), 3-540-36100-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098298/.

Janenko:1969:FM

[Jan69h] N. N. Janenko. Front matter. Lecture Notes in Mathematics, 91:N2–VIII, 1969. CODEN LNMAA2. ISBN 3-540-04610-0 (print), 3-540-36100-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-36100-8/1.pdf.

Janenko:1969:HSG

[Jan69i] N. N. Janenko. Homogene Schemata. (German) [Homogeneous schemata]. Lecture Notes in Mathematics, 91:1–16, 1969. CO-DEN LNMAA2. ISBN 3-540-04610-0 (print), 3-540-36100-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0098291/.

Janenko:1969:IIA

[Jan69j] N. N. Janenko. Integrodifferentialgleichungen, Integralgleichungen und algebraische Gleichungen. (German) [ntegrodifferential, integral equations, and algebraic equations]. Lecture Notes in Mathematics, 91:119–123, 1969. CODEN LNMAA2. ISBN 3-540-04610-0 (print), 3-540-36100-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098297/.

Janenko:1969:REG

[Jan69k] N. N. Janenko. Randwertaufgaben der Elastizitätstheorie. (German) [Boundary value problems of elasticity theory]. Lecture Notes in Mathematics, 91:97–110, 1969. CODEN LNMAA2. ISBN 3-540-04610-0 (print), 3-540-36100-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098295/.

Janenko:1969:SEG

[Jan69l] N. N. Janenko. Schemata erhöhter Genauigkeit. (German) [Schemes of increased accuracy]. Lecture Notes in Mathematics, 91:110–119, 1969. CODEN LNMAA2. ISBN 3-540-04610-0 (print), 3-540-36100-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098296/.

Janenko:1969:ZLM

[Jan69m] N. N. (Nikolaĭ Nikolaevich) Janenko. Die Zwischenschrittmethode zur Lösung mehrdimensionaler Probleme der mathematischen Physik. (German) [The intermediate step method for the solution of multidimensional problems of mathematical physics], volume 91 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CO-DEN LNMAA2. ISBN 3-540-04610-0 (paperback), 3-540-36100-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 194 pp. LCCN QA3 .L28 no. 91. URL http://link.springer.

com/book/10.1007/BFb0098290. Translation from Russian by K.

Roesner.

Jensen:1967:ADG

[Jen67a] Ronald Björn Jensen. Absolutheit und Definierbarkeit. (German) [Absoluteness and definability]. Lecture Notes in Mathematics, 37:24–44, 1967. CODEN LNMAA2. ISBN 3-540-03908-2 (print), 3-540-34976-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073885/.

Jensen:1967:WMG

[Jen67b] Ronald Björn Jensen. B-wertige Modelle. (German) [B-valued models]. Lecture Notes in Mathematics, 37:104–110, 1967. CO-DEN LNMAA2. ISBN 3-540-03908-2 (print), 3-540-34976-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0073891/.

Jensen:1967:KMG

[Jen67c] Ronald Björn Jensen. Das konstruktible Modell. (German) [The constructible model]. Lecture Notes in Mathematics, 37:67– 82, 1967. CODEN LNMAA2. ISBN 3-540-03908-2 (print), 3-540-34976-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073887/.

Jensen:1967:SFS

[Jen67d] Ronald Björn Jensen. Das System ZF Die Formale Sprache. (German) [The ZF system: The formal language]. Lecture Notes in Mathematics, 37:1–15, 1967. CODEN LNMAA2. ISBN 3-540-03908-2 (print), 3-540-34976-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073882/.

Jensen:1967:CEG

[Jen67e] Ronald Björn Jensen. Die Cohensche Erzwingungsmethode. (German) [The Cohen enforcement method]. Lecture Notes in Mathematics, 37:83–96, 1967. CODEN LNMAA2. ISBN 3-540-03908-2 (print), 3-540-34976-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073888/.

Jensen:1967:EBW

[Jen67f] Ronald Björn Jensen. Die Erzwingungsbeziehung als Booleschwertige Wahrheitsdefinition. (German) [The enforcement relationship as Boolean valued truth definition]. Lecture Notes in Mathematics, 37:100–103, 1967. CODEN LNMAA2. ISBN 3-540-03908-2 (print), 3-540-34976-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073890/.

Jensen:1967:MGMa

[Jen67g] Ronald Björn Jensen. Die Metasprache. (German) [The metalanguage]. Lecture Notes in Mathematics, 37:16–18, 1967. CODEN LNMAA2. ISBN 3-540-03908-2 (print), 3-540-34976-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073883/.

Jensen:1967:EMG

[Jen67h] Ronald Björn Jensen. Einbettungssatz für ZF -Modelle. (German) [Embedding theorem for ZF models]. Lecture Notes in Mathematics, 37:155–158, 1967. CODEN LNMAA2. ISBN 3-540-03908-2 (print), 3-540-34976-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073896/.

Jensen:1967:GMG

[Jen67i] Ronald Björn Jensen. Generische Modelle. (German) [Generic models]. Lecture Notes in Mathematics, 37:111–117, 1967. CO-DEN LNMAA2. ISBN 3-540-03908-2 (print), 3-540-34976-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0073892/.

Jensen:1967:IMG

[Jen67j] Ronald Björn Jensen. Innere Modelle. (German) [Inner models]. Lecture Notes in Mathematics, 37:45–66, 1967. CODEN LNMAA2. ISBN 3-540-03908-2 (print), 3-540-34976-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073886/.

Jensen:1967:LFA

[Jen67k] Ronald Björn Jensen. Lösung der Frage von Addison. (German) [Solution to the question of Addison]. Lecture Notes in Mathematics, 37:97–99, 1967. CODEN LNMAA2. ISBN 3-540-03908-2 (print), 3-540-34976-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073889/.

Jensen:1967:MII

[Jen67l] Ronald Björn Jensen. Maximale innere ib-modelle. (German) [Maximal internal IB models]. Lecture Notes in Mathematics, 37: 118–133, 1967. CODEN LNMAA2. ISBN 3-540-03908-2 (print), 3-540-34976-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073893/.

Jensen:1967:MMW

[Jen67m] Ronald Björn Jensen. Modelle der Mengenlehre: Widerspruchsfreiheit und Unabhängigkeit der Kontinuum-Hypothese und des Auswahlaxioms. (German) [Models of set theory: consistency and independence of the Continuum Hypothesis and the Axiom of Choice], volume 37 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. CODEN LNMAA2. ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 176 pp. LCCN QA3 .L28 no. 37. URL http:/

/link.springer.com/book/10.1007/BFb0073881. Further devel-

oped by Franz Josef Leven.

Jensen:1967:MGMb

[Jen67n] Ronald Björn Jensen. Modelle. (German) [Models]. Lecture Notes in Mathematics, 37:19–23, 1967. CODEN LNMAA2. ISBN 3-540-03908-2 (print), 3-540-34976-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073884/.

Jensen:1967:UKH

[Jen67o] Ronald Björn Jensen. Unabhängigkeit der Kontinuum-Hypothese. (German) [Independence of the continuum hypothesis]. Lecture Notes in Mathematics, 37:145–154, 1967. CODEN LNMAA2. ISBN 3-540-03908-2 (print), 3-540-34976-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073895/.

Jensen:1967:UAG

[Jen67p] Ronald Björn Jensen. Unabhängigkeit des Auswahlaxioms. (German) [Independence of the axiom of choice]. Lecture Notes in Mathematics, 37:159–167, 1967. CODEN LNMAA2. ISBN 3-540-03908-2 (print), 3-540-34976-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073897/.

Jensen:1967:UGI

[Jen67q] Ronald Björn Jensen. Unabhängigkeit von V = L. (German) [Independence of V = L]. Lecture Notes in Mathematics, 37:134– 144, 1967. CODEN LNMAA2. ISBN 3-540-03908-2 (print), 3-540-34976-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073894/.

Jones:1968:ECP

[Jon68a] G. Stephen Jones. The existence of critical points in generalized dynamical systems. Lecture Notes in Mathematics, 60:7–19, 1968. CODEN LNMAA2. ISBN 3-540-04230-X (print), 3-540-35862-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077122/.

Jones:1968:SDE

[Jon68b] G. Stephen Jones, editor. Seminar on Differential Equations and Dynamical Systems, volume 60 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. CODEN LNMAA2. ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 106 pp. LCCN QA3 .L28 no. 60. URL

http://link.springer.com/book/10.1007/BFb0077120.

Kac:1969:SMP

[Kac69] M. Kac. Some mathematical problems in statistical mechanics. Lecture Notes in Mathematics, 89:106–124, 1969. CODEN LN-MAA2. ISBN 3-540-04608-9 (print), 3-540-36098-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079121/.

Kappos:1967:RAS

[Kap67a] Demetrios A. Kappos. Representation of abstract L-spaces. Lecture Notes in Mathematics, 31:136–145, 1967. CODEN LNMAA2. ISBN 3-540-03902-3 (print), 3-540-34970-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061114/.

Kappos:1967:SPM

[Kap67b] Dēmētrios Andreou Kappos, editor. Symposium on Probability Methods in Analysis, volume 31 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. CODEN LNMAA2. ISBN 3-540-03902-3 (paperback), 3-540-34970-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 31. URL http://

link.springer.com/book/10.1007/BFb0061101.

Karp:1968:APB

[Kar68] Carol Karp. An algebraic proof of the barwise compactness theorem. Lecture Notes in Mathematics, 72:80–95, 1968. CODEN LNMAA2. ISBN 3-540-04242-3 (print), 3-540-35900-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079684/.

Karoubi:1969:FDT

[Kar69] Max Karoubi. Foncteurs dèrivès et K-thèorie. (French) [Derived functors and K-theory]. Lecture Notes in Mathematics, 99:365– 383, 1969. CODEN LNMAA2. ISBN 3-540-04618-6 (print), 3-540-36140-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081967/.

Kato:1968:AE

[Kat68a] Junji Kato. Asymptotic equivalence. Lecture Notes in Mathematics, 60:26–30, 1968. CODEN LNMAA2. ISBN 3-540-04230-X (print), 3-540-35862-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077124/.

Kato:1968:FML

[Kat68b] Junji Kato. Fundamental matrix in linear functional differential equations. Lecture Notes in Mathematics, 60:55–64, 1968. CO-DEN LNMAA2. ISBN 3-540-04230-X (print), 3-540-35862-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0077128/.

Kato:1968:RRS

[Kat68c] Junji Kato. A remark on a result of Strauss. Lecture Notes in Mathematics, 60:83–91, 1968. CODEN LNMAA2. ISBN 3-540-04230-X (print), 3-540-35862-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077131/.

Keisler:1968:FLO

[Kei68] H. J. Keisler. Formulas with linearly ordered quantifiers. Lecture Notes in Mathematics, 72:96–130, 1968. CODEN LNMAA2. ISBN 3-540-04242-3 (print), 3-540-35900-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079685/.

Kellerer:1967:ESP

[Kel67] Hans G. Kellerer. Extension of stationary processes. Lecture Notes in Mathematics, 31:146, 1967. CODEN LNMAA2. ISBN 3-540-03902-3 (print), 3-540-34970-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061115.

Kelly:1969:AEC

[Kel69] G. M. Kelly. Adjunction for enriched categories. Lecture Notes in Mathematics, 106:166–177, 1969. CODEN LNMAA2. ISBN 3-540-04625-9 (print), 3-540-36150-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059145/.

Kemperman:1969:ORT

[Kem69] J. H. B. Kemperman. On the optimum rate of transmitting information. Lecture Notes in Mathematics, 89:126–169, 1969. CODEN LNMAA2. ISBN 3-540-04608-9 (print), 3-540-36098-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079123/.

Kendall:1967:RST

[Ken67] David. G. Kendall. Renewal sequences and their arithmetic. Lecture Notes in Mathematics, 31:147–175, 1967. CODEN LN-MAA2. ISBN 3-540-03902-3 (print), 3-540-34970-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0061116/.

Kennison:1969:CMW

[Ken69] J. F. Kennison. Coreflection maps which resemble universal coverings. Lecture Notes in Mathematics, 86:46–75, 1969. CODEN LNMAA2. ISBN 3-540-04605-4 (print), 3-540-36095-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079384/.

Kervaire:1967:TBM

[Ker67] M. Kervaire. Le théorème de Barden–Mazur–Stallings. (French) [The Barden–Mazur–Stallings theorem]. Lecture Notes in Mathematics, 48:83–95, 1967. CODEN LNMAA2. ISBN 3-540-03919-8 (print), 3-540-35547-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073935/.

Kiselman:1968:SFE

[Kis68] C. O. Kiselman. Supports des fonctionnelles sur un espace de solutions d’une équation aux dérivées partielles à coefficients constants. (French) [Supports of functionals on a space of solutions of a partial differential equations with constant coefficients]. Lecture Notes in Mathematics, 71:118–126, 1968. CODEN LNMAA2. ISBN 3-540-04241-5 (print), 3-540-35889-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080103/.

Kambo:1969:ABA

[KK69] N. S. Kambo and S. Kotz. Asymptotic behaviour of the average probability of error for low rates of information transmission. Lecture Notes in Mathematics, 89:125, 1969. CODEN LNMAA2. ISBN 3-540-04608-9 (print), 3-540-36098-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079122.

Kopperman:1968:SPG

[KM68] R. D. Kopperman and A. R. D. Mathias. Some problems in group theory. Lecture Notes in Mathematics, 72:131–138, 1968. CODEN LNMAA2. ISBN 3-540-04242-3 (print), 3-540-35900-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079686/.

Kristensen:1967:FTI

[KMP67] Povl Kristensen, Lars Mejlbo, and Ebbe Thue Poulsen. On a Fourier transform in infinitely many dimensions. Lecture Notes in Mathematics, 31:187–196, 1967. CODEN LNMAA2. ISBN 3-540-03902-3 (print), 3-540-34970-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061118/.

Kneser:1969:NSI

[Kne69] Martin Kneser. Normal subgroups of integral orthogonal groups. Lecture Notes in Mathematics, 108:67–71, 1969. CODEN LN-MAA2. ISBN 3-540-04627-5 (print), 3-540-36156-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0059991/.

Knighten:1968:ACF

[Kni68] R. L. Knighten. An application of categories of fractions to homotopy theory. Lecture Notes in Mathematics, 61:62–68, 1968. CODEN LNMAA2. ISBN 3-540-04231-8 (print), 3-540-35863-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077117/.

Knus:1969:AGG

[Knu69] Max A. Knus. Algebras graded by a group. Lecture Notes in Mathematics, 92:117–133, 1969. CODEN LNMAA2. ISBN 3-540-04611-9 (print), 3-540-36101-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080768/.

Kohn:1968:HRB

[Köh68] Johannes Köhn. Harmonische Räume mit einer Basis semiregulärer Mengen. (German) [Harmonic space with a basis of semiregular measures]. Lecture Notes in Mathematics, 69:1–12, 1968. CODEN LNMAA2. ISBN 3-540-04239-3 (print), 3-540-35884-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080474/.

Kolzow:1968:BM

[Köl68a] Dietrich Kölzow. Back matter. Lecture Notes in Mathematics, 65:99–104, 1968. CODEN LNMAA2. ISBN 3-540-04235-0 (print), 3-540-35875-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-35875-6/1.pdf.

Kolzow:1968:BGG

[Köl68b] Dietrich Kölzow. Bezeichnungen und Grundbegriffe. (German) [Terms and basic concepts]. Lecture Notes in Mathematics, 65: 1–2, 1968. CODEN LNMAA2. ISBN 3-540-04235-0 (print), 3-540-35875-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077211.

Kolzow:1968:DLG

[Köl68c] Dietrich Kölzow. Differentiation und Liftings. (German) [Differentiation and liftings]. Lecture Notes in Mathematics, 65:3–64, 1968. CODEN LNMAA2. ISBN 3-540-04235-0 (print), 3-540-35875-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0077212/.

Kolzow:1968:DMH

[Köl68d] Dietrich Kölzow. Differentiation von Maßen: Habilitationsschrift, Saarbrücken 1967. (German) [Differentiation of dimensions: Habilitation thesis, Saarbrücken 1967], volume 65 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. CODEN LNMAA2. ISBN 3-540-04235-0 (paperback), 3-540-35875-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xi + 102 pp. LCCN QA3

.L28 no. 65. URL http://link.springer.com/book/10.1007/
BFb0077210.

Kolzow:1968:FM

[Köl68e] Dietrich Kölzow. Front matter. Lecture Notes in Mathematics, 65:I–XII, 1968. CODEN LNMAA2. ISBN 3-540-04235-0 (print), 3-540-35875-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-35875-6/1.pdf.

Kolzow:1968:SGS

[Köl68f] Dietrich Kölzow. Sonderfälle. (German) [Special cases]. Lecture Notes in Mathematics, 65:70–98, 1968. CODEN LNMAA2. ISBN 3-540-04235-0 (print), 3-540-35875-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077214/.

Kolzow:1968:VAG

[Köl68g] Dietrich Kölzow. Verschärfungen des Ableitungsbegriffs. (German) [Tightening of the derivative term]. Lecture Notes in Mathematics, 65:65–69, 1968. CODEN LNMAA2. ISBN 3-540-04235-0 (print), 3-540-35875-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077213/.

Kounias:1967:OBC

[Kou67] Eustratios Kounias. Optimal bounded control with linear stochastic equations and quadratic cost. Lecture Notes in Mathematics, 31:176–186, 1967. CODEN LNMAA2. ISBN 3-540-03902-3 (print), 3-540-34970-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061117/.

Kreisel:1968:CIL

[Kre68] G. Kreisel. Choice of infinitary languages by means of definability criteria; generalized recursion theory. Lecture Notes in Mathematics, 72:139–151, 1968. CODEN LNMAA2. ISBN 3-540-04242-3 (print), 3-540-35900-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079687/.

Krengel:1969:NSC

[Kre69] Ulrich Krengel. A necessary and sufficient condition for the validity of the local ergodic theorem. Lecture Notes in Mathematics, 89:170–177, 1969. CODEN LNMAA2. ISBN 3-540-04608-9 (print), 3-540-36098-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079124/.

Krickeberg:1969:RRM

[Kri69] K. Krickeberg. Recent results on mixing in topological measure spaces. Lecture Notes in Mathematics, 89:178–185, 1969. CODEN LNMAA2. ISBN 3-540-04608-9 (print), 3-540-36098-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079125/.

Kronk:1969:VTP

[Kro69] Hudson V. Kronk. Variations on a theorem of Pósa. Lecture Notes in Mathematics, 110:193–197, 1969. CODEN LNMAA2. ISBN 3-540-04629-1 (print), 3-540-36161-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060117/.

Kunzi:1966:DGD

[KT66a] H. P. Künzi and S. T. Tan. Dekompositionsmethoden. (German) [Decomposition methods]. Lecture Notes in Mathematics, 27:71–101, 1966. CODEN LNMAA2. ISBN 3-540-03609-1 (print), 3-540-34853-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071087/.

Kunzi:1966:RSD

[KT66b] H. P. Künzi and S. T. Tan. Die revidierten Simplexverfahren und das duale Simplexverfahren. (German) [The revised simplex method and the dual simplex method]. Lecture Notes in Mathematics, 27:24–42, 1966. CODEN LNMAA2. ISBN 3-540-03609-1 (print), 3-540-34853-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071085/.

Kunzi:1966:MGO

[KT66c] H. P. Künzi and S. T. Tan. Mathematische Grundlagen zur Optimierungstheorie. (German) [Mathematical foundation of optimization theory]. Lecture Notes in Mathematics, 27:1–23, 1966. CO-DEN LNMAA2. ISBN 3-540-03609-1 (print), 3-540-34853-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0071084/.

Kunzi:1966:MDG

[KT66d] H. P. Künzi and S. T. Tan. Mehrphasen- und Duoplexmethode. (German) [Multiphase and duoplex method]. Lecture Notes in Mathematics, 27:43–70, 1966. CODEN LNMAA2. ISBN 3-540-03609-1 (print), 3-540-34853-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071086/.

Kunzi:1966:LOG

[KT66e] Hans Paul Künzi and Soei Tien Tan. Lineare Optimierung großer Systeme. (German) [Nonlinear Optimization of Large Systems], volume 27 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1966. CODEN LNMAA2. ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 121 pp. LCCN QA3 .L28 no. 27. URL http://link.

springer.com/book/10.1007/BFb0071083.

Kamber:1968:CCF

[KT68a] Franz Kamber and Philippe Tondeur. Characteristic classes of flat bundles. Lecture Notes in Mathematics, 67:14–24, 1968. CODEN LNMAA2. ISBN 3-540-04237-7 (print), 3-540-35879-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076914/.

Kamber:1968:CCE

[KT68b] Franz Kamber and Philippe Tondeur. The Chern character of elliptic symbols associated to a flat G-structure. Lecture Notes in Mathematics, 67:42–45, 1968. CODEN LNMAA2. ISBN 3-540-04237-7 (print), 3-540-35879-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076917/.

Kamber:1968:FB

[KT68c] Franz Kamber and Philippe Tondeur. Flat bundles. Lecture Notes in Mathematics, 67:9–13, 1968. CODEN LNMAA2. ISBN 3-540-04237-7 (print), 3-540-35879-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076913/.

Kamber:1968:FSM

[KT68d] Franz Kamber and Philippe Tondeur. Flat G-structures on manifolds and the Euler characteristic. Lecture Notes in Mathematics, 67:37–42, 1968. CODEN LNMAA2. ISBN 3-540-04237-7 (print), 3-540-35879-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076916/.

Kamber:1968:FMa

[KT68e] Franz Kamber and Philippe Tondeur. Flat manifolds. Lecture Notes in Mathematics, 67:2–4, 1968. CODEN LNMAA2. ISBN 3-540-04237-7 (print), 3-540-35879-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076911/.

Kamber:1968:FMb

[KT68f] Franz Kamber and Philippe Tondeur. Flat Manifolds, volume 67 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. CO-DEN LNMAA2. ISBN 3-540-04237-7 (paperback), 3-540-35879-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 53

pp. LCCN QA3 .L28 no. 67. URL http://link.springer.com/
book/10.1007/BFb0076909.

Kamber:1968:FMP

[KT68g] Franz Kamber and Philippe Tondeur. Flat manifolds with parallel torsion. Lecture Notes in Mathematics, 67:4–9, 1968. CODEN LNMAA2. ISBN 3-540-04237-7 (print), 3-540-35879-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076912/.

Kamber:1968:IEC

[KT68h] Franz Kamber and Philippe Tondeur. The index of elliptic complexes associated to a flat G-structure. Lecture Notes in Mathematics, 67:45–47, 1968. CODEN LNMAA2. ISBN 3-540-04237-7 (print), 3-540-35879-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076918/.

Kamber:1968:I

[KT68i] Franz Kamber and Philippe Tondeur. Introduction. Lecture Notes in Mathematics, 67:1, 1968. CODEN LNMAA2. ISBN 3-540-04237-7 (print), 3-540-35879-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076910.

Kamber:1968:P

[KT68j] Franz Kamber and Philippe Tondeur. Problems. Lecture Notes in Mathematics, 67:47, 1968. CODEN LNMAA2. ISBN 3-540-04237-7 (print), 3-540-35879-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076919.

Kamber:1968:SFV

[KT68k] Franz Kamber and Philippe Tondeur. Stably flat vector bundles on complexes covered by a homotopy sphere. Lecture Notes in Mathematics, 67:24–37, 1968. CODEN LNMAA2. ISBN 3-540-04237-7 (print), 3-540-35879-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076915/.

Kueker:1968:DAI

[Kue68] David W. Kueker. Definability, automorphisms, and infinitary languages. Lecture Notes in Mathematics, 72:152–165, 1968. CODEN LNMAA2. ISBN 3-540-04242-3 (print), 3-540-35900-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079688/.

Kuramochi:1968:BFT

[Kur68] Zenjiro Kuramochi. On Beurling’s and Fatou’s theorems. Lecture Notes in Mathematics, 58:43–69, 1968. CODEN LNMAA2. ISBN 3-540-04228-8 (print), 3-540-35859-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098445/.

Lambek:1966:CWS

[Lam66a] Joachim Lambek. The completion when A is not small. Lecture Notes in Mathematics, 24:27–34, 1966. CODEN LNMAA2. ISBN 3-540-03607-5 (print), 3-540-34840-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077271/.

Lambek:1966:CCA

[Lam66b] Joachim Lambek. Completions of categories of algebras. Lecture Notes in Mathematics, 24:55–57, 1966. CODEN LNMAA2. ISBN 3-540-03607-5 (print), 3-540-34840-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077275/.

Lambek:1966:CCM

[Lam66c] Joachim Lambek. Completions of categories of modules. Lecture Notes in Mathematics, 24:58–68, 1966. CODEN LNMAA2. ISBN 3-540-03607-5 (print), 3-540-34840-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077276/.

Lambek:1966:CCS

[Lam66d] Joachim Lambek, editor. Completions of Categories: Seminar lectures given 1966 in Zürich, volume 24 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1966. CODEN LNMAA2. ISBN 3-540-03607-5 (paperback), 3-540-34840-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 24. URL http://

link.springer.com/book/10.1007/BFb0077265.

Lambek:1966:CG

[Lam66e] Joachim Lambek. Completions of groups. Lecture Notes in Mathematics, 24:51–54, 1966. CODEN LNMAA2. ISBN 3-540-03607-5 (print), 3-540-34840-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077274/.

Lambek:1966:GSD

[Lam66f] Joachim Lambek. Generating and sup-dense subcategories. Lecture Notes in Mathematics, 24:10–16, 1966. CODEN LNMAA2. ISBN 3-540-03607-5 (print), 3-540-34840-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077268/.

Lambek:1966:I

[Lam66g] Joachim Lambek. Introduction. Lecture Notes in Mathematics, 24:2–5, 1966. CODEN LNMAA2. ISBN 3-540-03607-5 (print), 3-540-34840-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077266/.

Lambek:1966:LPF

[Lam66h] Joachim Lambek. Limit preserving functors. Lecture Notes in Mathematics, 24:17–23, 1966. CODEN LNMAA2. ISBN 3-540-03607-5 (print), 3-540-34840-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077269/.

Lambek:1966:RBD

[Lam66i] Joachim Lambek. The relationship between different forms of completeness. Lecture Notes in Mathematics, 24:35–41, 1966. CODEN LNMAA2. ISBN 3-540-03607-5 (print), 3-540-34840-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077272/.

Lambek:1966:SCS

[Lam66j] Joachim Lambek. A sup-complete sup-dense, sup-preserving extension. Lecture Notes in Mathematics, 24:24–26, 1966. CODEN LNMAA2. ISBN 3-540-03607-5 (print), 3-540-34840-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077270/.

Lambek:1966:T

[Lam66k] Joachim Lambek. Terminology. Lecture Notes in Mathematics, 24:6–9, 1966. CODEN LNMAA2. ISBN 3-540-03607-5 (print), 3-540-34840-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077267/.

Lambek:1966:TPC

[Lam66l] Joachim Lambek. Theorems without properness conditions. Lecture Notes in Mathematics, 24:42–50, 1966. CODEN LN-MAA2. ISBN 3-540-03607-5 (print), 3-540-34840-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077273/.

Lambek:1969:DSC

[Lam69] Joachim Lambek. Deductive systems and categories II. Standard constructions and closed categories. Lecture Notes in Mathematics, 86:76–122, 1969. CODEN LNMAA2. ISBN 3-540-04605-4 (print), 3-540-36095-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079385/.

Larsen:1969:GTM

[Lar69a] Ronald Larsen. The general theory of multipliers. Lecture Notes in Mathematics, 105:19–105, 1969. CODEN LNMAA2. ISBN 3-540-04624-0 (print), 3-540-36149-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059713/.

Larsen:1969:MP

[Lar69b] Ronald Larsen. The Multiplier Problem, volume 105 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CODEN LNMAA2. ISBN 3-540-04624-0 (paperback), 3-540-36149-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 284 pp. LCCN

QA3 .L28 no. 105. URL http://link.springer.com/book/10.
1007/BFb0059711.

Larsen:1969:M

[Lar69c] Ronald Larsen. The multipliers for L_{p}(G). Lecture Notes in Mathematics, 105:141–170, 1969. CODEN LNMAA2. ISBN 3-540-04624-0 (print), 3-540-36149-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059716/.

Larsen:1969:MCH

[Lar69d] Ronald Larsen. The multipliers for commutative h*-algebras. Lecture Notes in Mathematics, 105:106–116, 1969. CODEN LN-MAA2. ISBN 3-540-04624-0 (print), 3-540-36149-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0059714/.

Larsen:1969:MFF

[Lar69e] Ronald Larsen. The multipliers for functions with Fourier transforms in L_{p}(Ĝ). Lecture Notes in Mathematics, 105:234–269, 1969. CODEN LNMAA2. ISBN 3-540-04624-0 (print), 3-540-36149-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0059718/.

Larsen:1969:MPG

[Lar69f] Ronald Larsen. The multipliers for the pair (L_{p}(G), L_{q}(G))(1 ≤ p, q ≤ ∞). Lecture Notes in Mathematics, 105:171–233, 1969. CO-DEN LNMAA2. ISBN 3-540-04624-0 (print), 3-540-36149-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0059717/.

Larsen:1969:MTL

[Lar69g] Ronald Larsen. Multipliers for topological linear spaces of functions and measures. Lecture Notes in Mathematics, 105:117–140, 1969. CODEN LNMAA2. ISBN 3-540-04624-0 (print), 3-540-36149-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0059715/.

Larsen:1969:PMG

[Lar69h] Ronald Larsen. Prologue: The multipliers for L_{1}(G). Lecture Notes in Mathematics, 105:1–18, 1969. CODEN LNMAA2. ISBN 3-540-04624-0 (print), 3-540-36149-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059712/.

Lawvere:1968:SAP

[Law68] F. William Lawvere. Some algebraic problems in the context of functorial semantics of algebraic theories. Lecture Notes in Mathematics, 61:41–61, 1968. CODEN LNMAA2. ISBN 3-540-04231-8 (print), 3-540-35863-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077116/.

Lawvere:1969:DAC

[Law69a] F. William Lawvere. Diagonal arguments and Cartesian closed categories. Lecture Notes in Mathematics, 92:134–145, 1969. CO-DEN LNMAA2. ISBN 3-540-04611-9 (print), 3-540-36101-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0080769/.

Lawvere:1969:OSE

[Law69b] F. William Lawvere. Ordinal sums and equational doctrines. Lecture Notes in Mathematics, 80:141–155, 1969. CODEN LN-MAA2. ISBN 3-540-04601-1 (print), 3-540-36091-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0083085/.

Lehto:1969:HGD

[Leh69] Olli Lehto. Homeomorphisms with a given dilatation. Lecture Notes in Mathematics, 118:58–73, 1969. CODEN LNMAA2. ISBN 3-540-04907-X (print), 3-540-36246-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060252/.

Lelong:1968:FPE

[Lel68a] Pierre Lelong. Fonctions plurisousharmoniques dans les espaces vectoriels topologiques. (French) [Plurisubharmonic functions in topological vector spaces]. Lecture Notes in Mathematics, 71:167– 190, 1968. CODEN LNMAA2. ISBN 3-540-04241-5 (print), 3-540-35889-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080107/.

Lelong:1968:SPL

[Lel68b] Pierre Lelong, editor. Séminaire Pierre Lelong (Analyse) Année 1967–68, volume 71 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. CODEN LNMAA2. ISBN 3-540-04241-5 (paperback), 3-540-35889-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 71. URL http://link.springer.

com/book/10.1007/BFb0080091.

Lal:1969:NST

[LG69] Mohan Lal and Paul Gillard. Numerical solution of two differentialdifference equations of analytic theory of numbers. Lecture Notes in Mathematics, 109:179–187, 1969. CODEN LNMAA2. ISBN 3-540-04628-3 (print), 3-540-36158-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060026/.

Lick:1969:CMC

[Lic69] R. Don Lick. Critically and minimally n-connected graphs. Lecture Notes in Mathematics, 110:199–205, 1969. CODEN LN-MAA2. ISBN 3-540-04629-1 (print), 3-540-36161-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0060118/.

Liniger:1969:GAS

[Lin69a] W. Liniger. Global accuracy and A-stability of one- and twostep integration formulae for stiff ordinary differential equations. Lecture Notes in Mathematics, 109:188–193, 1969. CODEN LN-MAA2. ISBN 3-540-04628-3 (print), 3-540-36158-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0060027/.

Linton:1969:AFS

[Lin69b] F. E. J. Linton. Applied functorial semantics, II. Lecture Notes in Mathematics, 80:53–74, 1969. CODEN LNMAA2. ISBN 3-540-04601-1 (print), 3-540-36091-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083081/.

Linton:1969:CCA

[Lin69c] F. E. J. Linton. Coequalizers in categories of algebras. Lecture Notes in Mathematics, 80:75–90, 1969. CODEN LNMAA2. ISBN 3-540-04601-1 (print), 3-540-36091-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083082/.

Linton:1969:OFS

[Lin69d] F. E. J. Linton. An outline of functorial semantics. Lecture Notes in Mathematics, 80:7–52, 1969. CODEN LNMAA2. ISBN 3-540-04601-1 (print), 3-540-36091-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083080/.

Linton:1969:RFS

[Lin69e] F. E. J. Linton. Relative functorial semantics: Adjointness results. Lecture Notes in Mathematics, 99:384–418, 1969. CODEN LNMAA2. ISBN 3-540-04618-6 (print), 3-540-36140-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081968/.

Lob:1968:PSS

[Löb68] M. H. Löb, editor. Proceedings of the Summer School in Logic Leeds, 1967: N.A.T.O. Advanced Study Institute Meeting of the Association for Symbolic Logic, volume 70 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. CODEN LNMAA2. ISBN 3-540-04240-7 (paperback), 3-540-35886-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 70. URL http://

link.springer.com/book/10.1007/BFb0079093.

Lob:1969:CTH

[Löb69] Martin H. Löb, editor. Category Theory, Homology Theory and Their Applications III: Proceedings of the Conference held at the Seattle Research Center of the Battelle Memorial Institute, June 24–July, 19, 1968 Volume Three, volume 99 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CODEN LNMAA2. ISBN 3-540-04618-6 (paperback), 3-540-36140-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 99. URL

http://link.springer.com/book/10.1007/BFb0081959.

Loynes:1967:SPA

[Loy67] R. M. Loynes. Some problems arising from spectral analysis. Lecture Notes in Mathematics, 31:197–207, 1967. CODEN LN-MAA2. ISBN 3-540-03902-3 (print), 3-540-34970-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0061119/.

Lukacs:1967:AMP

[Luk67] Eugene Lukacs. Analytical methods in probability theory. Lecture Notes in Mathematics, 31:208–238, 1967. CODEN LNMAA2. ISBN 3-540-03902-3 (print), 3-540-34970-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061120/.

Lumer:1968:AFE

[Lum68a] G. Lumer, editor. Algèbres de fonctions et espaces de Hardy: Rédigé par l’auteur à l’Institut de Recherche Mathématique Avancée de l’Université de Strasbourg 1968. (French) [Algebras of functions and Hardy spaces: Written by the author at the Research Institute of Advanced Mathematics from the University of Strasbourg in 1968], volume 75 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. CODEN LNMAA2. ISBN 3-540-04600-3 (paperback), 3-540-36090-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 75. URL http://

link.springer.com/book/10.1007/BFb0096045.

Lumer:1968:AFF

[Lum68b] G. Lumer. Algèbres de fonctions. (French) [Function algebra]. Lecture Notes in Mathematics, 75:1–6, 1968. CODEN LNMAA2. ISBN 3-540-04600-3 (print), 3-540-36090-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096046/.

Lumer:1968:EHC

[Lum68c] G. Lumer. Espaces de Hardy: Conjugués abstraits. (French) [Hardy spaces: Abstract conjugates]. Lecture Notes in Mathematics, 75:28–34, 1968. CODEN LNMAA2. ISBN 3-540-04600-3 (print), 3-540-36090-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096049/.

Lumer:1968:EHT

[Lum68d] G. Lumer. Espaces de Hardy: théorème de modification de convergence. (French) [Hardy spaces: convergence-change theorem]. Lecture Notes in Mathematics, 75:17–27, 1968. CODEN LN-MAA2. ISBN 3-540-04600-3 (print), 3-540-36090-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0096048/.

Lumer:1968:EFS

[Lum68e] G. Lumer. Étude de E et D(∗). (French) [Study of E and D(∗)]. Lecture Notes in Mathematics, 75:43–46, 1968. CODEN LN-MAA2. ISBN 3-540-04600-3 (print), 3-540-36090-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0096051/.

Lumer:1968:EDC

[Lum68f] G. Lumer. Étude du cas où M_{γ} possède un nombre fini de générateurs. (French) [Case study where M_{γ} has a finite number of generators]. Lecture Notes in Mathematics, 75:57–64, 1968. CODEN LNMAA2. ISBN 3-540-04600-3 (print), 3-540-36090-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0096054/.

Lumer:1968:MOF

[Lum68g] G. Lumer. Mesures orthogonales. (French) [Orthogonal measures]. Lecture Notes in Mathematics, 75:47–50, 1968. CODEN LN-MAA2. ISBN 3-540-04600-3 (print), 3-540-36090-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0096052/.

Lumer:1968:TMP

[Lum68h] G. Lumer. Théorie de la mesure par rapport à un système. (French) [Measurement theory with respect to a system]. Lecture Notes in Mathematics, 75:7–16, 1968. CODEN LNMAA2. ISBN 3-540-04600-3 (print), 3-540-36090-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096047/.

Lumer:1968:TPP

[Lum68i] G. Lumer. Théorie de la prediction (processus discrèts simplement stationnaires). (French) [Prediction theory (simple-stationary discrete processes)]. Lecture Notes in Mathematics, 75:35–42, 1968.

CODEN LNMAA2. ISBN 3-540-04600-3 (print), 3-540-36090-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0096050/.

Lumer:1968:UMR

[Lum68j] G. Lumer. Unicité de measure representative. Algèbres de Dirichlet et logmodulaires. (French) [Uniqueness of representative measure. Algebras of Dirichlet and log modulars]. Lecture Notes in Mathematics, 75:51–56, 1968. CODEN LNMAA2. ISBN 3-540-04600-3 (print), 3-540-36090-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096053/.

Luneburg:1965:KDP

[Lün65a] Heinz Lüneburg. S(q) als kollineationsgruppe des 3-dimensionalen projektiven Raumes über GF(q). (German) [S(q) as a collineation group of 3-dimensional projective space over GF(q)]. Lecture Notes in Mathematics, 10:72–79, 1965. CODEN LNMAA2. ISBN 3-540-03353-X (print), 3-540-37137-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082221/.

Luneburg:1965:KEO

[Lün65b] Heinz Lüneburg. S(q) als Kollineationsgruppe einer Ebene der Ordnung q^{2}. (German) [S(q) as a collineation group of a plane of order q^{2}]. Lecture Notes in Mathematics, 10:96–108, 1965. CO-DEN LNMAA2. ISBN 3-540-03353-X (print), 3-540-37137-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0082225/.

Luneburg:1965:APE

[Lün65c] Heinz Lüneburg. Affine und projektive Ebenen. (German) [Affine and projective planes]. Lecture Notes in Mathematics, 10:47– 52, 1965. CODEN LNMAA2. ISBN 3-540-03353-X (print), 3-540-37137-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082217/.

Luneburg:1965:ESG

[Lün65d] Heinz Lüneburg. Die Einfacheit der Suzukigruppen. (German) [The plainness of the Suzuki group]. Lecture Notes in Mathematics, 10:11–16, 1965. CODEN LNMAA2. ISBN 3-540-03353-X (print), 3-540-37137-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082213/.

Luneburg:1965:EBK

[Lün65e] Heinz Lüneburg. Die explizite Bestimmung der Kongruenz. (German) [The explicit determination of the congruence]. Lecture Notes in Mathematics, 10:91–95, 1965. CODEN LNMAA2. ISBN 3-540-03353-X (print), 3-540-37137-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082224/.

Luneburg:1965:GGG

[Lün65f] Heinz Lüneburg. Die Gruppen G. (German) [The group G]. Lecture Notes in Mathematics, 10:1–10, 1965. CODEN LNMAA2. ISBN 3-540-03353-X (print), 3-540-37137-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082212/.

Luneburg:1965:SIG

[Lün65g] Heinz Lüneburg, editor. Die Suzukigruppen und ihre Geometrien: Vorlesung Sommersemester 1965 in Mainz. (German) [The Suzuki Group and their geometries: Lectures of the summer semester 1965 in Mainz], volume 10 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1965. CODEN LNMAA2. ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 10. URL http://link.

springer.com/book/10.1007/BFb0082211.

Luneburg:1965:USG

[Lün65h] Heinz Lüneburg. Die Untergruppen der Suzukigruppen. (German) [The subgroups of the Suzuki group]. Lecture Notes in Mathematics, 10:26–38, 1965. CODEN LNMAA2. ISBN 3-540-03353-X (print), 3-540-37137-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082215/.

Luneburg:1965:SGM

[Lün65i] Heinz Lüneburg. Die zu den Suzukigruppen gehörigen Möbiusebenen. (German) [The Möbius planes belonging to the Suzuki group]. Lecture Notes in Mathematics, 10:68–71, 1965. CODEN LN-MAA2. ISBN 3-540-03353-X (print), 3-540-37137-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082220/.

Luneburg:1965:SGT

[Lün65j] Heinz Lüneburg. Die zu den Suzukigruppen gehörigen Translationsebenen. (German) [The translation planes belonging to the Suzuki group]. Lecture Notes in Mathematics, 10:86–90, 1965. CODEN LNMAA2. ISBN 3-540-03353-X (print), 3-540-37137-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082223/.

Luneburg:1965:KGG

[Lün65k] Heinz Lüneburg. Eine Kennzeichnung der (ZT)-Gruppen. (German) [A marking of (ZT)-groups]. Lecture Notes in Mathematics, 10:17–25, 1965. CODEN LNMAA2. ISBN 3-540-03353-X (print), 3-540-37137-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082214/.

Luneburg:1965:IGI

[Lün65l] Heinz Lüneburg. Inzidenzstrukturen. (German) [Incidence structures]. Lecture Notes in Mathematics, 10:39–46, 1965. CODEN LNMAA2. ISBN 3-540-03353-X (print), 3-540-37137-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082216/.

Luneburg:1965:MGM

[Lün65m] Heinz Lüneburg. Möbiusebenen. (German) [The Möbius plane]. Lecture Notes in Mathematics, 10:59–67, 1965. CODEN LN-MAA2. ISBN 3-540-03353-X (print), 3-540-37137-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082219/.

Luneburg:1965:PPE

[Lün65n] Heinz Lüneburg. Perspektivitäten von projektiven Ebenen. (German) [Perspectives on the projective plane]. Lecture Notes in Mathematics, 10:53–58, 1965. CODEN LNMAA2. ISBN 3-540-03353-X (print), 3-540-37137-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082218/.

Luneburg:1965:TGT

[Lün65o] Heinz Lüneburg. Translationsebenen. (German) [Translation planes]. Lecture Notes in Mathematics, 10:80–85, 1965. CODEN LNMAA2. ISBN 3-540-03353-X (print), 3-540-37137-0 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082222/.

Luneburg:1969:GMG

[Lün69a] Heinz Lüneburg. Die großen Mathieu-gruppen. (German) [The large Mathieu groups]. Lecture Notes in Mathematics, 84:86– 92, 1969. CODEN LNMAA2. ISBN 3-540-04603-8 (print), 3-540-36093-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077147/.

Luneburg:1969:HSG

[Lün69b] Heinz Lüneburg. Die Higman–Sims Gruppe. (German) [The Higman–Sims group]. Lecture Notes in Mathematics, 84:104– 107, 1969. CODEN LNMAA2. ISBN 3-540-04603-8 (print), 3-540-36093-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077150/.

Luneburg:1969:KMG

[Lün69c] Heinz Lüneburg. Die kleinen Mathieugruppen. (German) [The small Mathieu groups]. Lecture Notes in Mathematics, 84:35– 43, 1969. CODEN LNMAA2. ISBN 3-540-04603-8 (print), 3-540-36093-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077142/.

Luneburg:1969:EBG

[Lün69d] Heinz Lüneburg. Erweiterungen von t-Blockplänen. (German) [Extensions of t-block planes]. Lecture Notes in Mathematics, 84: 18–20, 1969. CODEN LNMAA2. ISBN 3-540-04603-8 (print), 3-540-36093-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077138/.

Luneburg:1969:IGI

[Lün69e] Heinz Lüneburg. Inzidenzstrukturen. (German) [Incidence structures]. Lecture Notes in Mathematics, 84:1–9, 1969. CODEN LNMAA2. ISBN 3-540-04603-8 (print), 3-540-36093-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077136/.

Luneburg:1969:KKE

[Lün69f] Heinz Lüneburg. Kollineationen und Kollineationsgruppen von endlichen projektiven und affinen Räumen. (German) [Collineations and collineation groups of finite projective and affine spaces]. Lecture Notes in Mathematics, 84:10–17, 1969. CODEN LN-MAA2. ISBN 3-540-04603-8 (print), 3-540-36093-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077137/.

Luneburg:1969:NTE

[Lün69g] Heinz Lüneburg. Nichtexistenz transitiver Erweiterungen von Gruppen vom Suzuki-Typ. (German) [Nonexistence of transitive extensions of Suzuki-type groups]. Lecture Notes in Mathematics, 84:27–34, 1969. CODEN LNMAA2. ISBN 3-540-04603-8 (print), 3-540-36093-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077141/.

Luneburg:1969:SCJ

[Lün69h] Heinz Lüneburg. Sätze von C. Jordan, Gorenstein–Hughes und M. Hall. (German) [Theorems of C. Jordan, Gorenstein–Hughes und M. Hall]. Lecture Notes in Mathematics, 84:44–60, 1969. CO-DEN LNMAA2. ISBN 3-540-04603-8 (print), 3-540-36093-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0077143/.

Luneburg:1969:HPG

[Lün69i] Heinz Lüneburg. t-Homogene Permutationsgruppen. (German) [t-Homogeneous permutation groups]. Lecture Notes in Mathematics, 84:108–117, 1969. CODEN LNMAA2. ISBN 3-540-04603-8 (print), 3-540-36093-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077151/.

Luneburg:1969:TEE

[Lün69j] Heinz Lüneburg. Transitive Erweiterungen endlicher Permutationsgruppen. (German) [Transitive extensions of finite permutation groups], volume 84 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CODEN LNMAA2. ISBN 3-540-04603-8 (paperback), 3-540-36093-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 119 pp. LCCN QA3 .L28 no. 84. URL http://link.

springer.com/book/10.1007/BFb0077135.

Luneburg:1969:TEA

[Lün69k] Heinz Lüneburg. Transitive Erweiterungen von Automorphismengruppen von t-Blockplänen. (German) [Transitive extensions of automorphism groups of t-block planes]. Lecture Notes in Mathematics, 84:21–24, 1969. CODEN LNMAA2. ISBN 3-540-04603-8 (print), 3-540-36093-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077139/.

Luneburg:1969:ETE

[Lün69l] Heinz Lüneburg. Über die Nicht-Existenz transitiver Erweiterungen gewisser Kollineationsgruppen. (German) [On the nonexistence of transitive extensions of certain collineation groups]. Lecture Notes in Mathematics, 84:25–26, 1969. CODEN LNMAA2. ISBN 3-540-04603-8 (print), 3-540-36093-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077140.

Luneburg:1969:UPEb

[Lün69m] Heinz Lüneburg. Unitale in der projektiven ebene der ordnung 4. (German) [Unitals in the projective plane of order 4]. Lecture Notes in Mathematics, 84:81–85, 1969. CODEN LNMAA2. ISBN 3-540-04603-8 (print), 3-540-36093-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077146/.

Luneburg:1969:UPEa

[Lün69n] Heinz Lüneburg. Unitäre polaritäten endlicher projektiver Ebenen. (German) [Unitary polarities of finite projective planes]. Lecture Notes in Mathematics, 84:69–80, 1969. CODEN LNMAA2. ISBN 3-540-04603-8 (print), 3-540-36093-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077145/.

Luneburg:1969:WEG

[Lün69o] Heinz Lüneburg. Weitere Eigenschaften von L_{22}. (German) [Other properties of L_{22}]. Lecture Notes in Mathematics, 84:96–103, 1969. CODEN LNMAA2. ISBN 3-540-04603-8 (print), 3-540-36093-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077149/.

Luneburg:1969:GPE

[Lün69p] Heinz Lüneburg. Zur geometrie der 21-punkte ebene. (German) [Toward a geometry of the 21-point plane]. Lecture Notes in Mathematics, 84:61–68, 1969. CODEN LNMAA2. ISBN 3-540-04603-8 (print), 3-540-36093-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077144/.

Luneburg:1969:SMG

[Lün69q] Heinz Lüneburg. Zur Struktur der Mathieurgruppen. (German) [On the structure of Mathieu groups]. Lecture Notes in Mathematics, 84:93–95, 1969. CODEN LNMAA2. ISBN 3-540-04603-8 (print), 3-540-36093-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077148/.

Lyche:1969:OOM

[Lyc69] Tom Lyche. Optimal order multistep methods with an arbitrary number of nonsteppoints. Lecture Notes in Mathematics, 109:194–199, 1969. CODEN LNMAA2. ISBN 3-540-04628-3 (print), 3-540-36158-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060028/.

MacLane:1969:FCS

[Mac69a] Saunders Mac Lane. Foundations for categories and sets. Lecture Notes in Mathematics, 92:146–164, 1969. CODEN LNMAA2. ISBN 3-540-04611-9 (print), 3-540-36101-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080770/.

MacLane:1969:OUF

[Mac69b] Saunders Mac Lane. One universe as a foundation for category theory. Lecture Notes in Mathematics, 106:192–200, 1969. CO-DEN LNMAA2. ISBN 3-540-04625-9 (print), 3-540-36150-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0059147/.

MacLane:1969:PPC

[Mac69c] Saunders Mac Lane. Possible programs for categorists. Lecture Notes in Mathematics, 86:123–131, 1969. CODEN LNMAA2. ISBN 3-540-04605-4 (print), 3-540-36095-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079386/.

Machover:1969:GI

[Mac69d] M. Machover. General introduction. Lecture Notes in Mathematics, 94:1–47, 1969. CODEN LNMAA2. ISBN 3-540-04613-5 (print), 3-540-36129-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101448/.

Maeda:1968:IKB

[Mae68a] Fumi-Yuki Maeda. Introduction to the Kuramochi boundary. Lecture Notes in Mathematics, 58:1–9, 1968. CODEN LNMAA2. ISBN 3-540-04228-8 (print), 3-540-35859-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098442/.

Maeda:1968:FSF

[Mae68b] Fumi-Yuki Maeda. On full-superharmonic functions. Lecture Notes in Mathematics, 58:10–29, 1968. CODEN LNMAA2. ISBN 3-540-04228-8 (print), 3-540-35859-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098443/.

Malitz:1968:HNC

[Mal68] Jerome Malitz. The Hanf number for complete L_{ω_{1},ω} sentences. Lecture Notes in Mathematics, 72:166–181, 1968. CODEN LN-MAA2. ISBN 3-540-04242-3 (print), 3-540-35900-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079689/.

Manes:1969:MSD

[Man69a] Ernest Manes. Minimal subalgebras for dynamic triples. Lecture Notes in Mathematics, 99:419–447, 1969. CODEN LNMAA2. ISBN 3-540-04618-6 (print), 3-540-36140-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081969/.

Manes:1969:TTC

[Man69b] Ernest Manes. A triple theoretic construction of compact algebras. Lecture Notes in Mathematics, 80:91–118, 1969. CODEN LNMAA2. ISBN 3-540-04601-1 (print), 3-540-36091-3 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083083/.

Manvel:1969:RG

[Man69c] Bennet Manvel. On reconstruction of graphs. Lecture Notes in Mathematics, 110:207–214, 1969. CODEN LNMAA2. ISBN 3-540-04629-1 (print), 3-540-36161-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060119/.

Martens:1969:CTJ

[Mar69] Henrik H. Martens. From the classical theory of Jacobian varieties. Lecture Notes in Mathematics, 118:74–98, 1969. CODEN LNMAA2. ISBN 3-540-04907-X (print), 3-540-36246-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060253/.

Matsumoto:1968:CEP

[Mat68] Kikuji Matsumoto. A condition for each point of the Kuramochi boundary to be of harmonic measure zero. Lecture Notes in Mathematics, 58:88–96, 1968. CODEN LNMAA2. ISBN 3-540-04228-8 (print), 3-540-35859-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098447/.

Maumary:1967:TMF

[Mau67a] S. Maumary. théorème de Mazur. (French) [Mazur’s theorem]. Lecture Notes in Mathematics, 48:65–73, 1967. CODEN LN-MAA2. ISBN 3-540-03919-8 (print), 3-540-35547-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0073933/.

Maumary:1967:TSHa

[Mau67b] S. Maumary. Type simple d’homotopie (théorie algébrique). (French) [Simple homotopy type (algebraic theory)]. Lecture Notes in Mathematics, 48:37–54, 1967. CODEN LNMAA2. ISBN 3-540-03919-8 (print), 3-540-35547-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073931/.

Maumary:1967:TSHb

[Mau67c] S. Maumary. Type simple d’homotopie (théorie géométrique). (French) [Simple homotopy type (geometric theory)]. Lecture Notes in Mathematics, 48:55–64, 1967. CODEN LNMAA2. ISBN 3-540-03919-8 (print), 3-540-35547-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073932/.

May:1969:CSI

[May69a] J. Peter May. Categories of spectra and infinite loop spaces. Lecture Notes in Mathematics, 99:448–479, 1969. CODEN LN-MAA2. ISBN 3-540-04618-6 (print), 3-540-36140-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081970/.

Mayer:1969:BGB

[May69b] K. H. Mayer. Bordismusgruppen. (German) [Bordism groups]. Lecture Notes in Mathematics, 111:5–18, 1969. CODEN LN-MAA2. ISBN 3-540-04630-5 (print), 3-540-36163-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0060161/.

Mayer:1969:DGD

[May69c] K. H. Mayer. Die Definition von W (X). (German) [The definition of W (X)]. Lecture Notes in Mathematics, 111:41–49, 1969. CODEN LNMAA2. ISBN 3-540-04630-5 (print), 3-540-36163-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060165/.

Mayer:1969:HGH

[May69d] K. H. Mayer. Die Homologie von W (BU(1)). (German) [The homology of W (BU(1))]. Lecture Notes in Mathematics, 111:57– 67, 1969. CODEN LNMAA2. ISBN 3-540-04630-5 (print), 3-540-36163-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060167/.

Mayer:1969:RZCa

[May69e] K. H. Mayer. Die Relationen zwischen den charakteristischen Zahlen einer SU-Mannigfaltigkeit und eines U(k)-Bündels. (German) [The relations between the characteristic numbers of an SU-manifold and a U(k) sheaf]. Lecture Notes in Mathematics, 111:70–75, 1969. CODEN LNMAA2. ISBN 3-540-04630-5 (print), 3-540-36163-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060169/.

Mayer:1969:TGT

[May69f] K. H. Mayer. Die Torsion von Ω_{∗}^{SU}(BU(1)). (German) [The torsion of Ω_{∗}^{SU}(BU(1))]. Lecture Notes in Mathematics, 111:68–69, 1969. CODEN LNMAA2. ISBN 3-540-04630-5 (print), 3-540-36163-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0060168.

Mayer:1969:EGR

[May69g] K. H. Mayer. Ein Ergebnis für Ω_{8n+4}^{Spin}(BSO(2k + 1)). (German) [A result for Ω_{8n+4}^{Spin}(BSO(2k + 1))]. Lecture Notes in Mathematics, 111:76–80, 1969. CODEN LNMAA2. ISBN 3-540-04630-5 (print), 3-540-36163-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060170/.

Mayer:1969:EGI

[May69h] K. H. Mayer. Einleitung. (German) [Introduction]. Lecture Notes in Mathematics, 111:1–4, 1969. CODEN LNMAA2. ISBN 3-540-04630-5 (print), 3-540-36163-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060160/.

Mayer:1969:ESB

[May69i] K. H. Mayer. Exakte Sequenzen zur Bestimmung von Ω_{∗}^{SU}(X). (German) [Exact sequences for the determination of Ω_{∗}^{SU}(X)]. Lecture Notes in Mathematics, 111:34–40, 1969. CODEN LN-MAA2. ISBN 3-540-04630-5 (print), 3-540-36163-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0060164/.

Mayer:1969:IGI

[May69j] K. H. Mayer. Identifikation von W (X). (German) [Identification of W (X)]. Lecture Notes in Mathematics, 111:50–56, 1969. CO-DEN LNMAA2. ISBN 3-540-04630-5 (print), 3-540-36163-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0060166/.

Mayer:1969:KVS

[May69k] K. H. Mayer. Komplexe Vektorraumbündel über schwach fastkomplexen und orientierten Mannigfaltigkeiten. (German) [Complex vector sheaves over weakly almost-complex and oriented manifolds]. Lecture Notes in Mathematics, 111:27–33, 1969. CODEN LNMAA2. ISBN 3-540-04630-5 (print), 3-540-36163-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060163/.

Mayer:1969:RVM

[May69l] K. H. Mayer. Reelle Vektorraumbündel über SU-Mannigfaltigkeiten. (German) [Real vector sheaf over SU-manifolds]. Lecture Notes in Mathematics, 111:96–97, 1969. CODEN LNMAA2. ISBN 3-540-04630-5 (print), 3-540-36163-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060172.

Mayer:1969:RVS

[May69m] K. H. Mayer. Reelle Vektorraumbündel über schwach fastkomplexen und orientierten Mannigfaltigkeiten. (German) [Real vector sheaf over weak almost-complex and oriented manifolds]. Lecture Notes in Mathematics, 111:81–95, 1969. CODEN LN-MAA2. ISBN 3-540-04630-5 (print), 3-540-36163-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0060171/.

Mayer:1969:RZCb

[May69n] K. H. Mayer. Relationen zwischen charakteristischen Zahlen. (German) [Relations between characteristic numbers], volume 111 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CODEN LN-MAA2. ISBN 3-540-04630-5 (paperback), 3-540-36163-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 99 pp. LCCN

QA3 .L28 no. 111. URL http://link.springer.com/book/10.
1007/BFb0060159.

Mayer:1969:VBG

[May69o] K. H. Mayer. Vorbereitungen und Bezeichnungen. (German) [Preparations and designations]. Lecture Notes in Mathematics, 111:19–26, 1969. CODEN LNMAA2. ISBN 3-540-04630-5 (print), 3-540-36163-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060162/.

McKee:1969:ADM

[McK69] S. McKee. Alternating direction methods for parabolic equations in two and three space dimensions with mixed derivatives.

Lecture Notes in Mathematics, 109:200–206, 1969. CODEN LN-MAA2. ISBN 3-540-04628-3 (print), 3-540-36158-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0060029/.

Mead:1969:CRV

[MD69] K. O. Mead and L. M. Delves. On the convergence rates of variational methods. Lecture Notes in Mathematics, 109:207–213, 1969. CODEN LNMAA2. ISBN 3-540-04628-3 (print), 3-540-36158-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0060030/.

Metivier:1967:MVV

[Mét67] Michel Métivier. Martingales à valeurs vectorielles application à la dérivation. (French) [Vector-valued martingales: application to differentiation]. Lecture Notes in Mathematics, 31:239–255, 1967. CODEN LNMAA2. ISBN 3-540-03902-3 (print), 3-540-34970-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0061121/.

Meyer:1967:ACD

[Mey67a] P.-A. Meyer. Appendice: Construction des deux processus croissants associés à une martingale de carré intégrable. (French) [Appendix: Construction of two increasing processes associated with a square integrable martingale]. Lecture Notes in Mathematics, 39:88–94, 1967. CODEN LNMAA2. ISBN 3-540-03910-4 (print), 3-540-34978-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080549/.

Meyer:1967:ISF

[Mey67b] P.-A. Meyer. Intégrales stochastiques I. (French) [Stochastic integrals I]. Lecture Notes in Mathematics, 39:72–87, 1967. CODEN LNMAA2. ISBN 3-540-03910-4 (print), 3-540-34978-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080548/.

Meyer:1967:ISIa

[Mey67c] P.-A. Meyer. Intégrales stochastiques II. (French) [Stochastic integrals II]. Lecture Notes in Mathematics, 39:95–117, 1967. CODEN LNMAA2. ISBN 3-540-03910-4 (print), 3-540-34978-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080550/.

Meyer:1967:ISIb

[Mey67d] P.-A. Meyer. Intégrales stochastiques III. (French) [Stochastic integrals III]. Lecture Notes in Mathematics, 39:118–141, 1967. CO-DEN LNMAA2. ISBN 3-540-03910-4 (print), 3-540-34978-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0080551/.

Meyer:1967:ISIc

[Mey67e] P.-A. Meyer. Intégrales stochastiques IV. (French) [Stochastic integrals IV]. Lecture Notes in Mathematics, 39:142–162, 1967. CO-DEN LNMAA2. ISBN 3-540-03910-4 (print), 3-540-34978-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0080552/.

Meyer:1967:TDF

[Mey67f] P.-A. Meyer. Sur un théorème de Deny. (French) [On a theorem of Deny]. Lecture Notes in Mathematics, 39:163–165, 1967. CODEN LNMAA2. ISBN 3-540-03910-4 (print), 3-540-34978-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080553/.

Meyer:1967:BM

[Mey67g] Paul-André Meyer. Back matter. Lecture Notes in Mathematics, 26:??, 1967. CODEN LNMAA2. ISBN 3-540-03901-5 (print), 3-540-34969-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-34969-3/1.pdf.

Meyer:1967:FM

[Mey67h] Paul-André Meyer. Front matter. Lecture Notes in Mathematics, 26:??, 1967. CODEN LNMAA2. ISBN 3-540-03901-5 (print), 3-540-34969-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-34969-3/1.pdf.

Meyer:1967:PHP

[Mey67i] Paul-André Meyer. Processus de Hunt, processus standard. (French) [Hunt processes, standard processes]. Lecture Notes in Mathematics, 26:78–124, 1967. CODEN LNMAA2. ISBN 3-540-03901-5 (print), 3-540-34969-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075151/.

Meyer:1967:PMF

[Mey67j] Paul-André Meyer. Processus de Markov. (French) [Markov processes], volume 26 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. CODEN LNMAA2. ISBN 3-540-03901-5 (paperback), 3-540-34969-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 189 pp. LCCN QA3 .L28 no. 26. URL http://link.

springer.com/book/10.1007/BFb0075148.

Meyer:1967:RMH

[Mey67k] Paul-André Meyer. Réduites, mesures harmoniques. (French) [Harmonic measures reduced]. Lecture Notes in Mathematics, 26:125– 188, 1967. CODEN LNMAA2. ISBN 3-540-03901-5 (print), 3-540-34969-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075152/.

Meyer:1967:SGF

[Mey67l] Paul-André Meyer. Semi-groupes de Feller. (French) [Feller semigroups]. Lecture Notes in Mathematics, 26:24–77, 1967. CODEN LNMAA2. ISBN 3-540-03901-5 (print), 3-540-34969-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075150/.

Meyer:1967:TEP

[Mey67m] Paul-André Meyer. Théorie élémentaire des processus de Markov. (French) [Elementary theory of Markov processes]. Lecture Notes in Mathematics, 26:1–23, 1967. CODEN LNMAA2. ISBN 3-540-03901-5 (print), 3-540-34969-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075149/.

Meyer:1968:CAR

[Mey68a] P.-A. Meyer. Compactifications associées à une résolvante. (French) [Compactifications associated with a resolvent]. Lecture Notes in Mathematics, 51:175–199, 1968. CODEN LNMAA2. ISBN 3-540-04221-0 (print), 3-540-35827-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059186/.

Meyer:1968:GDT

[Mey68b] P.-A. Meyer. Guide détaillé de la théorie ‘générale’ des processus. (French) [Detailed guide of the ‘general’ theory of processes].

Lecture Notes in Mathematics, 51:140–165, 1968. CODEN LN-MAA2. ISBN 3-540-04221-0 (print), 3-540-35827-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0059183/.

Meyer:1968:RFF

[Mey68c] P.-A. Meyer. Les résolvantes fortement fellériennes d’après Mokobodzki. (French) [Highly Fellerian resolvents according to Mokobodzki]. Lecture Notes in Mathematics, 51:171–174, 1968. CO-DEN LNMAA2. ISBN 3-540-04221-0 (print), 3-540-35827-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0059185/.

Meyer:1968:MDP

[Mey68d] P.-A. Meyer. Une majoration du processus croissant naturel associé à une surmartingale. (French) [An increase in the natural growing process associated with a supermartingale]. Lecture Notes in Mathematics, 51:166–170, 1968. CODEN LNMAA2. ISBN 3-540-04221-0 (print), 3-540-35827-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059184/.

Meyer:1968:CMF

[Mey68e] Paul-André Meyer. Compactifications de Martin. (French) [Martin compactifications]. Lecture Notes in Mathematics, 77:72–94, 1968. CODEN LNMAA2. ISBN 3-540-04246-6 (print), 3-540-35919-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0101402/.

Meyer:1968:CFE

[Mey68f] Paul-André Meyer. Compléments sur les fonctions excessives. (French) [Complements on excessive functions]. Lecture Notes in Mathematics, 77:2–47, 1968. CODEN LNMAA2. ISBN 3-540-04246-6 (print), 3-540-35919-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101399/.

Meyer:1968:LKW

[Mey68g] Paul-André Meyer. L’hypothèse K-W. Quelques conséquences. (French) [The K-W hypothesis. Some consequences]. Lecture Notes in Mathematics, 77:48–61, 1968. CODEN LNMAA2. ISBN 3-540-04246-6 (print), 3-540-35919-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101400/.

Meyer:1968:PGA

[Mey68h] Paul-André Meyer. Potentiels de Green, applications. (French) [Green potentials, applications]. Lecture Notes in Mathematics, 77:62–71, 1968. CODEN LNMAA2. ISBN 3-540-04246-6 (print), 3-540-35919-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101401/.

Meyer:1968:PMA

[Mey68i] Paul-André Meyer. Processus de Markov [à]la frontière de Martin. (French) [Markov process on the Martin boundary]. Lecture Notes in Mathematics, 77:1, 1968. CODEN LNMAA2. ISBN 3-540-04246-6 (print), 3-540-35919-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101398.

Meyer:1968:PMF

[Mey68j] Paul-André Meyer. Processus de Markov: la frontière de Martin. (French) [Markov processes: the Martin boundary], volume 77 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. CO-DEN LNMAA2. ISBN 3-540-04246-6 (paperback), 3-540-35919-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 123

pp. LCCN QA3 .L28 no. 77. URL http://link.springer.com/
book/10.1007/BFb0101397.

Meyer:1968:PCM

[Mey68k] Paul-André Meyer. Processus sur un compactifié de Martin. (French) [Processes on a Martin compactification]. Lecture Notes in Mathematics, 77:95–120, 1968. CODEN LNMAA2. ISBN 3-540-04246-6 (print), 3-540-35919-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101403/.

Meyer:1969:PAI

[Mey69a] P.-A. Meyer. Processus à accroissements indépendants et positifs. (French) [Processes with independent and positive increments]. Lecture Notes in Mathematics, 88:175–189, 1969. CODEN LN-MAA2. ISBN 3-540-04607-0 (print), 3-540-36097-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081637/.

Meyer:1969:REA

[Mey69b] P.-A. Meyer. Rectifications à des exposés antérieurs. (French) [Corrections to previous presentations]. Lecture Notes in Mathematics, 88:160–162, 1969. CODEN LNMAA2. ISBN 3-540-04607-0 (print), 3-540-36097-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081635/.

Meyer:1969:RTD

[Mey69c] P.-A. Meyer. Un résultat de théorie du potentiel. (French) [A result of potentiel theory]. Lecture Notes in Mathematics, 88:144– 151, 1969. CODEN LNMAA2. ISBN 3-540-04607-0 (print), 3-540-36097-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081632/.

Meyer:1969:RET

[Mey69d] P.-A. Meyer. Un résultat élémentaire sur les temps d’arrêt. (French) [An elementary result on stopping time]. Lecture Notes in Mathematics, 88:152–154, 1969. CODEN LNMAA2. ISBN 3-540-04607-0 (print), 3-540-36097-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081633/.

Meyer:1969:NDT

[Mey69e] P.-A. Meyer. Une nouvelle démonstration des théorèmes de section. (French) [A new proof of theorems of section]. Lecture Notes in Mathematics, 88:155–159, 1969. CODEN LNMAA2. ISBN 3-540-04607-0 (print), 3-540-36097-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081634/.

Machover:1969:LNS

[MH69] Moshé Machover and Joram Hirschfeld. Lectures on Non-Standard Analysis, volume 94 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CODEN LNMAA2. ISBN 3-540-04613-5 (paperback), 3-540-36129-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 79 pp. LCCN QA3 .L28 no. 94. URL http://link.

springer.com/book/10.1007/BFb0101447.

Mitchell:1969:DOC

[Mit69] Barry Mitchell. On the dimension of objects and categories III Hochschild dimension. Lecture Notes in Mathematics, 92:165– 196, 1969. CODEN LNMAA2. ISBN 3-540-04611-9 (print), 3-540-36101-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080771/.

Morley:1968:PM

[Mor68] Michael Morley. Partitions and models. Lecture Notes in Mathematics, 70:109–158, 1968. CODEN LNMAA2. ISBN 3-540-04240-7 (print), 3-540-35886-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079095/.

Morando:1969:MAF

[Mor69a] Philippe Morando. Mesures aléatoires. (French) [Random measures]. Lecture Notes in Mathematics, 88:190–191, 1969. CODEN LNMAA2. ISBN 3-540-04607-0 (print), 3-540-36097-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081638.

Morando:1969:PMA

[Mor69b] Philippe Morando. Prolongements des mesures alèatoires. (French) [Extensions of random measures]. Lecture Notes in Mathematics, 88:192–216, 1969. CODEN LNMAA2. ISBN 3-540-04607-0 (print), 3-540-36097-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081639/.

Morando:1969:RMA

[Mor69c] Philippe Morando. Representation des mesures aléatoires positives à accroissement independants. (French) [Representation of positive random measures to independent increase]. Lecture Notes in Mathematics, 88:217–229, 1969. CODEN LNMAA2. ISBN 3-540-04607-0 (print), 3-540-36097-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081640/.

Morris:1969:CNS

[Mor69d] J. Ll. Morris, editor. Conference on the Numerical Solution of Differential Equations: Held in Dundee/Scotland, June 23–27, 1969, volume 109 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CODEN LNMAA2. ISBN 3-540-04628-3 (paperback), 3-540-36158-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 109. URL http://link.springer.

com/book/10.1007/BFb0060012.

Moss:1969:ATG

[MT69] Robert Michael Forrester Moss and C. B. (Charles Benedict) Thomas, editors. Algebraic K-Theory and its Geometric Applications, volume 108 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CODEN LNMAA2. ISBN 3-540-04627-5 (paperback), 3-540-36156-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 86 pp. LCCN QA3 .L28 no. 108. URL http://link.

springer.com/book/10.1007/BFb0059986.

Muller:1966:AT

[Mül66a] Claus Müller. Addition theorem. Lecture Notes in Mathematics, 17:9–11, 1966. CODEN LNMAA2. ISBN 3-540-03600-8 (print), 3-540-37174-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094779/.

Muller:1966:AAT

[Mül66b] Claus Müller. Applications of the addition theorem. Lecture Notes in Mathematics, 17:14–15, 1966. CODEN LNMAA2. ISBN 3-540-03600-8 (print), 3-540-37174-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094781.

Muller:1966:ALF

[Mül66c] Claus Müller. Associated Legendre functions. Lecture Notes in Mathematics, 17:22–29, 1966. CODEN LNMAA2. ISBN 3-540-03600-8 (print), 3-540-37174-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094785/.

Muller:1966:DE

[Mül66d] Claus Müller. Differential equations. Lecture Notes in Mathematics, 17:37–39, 1966. CODEN LNMAA2. ISBN 3-540-03600-8 (print), 3-540-37174-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094787/.

Muller:1966:ESH

[Mül66e] Claus Müller. Expansions in spherical harmonics. Lecture Notes in Mathematics, 17:40–44, 1966. CODEN LNMAA2. ISBN 3-540-03600-8 (print), 3-540-37174-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094788/.

Muller:1966:FHF

[Mül66f] Claus Müller. Funk–Hecke formula. Lecture Notes in Mathematics, 17:18–20, 1966. CODEN LNMAA2. ISBN 3-540-03600-8 (print), 3-540-37174-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094783/.

Muller:1966:GBN

[Mül66g] Claus Müller. General background and notation. Lecture Notes in Mathematics, 17:1–5, 1966. CODEN LNMAA2. ISBN 3-540-03600-8 (print), 3-540-37174-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094776/.

Muller:1966:IRS

[Mül66h] Claus Müller. Integral representations of spherical harmonics. Lecture Notes in Mathematics, 17:21–22, 1966. CODEN LN-MAA2. ISBN 3-540-03600-8 (print), 3-540-37174-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0094784.

Muller:1966:LF

[Mül66i] Claus Müller. Legendre functions. Lecture Notes in Mathematics, 17:7–9, 1966. CODEN LNMAA2. ISBN 3-540-03600-8 (print), 3-540-37174-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094778/.

Muller:1966:OT

[Mül66j] Claus Müller. Orthogonal transformations. Lecture Notes in Mathematics, 17:5–7, 1966. CODEN LNMAA2. ISBN 3-540-03600-8 (print), 3-540-37174-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094777/.

Muller:1966:PLF

[Mül66k] Claus Müller. Properties of the Legendre functions. Lecture Notes in Mathematics, 17:29–37, 1966. CODEN LNMAA2. ISBN 3-540-03600-8 (print), 3-540-37174-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094786/.

Muller:1966:RT

[Mül66l] Claus Müller. Reprensentation theorem. Lecture Notes in Mathematics, 17:11–13, 1966. CODEN LNMAA2. ISBN 3-540-03600-8 (print), 3-540-37174-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094780/.

Muller:1966:RF

[Mül66m] Claus Müller. Rodrigues’ formula. Lecture Notes in Mathematics, 17:16–18, 1966. CODEN LNMAA2. ISBN 3-540-03600-8 (print), 3-540-37174-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094782/.

Muller:1966:SH

[Mül66n] Claus Müller. Spherical Harmonics, volume 17 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1966. CODEN LNMAA2. ISBN 3-540-03600-8 (paperback), 3-540-37174-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 45 pp. LCCN QA3 .L28 no. 17.

URL http://link.springer.com/book/10.1007/BFb0094775.

Narasimhan:1966:CT

[Nar66a] Raghavan Narasimhan. Coherence theorems. Lecture Notes in Mathematics, 25:64–90, 1966. CODEN LNMAA2. ISBN 3-540-03608-3 (print), 3-540-34845-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077075/.

Narasimhan:1966:HMC

[Nar66b] Raghavan Narasimhan. Holomorphic mappings of complex spaces. Lecture Notes in Mathematics, 25:123–136, 1966. CODEN LN-MAA2. ISBN 3-540-03608-3 (print), 3-540-34845-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077078/.

Narasimhan:1966:ITA

[Nar66c] Raghavan Narasimhan. Introduction to the Theory of Analytic Spaces, volume 25 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1966. CODEN LNMAA2. ISBN 3-540-03608-3 (paperback), 3-540-34845-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 143 pp. LCCN QA3 .L28 no. 25. URL http://link.

springer.com/book/10.1007/BFb0077071.

Narasimhan:1966:LPA

[Nar66d] Raghavan Narasimhan. Local properties of analytic sets. Lecture Notes in Mathematics, 25:31–63, 1966. CODEN LNMAA2. ISBN 3-540-03608-3 (print), 3-540-34845-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077074/.

Narasimhan:1966:NT

[Nar66e] Raghavan Narasimhan. The normalization theorem. Lecture Notes in Mathematics, 25:110–122, 1966. CODEN LNMAA2. ISBN 3-540-03608-3 (print), 3-540-34845-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077077/.

Narasimhan:1966:P

[Nar66f] Raghavan Narasimhan. Preliminaries. Lecture Notes in Mathematics, 25:2–8, 1966. CODEN LNMAA2. ISBN 3-540-03608-3 (print), 3-540-34845-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077072/.

Narasimhan:1966:RAS

[Nar66g] Raghavan Narasimhan. Real analytic sets. Lecture Notes in Mathematics, 25:91–109, 1966. CODEN LNMAA2. ISBN 3-540-03608-3 (print), 3-540-34845-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077076/.

Narasimhan:1966:WPT

[Nar66h] Raghavan Narasimhan. The Weierstrass preparation theorem. Lecture Notes in Mathematics, 25:9–30, 1966. CODEN LN-MAA2. ISBN 3-540-03608-3 (print), 3-540-34845-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077073/.

Noble:1969:IWS

[Nob69] Ben Noble. Instability when solving Volterra integral equations of the second kind by multistep methods. Lecture Notes in Mathematics, 109:23–39, 1969. CODEN LNMAA2. ISBN 3-540-04628-3 (print), 3-540-36158-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060015/.

Norguet:1968:VAS

[Nor68] François Norguet. Variétés algébriques strictement Q-pseudoconvexes. (French) [Strictly Q-pseudoconvex algebraic varieties]. Lecture Notes in Mathematics, 71:94–98, 1968. CODEN LNMAA2. ISBN 3-540-04241-5 (print), 3-540-35889-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080100/.

Norsett:1969:SMA

[Nør69] Syvert P. Nørsett. An A-stable modification of the Adams– Bashforth methods. Lecture Notes in Mathematics, 109:214– 219, 1969. CODEN LNMAA2. ISBN 3-540-04628-3 (print), 3-540-36158-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060031/.

Noverraz:1968:GSR

[Nov68] Ph. Noverraz. Generators for some rings of analytic functions d’après L. Hörmander [0]. Lecture Notes in Mathematics, 71:113– 117, 1968. CODEN LNMAA2. ISBN 3-540-04241-5 (print), 3-540-35889-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080102/.

Ohtsuka:1968:KPP

[Oht68] Makoto Ohtsuka. On Kuramochi’s paper “Potentials on Riemann surfaces”. Lecture Notes in Mathematics, 58:70–87, 1968. CO-DEN LNMAA2. ISBN 3-540-04228-8 (print), 3-540-35859-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0098446/.

Olum:1969:HSF

[Olu69] Paul Olum. Homology of squares and factoring of diagrams. Lecture Notes in Mathematics, 99:480–489, 1969. CODEN LNMAA2. ISBN 3-540-04618-6 (print), 3-540-36140-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081971/.

Oort:1966:AGS

[Oor66a] F. Oort. Algebraic group schemes. Lecture Notes in Mathematics, 15:31–97, 1966. CODEN LNMAA2. ISBN 3-540-03598-2 (print), 3-540-37171-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097481/.

Oort:1966:CGS

[Oor66b] F. Oort. Commutative group schemes, volume 15 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1966. CODEN LNMAA2. ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 133 pp. LCCN QA3

.L28 no. 15; QA3 .L28 no. 5. URL http://link.springer.com/
book/10.1007/BFb0097479.

Oort:1966:DTA

[Oor66c] F. Oort. Duality theorems for Abelian schemes. Lecture Notes in Mathematics, 15:98–130, 1966. CODEN LNMAA2. ISBN 3-540-03598-2 (print), 3-540-37171-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097482/.

Oort:1966:P

[Oor66d] F. Oort. Preliminaries. Lecture Notes in Mathematics, 15:1– 30, 1966. CODEN LNMAA2. ISBN 3-540-03598-2 (print), 3-540-37171-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097480/.

Padmanabhan:1969:CPA

[Pad69] A. R. Padmanabhan. Convergence in probability and allied results. Lecture Notes in Mathematics, 89:186, 1969. CODEN LN-MAA2. ISBN 3-540-04608-9 (print), 3-540-36098-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079126.

Pare:1969:AC

[Par69a] Robert Paré. Absolute coequalizers. Lecture Notes in Mathematics, 86:132–145, 1969. CODEN LNMAA2. ISBN 3-540-04605-4 (print), 3-540-36095-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079387/.

Parthasarathy:1969:BML

[Par69b] K. R. Parthasarathy. Borel multipliers on a locally compact group. Lecture Notes in Mathematics, 93:11–21, 1969. CODEN LNMAA2. ISBN 3-540-04612-7 (print), 3-540-36102-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097433/.

Parthasarathy:1969:I

[Par69c] K. R. Parthasarathy. Introduction. Lecture Notes in Mathematics, 93:1–3, 1969. CODEN LNMAA2. ISBN 3-540-04612-7 (print), 3-540-36102-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097431/.

Parthasarathy:1969:MSS

[Par69d] K. R. Parthasarathy. Miltipliers on some special groups. Lecture Notes in Mathematics, 93:28–52, 1969. CODEN LNMAA2. ISBN 3-540-04612-7 (print), 3-540-36102-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097435/.

Parthasarathy:1969:MLG

[Par69e] K. R. Parthasarathy. Multipliers on Lie groups. Lecture Notes in Mathematics, 93:22–28, 1969. CODEN LNMAA2. ISBN 3-540-04612-7 (print), 3-540-36102-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097434/.

Parthasarathy:1969:MLC

[Par69f] K. R. Parthasarathy. Multipliers on Locally Compact Groups, volume 93 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CO-DEN LNMAA2. ISBN 3-540-04612-7 (paperback), 3-540-36102-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 54

pp. LCCN QA3 .L28 no. 93. URL http://link.springer.com/
book/10.1007/BFb0097430.

Parthasarathy:1969:SGR

[Par69g] K. R. Parthasarathy. Standard groups with a right invariant measure. Lecture Notes in Mathematics, 93:3–11, 1969. CODEN LN-MAA2. ISBN 3-540-04612-7 (print), 3-540-36102-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0097432/.

Peyerimhoff:1969:CMF

[Pey69a] A. Peyerimhoff. Cesàro means. Lecture Notes in Mathematics, 107:4–10, 1969. CODEN LNMAA2. ISBN 3-540-04626-7 (print), 3-540-36152-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060953/.

Peyerimhoff:1969:HNS

[Pey69b] A. Peyerimhoff. Hausdorff and Nörlund summability. Lecture Notes in Mathematics, 107:88–101, 1969. CODEN LNMAA2. ISBN 3-540-04626-7 (print), 3-540-36152-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060956/.

Peyerimhoff:1969:IR

[Pey69c] A. Peyerimhoff. Introductory remarks. Lecture Notes in Mathematics, 107:2–4, 1969. CODEN LNMAA2. ISBN 3-540-04626-7 (print), 3-540-36152-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060952/.

Peyerimhoff:1969:LS

[Pey69d] A. Peyerimhoff. Lectures on Summability, volume 107 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CODEN LNMAA2. ISBN 3-540-04626-7 (paperback), 3-540-36152-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 111 pp. LCCN

QA3 .L28 no. 107. URL http://link.springer.com/book/10.
1007/BFb0060951.

Peyerimhoff:1969:MT

[Pey69e] A. Peyerimhoff. Matrix transformations. Lecture Notes in Mathematics, 107:10–63, 1969. CODEN LNMAA2. ISBN 3-540-04626-7 (print), 3-540-36152-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060954/.

Peyerimhoff:1969:TT

[Pey69f] A. Peyerimhoff. Tauberian theorems. Lecture Notes in Mathematics, 107:63–88, 1969. CODEN LNMAA2. ISBN 3-540-04626-7 (print), 3-540-36152-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060955/.

Pimbley:1969:DBS

[Pim69a] George H. Pimbley, Jr. Development of branches of solutions for nonlinear equations near an exceptional point. Bifurcation theory. Lecture Notes in Mathematics, 104:18–28, 1969. CODEN LNMAA2. ISBN 3-540-04623-2 (print), 3-540-36147-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065822/.

Pimbley:1969:EBN

[Pim69b] George H. Pimbley, Jr. Eigenfunction Branches of Nonlinear Operators, and their Bifurcations, volume 104 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CODEN LNMAA2. ISBN 3-540-04623-2 (paperback), 3-540-36147-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). i + 128 pp. LCCN QA3

.L28 no. 104. URL http://link.springer.com/book/10.1007/
BFb0065819.

Pimbley:1969:EPH

[Pim69c] George H. Pimbley, Jr. The eigenvalue problem; Hammerstein operators; sublinear and superlinear operators; oscillation kernels. Lecture Notes in Mathematics, 104:43–57, 1969. CODEN LN-MAA2. ISBN 3-540-04623-2 (print), 3-540-36147-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0065824/.

Pimbley:1969:E

[Pim69d] George H. Pimbley, Jr. An example. Lecture Notes in Mathematics, 104:4–10, 1969. CODEN LNMAA2. ISBN 3-540-04623-2 (print), 3-540-36147-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065820/.

Pimbley:1969:ESR

[Pim69e] George H. Pimbley, Jr. The example of section 1, reconsidered. Lecture Notes in Mathematics, 104:85–91, 1969. CODEN LN-MAA2. ISBN 3-540-04623-2 (print), 3-540-36147-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0065827/.

Pimbley:1969:EBEb

[Pim69f] George H. Pimbley, Jr. Extension of branches of eigenfunctions of Hammerstein operators. Lecture Notes in Mathematics, 104:80–84, 1969. CODEN LNMAA2. ISBN 3-540-04623-2 (print), 3-540-36147-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065826/.

Pimbley:1969:EBS

[Pim69g] George H. Pimbley, Jr. The extension of branches of solutions for nonlinear equations in Banach spaces. Lecture Notes in Mathematics, 104:11–17, 1969. CODEN LNMAA2. ISBN 3-540-04623-2 (print), 3-540-36147-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065821/.

Pimbley:1969:EBEa

[Pim69h] George H. Pimbley, Jr. On the extension of branches of eigenfunctions; conditions preventing secondary bifurcation of branches. Lecture Notes in Mathematics, 104:58–79, 1969. CODEN LN-MAA2. ISBN 3-540-04623-2 (print), 3-540-36147-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0065825/.

Pimbley:1969:SBE

[Pim69i] George H. Pimbley, Jr. Solution of the bifurcation equation in the case n = 1; bifurcation at the origin. Lecture Notes in Mathematics, 104:29–42, 1969. CODEN LNMAA2. ISBN 3-540-04623-2 (print), 3-540-36147-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065823/.

Pimbley:1969:SCH

[Pim69j] George H. Pimbley, Jr. Summary; collection of hypotheses; unsettled questions. Lecture Notes in Mathematics, 104:102–113, 1969. CODEN LNMAA2. ISBN 3-540-04623-2 (print), 3-540-36147-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0065829/.

Pimbley:1969:TPB

[Pim69k] George H. Pimbley, Jr. A two-point boundary value problem. Lecture Notes in Mathematics, 104:92–101, 1969. CODEN LN-MAA2. ISBN 3-540-04623-2 (print), 3-540-36147-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0065828/.

Piotrowski:1969:SCC

[Pio69] Peter Piotrowski. Stability, consistency and convergence of variable K-step methods for numerical integration of large systems of ordinary differential equations. Lecture Notes in Mathematics, 109:221–227, 1969. CODEN LNMAA2. ISBN 3-540-04628-3 (print), 3-540-36158-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060032/.

Poly:1968:CCT

[Pol68] J.-B. Poly. d′′-cohomologie à croissance: Un théorème de décomposition sur un ouvert pseudo-convexe. (French) [d′′- cohomology to growth: A decomposition theorem on a pseudoconvex opening]. Lecture Notes in Mathematics, 71:72–80, 1968. CODEN LNMAA2. ISBN 3-540-04241-5 (print), 3-540-35889-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080098/.

Pfanzagl:1966:A

[PP66a] J. Pfanzagl and W. Pierlo. Approximation. Lecture Notes in Mathematics, 16:5–9, 1966. CODEN LNMAA2. ISBN 3-540-03599-0 (print), 3-540-37172-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078993/.

Pfanzagl:1966:CA

[PP66b] J. Pfanzagl and W. Pierlo. Compact approximation. Lecture Notes in Mathematics, 16:10–12, 1966. CODEN LNMAA2. ISBN 3-540-03599-0 (print), 3-540-37172-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078994/.

Pfanzagl:1966:CAT

[PP66c] J. Pfanzagl and W. Pierlo. Compact approximation in topological spaces. Lecture Notes in Mathematics, 16:13–21, 1966. CODEN LNMAA2. ISBN 3-540-03599-0 (print), 3-540-37172-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078995/.

Pfanzagl:1966:CSSa

[PP66d] J. Pfanzagl and W. Pierlo. Compact systems of sets. Lecture Notes in Mathematics, 16:2–4, 1966. CODEN LNMAA2. ISBN 3-540-03599-0 (print), 3-540-37172-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078992/.

Pfanzagl:1966:CSSb

[PP66e] J. Pfanzagl and W. Pierlo. Compact Systems of Sets, volume 16 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1966. CO-DEN LNMAA2. ISBN 3-540-03599-0 (paperback), 3-540-37172-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 48

pp. LCCN QA3 .L28 no. 16. URL http://link.springer.com/
book/10.1007/BFb0078990.

Pfanzagl:1966:EPM

[PP66f] J. Pfanzagl and W. Pierlo. Existence of product measures. Lecture Notes in Mathematics, 16:25–33, 1966. CODEN LNMAA2. ISBN 3-540-03599-0 (print), 3-540-37172-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078997/.

Pfanzagl:1966:ERC

[PP66g] J. Pfanzagl and W. Pierlo. Existence of regular conditional probability measures. Lecture Notes in Mathematics, 16:34–41, 1966. CODEN LNMAA2. ISBN 3-540-03599-0 (print), 3-540-37172-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0078998/.

Pfanzagl:1966:I

[PP66h] J. Pfanzagl and W. Pierlo. Introduction. Lecture Notes in Mathematics, 16:1, 1966. CODEN LNMAA2. ISBN 3-540-03599-0 (print), 3-540-37172-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078991.

Pfanzagl:1966:PM

[PP66i] J. Pfanzagl and W. Pierlo. Perfect measures. Lecture Notes in Mathematics, 16:22–24, 1966. CODEN LNMAA2. ISBN 3-540-03599-0 (print), 3-540-37172-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078996/.

Preller:1968:QA

[Pre68] A. Preller. Quantified algebras. Lecture Notes in Mathematics, 72:182–203, 1968. CODEN LNMAA2. ISBN 3-540-04242-3 (print), 3-540-35900-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079690/.

Pyke:1969:AAS

[Pyk69] Ronald Pyke. Applications of almost surely convergent constructions of weakly convergent processes. Lecture Notes in Mathematics, 89:187–200, 1969. CODEN LNMAA2. ISBN 3-540-04608-9 (print), 3-540-36098-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079127/.

D:1966:HMH

[Q. 66a] Q. E. D. Hausdorff measure and the Hausdorff metric. Lecture Notes in Mathematics, 19:21–29, 1966. CODEN LNMAA2. ISBN 3-540-03602-4 (print), 3-540-37194-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097740/.

D:1966:UPM

[Q. 66b] Q. E. D. The use of the proper mapping. Lecture Notes in Mathematics, 19:30–37, 1966. CODEN LNMAA2. ISBN 3-540-03602-4 (print), 3-540-37194-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097741/.

Quillen:1967:AHT

[Qui67a] Daniel G. Quillen. Axiomatic homotopy theory. Lecture Notes in Mathematics, 43:1–64, 1967. CODEN LNMAA2. ISBN 3-540-03914-7 (print), 3-540-35523-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097439/.

Quillen:1967:ESH

[Qui67b] Daniel G. Quillen. Examples of simplicial homotopy theories. Lecture Notes in Mathematics, 43:65–155, 1967. CODEN LN-MAA2. ISBN 3-540-03914-7 (print), 3-540-35523-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0097440/.

Quillen:1967:HA

[Qui67c] Daniel G. Quillen. Homotopical Algebra, volume 43 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. CODEN LNMAA2. ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 157 pp. LCCN

QA3 .L28 no. 43. URL http://link.springer.com/book/10.
1007/BFb0097438.

Ramis:1968:SEA

[Ram68] J.-P. Ramis. Sous-ensembles analytiques d’une variété analytique banachique. (French) [Analytic subsets of a Banach analytic manifold]. Lecture Notes in Mathematics, 71:140–164, 1968. CODEN LNMAA2. ISBN 3-540-04241-5 (print), 3-540-35889-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080105/.

Rickart:1969:ERS

[Ric69] C. E. Rickart. Extension of results from several complex variables to general function algebras. Lecture Notes in Mathematics, 103:44–59, 1969. CODEN LNMAA2. ISBN 3-540-04622-4 (print), 3-540-36145-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099984/.

Rodding:1968:KRF

[Röd68] D. Rödding. Klassen rekursiver funktionen. (German) [Classes of recursive functions]. Lecture Notes in Mathematics, 70:159– 222, 1968. CODEN LNMAA2. ISBN 3-540-04240-7 (print), 3-540-35886-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079096/.

Roos:1967:LDS

[Roo67] Jan-Erik Roos. Locally distributive spectral categories and strongly regular rings. Lecture Notes in Mathematics, 47:156–181, 1967. CODEN LNMAA2. ISBN 3-540-03918-X (print), 3-540-35545-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074303/.

Roos:1968:LAH

[Roo68] Guy Roos. L’image d’une application holomorphe d’une variété dans l’espace projectif (d’après W. Stoll). (French) [The image of a holomorphic map of a variety in projective space (after W. Stoll)]. Lecture Notes in Mathematics, 71:127–139, 1968. CODEN LNMAA2. ISBN 3-540-04241-5 (print), 3-540-35889-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080104/.

Roos:1969:LNC

[Roo69] Jan-Erik Roos. Locally Noetherian categories and generalized strictly linearly compact rings. Applications. Lecture Notes in Mathematics, 92:197–277, 1969. CODEN LNMAA2. ISBN 3-540-04611-9 (print), 3-540-36101-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080772/.

Rose:1968:FSL

[Ros68a] Alan Rose. Formalisations of some ℵ_{0} Łukasiewicz propositional calculi. Lecture Notes in Mathematics, 70:269–271, 1968. CO-DEN LNMAA2. ISBN 3-540-04240-7 (print), 3-540-35886-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0079100/.

Rosenstein:1968:TWC

[Ros68b] Joseph G. Rosenstein. Theories which are not χ_{o}-categorical. Lecture Notes in Mathematics, 70:273–278, 1968. CODEN LN-MAA2. ISBN 3-540-04240-7 (print), 3-540-35886-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079101/.

Rossi:1969:SPM

[Ros69] Hugo Rossi. Strongly pseudoconvex manifolds. Lecture Notes in Mathematics, 103:10–29, 1969. CODEN LNMAA2. ISBN 3-540-04622-4 (print), 3-540-36145-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099982/.

Schapira:1968:EDP

[Sch68] Pierre Schapira. Équations aux dérivées partielles dans l’espace des hyperfonctions. (French) [Partial differential equations in the space of hyperfunctions]. Lecture Notes in Mathematics, 71:38–45, 1968. CODEN LNMAA2. ISBN 3-540-04241-5 (print), 3-540-35889-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0080095/.

Schwartz:1969:EDT

[Sch69] L. Schwartz. Extension du théorème de Sazonov–Minlos. (French) [Extension of the Sazonov– Minlos theorem]. Lecture Notes in Mathematics, 88:1–23, 1969. CODEN LNMAA2. ISBN 3-540-04607-0 (print), 3-540-36097-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081624/.

Strassen:1969:CLT

[SD69] Volker Strassen and R. M. Dudley. The central limit theorem and ϵ-entropy. Lecture Notes in Mathematics, 89:224–231, 1969. CODEN LNMAA2. ISBN 3-540-04608-9 (print), 3-540-36098-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0079129/.

Selberg:1969:RDT

[Sel69] Atle Selberg. Recent developments in the theory of discontinuous groups of motions of symmetric spaces. Lecture Notes in Mathematics, 118:99–120, 1969. CODEN LNMAA2. ISBN 3-540-04907-X (print), 3-540-36246-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060254/.

Serre:1965:BM

[Ser65a] Jean-Pierre Serre. Back matter. Lecture Notes in Mathematics, 5:207–217, 1965. CODEN LNMAA2. ISBN 3-540-03349-1 (print), 3-662-21576-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-662-21576-0/1.pdf.

Serre:1965:CGP

[Ser65b] Jean-Pierre Serre. Cohomologie des groupes profinis. (French) [Cohomology of profinite groups]. Lecture Notes in Mathematics, 5: 1–86, 1965. CODEN LNMAA2. ISBN 3-540-03349-1 (print), 3-662-21576-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21576-0_1/. See erratum [Ser65h].

Serre:1965:CGCa

[Ser65c] Jean-Pierre Serre. Cohomologie galoisienne — cas commutatif. (French) [Galois cohomology — commutative case]. Lecture Notes in Mathematics, 5:87–136, 1965. CODEN LNMAA2. ISBN 3-540-03349-1 (print), 3-662-21576-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21576-0_2/. See [Ser65i].

Serre:1965:CGCb

[Ser65d] Jean-Pierre Serre, editor. Cohomologie Galoisienne: Cours au Collège de France, 1962–1963. (French) [Galois cohomology Courses at the Collège de France, 1962–1963], volume 5 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1965. CODEN LN-MAA2. ISBN 3-540-03349-1 (paperback), 3-662-21576-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28

no. 5; QA3 .L28 no. 5 1965. URL http://link.springer.com/
book/10.1007/978-3-662-21576-0.

Serre:1965:CGCc

[Ser65e] Jean-Pierre Serre, editor. Cohomologie Galoisienne: Cours au Collège de France, 1962–1963. (French) [Galois cohomology Courses at the Collège de France, 1962–1963], volume 5 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1965. CODEN LN-MAA2. ISBN 3-540-03349-1 (paperback), 3-662-21576-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28

no. 5; QA3 .L28 no. 5 1965. URL http://link.springer.com/
book/10.1007/978-3-662-21576-0.

Serre:1965:CGN

[Ser65f] Jean-Pierre Serre. Cohomologie galoisienne non commutative. (French) [Noncommutative Galois cohomology]. Lecture Notes in Mathematics, 5:137–181, 1965. CODEN LNMAA2. ISBN 3-540-03349-1 (print), 3-662-21576-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21576-0_3/. See erratum

[Ser65j].

Serre:1965:E

[Ser65g] Jean-Pierre Serre. Erratum. Lecture Notes in Mathematics, 5: 218, 1965. CODEN LNMAA2. ISBN 3-540-03349-1 (print), 3-662-21576-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21576-0_9.

Serre:1965:ECGa

[Ser65h] Jean-Pierre Serre. Erratum to: Cohomologie des Groupes Profinis. (French) [Erratum to: Cohomology of profinite groups]. Lecture Notes in Mathematics, 5:218, 1965. CODEN LNMAA2. ISBN 3-540-03349-1 (print), 3-662-21576-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21576-0_5. See [Ser65b].

Serre:1965:ECGb

[Ser65i] Jean-Pierre Serre. Erratum to: Cohomologie Galoisienne — Cas Commutatif. (French) [Erratum to: Noncommutative Galois cohomology]. Lecture Notes in Mathematics, 5:218, 1965. CODEN LNMAA2. ISBN 3-540-03349-1 (print), 3-662-21576-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-662-21576-0_6.

See [Ser65c].

Serre:1965:ECGc

[Ser65j] Jean-Pierre Serre. Erratum to: Cohomologie Galoisienne non Commutative. (French) [Noncommutative Galois cohomology]. Lecture Notes in Mathematics, 5:218, 1965. CODEN LNMAA2. ISBN 3-540-03349-1 (print), 3-662-21576-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21576-0_7. See [Ser65f].

Serre:1965:EDC

[Ser65k] Jean-Pierre Serre. Erratum to: Dualité dans la cohomologie des groupes profinis. (French) [Erratum to: Duality in the cohomology of profinite groups]. Lecture Notes in Mathematics, 5: 218, 1965. CODEN LNMAA2. ISBN 3-540-03349-1 (print), 3-662-21576-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21576-0_8. See [Ver65].

Serre:1965:FM

[Ser65l] Jean-Pierre Serre. Front matter. Lecture Notes in Mathematics, 5:N2–vii, 1965. CODEN LNMAA2. ISBN 3-540-03349-1 (print), 3-662-21576-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-662-21576-0/1.pdf.

Serre:1966:MF

[Ser66a] J.-P. Serre. Modular forms. Lecture Notes in Mathematics, 21:II–1–II–16, 1966. CODEN LNMAA2. ISBN 3-540-03604-0 (print), 3-540-34798-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091552/.

Serre:1966:SR

[Ser66b] J.-P. Serre. Statement of results. Lecture Notes in Mathematics, 21:I–1–I–8, 1966. CODEN LNMAA2. ISBN 3-540-03604-0 (print), 3-540-34798-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091551/.

Servien:1968:EFE

[Ser68] Claude Servien. Espaces de fonctions entières et fonctionnelles analytiques. (French) [Spaces of entire functions and analytic functionals]. Lecture Notes in Mathematics, 71:57–71, 1968. CODEN LNMAA2. ISBN 3-540-04241-5 (print), 3-540-35889-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080097/.

Serre:1965:ALM

[SG65] Jean-Pierre Serre and Peter Gabriel. Algèbre locale, multiplicités; cours au Collège de France, 1957–1958. (French) [Local algebra, multiplicities; courses at the Collège of France, 1957–1958], volume 11 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1965. CODEN LNMAA2. ISBN 3-540-07028-1 (paperback), 3-540-37123-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 156 pp. LCCN QA3 .L28 no. 11. URL http://link.

springer.com/book/10.1007/978-3-540-37123-6.

Shatz:1969:GT

[Sha69] Stephen S. Shatz. Galois theory. Lecture Notes in Mathematics, 86:146–158, 1969. CODEN LNMAA2. ISBN 3-540-04605-4 (print), 3-540-36095-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079388/.

Shimura:1968:AVS

[Shi68a] Goro Shimura. Abelian varieties and Siegel modular functions. Lecture Notes in Mathematics, 54:16–25, 1968. CODEN LN-MAA2. ISBN 3-540-04224-5 (print), 3-540-35832-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0071102/.

Shimura:1968:AFN

[Shi68b] Gorō Shimura. Automorphic Functions and Number Theory, volume 54 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. CO-DEN LNMAA2. ISBN 3-540-04224-5 (paperback), 3-540-35832-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 69

pp. LCCN QA3 .L28 no. 54. URL http://link.springer.com/
book/10.1007/BFb0071097.

Shimura:1968:AFU

[Shi68c] Goro Shimura. Automorphic functions on the upper half plane, especially modular functions. Lecture Notes in Mathematics, 54: 2–8, 1968. CODEN LNMAA2. ISBN 3-540-04224-5 (print), 3-540-35832-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071099/.

Shimura:1968:CFT

[Shi68d] Goro Shimura. The class-field-theoretical characterization of K′(ϕ(z)). Lecture Notes in Mathematics, 54:33–39, 1968. CO-DEN LNMAA2. ISBN 3-540-04224-5 (print), 3-540-35832-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0071104/.

Shimura:1968:ECF

[Shi68e] Goro Shimura. Elliptic curves and the fundamental theorems of the classical theory of complex multiplication. Lecture Notes in Mathematics, 54:8–13, 1968. CODEN LNMAA2. ISBN 3-540-04224-5 (print), 3-540-35832-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071100/.

Shimura:1968:ERA

[Shi68f] Goro Shimura. The endomorphism-ring of an Abelian variety; the field of moduli of an Abelian variety with many complex multiplications. Lecture Notes in Mathematics, 54:26–32, 1968. CODEN LNMAA2. ISBN 3-540-04224-5 (print), 3-540-35832-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071103/.

Shimura:1968:FGC

[Shi68g] Goro Shimura. Further generalization and concluding remarks. Lecture Notes in Mathematics, 54:62–65, 1968. CODEN LN-MAA2. ISBN 3-540-04224-5 (print), 3-540-35832-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0071108/.

Shimura:1968:FMC

[Shi68h] Goro Shimura. A further method of constructing class fields. Lecture Notes in Mathematics, 54:39–47, 1968. CODEN LN-MAA2. ISBN 3-540-04224-5 (print), 3-540-35832-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0071105/.

Shimura:1968:HZF

[Shi68i] Goro Shimura. The Hasse zeta function of an algebraic curve. Lecture Notes in Mathematics, 54:48–54, 1968. CODEN LN-MAA2. ISBN 3-540-04224-5 (print), 3-540-35832-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0071106/.

Shimura:1968:IGE

[Shi68j] Goro Shimura. Infinite Galois extensions with l-adic representations. Lecture Notes in Mathematics, 54:55–62, 1968. CODEN LNMAA2. ISBN 3-540-04224-5 (print), 3-540-35832-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071107/.

Shimura:1968:I

[Shi68k] Goro Shimura. Introduction. Lecture Notes in Mathematics, 54: 1–2, 1968. CODEN LNMAA2. ISBN 3-540-04224-5 (print), 3-540-35832-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071098.

Shimura:1968:RBP

[Shi68l] Goro Shimura. Relation between the points of finite order on an elliptic curve and the modular functions of higher level. Lecture Notes in Mathematics, 54:13–16, 1968. CODEN LNMAA2. ISBN 3-540-04224-5 (print), 3-540-35832-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071101/.

Siciak:1968:SAF

[Sic68] Jozef Siciak. Separately analytic functions and envelopes of holomorphy of some lower dimensional subsets of C^{n}. Lecture Notes in Mathematics, 71:21–32, 1968. CODEN LNMAA2. ISBN 3-540-04241-5 (print), 3-540-35889-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080093/.

Sieveking:1968:ISF

[Sie68] Malte Sieveking. Integraldarstellung superharmonischer Funktionen mit Anwendung auf parabolische Differentialgleichungen. (German) [Super integral representation of harmonic functions with application to parabolic differential equations]. Lecture Notes in Mathematics, 69:13–68, 1968. CODEN LNMAA2. ISBN 3-540-04239-3 (print), 3-540-35884-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080475/.

Simms:1968:CI

[Sim68a] D. J. Simms. SU(3): Charge and isospin. Lecture Notes in Mathematics, 52:76–85, 1968. CODEN LNMAA2. ISBN 3-540-04222-9 (print), 3-540-35829-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069924/.

Simms:1968:CRF

[Sim68b] D. J. Simms. Classification of the relativistic free particles. Lecture Notes in Mathematics, 52:56–71, 1968. CODEN LNMAA2. ISBN 3-540-04222-9 (print), 3-540-35829-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069922/.

Simms:1968:DE

[Sim68c] D. J. Simms. The Dirac equation. Lecture Notes in Mathematics, 52:72–75, 1968. CODEN LNMAA2. ISBN 3-540-04222-9 (print), 3-540-35829-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069923/.

Simms:1968:IR

[Sim68d] D. J. Simms. Induced representations. Lecture Notes in Mathematics, 52:42–47, 1968. CODEN LNMAA2. ISBN 3-540-04222-9 (print), 3-540-35829-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069920/.

Simms:1968:LAP

[Sim68e] D. J. Simms. Lie algebras and physical observables. Lecture Notes in Mathematics, 52:20–25, 1968. CODEN LNMAA2. ISBN 3-540-04222-9 (print), 3-540-35829-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069918/.

Simms:1968:LGQ

[Sim68f] D. J. Simms. Lie Groups and Quantum Mechanics, volume 52 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. CODEN LN-MAA2. ISBN 3-540-04222-9 (paperback), 3-540-35829-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 90 pp. LCCN

QA3 .L28 no. 52. URL http://link.springer.com/book/10.
1007/BFb0069914.

Simms:1968:LPR

[Sim68g] D. J. Simms. Lifting projective representations. Lecture Notes in Mathematics, 52:9–16, 1968. CODEN LNMAA2. ISBN 3-540-04222-9 (print), 3-540-35829-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069916/.

Simms:1968:RFP

[Sim68h] D. J. Simms. The relativistic free particle. Lecture Notes in Mathematics, 52:17–19, 1968. CODEN LNMAA2. ISBN 3-540-04222-9 (print), 3-540-35829-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069917/.

Simms:1968:RI

[Sim68i] D. J. Simms. Relativistic invariance. Lecture Notes in Mathematics, 52:1–8, 1968. CODEN LNMAA2. ISBN 3-540-04222-9 (print), 3-540-35829-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069915/.

Simms:1968:RSD

[Sim68j] D. J. Simms. Representations of semi-direct products. Lecture Notes in Mathematics, 52:48–55, 1968. CODEN LNMAA2. ISBN 3-540-04222-9 (print), 3-540-35829-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069921/.

Simms:1968:UEA

[Sim68k] D. J. Simms. Universal enveloping algebra. Lecture Notes in Mathematics, 52:26–41, 1968. CODEN LNMAA2. ISBN 3-540-04222-9 (print), 3-540-35829-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069919/.

Slomson:1968:MFP

[Slo68] A. Slomson. The monadic fragment of predicate calculus with the Chang quantifier and equality. Lecture Notes in Mathematics, 70:279–301, 1968. CODEN LNMAA2. ISBN 3-540-04240-7 (print), 3-540-35886-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079102/.

Smale:1969:GSQ

[Sma69] S. Smale. Global stability questions in dynamical systems. Lecture Notes in Mathematics, 103:150–158, 1969. CODEN LN-MAA2. ISBN 3-540-04622-4 (print), 3-540-36145-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0099989/.

Spitzer:1969:RPD

[Spi69] Frank Spitzer. Random processes defined through the interaction of an infinite particle system. Lecture Notes in Mathematics, 89:201–223, 1969. CODEN LNMAA2. ISBN 3-540-04608-9 (print), 3-540-36098-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079128/.

Szego:1969:ADI

[ST69a] G. P. Szegő and G. Treccani. Attrazione degli insiemi compatti. (Italian) [Attraction of compact sets]. Lecture Notes in Mathematics, 101:81–87, 1969. CODEN LNMAA2. ISBN 3-540-04620-8 (print), 3-540-36143-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100818/.

Szego:1969:CDF

[ST69b] G. P. Szegő and G. Treccani. Classificazione del flusso nell’intorno di un insieme compatto fortemente invariante. (Italian) [Classification of the flow in the neighborhood of a strongly-invariant compact set]. Lecture Notes in Mathematics, 101:93–96, 1969. CO-DEN LNMAA2. ISBN 3-540-04620-8 (print), 3-540-36143-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0100820/.

Szego:1969:CCD

[ST69c] G. P. Szegő and G. Treccani. Condizioni con le derivate semidefinite. (French) [Conditions with semidefinite derivatives]. Lecture Notes in Mathematics, 101:135–148, 1969. CODEN LN-MAA2. ISBN 3-540-04620-8 (print), 3-540-36143-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0100825/.

Szego:1969:FLP

[ST69d] G. P. Szegő and G. Treccani. Funzioni di Liapunov per flussi senza unicità. (Italian) [Liapunov functions for flows without uniqueness]. Lecture Notes in Mathematics, 101:97–105, 1969. CODEN LNMAA2. ISBN 3-540-04620-8 (print), 3-540-36143-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100821/.

Szego:1969:SDGa

[ST69e] G. P. Szegő and G. Treccani. I sistemi dinamici generati da un’equazione automoma. (Italian) [Dynamical systems generated by an autonomous equation]. Lecture Notes in Mathematics, 101: 7–11, 1969. CODEN LNMAA2. ISBN 3-540-04620-8 (print), 3-540-36143-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100808/.

Szego:1969:CTT

[ST69f] G. P. Szegő and G. Treccani. Il cono di traiettorie e le traiettorie di un sistema dinamico generalizzato. (Italian) [The cone of trajectories and the trajectories of a generalized dynamic system]. Lecture Notes in Mathematics, 101:26–34, 1969. CODEN LNMAA2. ISBN 3-540-04620-8 (print), 3-540-36143-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100811/.

Szego:1969:SDGb

[ST69g] G. P. Szegő and G. Treccani. Il sistema dinamico generalizzato generato dalla equazione autonoma ẋ = g(x). (Italian) [The generalized dynamical system generated by the autonomous equation ẋ = g(x)]. Lecture Notes in Mathematics, 101:35–39, 1969. CO-DEN LNMAA2. ISBN 3-540-04620-8 (print), 3-540-36143-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0100812/.

Szego:1969:ILI

[ST69h] G. P. Szegő and G. Treccani. Insiemi limite. (Italian) [Limit sets]. Lecture Notes in Mathematics, 101:46–62, 1969. CODEN LNMAA2. ISBN 3-540-04620-8 (print), 3-540-36143-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100814/.

Szego:1969:IMI

[ST69i] G. P. Szegő and G. Treccani. Insiemi minimi. (Italian) [Minimal sets]. Lecture Notes in Mathematics, 101:75, 1969. CODEN LN-MAA2. ISBN 3-540-04620-8 (print), 3-540-36143-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0100816.

Szego:1969:IIIa

[ST69j] G. P. Szegő and G. Treccani. Introduzione. (Italian) [Introduction]. Lecture Notes in Mathematics, 101:1–6, 1969. CODEN LNMAA2. ISBN 3-540-04620-8 (print), 3-540-36143-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100807/.

Szego:1969:IIIb

[ST69k] G. P. Szegő and G. Treccani. Invarianza. (Italian) [Invariance]. Lecture Notes in Mathematics, 101:40–45, 1969. CODEN LN-MAA2. ISBN 3-540-04620-8 (print), 3-540-36143-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0100813/.

Szego:1969:PEI

[ST69l] G. P. Szegő and G. Treccani. Prolongazioni ed insiemi limite prolongazionali. (Italian) [Prolongation and prolongated limit sets]. Lecture Notes in Mathematics, 101:63–74, 1969. CODEN LN-MAA2. ISBN 3-540-04620-8 (print), 3-540-36143-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0100815/.

Szego:1969:RLI

[ST69m] G. P. Szegő and G. Treccani. Risultati locali. (Italian) [Local results]. Lecture Notes in Mathematics, 101:106–112, 1969. CODEN LNMAA2. ISBN 3-540-04620-8 (print), 3-540-36143-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100822/.

Szego:1969:STMb

[ST69n] G. P. Szegő and G. Treccani. Semigruppi di Trasformazioni Multivoche. (Italian) [Semigroups of multi-vocal transformations], volume 101 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CO-DEN LNMAA2. ISBN 3-540-04620-8 (paperback), 3-540-36143-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 177

pp. LCCN QA3 .L28 no. 101. URL http://link.springer.com/
book/10.1007/BFb0100806.

Szego:1969:STMa

[ST69o] G. P. Szegő and G. Treccani. Semigruppi di trasformazioni multivoche o sistemi dinamici generalizzati-depinizioni. (Italian) [Semigroups of transformations or multi-vocal-depinizioni generalized dynamical systems]. Lecture Notes in Mathematics, 101: 20–25, 1969. CODEN LNMAA2. ISBN 3-540-04620-8 (print), 3-540-36143-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100810/.

Szego:1969:SDO

[ST69p] G. P. Szegő and G. Treccani. Sistemi dinamici ordinari. (Italian) [Ordinary dynamical systems]. Lecture Notes in Mathematics, 101:12–19, 1969. CODEN LNMAA2. ISBN 3-540-04620-8 (print), 3-540-36143-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100809/.

Szego:1969:SAD

[ST69q] G. P. Szegő and G. Treccani. Stabilità asintotica degli insiemi compatti. (Italian) [Asymptotic stability of compact sets]. Lecture Notes in Mathematics, 101:88–92, 1969. CODEN LNMAA2. ISBN 3-540-04620-8 (print), 3-540-36143-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100819/.

Szego:1969:SIC

[ST69r] G. P. Szegő and G. Treccani. Stabilità di insiemi compatti. (Italian) [Stability of compact sets]. Lecture Notes in Mathematics, 101:76–80, 1969. CODEN LNMAA2. ISBN 3-540-04620-8 (print), 3-540-36143-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100817/.

Szego:1969:SDC

[ST69s] G. P. Szegő and G. Treccani. Struttura delle curve di livello. (Italian) [Structure of contour lines]. Lecture Notes in Mathematics, 101:122–134, 1969. CODEN LNMAA2. ISBN 3-540-04620-8 (print), 3-540-36143-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100824/.

Szego:1969:TEL

[ST69t] G. P. Szegő and G. Treccani. Teoremi di estensione locale e globale. (Italian) [Local and global extension theorems]. Lecture Notes in Mathematics, 101:113–121, 1969. CODEN LNMAA2. ISBN 3-540-04620-8 (print), 3-540-36143-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100823/.

Szego:1969:TTR

[ST69u] G. P. Szegő and G. Treccani. Teoremi tipo rolle. (Italian) [Rolletype theorems]. Lecture Notes in Mathematics, 101:149–157, 1969. CODEN LNMAA2. ISBN 3-540-04620-8 (print), 3-540-36143-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0100826/.

Stauffer:1969:DFI

[Sta69] H. B. Stauffer. Derived functors without injectives. Lecture Notes in Mathematics, 86:159–166, 1969. CODEN LNMAA2. ISBN 3-540-04605-4 (print), 3-540-36095-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079389/.

Stolzenberg:1966:AVM

[Sto66a] Gabriel Stolzenberg. Analytic varieties minimize volume. Lecture Notes in Mathematics, 19:6–15, 1966. CODEN LNMAA2. ISBN 3-540-03602-4 (print), 3-540-37194-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097738/.

Stolzenberg:1966:I

[Sto66b] Gabriel Stolzenberg. Introduction. Lecture Notes in Mathematics, 19:1–5, 1966. CODEN LNMAA2. ISBN 3-540-03602-4 (print), 3-540-37194-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097737/.

Stolzenberg:1966:LLB

[Sto66c] Gabriel Stolzenberg. A local lower bound for the volume of an analytic variety. Lecture Notes in Mathematics, 19:16–20, 1966. CODEN LNMAA2. ISBN 3-540-03602-4 (print), 3-540-37194-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097739/.

Stolzenberg:1966:VLE

[Sto66d] Gabriel Stolzenberg. Volumes, Limits, and Extensions of Analytic Varieties, volume 19 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1966. CODEN LNMAA2. ISBN 3-540-03602-4 (paperback), 3-540-37194-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 45 pp. LCCN QA3 .L28 no. 19; QA3L4 v.19. URL

http://link.springer.com/book/10.1007/BFb0097736.

Strauss:1968:RRP

[Str68a] Aaron Strauss. Recent results in perturbation theory. Lecture Notes in Mathematics, 60:1–6, 1968. CODEN LNMAA2. ISBN 3-540-04230-X (print), 3-540-35862-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077121/.

Strauss:1968:ULF

[Str68b] Aaron Strauss. The use of Liapunov functions for global existence. Lecture Notes in Mathematics, 60:76–82, 1968. CODEN LNMAA2. ISBN 3-540-04230-X (print), 3-540-35862-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077130/.

Stummel:1969:ATR

[Stu69a] F. Stummel. Allgemeine Theorie der Randwertaufgaben. (German) [General theory of boundary value problems]. Lecture Notes in Mathematics, 102:52–107, 1969. CODEN LNMAA2. ISBN 3-540-04621-6 (print), 3-540-36144-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059063/.

Stummel:1969:BBO

[Stu69b] F. Stummel. Beschränkte Bilinearformen und Operatoren. (German) [Limited bilinear forms and operators]. Lecture Notes in Mathematics, 102:1–28, 1969. CODEN LNMAA2. ISBN 3-540-04621-6 (print), 3-540-36144-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059061/.

Stummel:1969:BOE

[Stu69c] F. Stummel. Bilinearformen, Operatoren und V -Elliptizität. (German) [Bilinear forms, operators, and V -ellipticity]. Lecture Notes in Mathematics, 102:134–176, 1969. CODEN LNMAA2. ISBN 3-540-04621-6 (print), 3-540-36144-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059065/.

Stummel:1969:RGS

[Stu69d] F. Stummel. Die Räume H^{∞} und H_{0}^{∞}. (German) [The spaces H^{∞} und H_{0}^{∞}]. Lecture Notes in Mathematics, 102:177–203, 1969. CODEN LNMAA2. ISBN 3-540-04621-6 (print), 3-540-36144-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0059066/.

Stummel:1969:DRS

[Stu69e] F. Stummel. Dirichletsche Randwertaufgaben, Sobolewsche Funktionenräume W^{m,2}(Ω; p) und semielliptische Differentialoperatoren. (German) [Dirichlet boundary value problems, Sobolev function spaces P W^{m,2}(Ω, p) and semi-elliptic differential operators]. Lecture Notes in Mathematics, 102:287–338, 1969. CODEN LN-MAA2. ISBN 3-540-04621-6 (print), 3-540-36144-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0059070/.

Stummel:1969:ESP

[Stu69f] F. Stummel. Elliptische Systeme partieller Differentialgleichungen für periodische Funktionen. (German) [Elliptic systems of partial differential equations for periodic functions]. Lecture Notes in Mathematics, 102:257–286, 1969. CODEN LNMAA2. ISBN 3-540-04621-6 (print), 3-540-36144-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059069/.

Stummel:1969:LSF

[Stu69g] F. Stummel. Lineare stetige Funktionale auf dem Testraum H_{0}^{∞}. (German) [Linear continuous functionals on the test space H_{0}^{∞}]. Lecture Notes in Mathematics, 102:204–227, 1969. CODEN LN-MAA2. ISBN 3-540-04621-6 (print), 3-540-36144-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0059067/.

Stummel:1969:RES

[Stu69h] F. Stummel. Rand- und Eigenwertaufgaben in Sobolewschen Räumen. (German) [Boundary and eigenvalue problems in Sobolev spaces], volume 102 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CODEN LNMAA2. ISBN 3-540-04621-6 (paperback), 3-540-36144-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 386 pp. LCCN QA3 .L28 no. 102. URL http://

link.springer.com/book/10.1007/BFb0059060.

Stummel:1969:RRG

[Stu69i] F. Stummel. Reguläre Randwertaufgaben bei gewöhnlichen Differentialgleichungen. (German) [Regular boundary value problems for ordinary differential equations]. Lecture Notes in Mathematics, 102:228–256, 1969. CODEN LNMAA2. ISBN 3-540-04621-6 (print), 3-540-36144-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059068/.

Stummel:1969:SRG

[Stu69j] F. Stummel. Sobolewsche Räume. (German) [Sobolev spaces]. Lecture Notes in Mathematics, 102:108–133, 1969. CODEN LN-MAA2. ISBN 3-540-04621-6 (print), 3-540-36144-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0059064/.

Stummel:1969:VBO

[Stu69k] F. Stummel. Vollstetige Bilinearformen und Operatoren. (German) [Completely continuous bilinear forms and operators]. Lecture Notes in Mathematics, 102:29–51, 1969. CODEN LNMAA2. ISBN 3-540-04621-6 (print), 3-540-36144-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059062/.

Sunyach:1968:TDG

[Sun68] C. Sunyach. Sur le théorème du graphe fermé. (French) [The closed graph theorem]. Lecture Notes in Mathematics, 71:33–37, 1968. CODEN LNMAA2. ISBN 3-540-04241-5 (print), 3-540-35889-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0080094/.

Swan:1968:Ka

[Swa68a] R. G. Swan. K_{1}(A). Lecture Notes in Mathematics, 76:193– 204, 1968. CODEN LNMAA2. ISBN 3-540-04245-8 (print), 3-540-35917-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080296/.

Swan:1968:Kb

[Swa68b] R. G. Swan. K_{2}(R). Lecture Notes in Mathematics, 76:204– 211, 1968. CODEN LNMAA2. ISBN 3-540-04245-8 (print), 3-540-35917-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080297/.

Swan:1968:GR

[Swa68c] R. G. Swan. K_{O} of graded rings. Lecture Notes in Mathematics, 76:124–131, 1968. CODEN LNMAA2. ISBN 3-540-04245-8 (print), 3-540-35917-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080289/.

Swan:1968:SRH

[Swa68d] R. G. Swan. Spec(R) and H(R). Lecture Notes in Mathematics, 76:132–145, 1968. CODEN LNMAA2. ISBN 3-540-04245-8 (print), 3-540-35917-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080290/.

Swan:1968:AT

[Swa68e] R. G. Swan. Algebraic K-theory, volume 76 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. CODEN LNMAA2. ISBN 3-540-04245-8 (paperback), 3-540-35917-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 262 pp. LCCN QA3

.L28 no. 76. URL http://link.springer.com/book/10.1007/
BFb0080281.

Swan:1968:BTR

[Swa68f] R. G. Swan. Basic topological remarks. Lecture Notes in Mathematics, 76:155–160, 1968. CODEN LNMAA2. ISBN 3-540-04245-8 (print), 3-540-35917-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080292/.

Swan:1968:CTb

[Swa68g] R. G. Swan. Cancellation theorems. Lecture Notes in Mathematics, 76:183–193, 1968. CODEN LNMAA2. ISBN 3-540-04245-8 (print), 3-540-35917-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080295/.

Swan:1968:CTa

[Swa68h] R. G. Swan. Category theory. Lecture Notes in Mathematics, 76:1–40, 1968. CODEN LNMAA2. ISBN 3-540-04245-8 (print), 3-540-35917-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080282/.

Swan:1968:CCN

[Swa68i] R. G. Swan. Chain complexes and the nilpotence of K̃_{0}(R). Lecture Notes in Mathematics, 76:161–169, 1968. CODEN LNMAA2. ISBN 3-540-04245-8 (print), 3-540-35917-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080293/.

Swan:1968:CB

[Swa68j] R. G. Swan. The connection between K_{O}(R) and G_{O}(R). Lecture Notes in Mathematics, 76:100–109, 1968. CODEN LNMAA2. ISBN 3-540-04245-8 (print), 3-540-35917-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080287/.

Swan:1968:DE

[Swa68k] R. G. Swan. Definition of G(R) and examples. Lecture Notes in Mathematics, 76:92–100, 1968. CODEN LNMAA2. ISBN 3-540-04245-8 (print), 3-540-35917-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080286/.

Swan:1968:DSE

[Swa68l] R. G. Swan. Definition of K_{O}(A) and some examples. Lecture Notes in Mathematics, 76:65–75, 1968. CODEN LNMAA2. ISBN 3-540-04245-8 (print), 3-540-35917-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080284/.

Swan:1968:ESK

[Swa68m] R. G. Swan. The exact sequence of K_{i}’s. Lecture Notes in Mathematics, 76:211–223, 1968. CODEN LNMAA2. ISBN 3-540-04245-8 (print), 3-540-35917-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080298/.

Swan:1968:FR

[Swa68n] R. G. Swan. Further results on K_{1} and K_{0}. Lecture Notes in Mathematics, 76:224–247, 1968. CODEN LNMAA2. ISBN 3-540-04245-8 (print), 3-540-35917-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080299/.

Swan:1968:KST

[Swa68o] R. G. Swan. Krull–Schmidt theorems and applications. Lecture Notes in Mathematics, 76:75–92, 1968. CODEN LNMAA2. ISBN 3-540-04245-8 (print), 3-540-35917-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080285/.

Swan:1968:LRB

[Swa68p] R. G. Swan. Localization and relation between G_{O}(R) and G_{O}(R_{S}). Lecture Notes in Mathematics, 76:109–124, 1968. CO-DEN LNMAA2. ISBN 3-540-04245-8 (print), 3-540-35917-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0080288/.

Swan:1968:PGD

[Swa68q] R. G. Swan. Picard group and the determinant. Lecture Notes in Mathematics, 76:146–155, 1968. CODEN LNMAA2. ISBN 3-540-04245-8 (print), 3-540-35917-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080291/.

Swan:1968:QC

[Swa68r] R. G. Swan. Quotient categories. Lecture Notes in Mathematics, 76:40–64, 1968. CODEN LNMAA2. ISBN 3-540-04245-8 (print), 3-540-35917-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080283/.

Swan:1968:RBA

[Swa68s] R. G. Swan. Relations between algebraic and topological K theory. Lecture Notes in Mathematics, 76:247–256, 1968. CODEN LNMAA2. ISBN 3-540-04245-8 (print), 3-540-35917-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080300/.

Swan:1968:ST

[Swa68t] R. G. Swan. Serre’s theorem. Lecture Notes in Mathematics, 76:170–183, 1968. CODEN LNMAA2. ISBN 3-540-04245-8 (print), 3-540-35917-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080294/.

Taam:1969:LMA

[TA69] C. T. Taam and Michael Francis Atiyah, editors. Lectures in Modern Analysis and Applications I, volume 103 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CODEN LNMAA2. ISBN 3-540-04622-4 (paperback), 3-540-36145-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 162 pp. LCCN QA3

.L28 no. 103. URL http://link.springer.com/book/10.1007/
BFb0099980.

Tait:1968:NDC

[Tai68] W. W. Tait. Normal derivability in classical logic. Lecture Notes in Mathematics, 72:204–236, 1968. CODEN LNMAA2. ISBN 3-540-04242-3 (print), 3-540-35900-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079691/.

Takeuti:1968:CSR

[Tak68a] Gaisi Takeuti. The Π_{1}^{1}-comprehension schema and ω-rules. Lecture Notes in Mathematics, 70:303–331, 1968. CODEN LNMAA2. ISBN 3-540-04240-7 (print), 3-540-35886-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079103/.

Takeuti:1968:DL

[Tak68b] Gaisi Takeuti. A determinate logic. Lecture Notes in Mathematics, 72:237–264, 1968. CODEN LNMAA2. ISBN 3-540-04242-3 (print), 3-540-35900-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079692/.

Tanaka:1968:RSM

[Tan68] Hiroshi Tanaka. Riemann surfaces with Martin and Kuramochi boundary points. Lecture Notes in Mathematics, 58:30–42, 1968. CODEN LNMAA2. ISBN 3-540-04228-8 (print), 3-540-35859-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0098444/.

Targonski:1967:CRO

[Tar67a] György I. Targonski. Concluding remarks. Open questions. Lecture Notes in Mathematics, 33:104–107, 1967. CODEN LNMAA2. ISBN 3-540-03904-X (print), 3-540-34972-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082500/.

Targonski:1967:CIC

[Tar67b] György I. Targonski. Continuous iteration. Commuting substitution operators. Lecture Notes in Mathematics, 33:31–44, 1967. CODEN LNMAA2. ISBN 3-540-03904-X (print), 3-540-34972-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082490/.

Targonski:1967:CT

[Tar67c] György I. Targonski. Convergence theorems. Lecture Notes in Mathematics, 33:92–98, 1967. CODEN LNMAA2. ISBN 3-540-03904-X (print), 3-540-34972-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082497/.

Targonski:1967:IOR

[Tar67d] György I. Targonski. Integral operators: Remarks and definitions. Lecture Notes in Mathematics, 33:61–69, 1967. CODEN LNMAA2. ISBN 3-540-03904-X (print), 3-540-34972-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082493/.

Targonski:1967:IS

[Tar67e] György I. Targonski. Introduction and summary. Lecture Notes in Mathematics, 33:1–4, 1967. CODEN LNMAA2. ISBN 3-540-03904-X (print), 3-540-34972-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082487/.

Targonski:1967:MTS

[Tar67f] György I. Targonski. Multiplication theorems and square theorems; Bourlet operators. Lecture Notes in Mathematics, 33:5–20, 1967. CODEN LNMAA2. ISBN 3-540-03904-X (print), 3-540-34972-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082488/.

Targonski:1967:OLB

[Tar67g] György I. Targonski. Operators of locally bounded range. Lecture Notes in Mathematics, 33:102–103, 1967. CODEN LNMAA2. ISBN 3-540-03904-X (print), 3-540-34972-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082499.

Targonski:1967:RRC

[Tar67h] György I. Targonski. Recent results on Carleman operators. Lecture Notes in Mathematics, 33:75–85, 1967. CODEN LN-MAA2. ISBN 3-540-03904-X (print), 3-540-34972-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082495/.

Targonski:1967:RBB

[Tar67i] György I. Targonski. Relations between Bourlet operators of the specific form. Lecture Notes in Mathematics, 33:45–55, 1967. CO-DEN LNMAA2. ISBN 3-540-03904-X (print), 3-540-34972-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0082491/.

Targonski:1967:SFO

[Tar67j] György I. Targonski, editor. Seminar on Functional Operators and Equations: Forschungsinstitut für Mathematik, ETH, Zürich October 1965–July 1966, volume 33 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. CODEN LNMAA2. ISBN 3-540-03904-X (paperback), 3-540-34972-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 33. URL http://

link.springer.com/book/10.1007/BFb0082486.

Targonski:1967:SCO

[Tar67k] György I. Targonski. Strong Carleman operators. Lecture Notes in Mathematics, 33:86–91, 1967. CODEN LNMAA2. ISBN 3-540-03904-X (print), 3-540-34972-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082496/.

Targonski:1967:SOS

[Tar67l] György I. Targonski. Substitution operators. The Schröder equation. Lecture Notes in Mathematics, 33:21–30, 1967. CODEN LNMAA2. ISBN 3-540-03904-X (print), 3-540-34972-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082489/.

Targonski:1967:SSO

[Tar67m] György I. Targonski. Superposition of substitution operators. Lecture Notes in Mathematics, 33:56–60, 1967. CODEN LN-MAA2. ISBN 3-540-03904-X (print), 3-540-34972-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082492/.

Targonski:1967:TWN

[Tar67n] György I. Targonski. The theorems of Weyl and von Neumann on Hermitean Carleman operators. Lecture Notes in Mathematics, 33:70–74, 1967. CODEN LNMAA2. ISBN 3-540-03904-X (print), 3-540-34972-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082494/.

Targonski:1967:TSK

[Tar67o] György I. Targonski. Transformation of strong kernels. Lecture Notes in Mathematics, 33:99–101, 1967. CODEN LNMAA2. ISBN 3-540-03904-X (print), 3-540-34972-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082498/.

Thomas:1966:CMP

[Tho66a] Emery Thomas. The “classical” (Moore–Postnikov) method of decomposing a fibration. Lecture Notes in Mathematics, 13:12– 18, 1966. CODEN LNMAA2. ISBN 3-540-03596-6 (print), 3-540-37168-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097867/.

Thomas:1966:CPI

[Tho66b] Emery Thomas. Computing Postnikov invariants. Lecture Notes in Mathematics, 13:35–44, 1966. CODEN LNMAA2. ISBN 3-540-03596-6 (print), 3-540-37168-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097871/.

Thomas:1966:Ib

[Tho66c] Emery Thomas. An illustration. Lecture Notes in Mathematics, 13:29–34, 1966. CODEN LNMAA2. ISBN 3-540-03596-6 (print), 3-540-37168-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097870/.

Thomas:1966:Ia

[Tho66d] Emery Thomas. Introduction. Lecture Notes in Mathematics, 13: 1–6, 1966. CODEN LNMAA2. ISBN 3-540-03596-6 (print), 3-540-37168-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097865/.

Thomas:1966:KHG

[Tho66e] Emery Thomas. Killing homotopy groups. Lecture Notes in Mathematics, 13:23–28, 1966. CODEN LNMAA2. ISBN 3-540-03596-6 (print), 3-540-37168-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097869/.

Thomas:1966:OSB

[Tho66f] Emery Thomas. Oriented sphere bundles. Lecture Notes in Mathematics, 13:19–22, 1966. CODEN LNMAA2. ISBN 3-540-03596-6 (print), 3-540-37168-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097868/.

Thomas:1966:PFS

[Tho66g] Emery Thomas. Principal fiber spaces. Lecture Notes in Mathematics, 13:7–11, 1966. CODEN LNMAA2. ISBN 3-540-03596-6 (print), 3-540-37168-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097866/.

Thomas:1966:SFS

[Tho66h] Emery Thomas, editor. Seminar on Fiber Spaces: Lectures delivered in 1964 in Berkeley and 1965 in Zürich Berkeley, volume 13 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1966. CODEN LN-MAA2. ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN

QA3 .L28 no. 13. URL http://link.springer.com/book/10.

1007/BFb0097864. Berkeley notes edited by J. F. McClendon.

Thomas:1969:OGA

[Tho69] C. B. Thomas. Obstructions for group actions on S^{2n−1}. Lecture Notes in Mathematics, 108:78–86, 1969. CODEN LNMAA2. ISBN 3-540-04627-5 (print), 3-540-36156-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059993/.

Tierney:1969:ADC

[Tie69a] Myles Tierney. Adding directed colimits to a subcategory. Lecture Notes in Mathematics, 87:10–29, 1969. CODEN LNMAA2. ISBN 3-540-04606-2 (print), 3-540-36096-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101428/.

Tierney:1969:ACM

[Tie69b] Myles Tierney. Autonomous categories with models. Lecture Notes in Mathematics, 106:130–165, 1969. CODEN LNMAA2. ISBN 3-540-04625-9 (print), 3-540-36150-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059144/.

Tierney:1969:CCS

[Tie69c] Myles Tierney, editor. Categorical Constructions in Stable Homotopy Theory: A Seminar given at the ETH, Zürich, in 1967, volume 87 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CO-DEN LNMAA2. ISBN 3-540-04606-2 (paperback), 3-540-36096-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN

QA3 .L28 no. 87. URL http://link.springer.com/book/10.
1007/BFb0101425.

Tierney:1969:FS

[Tie69d] Myles Tierney. FD-spectra. Lecture Notes in Mathematics, 87: 47–63, 1969. CODEN LNMAA2. ISBN 3-540-04606-2 (print), 3-540-36096-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101430/.

Tierney:1969:I

[Tie69e] Myles Tierney. Introduction. Lecture Notes in Mathematics, 87: 1–2, 1969. CODEN LNMAA2. ISBN 3-540-04606-2 (print), 3-540-36096-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101426.

Tierney:1969:IE

[Tie69f] Myles Tierney. Inverting an endomorphism. Lecture Notes in Mathematics, 87:2–9, 1969. CODEN LNMAA2. ISBN 3-540-04606-2 (print), 3-540-36096-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101427/.

Tierney:1969:SS

[Tie69g] Myles Tierney. Simplicial spectra. Lecture Notes in Mathematics, 87:30–47, 1969. CODEN LNMAA2. ISBN 3-540-04606-2 (print), 3-540-36096-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101429/.

Tits:1967:BGD

[Tit67a] Jacques Tits. Bezeichnungen. (German) [Designations]. Lecture Notes in Mathematics, 40:23–51, 1967. CODEN LNMAA2. ISBN 3-540-03911-2 (print), 3-540-35169-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080327/.

Tits:1967:KAG

[Tit67b] Jacques Tits. Komplex-analytische Gruppen. (German) [Complexanalytic groups]. Lecture Notes in Mathematics, 40:1–13, 1967. CODEN LNMAA2. ISBN 3-540-03911-2 (print), 3-540-35169-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0080325/.

Tits:1967:RAG

[Tit67c] Jacques Tits. Reell-analytische gruppen. (German) [Real analytical groups]. Lecture Notes in Mathematics, 40:13–22, 1967. CO-DEN LNMAA2. ISBN 3-540-03911-2 (print), 3-540-35169-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0080326/.

Tits:1967:TEL

[Tit67d] Jacques Tits. Tabellen zu den einfachen Lie Gruppen und ihren Darstellungen. (German) [Tables of the simple Lie groups and their representations], volume 40 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. CODEN LNMAA2. ISBN 3-540-03911-2 (paperback), 3-540-35169-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 53 pp. LCCN QA3 .L28 no. 40. URL

http://link.springer.com/book/10.1007/BFb0080324.

Tondeur:1965:EML

[Ton65a] Philippe Tondeur. The exponential map of a Lie group. Lecture Notes in Mathematics, 7:103–127, 1965. CODEN LNMAA2. ISBN 3-540-04599-6 (print), 3-540-36089-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097835/.

Tondeur:1965:M

[Ton65b] Philippe Tondeur. G-manifolds. Lecture Notes in Mathematics, 7:34–39, 1965. CODEN LNMAA2. ISBN 3-540-04599-6 (print), 3-540-36089-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097832/.

Tondeur:1965:O

[Ton65c] Philippe Tondeur. G-objects. Lecture Notes in Mathematics, 7: 1–27, 1965. CODEN LNMAA2. ISBN 3-540-04599-6 (print), 3-540-36089-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097830/.

Tondeur:1965:S

[Ton65d] Philippe Tondeur. G-spaces. Lecture Notes in Mathematics, 7: 28–33, 1965. CODEN LNMAA2. ISBN 3-540-04599-6 (print), 3-540-36089-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097831/.

Tondeur:1965:GA

[Ton65e] Philippe Tondeur. Groups of automorphisms. Lecture Notes in Mathematics, 7:160–169, 1965. CODEN LNMAA2. ISBN 3-540-04599-6 (print), 3-540-36089-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097837/.

Tondeur:1965:ILG

[Ton65f] Philippe Tondeur. Introduction to Lie Groups and Transformation Groups, volume 7 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1965. CODEN LNMAA2. ISBN 3-540-04599-6 (paperback), 3-540-36089-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 176 pp. LCCN QA3 .L28 no. 7. URL http://link.

springer.com/book/10.1007/BFb0097829.

Tondeur:1965:SS

[Ton65g] Philippe Tondeur. Subgroups and subalgebras. Lecture Notes in Mathematics, 7:128–159, 1965. CODEN LNMAA2. ISBN 3-540-04599-6 (print), 3-540-36089-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097836/.

Tondeur:1965:V

[Ton65h] Philippe Tondeur. Vectorfields. Lecture Notes in Mathematics, 7:40–65, 1965. CODEN LNMAA2. ISBN 3-540-04599-6 (print), 3-540-36089-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097833/.

Tondeur:1965:VPG

[Ton65i] Philippe Tondeur. Vectorfields and 1-parameter groups of transformations. Lecture Notes in Mathematics, 7:66–102, 1965. CODEN LNMAA2. ISBN 3-540-04599-6 (print), 3-540-36089-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097834/.

Tondeur:1969:ILG

[Ton69] Philippe Tondeur. Introduction to Lie Groups and Transformation Groups, volume 7 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1969. CODEN LNMAA2. ISBN 3-540-04599-6 (paperback), 3-540-36089-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 176 pp. LCCN QA3 .L28 no. 7;

QA3 .L28 no. 7, 1969. URL http://link.springer.com/book/
10.1007/BFb0097829.

Treves:1969:LCP

[Tre69] F. Treves. Local Cauchy problem for linear partial differential equations with analytic coefficients. Lecture Notes in Mathematics, 103:72–100, 1969. CODEN LNMAA2. ISBN 3-540-04622-4 (print), 3-540-36145-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099986/.

Troelstra:1969:ACP

[Tro69a] A. S. Troelstra. Applications of the continuity principles and the fan theorem. Lecture Notes in Mathematics, 95:71–75, 1969. CO-DEN LNMAA2. ISBN 3-540-04614-3 (print), 3-540-36130-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0080656/.

Troelstra:1969:BTC

[Tro69b] A. S. Troelstra. Brouwer’s theory of the creative subject. Lecture Notes in Mathematics, 95:95–107, 1969. CODEN LNMAA2. ISBN 3-540-04614-3 (print), 3-540-36130-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080659/.

Troelstra:1969:CS

[Tro69c] A. S. Troelstra. Choice sequences. Lecture Notes in Mathematics, 95:44–56, 1969. CODEN LNMAA2. ISBN 3-540-04614-3 (print), 3-540-36130-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080653/.

Troelstra:1969:CLA

[Tro69d] A. S. Troelstra. Constructive or lawlike analysis. Lecture Notes in Mathematics, 95:29–33, 1969. CODEN LNMAA2. ISBN 3-540-04614-3 (print), 3-540-36130-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080651/.

Troelstra:1969:EA

[Tro69e] A. S. Troelstra. Elementary arithmetic. Lecture Notes in Mathematics, 95:12–13, 1969. CODEN LNMAA2. ISBN 3-540-04614-3 (print), 3-540-36130-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080646.

Troelstra:1969:ETR

[Tro69f] A. S. Troelstra. Elementary theory of real numbers. Lecture Notes in Mathematics, 95:22–26, 1969. CODEN LNMAA2. ISBN 3-540-04614-3 (print), 3-540-36130-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080649/.

Troelstra:1969:I

[Tro69g] A. S. Troelstra. Introduction. Lecture Notes in Mathematics, 95: 2–4, 1969. CODEN LNMAA2. ISBN 3-540-04614-3 (print), 3-540-36130-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080644/.

Troelstra:1969:LSN

[Tro69h] A. S. Troelstra. Lawless sequences of natural numbers. Lecture Notes in Mathematics, 95:34–43, 1969. CODEN LNMAA2. ISBN 3-540-04614-3 (print), 3-540-36130-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080652/.

Troelstra:1969:L

[Tro69i] A. S. Troelstra. Logic. Lecture Notes in Mathematics, 95:5– 11, 1969. CODEN LNMAA2. ISBN 3-540-04614-3 (print), 3-540-36130-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080645/.

Troelstra:1969:ORO

[Tro69j] A. S. Troelstra. Ordering relations and order on the real line. Lecture Notes in Mathematics, 95:26–28, 1969. CODEN LN-MAA2. ISBN 3-540-04614-3 (print), 3-540-36130-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080650/.

Troelstra:1969:SCL

[Tro69k] A. S. Troelstra. Sequences and constructive (lawlike) objects. Lecture Notes in Mathematics, 95:16–21, 1969. CODEN LN-MAA2. ISBN 3-540-04614-3 (print), 3-540-36130-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080648/.

Troelstra:1969:S

[Tro69l] A. S. Troelstra. Species. Lecture Notes in Mathematics, 95:14– 16, 1969. CODEN LNMAA2. ISBN 3-540-04614-3 (print), 3-540-36130-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080647/.

Troelstra:1969:SRR

[Tro69m] A. S. Troelstra. Species revisited; the role of the comprehension principle. Lecture Notes in Mathematics, 95:91–94, 1969. CO-DEN LNMAA2. ISBN 3-540-04614-3 (print), 3-540-36130-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0080658/.

Troelstra:1969:STR

[Tro69n] A. S. Troelstra. Spreads and a theory of real numbers. Lecture Notes in Mathematics, 95:57–64, 1969. CODEN LNMAA2. ISBN 3-540-04614-3 (print), 3-540-36130-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080654/.

Troelstra:1969:TSM

[Tro69o] A. S. Troelstra. Topology; separable metric spaces. Lecture Notes in Mathematics, 95:64–70, 1969. CODEN LNMAA2. ISBN 3-540-04614-3 (print), 3-540-36130-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080655/.

Troelstra:1969:WOO

[Tro69p] A. S. Troelstra. Well-orderings and ordinals. Lecture Notes in Mathematics, 95:76–90, 1969. CODEN LNMAA2. ISBN 3-540-04614-3 (print), 3-540-36130-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080657/.

Troelstra:1969:PIL

[Tro69q] Anne Sjerp Troelstra, editor. Principles of Intuitionism: Lectures presented at the summer conference on Intuitionism and Proof theory (1968) at SUNY at Buffalo, N.Y., volume 95 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CODEN LNMAA2. ISBN 3-540-04614-3 (paperback), 3-540-36130-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 95. URL

http://link.springer.com/book/10.1007/BFb0080643.

Tierney:1969:SDF

[TV69] Myles Tierney and Wolfgang Vogel. Simplicial derived functors. Lecture Notes in Mathematics, 86:167–180, 1969. CODEN LN-MAA2. ISBN 3-540-04605-4 (print), 3-540-36095-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079390/.

Ulmer:1969:AMK

[Ulm69a] Friedrich Ulmer. Acyclic models and Kan extensions. Lecture Notes in Mathematics, 86:181–204, 1969. CODEN LNMAA2. ISBN 3-540-04605-4 (print), 3-540-36095-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079391/.

Ulmer:1969:KEC

[Ulm69b] Friedrich Ulmer. Kan extensions, cotriples and André (co) homology. Lecture Notes in Mathematics, 92:278–308, 1969. CODEN LNMAA2. ISBN 3-540-04611-9 (print), 3-540-36101-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080773/.

Ulmer:1969:CAC

[Ulm69c] Friedrich Ulmer. On cotriple and André (co)homology, their relationship with classical homological algebra. Lecture Notes in Mathematics, 80:376–398, 1969. CODEN LNMAA2. ISBN 3-540-04601-1 (print), 3-540-36091-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083090/.

Urabe:1969:NSB

[Ura69] Minoru Urabe. Numerical solution of boundary value problems in Chebyshev series — a method of computation and error estimation. Lecture Notes in Mathematics, 109:40–86, 1969. CODEN LNMAA2. ISBN 3-540-04628-3 (print), 3-540-36158-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060016/.

Urbanik:1967:HT

[Urb67a] K. Urbanik. Herglotz’ theorem. Lecture Notes in Mathematics, 44:2–6, 1967. CODEN LNMAA2. ISBN 3-540-03915-5 (print), 3-540-35524-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074244/.

Urbanik:1967:LPT

[Urb67b] K. Urbanik, editor. Lectures on Prediction Theory: Delivered at the University Erlangen–Nürnberg 1966, volume 44 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. CODEN LNMAA2. ISBN 3-540-03915-5 (paperback), 3-540-35524-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN

QA3 .L28 no. 44. URL http://link.springer.com/book/10.

1007/BFb0074242. Prepared for publication by J. Rosemüller.

Urbanik:1967:PDS

[Urb67c] K. Urbanik. Positive definite sequences. Lecture Notes in Mathematics, 44:1–2, 1967. CODEN LNMAA2. ISBN 3-540-03915-5 (print), 3-540-35524-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074243.

Urbanik:1967:PP

[Urb67d] K. Urbanik. Prediction problems. Lecture Notes in Mathematics, 44:40–50, 1967. CODEN LNMAA2. ISBN 3-540-03915-5 (print), 3-540-35524-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074251/.

Urbanik:1967:PCI

[Urb67e] K. Urbanik. Predictors for completely indetermined sequences. Lecture Notes in Mathematics, 44:36–40, 1967. CODEN LN-MAA2. ISBN 3-540-03915-5 (print), 3-540-35524-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074250/.

Urbanik:1967:SR

[Urb67f] K. Urbanik. Series representation. Lecture Notes in Mathematics, 44:11–17, 1967. CODEN LNMAA2. ISBN 3-540-03915-5 (print), 3-540-35524-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074247/.

Urbanik:1967:SDE

[Urb67g] K. Urbanik. Spectral decomposition and ergodic theorem. Lecture Notes in Mathematics, 44:9–11, 1967. CODEN LNMAA2. ISBN 3-540-03915-5 (print), 3-540-35524-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074246/.

Urbanik:1967:SKK

[Urb67h] K. Urbanik. The Szegő–Krein–Kolmogorov theorem. Lecture Notes in Mathematics, 44:24–36, 1967. CODEN LNMAA2. ISBN 3-540-03915-5 (print), 3-540-35524-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074249/.

Urbanik:1967:VVM

[Urb67i] K. Urbanik. Vector valued measures. Lecture Notes in Mathematics, 44:6–9, 1967. CODEN LNMAA2. ISBN 3-540-03915-5 (print), 3-540-35524-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074245/.

Urbanik:1967:WD

[Urb67j] K. Urbanik. Wold decomposition. Lecture Notes in Mathematics, 44:17–24, 1967. CODEN LNMAA2. ISBN 3-540-03915-5 (print), 3-540-35524-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074248/.

vonBliedtner:1968:HGH

[vB68] Jürgen von Bliedtner. Harmonische Gruppen und Huntsche Faltungskerne. (German) [Harmonic groups and Hunt convolution kernels]. Lecture Notes in Mathematics, 69:69–102, 1968. CODEN LNMAA2. ISBN 3-540-04239-3 (print), 3-540-35884-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080476/.

Verdier:1965:DCG

[Ver65] Jean-Louis Verdier. Dualité dans la cohomologie des groupes profinis. (French) [Duality in the cohomology of profinite groups]. Lecture Notes in Mathematics, 5:183–206, 1965. CODEN LN-MAA2. ISBN 3-540-03349-1 (print), 3-662-21576-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-662-21576-0_4/. See

[Ser65b].

Verner:1969:IMI

[Ver69] J. H. Verner. Implicit methods for implicit differential equations. Lecture Notes in Mathematics, 109:261–266, 1969. CODEN LN-MAA2. ISBN 3-540-04628-3 (print), 3-540-36158-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

vonFinckenstein:1969:NTH

[vF69] Karl Graf Finck von Finckenstein. On the numerical treatment of hyperbolic differential equations with constant coefficients, particularly the n-dimensional wave equation. Lecture Notes in Mathematics, 109:154–159, 1969. CODEN LNMAA2. ISBN 3-540-04628-3 (print), 3-540-36158-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060022/.

vonHansen:1968:PHK

[vH68] Wolfhard von Hansen. Potentialtheorie harmonischer Kerne. (German) [Potential theory of harmonic kernels]. Lecture Notes in Mathematics, 69:103–159, 1968. CODEN LNMAA2. ISBN 3-540-04239-3 (print), 3-540-35884-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080477/.

Vitasek:1969:NSS

[Vit69] Emil Vitasek. The numerical stability in solution of differential equations. Lecture Notes in Mathematics, 109:87–111, 1969. CO-DEN LNMAA2. ISBN 3-540-04628-3 (print), 3-540-36158-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0060017/.

Walters:1969:AAU

[Wal69] R. F. C. Walters. An alternative approach to universal algebra. Lecture Notes in Mathematics, 106:64–73, 1969. CODEN LN-MAA2. ISBN 3-540-04625-9 (print), 3-540-36150-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0059142/.

Weil:1967:RDF

[Wei67a] M. Weil. Résolvantes en dualité. (French) [Resolvents in duality]. Lecture Notes in Mathematics, 39:177–189, 1967. CODEN LNMAA2. ISBN 3-540-03910-4 (print), 3-540-34978-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080555/.

Weil:1967:RDT

[Wei67b] M. Weil. Retournement du temps dans les processus markoviens. (French) [Turnaround time in th Markov processes]. Lecture Notes in Mathematics, 39:166–176, 1967. CODEN LNMAA2. ISBN 3-540-03910-4 (print), 3-540-34978-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080554/.

Weinstein:1968:PUM

[Wei68] Joseph Weinstein. (ω_{1}, ω) properties of unions of models. Lecture Notes in Mathematics, 72:265–268, 1968. CODEN LNMAA2. ISBN 3-540-04242-3 (print), 3-540-35900-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079693/.

Wermer:1964:BAF

[Wer64a] John Wermer. Beschränkte analytische Funktionen. (German) [Restricted analytic functions]. Lecture Notes in Mathematics, 1:26– 28, 1964. CODEN LNMAA2. ISBN 3-540-03178-2 (print), 3-540-37091-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099644/.

Wermer:1964:SMG

[Wer64b] John Wermer. Der Satz von Mergelyan. (German) [The Mergelyan set]. Lecture Notes in Mathematics, 1:21–22, 1964. CODEN LN-MAA2. ISBN 3-540-03178-2 (print), 3-540-37091-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0099642.

Wermer:1964:AGA

[Wer64c] John Wermer. Die Algebren P (X). (German) [The P (X) algebras]. Lecture Notes in Mathematics, 1:16–20, 1964. CODEN LNMAA2. ISBN 3-540-03178-2 (print), 3-540-37091-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099641/.

Wermer:1964:KGC

[Wer64d] John Wermer. Die Klassen für P (X). (German) [The classes for P (X)]. Lecture Notes in Mathematics, 1:23–25, 1964. CODEN LNMAA2. ISBN 3-540-03178-2 (print), 3-540-37091-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099643/.

Wermer:1964:RGS

[Wer64e] John Wermer. Die Räume H^{P} (λ). (German) [The space H^{P} (λ)]. Lecture Notes in Mathematics, 1:9–12, 1964. CODEN LN-MAA2. ISBN 3-540-03178-2 (print), 3-540-37091-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0099639/.

Wermer:1964:FMG

[Wer64f] John Wermer. Eine Formel für Masse in A^{⊥}. (German) [A formula for the dimension in A^{⊥}]. Lecture Notes in Mathematics, 1:13–15, 1964. CODEN LNMAA2. ISBN 3-540-03178-2 (print), 3-540-37091-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099640/.

Wermer:1964:EGI

[Wer64g] John Wermer. Einführung. (German) [Introduction]. Lecture Notes in Mathematics, 1:1–3, 1964. CODEN LNMAA2. ISBN 3-540-03178-2 (print), 3-540-37091-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099637/.

Wermer:1964:LZM

[Wer64h] John Wermer. Lebesguesche Zerlegung von Massen aus A^{⊥}. (German) [Lebesgue decomposition of dimensions of A^{⊥}]. Lecture Notes in Mathematics, 1:4–8, 1964. CODEN LNMAA2. ISBN 3-540-03178-2 (print), 3-540-37091-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099638/.

Wermer:1964:LGR

[Wer64i] John Wermer. Literatur. (German) [References]. Lecture Notes in Mathematics, 1:29–30, 1964. CODEN LNMAA2. ISBN 3-540-03178-2 (print), 3-540-37091-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099645.

Wermer:1964:SFA

[Wer64j] John Wermer, editor. Seminar über Funktionen-Algebren: Eidg. Technische Hochschule, Zürich Forschungsinstitut für Mathematik. (German) [Seminar on function algebras: Swiss Federal Institute of Technology, Zurich, Research Institute for Mathematics], volume 1 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1964. CO-DEN LNMAA2. ISBN 3-540-03178-2 (paperback), 3-540-37091-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN

QA3 .L28 no. 1. URL http://link.springer.com/book/10.
1007/BFb0099636.

Werner:1966:ARF

[Wer66a] Helmut Werner. Approximation mit rationalen Funktionen. (German) [Approximation with rational functions]. Lecture Notes in Mathematics, 14:19–29, 1966. CODEN LNMAA2. ISBN 3-540-03597-4 (print), 3-540-37169-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074615/.

Werner:1966:AUG

[Wer66b] Helmut Werner. Asymptotische Untersuchungen. (German) [Asymptotic studies]. Lecture Notes in Mathematics, 14:149– 155, 1966. CODEN LNMAA2. ISBN 3-540-03597-4 (print), 3-540-37169-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074630/.

Werner:1966:BGE

[Wer66c] Helmut Werner. Beispiele. (German) [Examples]. Lecture Notes in Mathematics, 14:95–101, 1966. CODEN LNMAA2. ISBN 3-540-03597-4 (print), 3-540-37169-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074622/.

Werner:1966:CAN

[Wer66d] Helmut Werner. Charakterisierung der Approximierenden in der L_{γ}-Norm bei linearem Ansatz. (German) [Characterization of the approximating in the L_{γ}-norm with linear shape]. Lecture Notes in Mathematics, 14:38–55, 1966. CODEN LNMAA2. ISBN 3-540-03597-4 (print), 3-540-37169-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074617/.

Werner:1966:CTA

[Wer66e] Helmut Werner. Charakterisierung der Tschebyscheff-Approximation. (German) [Characterization of the Chebyshev approximation]. Lecture Notes in Mathematics, 14:85–94, 1966. CODEN LN-MAA2. ISBN 3-540-03597-4 (print), 3-540-37169-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074621/.

Werner:1966:AGA

[Wer66f] Helmut Werner. Das Approximationsproblem. (German) [The approximation problem]. Lecture Notes in Mathematics, 14:16– 18, 1966. CODEN LNMAA2. ISBN 3-540-03597-4 (print), 3-540-37169-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074614/.

Werner:1966:AVA

[Wer66g] Helmut Werner. Das asymptotische Verhalten der Approximationen analytischer Funktionen. (German) [The asymptotic behavior of the approximation of analytic functions]. Lecture Notes in Mathematics, 14:156–162, 1966. CODEN LNMAA2. ISBN 3-540-03597-4 (print), 3-540-37169-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074631/.

Werner:1966:PHG

[Wer66h] Helmut Werner. Das Problem von Haar. (German) [The Haar problem]. Lecture Notes in Mathematics, 14:132–138, 1966. CODEN LNMAA2. ISBN 3-540-03597-4 (print), 3-540-37169-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074627/.

Werner:1966:DLN

[Wer66i] Helmut Werner. Definition des linearen, normierten Raumes, Beispiele. (German) [Definition of the linear normed space, examples]. Lecture Notes in Mathematics, 14:7–15, 1966. CODEN LNMAA2. ISBN 3-540-03597-4 (print), 3-540-37169-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074613/.

Werner:1966:TAM

[Wer66j] Helmut Werner. Die Tschebyscheff-Approximation bei mehreren Veränderlichen. (German) [The Chebyshev approximation on several variables]. Lecture Notes in Mathematics, 14:139–145, 1966. CODEN LNMAA2. ISBN 3-540-03597-4 (print), 3-540-37169-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0074628/.

Werner:1966:DGD

[Wer66k] Helmut Werner. Differenzenquotient. (German) [Difference quotient]. Lecture Notes in Mathematics, 14:72–84, 1966. CODEN LNMAA2. ISBN 3-540-03597-4 (print), 3-540-37169-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074620/.

Werner:1966:DKG

[Wer66l] Helmut Werner. Diskretisierung und Konvergenz. (German) [Discretization and convergence]. Lecture Notes in Mathematics, 14: 126–131, 1966. CODEN LNMAA2. ISBN 3-540-03597-4 (print), 3-540-37169-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074626/.

Werner:1966:EAG

[Wer66m] Helmut Werner. Eindeutigkeit bei L_{1}-Approximation. (German) [Uniqueness in L_{1} approximation]. Lecture Notes in Mathematics, 14:65–71, 1966. CODEN LNMAA2. ISBN 3-540-03597-4 (print), 3-540-37169-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074619/.

Werner:1966:EBG

[Wer66n] Helmut Werner. Einführung und Beispiele. (German) [Introduction and examples]. Lecture Notes in Mathematics, 14:1–6, 1966. CODEN LNMAA2. ISBN 3-540-03597-4 (print), 3-540-37169-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0074612/.

Werner:1966:FM

[Wer66o] Helmut Werner. Front matter. Lecture Notes in Mathematics, 14: i–iv, 1966. CODEN LNMAA2. ISBN 3-540-03597-4 (print), 3-540-37169-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-37169-4/1.pdf.

Werner:1966:NGN

[Wer66p] Helmut Werner. Normalität. (German) [Normality]. Lecture Notes in Mathematics, 14:102–105, 1966. CODEN LNMAA2. ISBN 3-540-03597-4 (print), 3-540-37169-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074623/.

Werner:1966:QFS

[Wer66q] Helmut Werner. Quantitative Fassung der Stetigkeit der Tschebyscheff-Approximation T [f]. (German) [Quantitative version of the continuity of the Chebyshev approximation T [f]]. Lecture Notes in Mathematics, 14:121–125, 1966. CODEN LN-MAA2. ISBN 3-540-03597-4 (print), 3-540-37169-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074625/.

Werner:1966:SAT

[Wer66r] Helmut Werner. Stetige Abhängigkeit der Tschebyscheff-Approximation von der Funktion. (German) [Continuous dependence of the Chebyshev approximation of functions]. Lecture Notes in Mathematics, 14:106–120, 1966. CODEN LNMAA2. ISBN 3-540-03597-4 (print), 3-540-37169-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074624/.

Werner:1966:SKN

[Wer66s] Helmut Werner. Strikt konvexe Normen und Eindeutigkeit des linearen Approximationsproblems. (German) [Strictly convex norms and uniqueness of the linear approximation problem]. Lecture Notes in Mathematics, 14:30–37, 1966. CODEN LNMAA2. ISBN 3-540-03597-4 (print), 3-540-37169-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074616/.

Werner:1966:TAL

[Wer66t] Helmut Werner. Tschebyscheff-Approximation und lineare (konvexe) Programmierung. (German) [Chebyshev approximation and linear (convex) programming]. Lecture Notes in Mathematics, 14: 146–148, 1966. CODEN LNMAA2. ISBN 3-540-03597-4 (print), 3-540-37169-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074629/.

Werner:1966:TSG

[Wer66u] Helmut Werner. Tschebyscheff-Systeme. (German) [Chebyshev systems]. Lecture Notes in Mathematics, 14:56–64, 1966. CO-DEN LNMAA2. ISBN 3-540-03597-4 (print), 3-540-37169-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0074618/.

Werner:1966:VAU

[Wer66v] Helmut Werner, editor. Vorlesung über Approximationstheorie: Universität Münster Institut für Numerische und Instrumentelle Mathematik Sommer-Semester 1964. (German) [Lectures on Approximation Theory: University of Muenster Institute for Computational and Instrumental Mathematics Summer semester 1964], volume 14 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1966. CODEN LNMAA2. ISBN 3-540-03597-4 (paperback), 3-540-37169-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

LCCN QA3 .L28 no. 14. URL http://link.springer.com/book/
10.1007/BFb0074611.

Wermer:1968:AUF

[Wer68] John Wermer. Approximation uniforme dans C^{n}. (French) [Uniform approximation in C^{n}]. Lecture Notes in Mathematics, 71: 165–166, 1968. CODEN LNMAA2. ISBN 3-540-04241-5 (print), 3-540-35889-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080106.

Wermer:1969:PA

[Wer69a] J. Wermer. Polynomial approximation. Lecture Notes in Mathematics, 118:133–162, 1969. CODEN LNMAA2. ISBN 3-540-04907-X (print), 3-540-36246-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060256/.

Wermer:1969:BAU

[Wer69b] John Wermer. Banach algebras and uniform approximation. Lecture Notes in Mathematics, 103:30–43, 1969. CODEN LN-MAA2. ISBN 3-540-04622-4 (print), 3-540-36145-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0099983/.

Widlund:1969:ESP

[Wid69] Olaf B. Widlund. On the effects of scaling of the Peaceman– Rachford method. Lecture Notes in Mathematics, 109:113–132, 1969. CODEN LNMAA2. ISBN 3-540-04628-3 (print), 3-540-36158-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060018/.

Wilansky:1967:CT

[Wil67a] Albert Wilansky. Comparison of topologies. Lecture Notes in Mathematics, 45:1–23, 1967. CODEN LNMAA2. ISBN 3-540-03916-3 (print), 3-540-35525-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079845/.

Wilansky:1967:CCS

[Wil67b] Albert Wilansky. Coregular and conull F K spaces. Lecture Notes in Mathematics, 45:89–100, 1967. CODEN LNMAA2. ISBN 3-540-03916-3 (print), 3-540-35525-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079851/.

Wilansky:1967:FBS

[Wil67c] Albert Wilansky. FH and BH spaces. Lecture Notes in Mathematics, 45:62–76, 1967. CODEN LNMAA2. ISBN 3-540-03916-3 (print), 3-540-35525-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079848/.

Wilansky:1967:GCT

[Wil67d] Albert Wilansky. Garling’s completeness theorem. Lecture Notes in Mathematics, 45:77–82, 1967. CODEN LNMAA2. ISBN 3-540-03916-3 (print), 3-540-35525-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079849/.

Wilansky:1967:IL

[Wil67e] Albert Wilansky. Inductive limits. Lecture Notes in Mathematics, 45:24–43, 1967. CODEN LNMAA2. ISBN 3-540-03916-3 (print), 3-540-35525-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079846/.

Wilansky:1967:MTT

[Wil67f] Albert Wilansky. Mixed topologies and two norm spaces. Lecture Notes in Mathematics, 45:83–88, 1967. CODEN LNMAA2. ISBN 3-540-03916-3 (print), 3-540-35525-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079850/.

Wilansky:1967:SPL

[Wil67g] Albert Wilansky. Some properties of linear topological spaces. Lecture Notes in Mathematics, 45:44–61, 1967. CODEN LN-MAA2. ISBN 3-540-03916-3 (print), 3-540-35525-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079847/.

Wilansky:1967:TFA

[Wil67h] Albert Wilansky. Topics in Functional Analysis, volume 45 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. CODEN LN-MAA2. ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 102

pp. LCCN QA3 .L28 no. 45. URL http://link.springer.com/
book/10.1007/BFb0079844. Notes by W. D. Laverell.

Wloka:1969:EGE

[Wlo69a] Joseph Wloka. Die Einbettungstheoreme. (German) [The embedding theorem]. Lecture Notes in Mathematics, 82:1–34, 1969. CO-DEN LNMAA2. ISBN 3-540-04250-4 (print), 3-540-35931-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0075162/.

Wloka:1969:GVF

[Wlo69b] Joseph Wloka. Grundräume und verallgemeinerte Funktionen. (German) [Basic spaces and generalized functions], volume 82 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CODEN LN-MAA2. ISBN 3-540-04250-4 (paperback), 3-540-35931-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 131 pp.

LCCN QA3 .L28 no. 82. URL http://link.springer.com/book/
10.1007/BFb0075161.

Wloka:1969:IRGb

[Wlo69c] Joseph Wloka. Induktive S(m)-Räume. (German) [Inductive S(m) spaces]. Lecture Notes in Mathematics, 82:124–129, 1969. CODEN LNMAA2. ISBN 3-540-04250-4 (print), 3-540-35931-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075169/.

Wloka:1969:IRGa

[Wlo69d] Joseph Wloka. Induktive A-Räume. (German) [Inductive A spaces]. Lecture Notes in Mathematics, 82:97–124, 1969. CO-DEN LNMAA2. ISBN 3-540-04250-4 (print), 3-540-35931-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0075168/.

Wloka:1969:ORG

[Wlo69e] Joseph Wloka. Operationen auf den W -, A- und S-Räumen. (German) [Operations on the W -, A- and S-spaces]. Lecture Notes in Mathematics, 82:34–51, 1969. CODEN LNMAA2. ISBN 3-540-04250-4 (print), 3-540-35931-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075163/.

Wloka:1969:PRGc

[Wlo69f] Joseph Wloka. Projektive S(m)-Räume. (German) [Projective S(m) spaces]. Lecture Notes in Mathematics, 82:89–96, 1969. CO-DEN LNMAA2. ISBN 3-540-04250-4 (print), 3-540-35931-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0075167/.

Wloka:1969:PRGb

[Wlo69g] Joseph Wloka. Projektive A-Räume. (German) [Projective A spaces]. Lecture Notes in Mathematics, 82:76–89, 1969. CODEN LNMAA2. ISBN 3-540-04250-4 (print), 3-540-35931-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075166/.

Wloka:1969:PGG

[Wlo69h] Joseph Wloka. Projektive Grundräume. (German) [Projective basis spaces]. Lecture Notes in Mathematics, 82:51–57, 1969. CO-DEN LNMAA2. ISBN 3-540-04250-4 (print), 3-540-35931-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0075164/.

Wloka:1969:PRGa

[Wlo69i] Joseph Wloka. Projektive W -Räume. (German) [Projective W spaces]. Lecture Notes in Mathematics, 82:57–75, 1969. CODEN LNMAA2. ISBN 3-540-04250-4 (print), 3-540-35931-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075165/.

Weiss:1969:MPE

[WW69] L. Weiss and J. Wolfowitz. Maximum probability estimators with a general loss function. Lecture Notes in Mathematics, 89:232– 256, 1969. CODEN LNMAA2. ISBN 3-540-04608-9 (print), 3-540-36098-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079130/.

Maeda:1968:KBR

[yMO68] Fumi yuki Maeda and Makoto Otsuka, editors. Kuramochi Boundaries of Riemann Surfaces: A Symposium held at the Research Institute for Mathematical Sciences, Kyoto University, October [11– 13,] 1965, volume 58 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. CODEN LNMAA2. ISBN 3-540-04228-8 (paperback), 3-540-35859-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 58. URL http://link.springer.

com/book/10.1007/BFb0098441.

Yorke:1968:ASF

[Yor68a] James A. Yorke. Asymptotic stability for functional differential equations. Lecture Notes in Mathematics, 60:65–75, 1968. CO-DEN LNMAA2. ISBN 3-540-04230-X (print), 3-540-35862-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0077129/.

Yorke:1968:ELS

[Yor68b] James A. Yorke. Extending Liapunov’s second method to non-Lipschitz Liapunov functions. Lecture Notes in Mathematics, 60: 31–36, 1968. CODEN LNMAA2. ISBN 3-540-04230-X (print), 3-540-35862-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077125/.

Yorke:1968:ECI

[Yor68c] James A. Yorke. An extension of Chetaev’s instability theorem using invariant sets and an example. Lecture Notes in Mathematics, 60:100–106, 1968. CODEN LNMAA2. ISBN 3-540-04230-X (print), 3-540-35862-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077133/.

Yorke:1968:LFE

[Yor68d] James A. Yorke. Liapunov functions and the existence of solutions tending to 0. Lecture Notes in Mathematics, 60:48–54, 1968. CODEN LNMAA2. ISBN 3-540-04230-X (print), 3-540-35862-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077127/.

Yoshizawa:1968:SEP

[Yos68] Taro Yoshizawa. Stability and existence of periodic and almost periodic solutions. Lecture Notes in Mathematics, 60:20–25, 1968. CODEN LNMAA2. ISBN 3-540-04230-X (print), 3-540-35862-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077123/.

Zalcman:1968:ACb

[Zal68a] Lawrence Zalcman. AC capacity. Lecture Notes in Mathematics, 50:77–85, 1968. CODEN LNMAA2. ISBN 3-540-04220-2 (print), 3-540-35825-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070667/.

Zalcman:1968:ACa

[Zal68b] Lawrence Zalcman. Analytic capacity. Lecture Notes in Mathematics, 50:11–22, 1968. CODEN LNMAA2. ISBN 3-540-04220-2 (print), 3-540-35825-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070660/.

Zalcman:1968:ACR

[Zal68c] Lawrence Zalcman. Analytic Capacity and Rational Approximation, volume 50 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. CODEN LNMAA2. ISBN 3-540-04220-2 (paperback), 3-540-35825-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 155 pp. LCCN QA3 .L28 no. 50. URL http://link.

springer.com/book/10.1007/BFb0070657.

Zalcman:1968:A

[Zal68d] Lawrence Zalcman. Applications. Lecture Notes in Mathematics, 50:57–64, 1968. CODEN LNMAA2. ISBN 3-540-04220-2 (print), 3-540-35825-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070665/.

Zalcman:1968:AVT

[Zal68e] Lawrence Zalcman. Applications of Vitushkin’s theorem. Lecture Notes in Mathematics, 50:108–111, 1968. CODEN LNMAA2. ISBN 3-540-04220-2 (print), 3-540-35825-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070670/.

Zalcman:1968:EIG

[Zal68f] Lawrence Zalcman. Estimates for integrals. Lecture Notes in Mathematics, 50:27–44, 1968. CODEN LNMAA2. ISBN 3-540-04220-2 (print), 3-540-35825-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070662/.

Zalcman:1968:FAM

[Zal68g] Lawrence Zalcman. Function algebra methods. Lecture Notes in Mathematics, 50:117–129, 1968. CODEN LNMAA2. ISBN 3-540-04220-2 (print), 3-540-35825-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070672/.

Zalcman:1968:FR

[Zal68h] Lawrence Zalcman. Further results. Lecture Notes in Mathematics, 50:52–56, 1968. CODEN LNMAA2. ISBN 3-540-04220-2 (print), 3-540-35825-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070664/.

Zalcman:1968:GC

[Zal68i] Lawrence Zalcman. Geometric conditions. Lecture Notes in Mathematics, 50:112–116, 1968. CODEN LNMAA2. ISBN 3-540-04220-2 (print), 3-540-35825-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070671/.

Zalcman:1968:I

[Zal68j] Lawrence Zalcman. Introduction. Lecture Notes in Mathematics, 50:1–3, 1968. CODEN LNMAA2. ISBN 3-540-04220-2 (print), 3-540-35825-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070658/.

Zalcman:1968:MT

[Zal68k] Lawrence Zalcman. Melnikov’s theorem. Lecture Notes in Mathematics, 50:45–51, 1968. CODEN LNMAA2. ISBN 3-540-04220-2 (print), 3-540-35825-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070663/.

Zalcman:1968:PP

[Zal68l] Lawrence Zalcman. Peak points. Lecture Notes in Mathematics, 50:4–10, 1968. CODEN LNMAA2. ISBN 3-540-04220-2 (print), 3-540-35825-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070659/.

Zalcman:1968:PRA

[Zal68m] Lawrence Zalcman. The problem of rational approximation. Lecture Notes in Mathematics, 50:65–76, 1968. CODEN LN-MAA2. ISBN 3-540-04220-2 (print), 3-540-35825-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0070666/.

Zalcman:1968:SA

[Zal68n] Lawrence Zalcman. A scheme for approximation. Lecture Notes in Mathematics, 50:86–99, 1968. CODEN LNMAA2. ISBN 3-540-04220-2 (print), 3-540-35825-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070668/.

Zalcman:1968:SOQ

[Zal68o] Lawrence Zalcman. Some open questions. Lecture Notes in Mathematics, 50:130–131, 1968. CODEN LNMAA2. ISBN 3-540-04220-2 (print), 3-540-35825-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070673.

Zalcman:1968:SUF

[Zal68p] Lawrence Zalcman. Some useful facts. Lecture Notes in Mathematics, 50:23–26, 1968. CODEN LNMAA2. ISBN 3-540-04220-2 (print), 3-540-35825-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070661/.

Zalcman:1968:VT

[Zal68q] Lawrence Zalcman. Vitushkin’s theorem. Lecture Notes in Mathematics, 50:100–107, 1968. CODEN LNMAA2. ISBN 3-540-04220-2 (print), 3-540-35825-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070669/.

Zariski:1969:ITAb

[Zar69a] Oscar Zariski, editor. An Introduction to the Theory of Algebraic Surfaces, volume 83 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. CODEN LNMAA2. ISBN 3-540-04602-X (paperback), 3-540-36092-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 83; QA3 .L28 no. 83 QA571. URL

http://link.springer.com/book/10.1007/BFb0082246. Notes

by James Alan Cohn, Harvard University, 1957–58.

Zariski:1969:AG

[Zar69b] Oscar Zariski. The arithemetic genus. Lecture Notes in Mathematics, 83:49–51, 1969. CODEN LNMAA2. ISBN 3-540-04602-X (print), 3-540-36092-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082250/.

Zariski:1969:CSV

[Zar69c] Oscar Zariski. The canonical system on a variety V. Lecture Notes in Mathematics, 83:29–34, 1969. CODEN LNMAA2. ISBN 3-540-04602-X (print), 3-540-36092-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082263/.

Zariski:1969:CRI

[Zar69d] Oscar Zariski. Coordinate rings of irreducible varieties. Lecture Notes in Mathematics, 83:3–4, 1969. CODEN LNMAA2. ISBN 3-540-04602-X (print), 3-540-36092-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082256.

Zariski:1969:DGD

[Zar69e] Oscar Zariski. Differentials. Lecture Notes in Mathematics, 83: 24–28, 1969. CODEN LNMAA2. ISBN 3-540-04602-X (print), 3-540-36092-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082262/.

Zariski:1969:DCN

[Zar69f] Oscar Zariski. Divisorial cycles on a normal projective variety V/k(dim(V) = r ≥ 1). Lecture Notes in Mathematics, 83:6– 11, 1969. CODEN LNMAA2. ISBN 3-540-04602-X (print), 3-540-36092-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082258/.

Zariski:1969:DAV

[Zar69g] Oscar Zariski. Divisors on an arbitrary variety V. Lecture Notes in Mathematics, 83:15–19, 1969. CODEN LNMAA2. ISBN 3-540-04602-X (print), 3-540-36092-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082260/.

Zariski:1969:HCF

[Zar69h] Oscar Zariski. The Hilbert characteristic function and the arithmetic genus of a variety. Lecture Notes in Mathematics, 83:63– 71, 1969. CODEN LNMAA2. ISBN 3-540-04602-X (print), 3-540-36092-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082252/.

Zariski:1969:HNH

[Zar69i] Oscar Zariski. Homogeneous and non-homogeneous point coordinates. Lecture Notes in Mathematics, 83:1–3, 1969. CODEN LNMAA2. ISBN 3-540-04602-X (print), 3-540-36092-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082248/.

Zariski:1969:ITAa

[Zar69j] Oscar Zariski. Intersection theory on algebraic surfaces (k algebraically closed). Lecture Notes in Mathematics, 83:19–23, 1969. CODEN LNMAA2. ISBN 3-540-04602-X (print), 3-540-36092-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082261/.

Zariski:1969:LS

[Zar69k] Oscar Zariski. Linear systems. Lecture Notes in Mathematics, 83:11–15, 1969. CODEN LNMAA2. ISBN 3-540-04602-X (print), 3-540-36092-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082259/.

Zariski:1969:NV

[Zar69l] Oscar Zariski. Normal varieties. Lecture Notes in Mathematics, 83:4–5, 1969. CODEN LNMAA2. ISBN 3-540-04602-X (print), 3-540-36092-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082257.

Zariski:1969:NCS

[Zar69m] Oscar Zariski. Normalization and complete systems. Lecture Notes in Mathematics, 83:52–62, 1969. CODEN LNMAA2. ISBN 3-540-04602-X (print), 3-540-36092-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082251/.

Zariski:1969:PFL

[Zar69n] Oscar Zariski. Proof of the fundamental lemma. Lecture Notes in Mathematics, 83:95–100, 1969. CODEN LNMAA2. ISBN 3-540-04602-X (print), 3-540-36092-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082255/.

Zariski:1969:RRT

[Zar69o] Oscar Zariski. The Riemann–Roch theorem. Lecture Notes in Mathematics, 83:72–80, 1969. CODEN LNMAA2. ISBN 3-540-04602-X (print), 3-540-36092-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082253/.

Zariski:1969:SPG

[Zar69p] Oscar Zariski. Subadjoint polynomials. Lecture Notes in Mathematics, 83:80–95, 1969. CODEN LNMAA2. ISBN 3-540-04602-X (print), 3-540-36092-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082254/.

Zariski:1969:TD

[Zar69q] Oscar Zariski. Trace of a differential. Lecture Notes in Mathematics, 83:34–48, 1969. CODEN LNMAA2. ISBN 3-540-04602-X (print), 3-540-36092-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082249/.

Zisman:1969:DCP

[Zis69] M. Zisman. Derived category and Poincaré duality. Lecture Notes in Mathematics, 86:205–216, 1969. CODEN LNMAA2. ISBN 3-540-04605-4 (print), 3-540-36095-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079392/.